A Complete Bibliography of Publications in Linear Algebra and its Applications: 2020–2024

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 April 2024 Version 1.43

Title word cross-reference

(−1, 1) [Obo24]. (0, 1) [Loe21, Shi21c, DGS20, EJS24, GS23b]. (0, ±1) [DGS23]. (1, −1) [MSX22]. (α, β) [SWW22]. (Au, A^{r}v) = (u, v) [BKS^{+}21]. (Au, v) = (u, A^{r}v) [BKS^{+}21]. (b, c) [Dra23]. (∞, ∞) [GOZ21]. (k + 1) [OdFVM21]. (n + k) × n [BDNO20]. (n, m) [GZZ21]. (p, q) [AdST22, MMV22]. (t, t) [HW21b]. ∗ [BN23]. −1 [WLL23, ZW22, BD21b]. {−1, +1} [AdL21]. −3 [LLL22]. 0 [ABW22, BD21b, CHY20, HL20, Lyu20]. {0, 1} [WW22]. 1 [ABW22, BD21b, CHY20, HL20, KLR22, Lyu20, ZLLZ22]. 1/2 [GK20, WQP23]. 12 ≤ α ≤ 1 [LQ21]. 2 [Alf20b, BS21a, Ban22, BEF20, BDV21, GLT23, LNN^{+}20, SV21, XZWT20, ZG22]. 2 + 2 [LWB20]. 2^{w} × 2^{w} [VB20]. 2θ(G) + ρ(G) − 1 [ZW22]. 2 × 2 [BDK21, WZL21, Xie21]. 3 [BC24, IKPR20, MV24, SS23c, ZZJ23, Zhe21]. 3 × 3 [CLW22, CdSV22]. 4 [CN23a, CN23b, Ere21, FN23, TCC21]. 4 × 4 [CRSY22, GHL^{+}24]. 5 [JMP21b, TCC21]. 5 × 5 [LV22b]. 6 [ZWF20]. [a, b] [WZC23]. ^{⋆} [CAEH^{+}22]. _{α} [CLM22]. A [BKPS23, CLM22, Djo23]. A(R, S) [FdCP20]. a_{2} − c_{3} [IKPR20]. A_{α} [CT23, CLM20, FW22, HLS22, LQ21, LTYZ23, LY23, LMG20, LDSS20, WWT20, WFGT20, XWT20, YBW24, ZZ21, ZHW21]. α [CDG21, DLVV21, LL21c, TCC21, WSW20, XWT20]. α(M) [FQ20]. ∗ [AG15, AG20]. AX^{ε} + BX^{δ} = 0 [DL23].

AXA = XAX [Djo23]. b [ZL23c]. ⊗ [Mor22]. c [GS21a, vdMvSW24]. C^{∗} [CMM20]. C^{n} [EJJL20]. C^{⋆} [BH24, CN20a, CDR23, EE23, FPS22, Ham22, MS24, Sin21, XY20b]. C_{2} [RT21]. C_{3} [LMX23]. C_{3}^{−} [WHL24]. C_{5} [MLH22]. C_{6} [MLH22]. T [CN20a]. D [KP21, Kus23, Zha21, vED22]. D_{n} [Sim21, WWG22, Wan23]. det^{S^{3}} [LS24a]. ℓ^{∞} [FISW21]. ℓ^{p} [HJ24]. η(G, σ) = |V (G)| − 2m(G) + 2c(G) − 1 [LW21d]. F_{q^{6}} × F_{q^{6}} [BCM21]. G [CDD21, CL22, GG21, AAMR20, DY24, ZWT21]. G^{⋆}JG = J [EJ23]. Gr_{3}(C^{6}) [HPSZ21]. H [AMO24, CGP22, CS21, Gor24, GvRR21a, GvRR^{+}21b, SMA21, WES21, vRvSTT21, MMV21]. H(3, 4) [DG20]. i [GZZ21]. J_{q}(N, D) [LIW20]. K [FSRVSST22, LL21d, AdLN23, BCE21, Fio20, HJBKR24, HL20, HLS21b, KLSP24, Le22, MP20a, OdFVM21, SSP22, SWW22, SSvW20, VTBC20, WSW20, Yaz21, Zhe20, ZBWL20]. K_{1,t} [ZL22]. K_{2,t} [LLW23]. K_{3,3}/K_{2,4} [WCF21]. K_{5}^{−} [Wan24]. {K_{k+1}, L_{s}} [ZY24]. L [BM22, JLW22]. l^{∞} [Yu20]. L_{+}^{n} [HS23b]. L_{1} [NNSS20]. λ [Mif21]. ⌈g⌉/2 [DY24]. ⌈g⌉/2 − 1 [DY24]. ≤ 4 [Com21]. LU [LY20]. M [CS21, DMS20, KLS21, Men22, Beh21, BCG24, EJJL20, GvRR^{+}21b, JJS20a, SZH22, YZY24]. M_{n} [BSC23]. M_{n}(E) [Fid22]. M_{n}(F_{2}) [Šte22b]. CP^{n} [SS22]. DL(P) [DN23]. F_{q}^{n} [AS20a]. Q [GvRR^{+}23]. T [Kha24]. Z [dAGK21]. Z_{n} [CPS20]. AN [GOUY23]. B(H) [AGK^{+}21]. CX ⊗ BH [EE23]. D_{α} [DPR20]. L(C^{n}) [MRZ20]. L(H) [Wój22b]. L(l_{2}^{2}) [Wój22b]. ∫↕(2, C) [JL22]. ∫≀_{4} [Mor22]. L [BES21]. N [CT20, CK21, JKX20, BDNO20, BCE21, CW23a, CW23b, HHS21, LP21, SZH22, ZWT21]. n(G) − g(G) − 1 [CL22]. n − 3 [TWW20, TW21]. n − g [ZWT21]. n = 2, 3, 4 [SZH22]. n ≡ 0 (mod 4) [Wan23]. n × n [DZCY22, SSvW20]. O(n) [TP23]. O(q, k) [HHS20b]. ω [MPS24]. Ω(2^{n}/n) [Sha22]. P [CK21, MC21b, ME21, AAFM23, AdST22, BM22, CI21, KLLW21, LL21a, dSP20a, ZKLS20]. p(λ) [GvRR^{+}23]. p(X) = A [GvRR^{+}23]. ±1 [HT23]. ψ_{S} [AP20]. PSL(3, C) [BCNS22]. Q [HLN24, JLW22, LMX23, OPPS22, Sud22, Tan23, ZL22, AKVS23, BPS23, CKOS22, Hom20, KAGH24, Lee20, LHZ21, Ner20, ZZ20]. q + 1 [Hom20]. q + 2 [Hom20]. Q_{0} [CESS21, SNK22]. r [LC20]. R^{d} [DO22]. R_{0} [ME21]. ρ [KZ24]. S [DP23, AK20, BNS22]. S_{n} [HMS21]. s ≥ 3 [BNS22]. sl_{n} [Bar22]. SR [FRS21]. ⋆ [Ham22]. SU(2) [Soł20]. T [MQW20, BNS22, FdCP20, WXY24]. t ≥ 2s − 2 [BNS22]. D̃_{n} [MP23b]. × [CL20].

U_{q}(ˆ∫↕^{2}) [LIW20]. W [ME21]. W (B_{n})

[Ces20]. x^{q} + x^{q^{3}} + cx^{q^{5}} ∈ F_{q^{6}}[x] [MMZ20]. z [DLVV21]. Z_{2} [FGGG24]. Z_{q} [GFD21].

*congruence [Sta21].

{-1 [AdLN23]. -actions [Gor24]. -algebra [CDR23]. -algebras [CMM20, CN20a, CAEH^{+}22, EE23, FPS22, Ham22, MS24]. -analogue [MMV22, Ner20]. -analysis [AKVS23]. -Bernstein [MMV21]. -Bures [DLVV21]. -by- [CN23a, CN23b]. -chromatic [Ere21]. -circulant [AAMR20]. -commutant [BPS23]. -commuting [BPS23, LP21]. -complex [EJJL20]. -connected [HLS21b, ZG22, ZBWL20]. -contractivity [BCE21]. -Core [Mif21]. -covered [WZC23]. -cut [AdLN23]. -cycle [TCC21, ZWF20]. -deformed [CKOS22]. -designs [BNS22, HW21b]. -dimensional [BS21a, Ban22, FN23, GLT23, LNN^{+}20, SZH22, SS23c, ZZJ23]. -distance [BNS22, KAGH24]. -domination [AAFM23]. -double [AdST22, ZZ20]. -eigenvalue [CLM20, WFGT20]. -eigenvalues [Zha21]. -extreme [BH24]. -factor [ZLLZ22]. -factors [ZL23c]. -forest [BEF20]. -frames [LL21d]. -free [BES21, LMX23, MLH22, WHL24, Wan24, ZY24]. [?]free/Min:2022:SUB. -functions [Beh21]. -gain [Kha24]. -graded [GFD21, dAGK21, RT21]. -gradings [FGGG24]. -graph [GZZ21]. -graphs [LC20]. -Hahn [Lee20]. -hypercontractivity [JKX20]. -hyponormality [BCE21]. -idempotent [HL20]. -independence [Fio20]. -index [TCC21, CT23, JLW22, LTYZ23, LMX23, ZL22]. -infinity [WES21]. -integral [OPPS22]. -intersecting [WXY24]. -invariant [GG21, TP23]. -inverses [Dra23]. -Isometric [JJS20a]. -isometries [BCG24]. -join [CGP22, SMA21]. -labelling [SWW22]. -line [OdFVM21]. -linear [FSRVSST22]. -Majorization [vED22]. -matrices [DGS20, DGS23, EJS24, GS23b, MSX22, WW22, CS21, CT20, CK21, KLS21, Men22, Tan23]. -matrix [BM22, DMS20, LDSS20, SNK22]. -minimizer [HLN24]. -minor [WCF21, ZL22]. -minor-free [LLW23]. -module [XY20b]. -nearest [Yu20]. -nonnegative [vRvSTT21]. -norm [BM22, KLSP24, NNSS20]. -normal [LL21c]. -norms [CI21, FISW21]. -numerical [vdMvSW24]. -operator [KZ24]. -orbit [HMS21]. -order [SZH22]. -orthogonal [AP20]. -palindromic [BN23]. -phases [CDD21]. -polynomial [Sud22]. -positive [Le22]. -power [LL21a]. -product [MQW20]. -property [GOUY23, MC21b]. -rings [Ham22]. -Schröder [YZY24]. -selfadjoint [AMO24, GvRR^{+}21b]. -semigroups [AG15, AG20]. -semipositive [HS23b]. -skeleton [KLR22]. -smooth [MP20a]. -spanning [HJBKR24]. -specified [AK20]. -spectra [JMP21b, DP23, Zha21]. -spectral [GS21a, KLLW21, WSW20, ZKLS20, BKPS23, DPR20, FW22, HLS22, LQ21, OPPS22, ZZ21, ZHW21]. -spread [LMG20]. -stability [KP21]. -stable [Kus23]. -step [BDV21]. -sun [AdST22].

-symmetries [AP20]. -tensors [BC24]. -term [FdCP20]. -th [CW23a, CW23b, GZZ21, GvRR^{+}21b, dSP20a]. -trees [BEF20, OdFVM21]. -Uniform [MV24, SSP22, SWW22, WSW20, Zhe20, Zhe21]. -unitary [GvRR21a].

1} [AdLN23, Ano22-57, Ano22-69, Ano22-65, Ano22-59, Ano22-52, Ano22-58, Ano22-63, Ano22-60, Ano23-52, Ano23-53, Ano23-61, Ano23-63, Ano23-51, Ano23-54, Ano23-49, Ano23-65, Ano23-68, Ano23-69, Ano23-58, Ano24-27, Ano24-38, Ano24-39, Ano24-30, Ano24-36, Ano24-28, Ano24-33]. 1-350 [Ano23-72]. 15 [Ano22-62, Ano22-49, Ano22-66, Ano22-61, Ano22-64, Ano22-53, Ano22-72, Ano22-70, Ano22-71, Ano23-57, Ano23-50, Ano23-62, Ano23-71, Ano23-60, Ano23-56, Ano23-55, Ano23-59, Ano23-66, Ano23-64, Ano23-67, Ano24-34, Ano24-29, Ano24-31, Ano24-32, Ano24-37].

2-Cayley [WFY24]. 2-Steiner [AS22]. 2022 [Ano22-62, Ano22-57, Ano22-49, Ano22-69, Ano22-66, Ano22-65, Ano22-61, Ano22-59, Ano22-64, Ano22-52, Ano22-53, Ano22-58, Ano22-72, Ano22-63, Ano22-70, Ano22-60, Ano22-71]. 2023 [Ano23-52, Ano23-57, Ano23-53, Ano23-50, Ano23-61, Ano23-62, Ano23-63, Ano23-71, Ano23-60, Ano23-51, Ano23-56, Ano23-54, Ano23-55, Ano23-49, Ano23-59, Ano23-65, Ano23-66, Ano23-68, Ano23-64, Ano23-69, Ano23-58, Ano23-67]. 2024 [Ano24-27, Ano24-38, Ano24-34, Ano24-39, Ano24-29, Ano24-30, Ano24-31, Ano24-36, Ano24-28, Ano24-32, Ano24-33, Ano24-37]. 22nd [AHKT21].

436 [CdGS20]. 479 [BKSP23]. 486 [AG20].

520 [FJS21]. 532 [Hou20]. 553 [AAH^{+}22]. 561 [PW20]. 577 [Ali20]. 582 [KAPS20]. 585 [AD20a]. 586 [JLM20a]. 594 [Sło21a].

595 [MH21].

610 [FHMP22]. 611 [SD21a]. 630 [BP23b]. 646 [BDHS23b].

ABC [LW20b, YD21, GLHNW20]. Abel [CN21a, CN22, CN23b]. abelian [BF24a, BL23, CW22, CK23a, KB22, MDH23, BFLT23, CL21]. absolute [StW22]. absorbing [CR21]. abstract [AFS^{+}23, BDN24]. accelerated [OPA23]. Accelerating [BHD^{+}21]. Acceptability [Pan23]. Accuracy [GNR20]. Accurate [MPR22, MPR24, MMV22, OP21, Yan21]. achieve [Sha22]. acting [BMR23]. actions [CXZ22, Gor24]. acyclic [DdF23]. Addendum [AAH^{+}22]. addition [Arl20]. additive [GS20b, LP21, XX20]. ADI [Rob20]. Adjacency [CDVPT22, CL24, Coo20, HS23a, LL24a, AdL21, BGGA22, BS23a, CGP22, Du23, LW21b, WZFW23, XW21]. adjoint [DH21, GLT21, PPŻ21, Pan22, PPŻ23, SZ20]. adjointable [AG15, AG20]. adjointness [AC21]. adjugate [CZ22b, RMG24]. Admissibility [NS22]. admitting [BL23]. Affine [BS20f, Rub22, BFO23, Hom20, LIW20, Yan22]. affine-plane [Hom20]. after [Bar20, JST21]. Algebra [AAH^{+}22, AD20a, Ali20, AG20, BP23b, BDHS23b, BKSP23, CdGS20, FJS21, FHMP22, Hou20, JLM20a, KAPS20, MH21, PW20, Sło21a, SD21a, AOCM20, BK22, BE20, CW23a, CW23b, CCS22, CFR23, CDR23, CdSV22, DRG24, FM20a, FGGG24, FP22, FT23, GFD21, dAGK21, Jon21, KS24, KZPH22, LIW20, LP22, MPS22, Mor22, Mur22, PGST23, Sch24, SlW22, Vit21, WFSS20, Yas23, AHKT21]. Algebraic [AM21, HS23b, Tar20, AM23b, BEHK22, BF24b, Bri22, CF24, Das22, KS22a, KW20, MS24, BCP22]. algebraically [BHBO21]. algebras [ABCO22, AM21, AEGV21, Alt22, AOCM20, ABC^{+}22, BS20b, BB22, BFLT23, BF24a, BRL23, BCS24, Bez22, BDHS22, BDHS23b, BR24, Bla20, BDvW24, BS23a, BE20, BDV21, CGG^{+}21, CMM20, CWW24, CN20a, CLW22, CdGS12, CdGS20, CZ20a, CMR20, CDG22b, Cos20b, Cra21, CMR21, CAEH^{+}22, DdR20, DP21, DGdSS23, DHR20, DM24, EE23, Eld22, EGL21, FK23, FPS22, FGDY20, FKM22, FN23, GLNW24, GK22a, GT23, GKS21, GS20a, Gow20, GMP22, Ham22, HKS23, JJL20b, KK23a, KK23b, KY21, Kol21, KLL22, Kye23, LM22a, LNN^{+}20, LS20b, MT21a, Mar21, Mig21, MRT20, MDH23, MS24, Rai23, Sas21, Sos23, SZ22b, Uru22, Vol23, WZL21, WZ23b, XY20a, XZ21, Yan22, ZG21, dMdSS23]. algorithm [Bär21a, BC22, BT20a, BHD^{+}21, CT24, CLLK21, CZ20b, FM22a, KPP21, NPkT21, YOV21]. Algorithmic [CT20, VV22]. Algorithms [Boj21, BF24b, CLRT22, HNS21, fLLV20, LY20, Sch23, SZ21, SZ22a, VD22]. ALIF [BC22]. allow [BFF^{+}23b, BFF^{+}23a, Das22]. allowed [BBC^{+}22b]. Almost [BDV21, AK23, BF24a, CDdM22, CD23a, CK21, DYLL23, FMR20, Roc20, Sło21b]. almost-principal [FMR20]. almost-Riordan [AK23, Sło21b]. Alpha [Min22]. Alternating [OJQ20, OQ22, EN22, KKS21, RZ24]. alternative [GvRR21a]. Aluthge [GOUY23, Yam21]. amenable [CN20a]. among [DJ23, LW21b]. amplitudes [JS22]. analogue [MMV22, Ner20]. analysis [AFSS20, AK20, AKVS23, AV23, BEFSC22, BDFF23, CW21, FKM20, GM20, GFW24, HH23a, LS24b, LHZ21, MMV21, Mis20, RRvW24, SDTR23, Szc21, WvdVvdH24]. analytic [AV23]. Analyticity [LL21b]. Anderson [BKPS23]. Ando [ACM20, BMS20, Gye23, LS20a, PV21a]. angle [DO22]. angled [FKK21a]. angles [BKRS20, BMPS23, ŠH21a]. annihilated [LTWW23]. annihilating [FG22]. anomaly [Dax23]. Anti [FP22, AFSS20, BW22, FN23, TT21]. anti-commutative [FN23]. anti-diagonal [BW22]. anti-dual [TT21]. Anti-isomorphisms [FP22]. anti-regular [AFSS20]. Anticodes [BGLR22]. Anticommutative [Vol23]. antilinear [RT21, SZ20]. antinorm [GZ20b]. any [FJ24]. Apéry [MRT20]. Appl [AAH^{+}22, AD20a, Ali20, AG20, BP23b, BDHS23b, BKSP23, CdGS20, FJS21, FHMP22, Hou20, JLM20a, KAPS20, MH21, PW20, Sło21a, SD21a]. application [AFK^{+}23, ACM22, AP23, CKOS23, DMQD20, Fio20, Miy24, PŠ23a, PS22, Pra20, RSS21, SMP24, ZS22a]. Applications [CC22b, Kak20, KK20a, SZ22a, AWvW24, AAMR20, CST21, CCL21, Djo21, DL22, EGJP21, FISW21, GSS24, GCOP24, HCZ20, HLYZ23, HVX23, KS23c, KP21, LHFL20, MMV21, MP22, MG21, NN24, SWW22, SS24b, Wój22a]. Applying [LNN^{+}20]. approach [AKVS23, AC22, Djo21, DFH^{+}20, FGM21, GdM23, HT22, IB22, LT24, LHK20, Lin23, SK21a, SKS22, Ste21a, VS23, WZFW23]. approaches [Ghi20]. Approximate [AKN23, Wój22a, GNR20, HLW23, ORŠ23, dAdCG^{+}21]. Approximating [AGKS23, Bar21b]. Approximation [FKMM21, RKBA21, ZRQK22, Bot21, CEF^{+}22, CER^{+}23, CKM20, FGM21, KKM22a, KPP21, KST22, MSZ23, Rao21, Rod21, SK21a, Wój22b]. approximations [CHHS20, DMQD20, GZ20b, LV22b, dHH23]. April [Ano20-63, Ano20-65, Ano21-53, Ano22-55, Ano22-62, Ano23-52, Ano23-57, Ano24-27, Ano24-35]. arbitrary [ABH^{+}23, ADHM23, BDRS20, CGP22, FISW21, MMR20, VB20, WWJ20, WFGT20, ZHSX20]. arboricity [HXLL21]. arcs [KK20b]. argument [MB20]. arise [CGP21]. arising [BBS23, CER^{+}23, KMZ20, MMZ20, Oli21]. arithmetic [BLS20, ES23]. arithmetical [DLL23, VV22]. Arithmeticity [BSS23]. arrangements [Win21]. array [He20, YY23, ZZJ23]. arrays [AK23, KNiO23, MMW22, Pet21, Sło20, Sło21a, Sło21b, SS23c, YY21]. arrowhead [CRSY22]. Artin [FKK21a]. Artinian [Alt22]. ascent [HP23b]. Askey [SS20]. aspects [Got20, VV22]. assignment [ZK21]. assignments [NMS20]. associated [AG15, AG20, BR22, BLS20, CCKM22, CA22, Chr24b, DZCY22, DM24, GLZ20, HCZ20, HP22, KL21a, KS22c, LW21a, MRT20, PPŻ21, RW21a, Sta23b]. association [BMS21, LTTW22, XTLK23]. associative [Bri22]. Asymptotic [LLW21, MPS22, BCR^{+}23, Du23, LLW20, RRvW24]. asymptotically [RT23, Roc20]. asynchronous [SDTR23]. attaining [BDSS21]. attainment [PSZ21]. Audenaert [DF20b]. August [Ano20-50, Ano20-52, Ano21-61, Ano22-57, Ano22-49, Ano23-53, Ano23-50]. automata [Phu24]. Automatic [PSW23]. automorphic [BKT23]. automorphism [Oli21, SZ22b]. Automorphisms [AM23a, Bez22, CGMP23, PPŻ23, DGET21, FN23, MDH23]. autoregression [FKMM21]. autoregressive [GZ20a]. averaging [MA24]. Avoiding [BMPS23, OPA23].

B [BGOO21]. B-splines [BGOO21]. back [MZZ24]. background [BCP22]. backtracking [GK21, KKS21, MZZ24]. Backward [Pra20, Hal20, Pik23, TV21, WW20]. balanced [BW21, DFS23, DK21, GT23]. ball [MAG21, Wój22b]. balls [FKKM22]. Banach [AEGV21, AHV20, BS21a, Ban22, BDI20, Djo21, FPR20, GGGD23, IKLP20, MP20a, Sas21, Sch24]. Band [GGGD23, CCCS21, EGV20]. Band-diagonal [GGGD23]. banded [OW23]. Bartholdi [Zhu21]. based [Bla20, BDFF23, CL21, FQ20, GSS24, GFW24, GMR22, KCML20, LLW20, LLW21, LW21b, MQW20, NPkT21, NPHT20, QJMW23, SGGZ21, WXL24, XJLL23]. bases [AMO24, ADMZ21, BC24, HJBKR24, NWW20, NV23a]. basic [VS21]. basis [GZ23, LT24, MMV22, TL21, Yos21]. Baxter [MDH23]. Bayes’ [PR22a]. Bayesian [DTA23]. behavior [GP20]. Bell [GZ20c, Li20a]. below [GK20]. Berezin [CF22]. Berge [KLLW21, ZKLS20]. Bernoulli [Kim20]. Bernstein [MMV21, MMV22]. Bessel [MPR22]. Best [Bot21, Rao21, Wój22b]. Bethe [FB21, YW22]. Better [OM20]. Between [PPW20, Ago24, BCE21, CMS21, CAEH^{+}22, FPR20, GKS21, GMR22, IOT21, KS22c, Kye23, LL24a, MP20a, Min22, Mon23, Sán21, Sta23b, WFL20, WWY23, ZXZW23]. betweenness [DF20b, DF22]. Beurling [ACSS22]. beyond [RS21b, XZ23]. Bhatia [Roy24]. Bhattacharya [CLwW22, GF20]. Bhattacharya-Mesner [GF20]. Bi [CR23]. Bi-isometries [CR23]. bicyclic [DGS^{+}21a, SWW21, YD21]. bidegreed [SSS21]. bidiagonal [BFMM23, MPR24, MMV21, TŽ23]. bifurcation [FHLS22]. bifurcations [Fik20]. Bilinear [SDS21, Bui21, Chr24b, KLLS22, Ros22, RSS21]. binary [BT20a, HP23a, May21, Sha22, SZH22, WJLW24]. Binomial [BCR20, Bay24]. bipartite [AD20a, AD20b, ADH^{+}24, BJP21, BS23b, CWYZ23, CP21, Du23, GLZ21, JMP21a, Kub22, LG21a, LMX23, OJQ20, Sim21, SZ21]. biregular [FM20a, HP22]. Biregularity [RS21b]. Birkhoff [AGK^{+}21, Shi22]. Bivariate [CCCS21, BCP24]. Blanchfield [BCP22]. Blind [GZ20a, KKM22b]. Block [CLRT22, HMPT20, AT24, BS20d, BJ23, CW21, CZ20b, CDG21, CDG22a, Dew21, DdR20, FGDY20, FLT20, FG20, HP23a, JGW24, LOŠ21a, LG21b, PŠ23a, RS21a, RKBA21, SZSMB20, VD22, XLS20, YOV21, ZL23a, ZL23b, dAdCG^{+}21, KK23a]. block-diagonal [BS20d]. block-Jacobi [YOV21]. Block-sparse [HMPT20]. block-triangular [FGDY20]. blocks [B– DK^{+}24, BL24, CDG22a, DGS^{+}21a, EGC21, KS20a, Sta20a]. blow [LZ21b, XHW23]. blow-up [LZ21b, XHW23]. Board [Ano20a, Ano20b, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y, Ano20z, Ano21a, Ano21b, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p]. Board [Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano22a, Ano22b, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano23a, Ano23b, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano23c, Ano23d, Ano23e, Ano23f]. Board [Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z]. Bogomolov [Rai23]. Bohemians [CCGV^{+}20]. book [TTV23]. Boolean [CKKR23, KCML20]. Bordered [Kli20a, CESS21]. Bordering [AFK^{+}23]. Borders [AHKT21]. Bose [Mar21]. Böttcher [CKOS23]. bound [AAFM23, AL23a, BT20b, FLRS20, Fio20, Hae21, LZZ23, LHZ21, MLH22, MT23, WG20a]. boundary [CEM21, KŠ22b]. Bounded [BDRS20, DS24, AGV20, BS20d, GK20, Lyu20, MS21, Nag23, RT23, WHM23]. Boundedness [SS23a, RAK21]. Bounding [CKOS22, ES23, GLS22, KL21b, ZS22a]. Bounds [DW21, GLHNW20, KKP21, LW21c, PP20, SK21b, WWT20, AOA23, ASS21, AGH20, AGNPSE23, BHHP21, BR22, BKPS23, Bhu24, CKTP20, CAJ21, CI21, DYB20, DK21, FQ20, FS23, Gan23, GMR21, IGW17, JZZ24, KL23, KCPMNS21, Kus23, MT21b, MP21, OM20, Ran23, Rum22, SZ23a, Skr20, Skr21, Sta20a, VCR21, WG20b, ZZ21, dHH23, CN24, GZ20b]. Braess’ [Cia20a]. braided [Soł20]. Brazil [AHKT21]. breaking [AKL^{+}21, BH24]. bridge [BCE21]. bridges [CGJvB22]. Brouwer [Coo21, GPRT20, Roc20, TT24]. Brualdi [YLCB21]. Brualdi-type [YLCB21]. Bruhat [FdCS20]. Brunn [LF20, ZJCA20]. bundles [DD23b]. Bures [DLVV21, Min22]. caching [FT23]. Cadzow [KPP21]. cages [HP22]. Calabi [GLNW24]. calculus [LMR21, NS21]. Calderón [GR22]. can [HMS21]. canonical [AMO24, GvRR21a, Li24, Sta21, SZ20]. cardinality [AGNPSE23, SS21b]. Cartan [CKL24, FKM22, Sim20]. Cartesian [TT24]. case [AS20a, CV21, GS21a, LTYZ23, Mań22, Zem22]. cases [dSP22, dSP23b]. Catalan [Pet21]. categories [Ago24]. Cauchy [Ner20, Yan21]. Cayley [AY23, ASG22, AS23, BL22, BL23, CW22, CK23a, Ces20, DF20a, ET20, Hua22, KB22, Kad23, LW21a, MT23, MST24, Mur22, PV21b, SSvW20, WFY24]. cellular [Phu24]. Cellularity [XY20a]. center [CDM20, Cia20b, PV21a]. Centers [HLYZ23]. Central [CNO23, GFD21]. centrality [BD21a]. centralizer [B– DK^{+}24, MMR20, XZ21]. centralizers [AM23b]. Centrosymmetric [BCP24, CCK21a, XY20a]. Certain [DY23a, AG21, BDNO20, BS20f, BMPS23, Cra21, CMR21, DO22, EMŠV22, GLZ20, GL20a, JMP21b, LW21a, Shi21c, TW21]. CG [CLS24]. chain [BHKW22, CM23b]. Chains [CGG^{+}21, AHV20, CCPKP20, CK23b, CR21, DS20, HX24, LL20, VS20]. Change [JST21, AS24, GZ23, SD21a, SD21d, TLC20]. change-of-basis [GZ23]. changes [KK22b]. channels [AKL^{+}21, GL21, LCH23]. Characterisation [Ram23]. Characteristic [JL22, Kad20, ABSV21, CB21, Cia20b, DvDW23, GM22, HY24, HHS20b, JR24, Kor21, KLLS22, LS20b, Pan23, YLZH21, Zhe21]. characteristic-like [GM22]. characteristics [AKPS19, BN23, KAPS20]. Characterization [AG21, MP20a, NYW22, TWW20, AMR20, AdST22, AK23, BP21b, BP23b, DMS20, DY23b, DdF23, FGGG24, MA22, MM21, NWW20, SC20a, SSS21, TW21, Lew21].

Characterizations [CMM20, Ime23, AAH^{+}22, MC21b, ME21, NPHT20, QJMW23, QQM24, ZBCI21]. Characterizing [AHZ22, KSSW20, Tan23, VTBC20]. Charlier [XJLL23]. Chebyshev [MSZ23, WCS20]. Chebyshev-like [WCS20]. checkerboards [EJS24]. Cheeger [KMP23]. Chief [Ano21-49, Bru21, Bru22, Bru24]. chiral [FP23]. Cholesky [Kim20]. chordal [AJM21, AJM24]. Chromatic [CK23a, Ere21, Gan23, PK21, WYF^{+}20, WYQ21]. Chung [YZY24]. circle [Adr22, BBCD21, FG20, FG23b]. Circulant [DN21, AAMR20, ACOM21, CEJM21, Che21, DS22, HP23a, LS23, – DMM^{+}22]. Circular [Lee22, TŽ23, GW20, PW21]. circularity [vdMvSW24]. CKV [CCL21]. CKV-type [CCL21]. Clarkson [BL20]. class [AR21, BCR^{+}23, BSC23, CHT20, CESS21, EGV20, Fio20, Guo22, GLT23, Hu21, KS23a, KW20, Loe21, PS23b, SL20a, SL21]. Classes [BP21a, ACD22, BC24, GLZ20, GK22b, KZPH22, ME21, MST24, OP21, PPŻ21, Pan22, PPŻ23, WZ23b, ZM21]. classical [CGP21, KW20, Pan23, Pet21]. Classification [AADT23, BKS^{+}21, DGL23b, FN23, SS23b, ABCO22, DGdSS23, HX24, Sim21]. classify [LNN^{+}20]. Classifying [BDN24, IdSSV21]. clean [Šte22b]. Cleaning [KS22a]. clique [DLL23, DW21, HLX20, HXLL21, JGZ21, WHM23, ZvD21]. cliques [DKL^{+}22, MLZ24]. close [MS21]. Closed [BKLT23, AT24, BHBO21, De 20, SL21]. Closed-range [BKLT23]. closest [Kov21]. Cluster [EGL21]. clustered [BDGY21, KN20]. clustering [AEGL21]. Clusters [CK23b]. CM [EL20]. CMV [FG20, GP20]. Co [VJ24, AV23, BPS23, ZBCI21]. Co-eigenvector [VJ24]. co-EP [ZBCI21]. co-extensions [BPS23]. co-isometric [AV23]. coalgebras [FMS21, KUZ22, Pan21]. cocharacters [CdSV22, Iop20]. coclique [SD21c]. code [KCML20]. coded [FT23]. Codes [BCR20, AGNPSE23, BZZ22, CM23b, CL21, CXZ22, GL20a, GLT23, LLSV21, MMZ20, NPkT21, NPR22, NPSE24, NPHT20, Ner20, Oli21, ZT23]. Codeterminantal [AAHV22]. codimension [Alt22, MRT20]. codimensions [Riz23]. coding [KS22a]. coefficient [GZZ21, GQZ20, LL21b]. coefficients [Bay24, Du23, GP21, MMA22, PP20, RS21a, VS23]. cograph [LL21e]. cographs [BR23a]. Cohen [FG22]. coherence [MB20]. coherent [Sud22]. cohomology [FKK21a]. coin [SMA20]. coincide [tHvdM21]. cokernel [CLS23]. Colin [MS20b]. Collatz [KL21a]. collection [MAG21]. collocation [MMV22]. Coloring [AMZ21, KKOO22]. Colorings [O.22]. colourings [OJQ20]. Column [Shi21b, RC23]. columns [FJ24]. Combinatorial [AD24, AS23, Kub23, LS21, CST21, Got20]. Combinatorics [DS20, Gol20]. common [LHK20]. communicability [HD22]. commutant [BPS23]. Commutative [BE20, FN23, Got20, Mig21, PR22a, SS22, Szc21]. Commutativity [Pan22, ORŠ23]. Commutator [CAJ21, CJ20, CKOS22]. commutators [BDHS22, BDHS23b, BR23b, VZ20, ZS22a]. Commuting [CMT21, BCG24, BPS23, BFPS21, BKS15, BKSP23, Fek20, LP21, SS20, Xie21]. Comon [ZHSX20]. compact [JLM20a, JLM20b, Pan22, Rao21, hB21]. compactly [AKPS19, KAPS20]. Comparative [ŠH21a]. Comparing [CTS23, RZ24]. Compatible [BKS15, BKSP23, Gor24, GFW24]. compensation [MM21]. competition [CKKR23]. complement [Dev22, MH20, MH21, TT21]. complementable [ACM20]. Complementarity [BCFT21, CD23a, DY23b, FQ20, MG21, MC21b, SS21b, SNK22, XX20]. Complementary [EGK^{+}24, VD22]. complements [ARS24, CMM22, LQL21, MK23, QLCM20]. Complete [AC21, FHMP21, FHMP22, BGOO21, CCPKP20, Coo20, IMT20, SC20a, Shi21b, SSS21, SD21a, SD21d, TLC20, Zhe20, Zhe21, tHvdM21]. Completely [BLS20, JS22, AKN23, BHMU20, BKT23, FOR22, GD20, LHFL20, SM23, ZFZ20]. completion [BT20a, BBS20, DAG21, Kim22, SL20a, TL21]. completions [CP21, EGC21]. Complex [BFLT23, CFK^{+}20, LWH21, XZWT20, ACLS20, CDM21, ES20, EJJL20, Fik20, FP21, Har22, HHD20, HW21a, Kni20, KP23b, LZZ23, LW21c, MRZ22, MvWC23, dSP20a, SK21b, WCKQ20]. complex-valued [MvWC23]. complexes [KMZ20]. Complexifications [IKLP20]. complexity [TTV23]. component [FKMM21]. components [Van24a, Yaz21]. composition [AOCM20, SS23a, SS24b]. Compositions [Kye23]. compound [FKMM21]. compressing [Cos22]. compression [Cra21, CMR21]. comprised [FS20]. Computation [SK21a, Bol21, CHHS20, DK21, MC21a, NV23b, OSC20, Var21, Zhe20]. computational [Bai24, Men22, Sim20]. computations [MPR22, MMV22, Mar21, Yan21]. computed [CKTP20]. computing [AHZ22, KK20a, NV23a]. Concavity [CZ22a, Shi20a, BW24, BO23, Nie24]. Concentration [HW20]. concentric [PW21]. concept [KP21, SG20]. concerning [CJ20, FdCS20, GAM20, Gye23, Iop20, Miy24, QJW20, Row20]. condition [AJNS24, BCR^{+}23, BS20e, CV21, DBV23, Gal21, GLZ21, fLLV20, LG21a, Pra20, Sch23, Van24b]. Conditional [BCE21, EM23]. Conditionally [JJLS23]. conditions [BP21b, BP23b, HMPT20, HWD23, Kub23, LBW23, Pri22, Tsa23, WZC23, ZWL20b, ZWL20a, ZBWL20, ZL23c]. cone [CIM22, GH23, GNR20, LV22b, MC21b]. cones [GKS21, HS23b, MS24, ST20]. Conference [AHKT21]. configuration [cRLY20]. configurations [CDG22b, Sud22]. conformal [BF24a]. Congruence [BFMS21, HWK21]. congruences [DY23a]. conic [LPV21]. conjecture [BD21a, BCM21, Buc20, CLL20, CJ20, CLwW22, Coo21, FdCS20, GPRT20, GdM23, Gau24, HLH23, HD22, KL21a, Loe23, LZ21b, PPW20, Roc20, SD20, SL20b, SD21a, SD21d, Tur21, WZ23a, XZ23, YW22, ZL23b, ZHSX20, TT24]. Conjectured [DID22, ELW24, GAMA20]. Conjectures [BC22, KK22a, BEHK22, GLLZ20, KK20b, SD21c]. conjugacy [PPŻ21, Pan22, PPŻ23]. Conjugate [EL20, CF24, ES20]. Connected [LTYZ23, AHLvdH21, CCPKP20, CLL20, CLT22, HLS21b, Kha24, dSP20a, SS21b, WFGT20, WG20a, WZW^{+}21, WLT22, ZG22, ZWL20b, ZWL20a, ZBWL20, ZWT21]. connection [VS23]. Connectivity [CDS21, HLX20, BEHK22, DPR20, HS22, HWD23, WYF^{+}20]. conquer [BT20a, XBSF23]. Consistent [Ais24, LG21b, SDTR23]. consisting [SZ20]. constant [JKS23, LL20, LLSV21, Rub22]. constants [CN24, Mło22, RT23]. constrained [fLLV20]. constraint [ZMP23]. Constraints [CFKO21, Coo21, BES21, PS22, SDS21]. Constructing [HLS24, Kin21, LL21d, HJBKR24, XBSF23]. Construction [APNP^{+}21, CL21, AO22, CT20, ET20, cRLY20, PM23, QJW20, VS21]. Constructions [BC24, CXZ22, AM21, BRL23, ZT23]. contain [Vit21]. containing [AS20a]. contingency [PP20]. Continuity [CG20, RW21a, Car21]. continuous [Zha22]. continuous-time [Zha22]. contour [FP21]. contracting [LC20]. contraction [AWvW24]. contractions [LZ21a]. contractivity [BCE21]. control [BPZ21, Dra22, Ost23, WES21, ZK21]. controllability [AD20a, AD20b, BPZ21]. controllable [DYLL23]. Convergence [CW21, ABCH23, BWW21, BDFF23, Har22, SDTR23, YOV21]. converters [GS21c]. Converting [DGS21b]. Convex [GKS21, MS21, JJL20a, Sán21, ST22, ST20]. Convexity [CF22, EMŠV22, San23]. Convolutional [HNS21, NPR22]. coordinates [CR23, FM22a, Sar21]. Coordination [FT23]. copositive [LV22b, LHFL20]. copositivity [Ans21]. coprime [Var21]. corank [EG20, Kop20, SZ21]. core [BCG24, WJLW24, Mif21]. core-nilpotent [WJLW24]. corners [MRZ20]. corona [LLH20]. corona-type [LLH20]. coronae [SC20b]. Correlation [BAPF23, GQZ20, Li20a]. correspondence [WvdVvdH24]. corresponding [BLS20, FKKM22, PPŻ23, RS21a, TV21]. Corrigendum [AD20a, Ali20, AG20, BP23b, BDHS23b, BKSP23, CdGS20, FJS21, FHMP22, Hou20, JLM20a, MH21, PW20, Sło21a, SD21a]. cosine [HJBKR24, SS24b]. Cospectral [ADHS23, GZ22, KSSW20, LWYL20, QJW20]. Cospectrality [MPRS21, Her20]. costly [OPA23]. countable [LL20, dSP23a]. counter [Sło20, Sło21a]. Counterexamples [CLwW22, CDPR23, FJS17, FJS21]. Counting [Rei23]. coupled [GP21, ZBdAF20]. couplings [OSC20]. Covariance [RRvW24, GZ20a]. covariant [LCH23]. cover [BK22]. covered [WZC23]. covering [AADT23]. Cowen [JKX20]. Coxeter [Sim20, Sim21, SZ21, SZ22a, dlPG21]. CP [ZBdAF20]. Cramer [NWW20]. criteria [SlW22]. criterion [MMW22]. Critical [DLL23, BM22]. Cross [DGET21, CKM20]. Crouzeix [Li20b, PPW20]. cubic [CH21]. Cuntz [ACSS22]. curvature [ÁVKP20b, Sic21]. curve [GQZ20, LPS23, SS23b]. curves [BdPS21, Hom20, NYW22, Sán21]. Cut [Mis22, Nic20, AdLN23, FYS^{+}23, MMP21]. cut-and-project [MMP21]. Cvetković [KKOO22]. Cvetković-type [KKOO22]. cycle [BAPF23, TCC21, ZWF20]. cycles [AAH^{+}22, AGNN20, ACDG21, CWZ22, CP21, Des24, GMSG22, JMP21a, KSY21, LW21a, LY23, LG21a, LWYL20, MLZ24, dSP23a, RDFM23, SMA20, SD21c, Tho22, ZLLZ22]. Cyclic [BDvW24, Bol22, CL21, CKLN23, DdR20, Gau24, Mig21, ZT23]. cyclicity [KCPMNS21]. cyclicly [KS23a]. cyclomatic [WWJ20].

D [SV21]. DAHA [Lee20]. damped [YGW20]. data [ADGK21, APLN^{+}23, CLS20, SL20a, XBSF23, YGW20, ZW20]. decay [BR22]. December [Ano20-55, Ano20-66, Ano21-62, Ano21-67, Ano22-69, Ano22-66, Ano23-61, Ano23-62]. decoding [NPkT21]. decoherence [FOR22]. Decomposable [PM23]. Decomposition [Kor21, ÁVKP20a, AOZ23, BES21, Bhu24, DGL23b, FRS21, Gal21, GSS24, JMP21a, Kim20, LY20, MMV21, MQW20, Pre20, SZH22, Ste21b, VB20, VCR21, WJLW24, Wój20, WX20]. decomposition-based [GSS24]. Decompositions [BDHS22, BDHS23b, BR23b, CCP22, DGL23a, DID22, DBV23, Ham23, HS23a, LDW21, MO20, OPA23, SDS21, ZBdAF20]. deconvolution [GZ20a]. decrease [KS23a]. defective [FFSW23]. deficient [DB23]. defined [KSY21]. definite [Bar21b, BL24, CRR20, CIM22, GT21, Har22, IB22, JJLS23, LWW22, Lim21, Min22, MS24, Sra21]. definiteness [BCE21]. definition [BEFSC22, ACM22]. deflating [ADGK21]. deformations [FKK^{+}21b]. deformed [CKOS22]. degenerate [CDdM22, GNP22, Lee20, VMB22]. Degree [SGGZ21, AIK^{+}24, BFF^{+}23b, CLL20, CLT22, GMR22, HHS21, Hom20, HLYZ21, LLW20, LLW21, LW21b, LLA23, LW21c, NdLO21, WG20a, WYF^{+}20, ZWW23, ZWL20a, dLMO20]. Degree-based [SGGZ21, LW21b]. degree-distance-based [LLW20, LLW21]. degrees [WWT20]. Delannoy [YZY24]. delay [WXL24, XJLL23]. Deletion [AWvW24, SD21a, SD21d, TLC20]. Deletion-contraction [AWvW24]. Delsarte [BMS21]. delta [FISW21]. demystified [CCCS21]. dendrimers [DFS23]. dendrite [XY20c]. dense [AM23a]. Density [BHKW22, BMR23]. dependence [CZ20a, Lim22b]. depth [SS21c]. derivations [Bez22, BDV21, HKS23, JLM20a, JLM20b, LS20b]. derivative [Sin21, ZX20]. derivatives [RMG24, Sch23, Ste22a]. derived [CIM22, LNN^{+}20]. descent [DAG21, HP23b]. descriptor [DYT22]. design [AKPS19, GCOP24, KAPS20]. designs [BMS21, BNS22, Gol20, HW21b, RC23]. detection [CT20]. determinant [ABW22, BT20b, Dom20, DGS^{+}21a, LS20c]. Determinantal [Tsa23, BR21, CN22, GAMA20, GAM20, Koz21, Sos23, SW21]. Determinants [Bri22, Guo22, Sha22, Gil20, Kus23, MSX22, OP21, Shi21c, dC21]. Determination [Sta23a, CWBB23]. determine [LWW22]. determined [AGNN20, AEGV21, BB22, CW23b, QWW21, Zha21]. determining [KK22a]. deterministic [CMS21, Ghi20, OW23]. Development [BP21b, BP23b]. developments [GPRT20]. Diagonal [BL24, JMPS24, APNP^{+}21, BS20d, Bot21, BW22, GGGD23, KS23a, KP21, MRZ20, MG21, Mil24, NS21, TJ21]. diagonalisable [CFK^{+}20]. Diagonalizability [AdlCPS20]. Diagonalizable [HLYZ21, AAF^{+}21, CFSiT20, FFSW23]. diagonalizably [AMS21]. diagonally [CKM20, JMPS24]. diameter [ABH^{+}23, FW22, HLS21b, HLS22, IKPR20, JLW22, LBW23, XZ23]. dichotomy [BCG24]. dichromatic [YBW24]. dictionaries [SC20a]. dicyclic [ASG22, BL22]. differ [Alf20a]. difference [CER^{+}23, FISW21, LH21, MH20, MH21, OJQ20]. difference-1 [OJQ20]. differences [BGOO21]. differentiability [Roy24]. Differential [Nar23, Riz23, GP21, LMR21, Shi20a]. diffusion [Est20]. digraph [BGGA22, CGMP23, GS23a, KKS21]. digraphs [AMR20, ADGKM21, BCE^{+}23, BCFT21, BCF^{+}20, CHSW22, FPS22, GMR21, GM23, Her20, KKOO22, LLWW23, LS24b, Moh20, Mor20, O.22, TFT20, YBW24]. dilations [FFSW23]. dimension [JMPS21, JCD24, LLSV21, dSP23a, Szc21, Yos21, ZG21]. dimensional [AKPS19, BS21a, Ban22, Bla20, BFPS21, CM23a, CG20, CdGS12, CdGS20, ĆD20, ĆD23b, DGdSS23, DJ22, FN23, GS21a, GLT23, Ime23, JL22, KAPS20, KY21, LNN^{+}20, Li20a, LM24, O’L23, SZH22, SS23c, ZZJ23]. Dimensionally [MKE^{+}20]. Diminimal [ABH^{+}23]. Diophantine [QSS22]. direct [AS20a, GNP22, HS23a, KS20a, MvWC23]. directed [BES21, Bol21, GM20, LL21c, Pik23]. directional [AKPS19, GS21c, KAPS20]. Dirichlet [dC21]. discontinuity [CG20, Mis22]. discontinuous [FSRVSST22]. discoordination [FT23]. discrepancy [ZS22a]. Discrete [BCNS22, Ruž21, UZ21, AV23, CF24, CMS21, DK21, EAHS21, Lew21, Mar24, WXL24, XJLL23, Zem22, Zha22]. discrete-time [CF24, CMS21, DK21, Lew21, Zha22]. disjoint [HO20, ZLLZ22]. Dispersing [MRZ22]. dispersionless [VS20]. Disproof [HLH23, YW22, SD21c]. dissipative [GS21b, MMW21]. Distance [AT20, AGNPSE23, IKPR20, KAGH24, LZ20, LL21e, MT21b, MMW21, WZFW23, ZLLZ22, Alf20a, Alf20b, ABG20, AS22, BBG21, BS20e, BJP21, BEF20, BNS22, Bud24, BLS20, CN24, Dev22, DGS^{+}21a, Ere21, Est23, FM20a, GK22b, Goe21, GS22, GLT23, HSS^{+}21, HP22, HD22, JGZ21, JT24, LLW20, LLW21, LQL21, MK23, MC21a, Mif21, Mor23, PGRU20, PK21, QLCM20, SL20a, SD21a, SD21d, TLC20, WZ23a, WFY24, XLS20, Zha22, ZvD23, ZTX24, MT21b]. distance-biregular [FM20a]. Distance-regular [IKPR20, KAGH24, GK22b]. distances [BMPS23, SL20a, TXF24]. distillation [QCCS21]. Distinct [KK24, AFK^{+}23, ABR21, Bar20, BFF^{+}23b, BFF^{+}23a, BBC^{+}22b, FM22b, GHL^{+}24, Kub22, MP21, Rei23, Row20, SSS21, Zha21]. distinction [BD21a]. Distinguishability [DYT22]. distinguishing [KMNP20]. distributed [BHD^{+}21]. Distribution [Lub22, ŠH21b, AAFT22, AADT23, Bou21, CM23b, WYF^{+}20, XZ23]. distributions [CT24, RW22]. divergence [DLVV21, PV21a]. divergences [LL21a, Sra21]. divide [BT20a, XBSF23]. divide-and-conquer [BT20a, XBSF23]. divided [BGOO21]. Divisibility [GR21, CWBB23]. division [BDHS22, BDHS23b, Bol22]. divisor [CPS20, CN23b]. do [CGJvB22]. does [LWB20, Sta20b]. domain [Gal21]. domains [Nag23]. dominance [KP21, MG21]. dominant [CKM20, JMPS24]. dominating [LTYZ23].

Domination [XLYS20, AAFM23, HLH23, LLX23, WYF^{+}20]. Double [ABHR22, AdST22, DvDW23, RS21a, Sta23a, SS23c, ZZ20]. double-structured [RS21a]. Doubly [Djo23, APNP^{+}21, CC22a, GK20, JJL20b, KK22a, PSAPNV20, RX21, SM23]. doubly-nonnegative [RX21]. Douglas [JKX20]. Drazin [BR24]. Dual [WJLW24, ANAS20, AP23, GKLZ20, JN22, KCML20, Lee20, LT24, SZ22a, TT21]. duality [LTTW22, Phu24]. due [SD21a, SD21d, TLC20]. DXZ [DID22]. Dynamic [AHNW22, YOV21]. Dynamical [AP23, CFKO21, GP20, KHU23, Rei23]. dynamics [CKOS23, GGGD23, Phu24, RMP21]. Dynkin [MP23b, Sim21, WWG22, Wan23].

Each [GZZ21]. economics [EGJP21]. Edge [HS22, LZ21a, AADT23, CW22, GMSG22, HWD23, RFP22, Sim21, SZ21, SD21a, SD21d, TLC20, Zak21]. edge-bipartite [Sim21, SZ21]. Edge-connectivity [HS22]. edge-free [Zak21]. Edge-matching [LZ21a]. edges [CB21, LC20, LQ21, XW21, ZXL20]. Editor [Bru21, Bru22, Bru24]. Editor-in-Chief [Bru21, Bru22, Bru24]. Editorial [Ano20a, Ano20b, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y, Ano20z, Ano21a, Ano21b, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p]. Editorial [Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano22a, Ano22b, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano23a, Ano23b, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano23c, Ano23d, Ano23e, Ano23f]. Editorial [Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z]. Editors [Ano21-49]. Editors-in-Chief [Ano21-49]. effect [ADGK21, EE23, LC20, TJ21, XW21]. Effective [Dev22, BBW20]. Efficiency [FJ24]. Efficient [CEF^{+}22, fLLV20, CLLK21]. efficiently [KHU23]. eigendecomposition [BN23]. eigenfunctions [GKK^{+}20, KKSV21]. Eigenpairs [Los21b, AK20]. eigenspaces [DF20a, WZFW23]. eigenspectra [MvWC23]. eigenstructure [Kli20b]. Eigenvalue [CH21, Fik20, FP21, GA22, AAFM23, AAFT22, AvDF20, AADT23, ABG20, ÁVKP20a, AAMR20, APLN^{+}23, AGV20, BP21a, Beh21, BLMT21, BW22, CLM20, DMQD20, EN22, FHLS22, Har22, IKPR20, JL20, JMP21b, JCD24, KL21b, KGYY23, KAGH24, KP22, LWB20, LOŠ21a, LOŠ23, LDSS20, LWH21, LLL22, MvWC23, NYW22, Obo22, Obo24, PS22, Ran23, SS21c, Sta20a, TWW20, TW21, TJ21, VD22, WWJ20, WFGT20, WYF^{+}20, WZW^{+}21, WCS20, WQP23, XWT20, XLYS20, XLS20, YGW20, YK21, YW23, YLCB21, YWF21, ZK21, ZW22]. Eigenvalues [CEM21, EGJP21, GMSG22, Hu21, WYQ21, AFK^{+}23, AJM21, AJM24, AFG^{+}21, APNP^{+}21, ABR21, Bar20, BFF^{+}23b, BFF^{+}23a, Bas21, BJ20b, BBC^{+}22b, CPS20, CZ22c, CvDB23, CP23, Csi22, ĆD20, ĆD23b, DN21, DJ23, DH22, DH23, ET20, FM22b, FR21, Fra20, GPRT20, GHL^{+}24, GS23a, HT23, HLX20, HXLL21, Jai21, KKSV21, KKP21, KK20b, KK22a, KK24, Kub22, LW21a, hLqJ21, LL21e, Lub22, Mel23, Mis20, Moh22, MP21, Nic20, PSAPNV20, QLCM20, Row20, RS21b, RB22, Sic21, SSS21, SMM24, Ste22a, SD21b, SLL22, TFT20, TW22, Tur21, VS20, VTBC20, WWT20, WLL23, Wei22, WZT20, WWY23, XZ23, YOV21, Yan21, Yaz21, Zha21, ZZW22, Zha24a]. Eigenvector [Van24a, CTS23, VJ24, MPRS21, StW22]. eigenvectors [AdL21, AdLN23, CCP22, CA22, Coh20, EGJP21, ET20, Hu21, StW22, YOV21]. eight [GS21a]. Eisenstein [KB22, Kad23]. elementary [VZ20]. Elements [Coh20, APNP^{+}21, BS21b, GLAdS23, Hom20, Pat20]. elementwise [TV21]. elimination [SG20]. Elliptic [CN23a, BMS21, CER^{+}23]. embedded [ÁVKP20a]. embedding [BS20a, CGP21, CN24, GA22, Mło22]. embeddings [DG20, UZ21]. enclosing [MAG21]. enclosure [Miy24]. encoders [Ner20]. end [VZ20]. endomorphism [AM23a, MMR20]. endomorphisms [DGL23a, Mur22]. endpoint [ZS22b]. energies [LLW21, MH20, MH21, SGGZ21]. Energy [AA21, Obo24, AGH20, AAAH22, AL23a, ABR21, AS24, BP20, CZG21, CDVPT22, GLHNW20, GMR22, JGZ21, Kni20, LLW20, LW20a, PGRU20, SK21b, SD21a, SD21d, TFT20, TLC20, UZ21]. ensemble [BBS23, Sch22]. Entanglement [AKL^{+}21, BH24, CJMP22, LS21, QCCS21]. entries [AdL21, BR22, Bib22, JMPS24, LLW20, LLW21, Lub22, Lyu20, MR21a, StW22, ZC23]. entropies [Shi21a]. Entropy [BCF^{+}20, CHHS20, FKMM21, Kak20, MKE^{+}20]. entry [Alf20a, TJ21]. Entrywise [Rum22]. Enumerating [AR21]. enveloping [CZ20a, Yas23]. EP [Dra23, ZBCI21]. EP-like [Dra23]. Epstein [CZ22a]. equal [GJV20, Kha24, KP23b, NdLO21, WLL23, WLT22]. equality [BP21b, BP23b, SDS21]. equation [BS21b, CF24, Djo21, DL23, EJ23, GJ23, GvRR^{+}23, Li24, Pra20, Ruž21, Sas21]. equations [ABCH23, BF21, BF24b, CLS23, GP21, GJ23, IB22, Lim22a, LH21, MPS24, MP20b, PS23b, QSS22, Rob20, RW22, Uch21]. equiangular [FS20, cRLY20, GIJM22]. equienergetic [PV21b]. equilateral [GL23]. Equitable [RFP22, AHZ22, LDW21]. equitably [Kad20]. Equivalence [AdF20, NPSE24, NPHT20, Oli21, FHMP21, FHMP22, GLT23, MPRS21, PR22b, Sch22, DD23b]. equivalences [MZ22]. Equivalent [BRL23, FPS22, Her20]. erasure [LM22b]. Erdos [Adr22]. ergodic [CR21, FOR22]. Erratum [KAPS20]. Error [CKTP20, MMV21, FQ20, Hal20, LHZ21, Pra20, TV21, VCR21, WW20, dHH23]. errors [Ais24]. essential [AHV20, WMCQ23, ZS22b, ĆD23b]. estimates [CWW24, CER^{+}23]. Estimating [Hal20]. estimation [Ais24, HT22, MN24, SL20a, ZBdAF20]. estimations [KZ23]. estimator [GQZ20]. estimators [CKTP20, Skr21]. Estrada [ALMZ^{+}20, BD21a, FYS^{+}23]. Euclidean [Alf20a, BLS20, Est23, GS20a, Gow20, JJL20b, SL20a, Sos23]. Euclidean/Log [Min22]. Euler [JR24]. evaluated [BHBO21]. EVD [YOV21]. even [BBG21, Des24, Goe21, LY23, Uru22]. eventown [AC22]. Eventual [GH23, JWZ22]. Every [AdlC20, BS21a, Ban22, Est23, AT21, Sha22]. evolution [CGG^{+}21, Pan21, SZ22b, WZ23b]. Exact [CZ20b, KST22, Li20a]. exactly [KAGH24, LWH21]. exactness [LV22b]. examples [BS20g, CLS24, KUZ22, Sło20, Sło21a]. exceed [LWB20, Sta20b]. Exceptional [dSP22, dSP23b]. exclusion [BS23b]. Existence [LS24a, GLZ21, LBW23, Mar24]. Expansion [Han21, Dom20, WXL24]. expansions [VCR21]. expansive [BS24b, MS21, Ran22]. expectations [EM23]. Explicit [ET20]. exponent [ZM21]. Exponential [KNiO23, Riz23]. Exponential-type [KNiO23]. exponentiated [CWW24]. Expressing [HLZ17, Hou20]. expression [IIM^{+}21]. Extended [SLTW22, KNiO23, MP23b, RRF23]. extendibility [CDPR23]. extension [Ali20, BM23, HAD19, KK20a, Pik23, SD21c, ZvD23, ZG21]. Extensions [BS22, LF20, Li22, Mao20, AT24, BPS23, CNO23, GvRR^{+}23, MDH23, XY20a]. exterior [Kor21, LL24a]. extrema [CRR20, LY23, WHM23, ZY24]. Extremal [ALMZ^{+}20, JT24, MR21b, SC23, WCF21, WHL24, ZKLS20, APLN^{+}23, ABR21, BS20g, CKLL20, CP23, DKL^{+}22, Gol20, JGW24, KLLW21, KKP21, KS20b, LW21b, LJSZ23, LZ21b, MLZ24, TCC21, XHW23]. Extremality [SlW22]. Extreme [Yu20, BH24, CCK21a, Shi22]. extremely [HH23a]. extremum [CWYZ23].

Faces [CC22a, May21]. factor [KCPMNS21, ZLLZ22]. factorability [DdR20]. factorial [RC23]. Factoring [Zha22]. Factorisatio [dC21]. factorisations [HLS21a]. factorizable [LR21]. factorization [ACLS20, BFMM23, GW24, GD20, JLX22, MPR24, NGK23]. factorizations [AKN23, BLS20, Chr24b, SDS21, Var21]. factors [CMP21, LLA23, ZL23c]. Faddeev [Bär21a]. Faddeev-LeVerrier [Bär21a]. False [Bas21]. False-twin-free [Bas21]. Families [ES20, ABH^{+}23, ABR^{+}22, Bou21, CGP22, VS21]. family [BF24b, FN23, GZ23, Oli21, QQM24, Shi22]. Fan [KLSP24]. Fast [CHHS20, MC21a, MN24, Bol21, MB20]. Faster [OSC20]. Faulkner [AO22]. feature [Dax23]. February [Ano20-68, Ano20-67, Ano21-71, Ano21-69, Ano22-54, Ano22-51, Ano23-63, Ano23-71, Ano24-38, Ano24-34]. Feedback [Sch22, WES21, ZK21]. feedbacks [GS21b]. Feller [YZY24]. few [ABSV21, ABR21, Zha21]. Fibonacci [ACOM21]. Fiedler [AD24, Cia20b, SMA21, Sos23]. Fiedler-type [Sos23]. field [BHBO21, CGMP23, FG22, GvRR^{+}23, HS23a, Hom20, HHS20b, MMR20, SZH22]. fields [Adr22, ADHM23, BDRS20, BFO23, BM23, BAPF23, CCH^{+}21, CLNY22, DG22, Fid22, GIJM22, GJKM21, LM24, PW19, PW20, Pri22, Wer22, ZHSX20]. filiform [BBNS22, dMdSS23]. filling [Hom20]. finder [TV21]. Finding [Kov21]. Fine [AOCM21, NZ24]. finitary [FP22, Vit21]. Finite [CM23a, Adr22, BM23, BL23, BHKW22, CCH^{+}21, CLZ22, CM23b, CV21, CLNY22, CA22, CER^{+}23, Coh20, CR21, ĆD20, ĆD23b, DR20, DG22, FGPW24, FMS21, Fra20, FG22, GNP22, Got20, GIJM22, HS23a, Hom20, IIM^{+}21, Ime23, KY21, LWW22, LM24, MA24, Mor20, NPR22, OQ22, O’L23, PPŻ21, PPŻ23, PSW23, dSP23a, PW19, PW20, Pri22, Rei23, Szc21]. Finite-dimensional [CM23a, ĆD20, ĆD23b, Ime23]. finite-rank [Got20, PPŻ21, PPŻ23]. finitely [CN20b, JL22]. First [Mis20, SMM24, AY23, YD21, ZLG20, Zha24a]. Fischer [BS23b]. five [Kub22]. Fixed [BR20a, Bas21, Bib22, BMM24, CT23, Cos24, DGL23a, DPR20, HXLL21, KGYY23, LTYZ23, LTWW23, Lim22b, YK21, ZG22]. Flag [AGNPSE23, NPSE24, Pat20]. Flattening [CST21]. Floquet [Szc21]. flow [DJ22]. flows [KLS23]. focal [ŠH21b]. Foguel [GWW21]. footprint [ACM22]. forbidden [FY23, LW24]. forbidding [CLZ20, LY23]. forcing [KL21b]. forest [BEF20]. forests [CLZ20, CWYZ23]. form [ACM22, Arl20, BS20d, CFSiT20, GvRR21a, HLS21a, HVX23, IGW17, Li24, MP23b, Nay23, PK22, RMG24, SZ23a, Sta21, SZ20, Wan21, Wan23, BS20e, CLLK21]. formal [Zha23]. formats [Buc20]. forms [AC20b, Chr24b, Com21, Dew21, Ger20, GCOP24, HLYZ21, HLYZ23, KLLS22, May21, MZ22, PSRS20, RSS21, dlCS21]. formula [BBG21, BS23b, CZ22b, CB21, CMM22, GS22, LV22a, RRF23]. formulae [MPS22]. formulas [CWW24, DN21, KS23c]. formulation [Min22]. Fornasini [BCR^{+}23]. Four [GLNW24, AJNS24, BS20e, BFF^{+}23b, GHL^{+}24, HLS24, LLW21, LBW23, SSS21]. four-point [BS20e]. Four-vertex [GLNW24]. Fourier [HW21a, LDW21]. fraction [HMS21]. Fractional [GZ22, HXLL21, LLA23, Mon23, CCD^{+}22, WZC23]. Fragments [Hom20]. frame [NS22]. framed [ABHR22]. framelets [AKPS19, KAPS20]. Frames [GIJM22, ANAS20, AP23, DO22, FS20, LL21d, LM22b, NS22]. framework [BF21]. frameworks [KNS20]. Fréchet [Sch23, Shi20a]. Fredholm [Djo21]. free [AM23a, BES21, Bas21, BHKW22, Csi22, Got20, LNP21, LLW23, LMX23, MT22, MLH22, WCF21, WHL24, Wan24, Zak21, ZY24]. free/ [MLH22]. Frequency [BW21]. Frequency- [BW21]. Friedland [CLwW22]. Frobenius [CKOS22, KZPH22, KPP21, Lin23, TXF24, XY20a, ZC23]. Fucík [Mar24]. Full [TW21, HH23a, LNP21]. fully [Tan23]. function [FKMM21, GvRR^{+}23, IIM^{+}21, KW20, Kli20a, KKS21, MQW20, WXL24, Zhu21]. functional [Bay24, EMŠV22, Rei23]. functionals [RRvW24]. Functions [BCR20, AS20b, AL21, AL23b, AV23, AdF20, Beh21, BO23, Car21, FSRVSST22, Guo22, GJKM21, JLM20a, JLM20b, Kad20, KS23b, Li20b, LLH20, LWW22, Mor20, Nag23, ST22, Sch23, Shi20a, Skr21, SC20b, Yaz21, Zak21]. Functors [Ago24]. Fundamentals [CCD^{+}22]. Further [DYLL23, GPRT20, Nie24, SZ23b, ZT23, ZBCI21].

Gabidulin [BZZ22, Ner20]. Gaddum [CLM20]. Gain [CDD21, BBCD21, BBC^{+}22a, HHD20, KKP21, Kha24, LW21c, LWH21, SK21b, XZWT20]. Gain-line [CDD21]. Galerkin [KHU23, Pra20]. gap [Ces20, CMS21]. Gateaux [Sin21]. Gau [CRSY22]. Gauss [BWW21]. Gaussian [BAPF23, SG20]. GCD [GR21]. general [BP23a, CEM21, DYB20, DW21, HCZ20, HBHZ24, JST21, PQ22, PV21a, TL21, AS20a, Pre20]. Generalization [AOZ23, KP21, CKOS23, De 20, MG21, SMA21, ZW21]. generalizations [BD21a, Fid22, HVX23, ZJCA20]. Generalized [BCR20, DF20b, GQZ20, Ham23, LWYL20, MQW20, PPŻ21, QJMW23, QQM24, Ste22a, ADHS23, AKVS23, AHV20, ABG20, AO22, APLN^{+}23, BGGA22, BDK21, BHBO21, BL22, Beh21, Bur21, CWW24, CV21, DG20, DLL23, FB21, Gor24, GZ20c, Har22, HMPT20, IIM^{+}21, JGW24, JLM20a, JLM20b, KNiO23, KMP23, KLR22, LL21c, MMP21, Ner20, NN24, PGRU20, Pra20, QWW21, Roy24, SMA20, SS21c, YW22, TT21]. generalizing [Cia20b, LWZL23, LL24b]. generated [BDHS23a, Kol21, LW21a, XTLK23]. Generating [BS21b, BS20d, Yaz21]. Generation [EV20, GNP22]. generators [CFKO21, dBS20]. generic [BBS20, Com21, KK20a]. genetic [Pan21]. genus [BDV21]. Geodesic [DF22]. Geometric [Got20, LS24b, CMM20, DY23b, EHS22, FH23, FJ24, GLT21, HS23b, LMW21, RRvW24, SKS22, TJ21]. Geometries [CIM22, Adr22, SZ21, SZ22a]. Geometry [Win21, AG24, BPZ21, CCH^{+}21, FK21, GIJM22, MS22, NXS23, LR23]. Gerstenhaber [dSP20b]. girth [AIK^{+}24, DY24, HLX20, HXLL21, Kha24, WLT22, ZWT21]. given [AMR20, AIK^{+}24, BP20, CZG21, EG20, FY23, FW22, HLX20, HLN24, HLS21b, HLS22, JLW22, LJSZ23, LZ20, LLX23, LW24, MLH22, Ros22, SWW21, XBSF23, YBW24]. global [CZ20a, MORŠ21]. globally [WW20]. GMRES [CLS24]. Godsil [BBCD21, QJW20]. going [MZZ24]. Gordon [FGPW24]. Gorenstein [Alt22, MRT20]. gr} [FK23]. gr}-PI [FK23]. Graded [CDdM22, DRG24, FGDY20, FKM22, Fid22, Gor24, BCS24, CdSV22, DdR20, GLNW24, GFD21, dAGK21, Iop20, IdSSV21, RT21, Yas23]. gradient [DAG21, FGM21]. grading [CdSV22, FGGG24, FKM22]. Gradings [AOCM20, GLAdS23, AOCM21, BDvW24, CDdM22, DGET21, DP21, Eld22, FGGG24, GdM23, Gor24]. grafting [LC20, XW21]. Gram [CLRT22, KK22b, May21]. graph [AAFM23, AFSS20, AAFT22, AD24, AA21, ACDG21, BBG21, Bol21, CLS20, CPS20, CLM20, CG22, CvDB23, CLL20, CLT22, CMP21, Csi22, DJ23, DAG21, ET20, FHLS22, FM22b, Fio20, GH24, GK21, GZZ21, HHD20, HLN24, IIM^{+}21, KKSV21, KIS21, KK22c, KK24, KS22c, LZ21a, LW20a, LMG20, LOŠ23, LW21c, LW21d, Mar21, MLZ24, MS20b, MH20, MH21, MP23b, MPRS21, PŠ23a, QJMW23, SK21b, SS21c, Sta23b, SD21a, SD21d, TW22, Van24a, WWT20, WWJ20, WFGT20, WG20a, WG20b, WYQ21, WZW^{+}21, WWG22, Wan23, WFL20, WZFW23, XWT20, XLS20, YWF21, ZXZW23, Zhu21]. graph-regularized [DAG21]. graph-related [Van24a]. graph-theoretic [Mar21]. graph-vectors [QJMW23]. Graphical [Gol20, NS21]. graphons [Mis22]. Graphs [ABSV21, ABR21, CL22, CT23, LWB20, WQP23, Zha21, ZL22, ZW22, AHZ22, AAHV22, AFK^{+}23, ADHS23, AJM21, AJM24, AAF^{+}21, ASS21, AFSS20, AIK^{+}24, AGH20, AFG^{+}21, AAAH22, AL23a, AdL21, AdLN23, AT20, Alf20b, ABG20, ALMZ^{+}20, ABR^{+}22, ARS24, ADvdH21, AHLvdH21, ADvdH23, ADH^{+}24, ASG22, BP21a, BV24, BHHP21, BEHK22, BFF^{+}23b, BFF^{+}23a, BES21, Bas21, BAPF23, BL22, BL23, BC21, BBCD21, BBC^{+}22a, BP20, BDVR21, BS23b, CZG21, CW22, CCPKP20, CGP22, CDD21, CK23a, Ces20, CFK^{+}20, CCD^{+}22, CDS21, CLZ20, CH21, CWZ22, CLM22, CHHS20, CN24, Cia20b, CP21, CDG21, CDG22a, DF20a, DDFM^{+}21, DS22, DKL^{+}22, DPR20, DGS^{+}21a, Du23, DdF23, DYLL23, DH23, DY24, Ere21, EGK^{+}24, EG20, EHS22, Est20, FCLP22, FY23, FLRS20, FW22, FM20a, FS23, Gan23, GPRT20, GR22, Ger20, GLHNW20, GK22b, GM20]. graphs [GL20b, Goe21, GKLZ20, GKK^{+}20, GLZ21, GLMM23, HO20, HT23, HSS^{+}21, Han21, HLX20, HXLL21, HP22, HBHZ22, HWD23, HBHZ24, HLS21b, HLS22, HD22, Hua22, IKPR20, IMT20, JMP21a, JGW24, JZZ24, Kad20, KB22, Kad23, Kak20, Kha24, KGYY23, KAGH24, KS23c, KMNP20, Kub22, Kub23, KLR22, LW21a, Lee20, LOŠ22, LW20b, LLW20, LC20, LLH20, LLW21, LQL21, LQ21, LTYZ23, LY23, LJSZ23, LOS20, LOŠ21a, LG21a, LLW23, LWYL20, LDSS20, LHH22, LMX23, LLX23, LW24, LZ21b, LLA23, LWH21, LLL22, LDW21, MA24, MMN22, VJ24, MLH22, Mło22, MR21b, Nic24, NdLO21, O.21, OJQ20, Obo22, Obo24, OPPS22, OdFVM21, PPŻ21, PPŻ23, PGRU20, PK21, PV21b, QLCM20, QJW20, QWW21, QJMW23, QQM24, RFP22, Rei23, RRvW24, RDFM23, Row20, RS21b, SMA21, SS21b, Sic21, Sim21, SZ21, SZSMB20, SSS21, Sta20b, SD20, SD21b, SLL22, SC23, Tan23]. graphs [TTV23, TLC20, TWW20, TCC21, TW21, TT24, Tur21, UZ21, VTBC20, WH20, WCF21, WHM23, WLL23, WZC23, WHL24, Wan24, WLT22, WFY24, XZWT20, YK21, YW23, YLZH21, YD21, Zak21, ZXL20, ZWF20, ZY24, Zha22, ZG22, ZWW23, Zha24a, ZHG21, ZHW21, ZL23b, ZWL20b, ZWL20a, ZBWL20, ZWT21, ZvD21, ZL23c, dLMO20, JLW22]. graphs/networks [Est20]. Grassmann [ADHS23, FGGG24, GFD21, dAGK21, Lee20, LIW20, PPŻ21]. Grassmannian [HPSZ21]. Grassmannians [Pan20]. greater [LLL22, NdLO21]. greedy [KHU23, LG21b]. Green [GJKM21]. grid [DL22]. Grothendieck [Chr24a]. group [ASG22, BMR23, BBCD21, BR24, BS21b, CEJM21, CDD21, Ces20, Cob22b, Coh20, DHR20, Eld22, EV20, Gor24, KLS21, LW21a, LWW22, Oli21, PR22b, SS24a, Sło21b]. groups [BSS23, BCNS22, BFO23, BL22, BL23, BDHS23a, CK23a, CXZ22, DF20a, DZCY22, DdR20, DLL23, Dra22, FKK21a, GG21, GLZ20, HWK21, JJL20a, KB22, LM24, Pan23, Pre20, RT21, SZ22b]. Grover [IMT20, Kub22, KS22c, SMA20]. Growth [Bui21, BCS24, GQZ20, IdSSV21, Riz23, Ros22]. Guarantees [MB20, HJBKR24].

H [CWBB23]. H-shape [CWBB23]. Haagerup [Ali20, HAD19]. Hackbusch [Buc20]. Hadamard [FJS21, AAF^{+}21, BBCK24, BS23b, CFK^{+}20, DD23a, FJS17, HW21a, HY20, HK20, Le22, LH20, LS20c, MSX22, RW21b, Tei22]. Hadamard-type [Le22]. Haemers [ACG^{+}21, De 20]. Hafnian [KKT22]. Hahn [FG23b, Lee20]. Half [He20]. Halin [CLM22]. Halmos [XY20b]. halves [SS23c, YY23, YZY24, ZZJ23]. Hamburger [FKKM22]. Hamilton [AS23, Mur22, SSvW20, ZWL20b, ZWL20a]. Hamilton-connected [ZWL20b, ZWL20a]. Hamiltonian [Ayd23, FCLP22, FGPW24, GS21b, LWYL20, MMW21, SS21a, ŠH21b, ZS22b, dlCP20]. Hamiltonians [Lim22b]. Hamiltonicity [FKK21a, ZBWL20]. Hamming [BMPS23]. hand [Hal20]. Hankel [Bou21, Dom20, DG22, FGM21, GW22, KPP21, SS20]. Harary [CC22b]. hard [DL22]. Harmonic [FS20]. Harmonious [XY20b]. Harnack [JL20, YZ20]. Harpe [Ali20, HAD19]. Hartfiel [Mao20]. Hausdorff [FKM20]. having [CDG22a, GZZ21, LNN^{+}20, TW21]. Hecke [Wie22]. Heisenberg [FKM22]. Hellinger [DF20b]. helm [Goe21]. hereditary [AT24, Yan22]. Hermite [CLLK21]. Hermitian [BN23, BLMT21, BR22, BDI20, CN22, FM20b, GM23, GM20, GL22b, IB22, Kli20a, LR23, MSZ21, MRV20, Moh20, Sta21, SZ20, ZFZ20]. Hermitians [DM24]. Heron [MS24]. Hessenberg [BFMM23, CCGV^{+}20, Lee22]. Hessian [Yaz21]. hierarchical [IB22]. High [ZBdAF20, BBS23, XJLL23]. High-order [ZBdAF20, XJLL23]. Higher [Kak20, RMG24, BdPS21, HX24, HLYZ21, PW21, SH20, Sch23]. higher-order [Sch23]. Higher-rank [Kak20, BdPS21]. Hilbert [ANAS20, BM22, DP23, FG23a, Fra22, MPR24, Min22, OM20, Pan20, XY20b, ZW21, Zam23, hB21]. Hilbertian [BKPS23, MXZ20]. Hilton [WXY24]. hitting [LV22a]. Hoffman [Hae21]. holds [Roc20]. Homogeneous [Lim22a, BGOO21, DG20, DR20, FGGG24, HCZ20]. homomorphic [Tsa23]. homotopy [KW20, NS22]. Hong [Sta21]. Horadam [ACOM21]. horizontal [MG21]. Horn [BJ20b, SZ22a, ZX20]. Householder [CP20]. HS [KB22, Kad23]. HS-integral [KB22, Kad23]. Hubs [Est20]. Hubs-repelling [Est20]. hulls [CK21, Cos22, JJL20a, KS20a]. Hurwitz [AOCM21, BBCK24]. hybrid [HH23a]. hyper [AL21, CR23]. hyper-positive [AL21]. hyper-ranges [CR23]. hyperbolic [AG24, BG20, SS24b]. Hyperbolicity [GL22a]. hypercontractivity [JKX20]. hyperdeterminant [AY23]. hypergeometric [BSS23, VS21]. Hypergraph [GFW24, CB21, CvDB23, CTS23, CZ20b, SWW22, WSW20, WS22, YYS^{+}21, Zhe20, Zhe21]. hypergraphs [AMZ21, Ban21, BP23a, CHT20, CDVPT22, CC22b, Coo20, CO24, DW21, FYS^{+}23, GRR22, HCZ20, KLLW21, LL21c, LYZ23, MV24, SSP22, SB20, SWW21, XW21, XHW23, ZLG20, ZZ21, ZKLS20]. hypermatrices [GF20]. hyperovals [DG20]. hyperplane [HMS21]. hyperstars [DvDW23]. hypersurface [ÁVKP20b, BR21]. hypersurfaces [DFH^{+}20]. hypertriangles [DvDW23]. Hypoenergetic [ADGKM21]. hyponormality [BCE21]. HZ [Har22].

ideal [PSW23]. idealization [FP22]. Ideals [VMB22, AT20, Alf20b, ABSV21, Bot21, CDG22b, DJ22, JLM20a, JLM20b, KLR22, LNN^{+}20, LP21, Mań22, Wer22]. Idempotent [BDSS21, BM20, Bri22, HL20, SZ22b, TXF24, WZ23b, XTLK23, ZTX24].

Idempotents [EAHS21, BR24]. identifiability [AC20b]. Identifying [HW21a, Mań22]. Identities [CdSV22, LP22, BC23, CDdM22, DdR20, DGdSS23, DRG24, FKM22, Fid22, GFD21, dAGK21, Nar23, Uru22, dMdSS23]. identity [SSvW20]. Ihara [IIM^{+}21, SC20b]. IHS [OPA23]. II [AS20a, CG20, Cra21, KS23b, KP22, LL24b, SS22, Sim21, SMM24, Wój22a]. III [dSP20b]. ILAS [AHKT21]. image [WZL21]. Images [Vit21, BHBO21, CLW22, dMdSS23]. Immanant [BS24a]. immanants [CK24, DGS21b]. imperfect [Los21b, Sar21]. Improved [AL23a, Bhu24, FQ20]. improvements [KZ23]. impulsive [SLTW22]. in-betweenness [DF20b, DF22]. In-evolution [Pan21]. Incidence [GRR22, CCKM22, DHR20, FP22, GK22a]. inclusion [BS23b, Cos24, YLCB21]. inclusion-exclusion [BS23b]. indefinite [Boj21, BKS^{+}21, Kov20, Ran22]. independence [ABC^{+}22, CDG21, Fio20, HH23b, HLN24, Nic24]. Independent [Pat20, GLP21]. Index [LH21, AA21, BL23, CT23, CHY20, CMR20, DY24, FYS^{+}23, GMR22, HD22, JKS23, JLW22, KL21c, LTYZ23, LMX23, ŠH21a, Šte22b, SSvW20, TCC21, ZL22]. indices [ADMZ21, ALMZ^{+}20, BPZ21, FLRS20, LW21b, WFL20]. induce [Kub23]. induced [GLAdS23, GOUY23, Gow20, ST20, Sra21, Yam21]. inducibility [MDH23]. Inequalities [JLM20a, JLM20b, ANBK22, Ala21, AHK23, BDK21, BDN24, BS22, BP21b, BP23b, BL20, CZ22a, CLM20, Chr24a, ES20, FH23, FLT20, GAM20, Gow20, HW20, HVX23, Ime23, JJL20b, Le22, LS20a, Li20a, LH20, Li22, MPS22, MJ24, MKE^{+}20, MXZ20, PLT23, RT23, SKS22, SZ23b, YZ20, dVLG23]. inequality [Ali20, BMS20, BS23b, CKOS23, DA23, De 20, Dru24, GOZ21, GAMA20, HAD19, JL20, KMP23, LF20, LS20c, Mao20, Naj20, PPW20, QCCS21, RW21b, Sos23, ZJCA20]. inertia [DY24, WFL20]. Inertias [BCOvdD23]. inexact [WCS20]. infinite [Bez22, CG20, Fid22, GJ23, HLZ17, Hou20, Men22, Rob20]. infinite-dimensional [CG20]. infinity [WES21]. Information [MS22, BS20b, FT23]. Inheritance [CF24]. injective [LCH23]. Injectivity [GRR22]. inner [BKR^{+}20, BNS22, BDV21, CS20, Fek20, Ran22]. inspired [BS20e]. Integer [AJM21, HLS21a, AJM24, Ans21, KS20b, Kop20, NPkT21, PV22, Sim20, VS20]. integers [CLS23]. Integrability [DGKP24]. Integral [BDVR21, Ger20, GKLZ20, PV21b, Sch23, WH20, ÁVKP20b, ABR^{+}22, BL23, KB22, Kad23, MZ22, NdLO21, OPPS22, Sha22, dLMO20]. integrality [WFY24]. integration [NS21]. interlacing [CZ22c, LZ21a]. interpolation [Bol22, GS20a]. interpretation [PV21a]. intersecting [DKL^{+}22, Des24, HLS24, MLZ24, WXY24, ZHG21]. intersection [BMR23, BFPS21, CVV21]. Interval [CK21, Rub21, AG21, Bib22, MP20b, MP24, NN24, Obo24, SL20a]. Intervals [CS21]. Introducing [ZS22a]. invariance [GH23]. Invariant [Lim21, ANBK22, AG15, AG20, Che22, CKLL20, CMP21, DLL20, GG21, GGGD23, KK20a, KLLS22, LR23, NNSS20, RMP21, TP23, WW20]. Invariants [GCOP24, ÁVKP20b, CL20, CN23a, Elm23, KY21, LM24, NPHT20, Oli21]. Inverse [CN22, DMS20, GK20, PSAPNV20, YGW20, AAMR20, APLN^{+}23, BBG21, Beh21, Bol21, CEJM21, CS21, CCK21b, CN21a, CESS21, FHLS22, FG23a, GR22, HLW23, JCD24, LOŠ21a, LOŠ23, MvWC23, RX21, SM23, SS21c, WCS20, dAdCG^{+}21]. inverses [Dra23, GZ20a, JMPS24, KLS21, MR21a]. Inversion [AHV20, PŠ23a, RS21a]. invertibility [BR24, Cos20b, Sch24]. invertible [FM22a, Loe21, dSP23b]. Involution [BDHS23a, BCS24, CdSV22, DRG24, HS20, Iop20, IdSSV21, PS20, Uru22]. involutions [AdlC20, BDHS23a, EV20, FGDY20, FP22, HLZ17, Hou20, HHS20b, RAK21, Sło21b]. Involutive [BW22]. involutory [CCK21b]. involving [AHK23, ACOM21, LLA23, ZWW23]. irreducibility [MG21]. Irreducible [CCU21, Wie22, Dra22, Hu21, Miy24, YLZH21, GHL^{+}24]. irregularity [LYZ23]. isolated [AHV20]. Isometric [CFSiT20, AV23, JJS20a]. isometries [BCG24, BG20, BKS15, BKSP23, CR23, CAEH^{+}22, GS21a, GL22b, IKLP20, IOT21, MS21, SMP24, Sán21]. isometry [BL20, Sat21]. Isomorphism [WZ23b, KK22b]. isomorphisms [CMM20, EE23, FP22]. isospectral [KSSW20, PV21b]. Isospectrality [PW21, vdMvSW24]. isotropic [Car21]. issue [GLT23, AHKT21]. iteration [KLS23, Rob20]. iterations [BKRS20, WW22]. Iterative [HH23a, ABCH23, DL22]. iteratively [CW21].

Jacobi [AM21, APLN^{+}23, CN22, IB22, SS20, VS20, XBSF23, YOV21, ZW20]. Jacobsthal [ACOM21]. James [AGK^{+}21]. Janeiro [AHKT21]. January [Ano20-69, Ano20-56, Ano21-68, Ano21-70, Ano22-67, Ano22-68, Ano23-60, Ano24-39, Ano24-29, Ano23-70]. JB [CWW24, CAEH^{+}22]. JB-algebras [CWW24]. Jensen [Sra21]. Jiang [WW20]. Johnson [ADHS23, AHM20, MB20]. join [AFK^{+}23, CGP22, SMA21]. joining [AS24]. Joins [SB20, – DMM^{+}22, LOŠ22]. Joint [Pik23, Bui22, CMS21]. jointly [BO23]. Jordan [AM21, AMO24, AHV20, Alt22, AO22, B– DK^{+}24, CCU21, CLNY22, DGdSS23, FP20, GS20a, Gow20, JJL20b, JJS20b, KT21, KS20a, Li24, PK22, RMG24, Sán21, Sos23]. Jordan* [CMM20]. Jordan*-isomorphisms [CMM20]. Jordan-Kronecker [KT21]. July [Ano20-58, Ano20-53, Ano21-63, Ano21-59, Ano22-59, Ano22-64, Ano23-54, Ano23-55, Ano24-36, AHKT21]. June [Ano20-59, Ano20-51, Ano21-51, Ano21-58, Ano22-65, Ano22-61, Ano23-51, Ano23-56, Ano24-30, Ano24-31]. just [Row20].

Kac [FKM22, Fik20, FP21]. Kaczmarz [BHD^{+}21, Dax23, LG21b, SLTW22, ZL23a]. Kahan [WW20]. Kahane [RT23]. Kantor [AOCM21]. Kaplansky [GdM23]. Karcher [GLP21, GL22a, GL23, LL21b]. Karpelevic [KŠ22b]. Karpelevich [KK20b]. Kato [AOZ23]. Katz [AHNW22]. Kemeny [JKS23, LL20]. kernel [DD23a, DY23b]. kernels [AG15, AG20, BFPS21, NWW20]. kind [Moh20, XX20]. Kippenhahn [BdPS21]. Kirchhoff [KL21c, RFP22]. Kiselman [Wie22]. known [KZ23]. Ko [Adr22]. Kraus [ST22]. Kronecker [FKMM21, KT21, MB20, PP20]. Krylov [CCS22, Rob20, SS21a, XJLL23]. Krylov-Lagrangian [SS21a]. Kubo [Gye23, LS20a, PV21a]. Kuiper [MT21a]. Ky [KLSP24].

labeling [LL21c]. labelling [SWW22]. Lagrange [Bur21]. Lagrangian [SS21a]. Laguerre [MPR22, WXL24]. Laman [BC21]. Lanczos [CT24, GP21]. Lanczos-like [GP21]. Landau [CLNY22]. Laplacian [AMZ21, AJM21, AJM24, ASS21, AAFT22, AvDF20, AADT23, AdL21, AWvW24, BR23a, BEHK22, BCE^{+}23, Bol21, CLS20, CPS20, CLZ20, CWZ22, CK23b, Coo21, DDFM^{+}21, ET20, Est20, FCLP22, GM22, GPRT20, GH24, GZZ21, GMSG22, HLW23, HLS21b, JT24, KKP21, LWB20, LLW21, LQL21, LHH22, MK23, MT21b, MZZ24, Nic20, Obo22, PK21, SSP22, SSS21, Sta20a, Sta23a, SD21b, SC23, TWW20, TW21, VTBC20, WYF^{+}20, WZW^{+}21, WMCQ23, XW21, XY20c, XZ23, XLYS20, ZXL20, ZG22, ZHG21, ZWL20b, ZWL20a]. Laplacian-type [LLW21]. Laplacians [Dev22, KLR22, LLWW23]. large [AvDF20, BFLT23, BW21, BF24b, CHHS20, cRLY20, MSX22, ZT23, ZWL20a]. large-scale [BW21, BF24b, CHHS20]. larger [dLMO20]. largest [ABG20, Csi22, LWB20, WWT20, WQP23, XLS20, YK21, YW23, YWF21]. lattice [ACDG21, KS22a, PGST23, Pol20, Ram23]. lattice-subspaces [Pol20]. lattice-theoretical [KS22a]. lattices [FK21, GGGD23]. law [KNiO23]. Lax [ACSS22]. LCM [GR21]. learning [ÁVKP20b]. least [Boj21, CW21, CKTP20, Ghi20, Hal20, LDSS20, RC23, SLTW22, SC20a, Sta20a]. least-squares [Hal20]. Leavitt [BS23a]. Lefkovitch [APNP^{+}21]. Lefschetz [MRT20]. left [AL23b, DHH^{+}20, LS23]. left-star [DHH^{+}20]. Leibniz [CNO23, LM22a, dMdSS23]. lemma [FHLS22, FGPW24, SMA21, KS22a]. length [JMP21a, RRvW24]. Lengths [Mig21]. Lenstra [FG22]. Leslie [PSAPNV20]. less [WQP23]. level [Sch23, Vol23, XX20]. level-2 [Sch23]. LeVerrier [Bär21a]. Levinger [AC20a]. liberation [LOŠ23]. Lidskii [ŠH21a]. Lie [CdGS20, HMS21, ABCO22, BBNS22, BFLT23, BF24a, BRL23, BE20, BDV21, CWW24, CdGS12, CZ20a, CMR20, DM24, Eld22, GG21, GJV20, KK23a, KY21, KLLS22, LNN^{+}20, LS20b, LP21, MDH23, PGST23, QSH22, Rai23, SSvW20, Yas23]. Lieb [AS20b]. lifting [BPS23]. liftings [MS21]. lifts [WZFW23]. like [AKVS23, ACOM21, CGJvB22, Dra23, ES20, GM22, GP21, WCS20]. limit [Bri22, Zem22]. limitations [BEFSC22]. limited [BW21, DK21]. limiting [RW22]. limits [BC21, CN20b, CR21, Mań22]. Lindbladians [Szc21]. Lindenstrauss [MB20]. Line [YW23, AAAH22, AL23a, CZG21, CLZ22, CDD21, De 20, HLS24, OdFVM21, Onn21, VZ20]. Lineability [FSRVSST22]. Linear [AAH^{+}22, AD20a, Ali20, AG20, BP23b, BDHS23b, BKSP23, CdGS20, Cos20a, Cos21, Cos22, Cos24, FJS21, FHMP22, GS21c, GJKM21, GS23b, Hou20, HH23b, JLM20a, KAPS20, KLSP24, LTWW23, MH21, PW20, QSS22, Sło21a, SD21a, ZMP23, tHvdM21, AM23b, Ais24, Ans21, Arl20, ABCH23, AR21, BPZ21, BEF20, BDN24, BDHS23a, Boj21, BKS^{+}21, BS21b, BK22, BS20g, CLZ22, CVV21, CD23a, CLZ20, CWYZ23, CI21, CMS21, CZ20a, Com21, CMM22, DN21, DJ23, FM22a, FQ20, FSRVSST22, FT23, GP21, GL20a, GJ23, GLT23, GS24, HLW23, HPSZ21, KS22a, KK20a, KHU23, KCML20, Kye23, LPS23, Lew21, LG21b, MAG21, MG21, MC21b, MKE^{+}20, MPS24, NPkT21, NPHT20, dSP23a, Phu24, PLT23, PS23b, Pre20, Rei23, RW22, SDTR23, Sch24, ŠH21b, SL21, SNK22, ZK21, AHKT21]. linear-algebraic [KS22a]. Linearity [CKL24, Sán21]. Linearization [VS23]. Linearizations [PQ22, ADMZ21, DMQD20]. linearized [MP22, Oli21, PZ20, dVLG23]. linearly [GLP21, Pat20]. lines [BS20f, FP21, GIJM22, HLS24, cRLY20, SS23b, Zal22]. Link [BBW20]. linked [BS20f]. Linking [Mor22]. Lipschitz [AMO24, Car21]. list [NPkT21]. Littlewood [KK20a]. lively [SMA20]. LLL [CLLK21]. Local [ÁVKP20a, CRR20, DMQD20, MORŠ21, BC21, BMM24, CZ20a, JJS20a, SMP24]. local-global [CZ20a]. localization [Che22, GP20]. Locally [AM23b, BF24a, FMS21, dSP23a, Ste21b]. Locating [RB22]. LOCC [KMNP20]. Log [Min22]. Log-Euclidean [Min22].

Log-Euclidean/Log-Hilbert [Min22]. Log-Hilbert [Min22]. logarithmic [Fra22]. long [CWZ22]. look [KS22a]. loops [IMT20]. loss [CPTP21, FKMM21]. Lovász [AC22]. Low [DJ22, TL21, ZRQK22, BF24b, Bud24, FGM21, JMPS21, JCD24, KKM22b, KST22, Li20a, LHZ21, MN24, MO20, SK21a]. low-dimensional [Li20a]. Low-rank [TL21, BF24b, Bud24, FGM21, KKM22b, KST22, LHZ21, MN24, SK21a]. Lower [ASS21, GMR21, MP21, Sta20a, ZZ21, AOA23, AGH20, AL23a, BS20f, Dew21, GZ20b, KL23, Lim22a, LZZ23, Loe21, MT21b, Rum22]. lower-triangular [Dew21]. Lu [GLLZ20]. Łukasiewicz [YY21]. Lupas [MMV22]. Lupas-type [MMV22]. Lvov [GdM23]. Lyapunov [BF21, GP20].

M [OPA23]. M-IHS [OPA23]. Macaulay [VD22]. Macphail [Tei22]. Magic [HHS21]. magnetic [FCLP22, KS23c]. main [AFG^{+}21, FB21, VTBC20, YW22]. majorisation [BS22]. Majorization [DGS20, DGS23, Gow20, GS23b, JJL20b, YZ20, vED22]. majorizations [GS21c]. Manifold [ÁVKP20b, GSS24]. manifolds [ÁVKP20a]. many [CN20b]. map [CM23a, CN22, CN23b, LS24a]. mapping [AFS^{+}23, GKS21]. mappings [dSP23a, Sch24, Shi20a]. Maps [BMM24, HP23b, BH24, BFMS21, BSC23, BHMU20, BKT23, BS20g, Bui21, CDPR23, CMT21, CL24, Cos20a, Cos20b, Cos21, Cos22, Cos24, DN21, FOR22, GJV20, GS20b, GMP22, HH23b, JJL20b, JS22, KCML20, KLSP24, Kye23, LV22a, LL24a, LL21b, LR21, LHFL20, LTWW23, LP21, ST20, SS23b, tHvdM21]. March [Ano20-70, Ano20-54, Ano21-64, Ano21-66, Ano22-50, Ano22-56, Ano23-49, Ano23-59, Ano24-28, Ano24-32]. Marchesini [BCR^{+}23]. Markov [BS20a, BAPF23, CK23b, CR21, EGJP21, HX24, LL20, RMP21, TTV23]. mass [VS20]. Master [KKT22]. Matching [FCLP22, GJ23, ZvD23, ADH^{+}24, HHD20, LZ21a, LLA23, WZW^{+}21, WHM23, WWY23, YYS^{+}21]. matching-minor [ADH^{+}24]. matchings [GLMM23, O.21, ZWW23, ZHW21]. mates [ADHS23, KK22b]. Matrices [FdCP20, KLS21, Mel23, AFK^{+}23, AG21, AP20, AKL^{+}21, ACG^{+}21, AMO24, AKVS23, ANBK22, AS20b, Ala21, AdL21, Alf20a, ADMZ21, AAMR20, ACOM21, ABW22, APNP^{+}21, APLN^{+}23, Ayd23, BS20b, BDNO20, BS20d, BBS23, BDK21, BPZ21, BN23, BEF20, BS20f, Bar21b, BDGY21, Bay24, BAPF23, BHBO21, BdPS21, BJ20a, BC21, BR22, BCOvdD23, Bez22, BJ23, Bib22, BDHS22, BDHS23b, BDvW24, B– DK^{+}24, Bol22, BEFSC22, BFPS21, Bou21, BMM24, BL24, BCP24, BFMM23, BM20, BR23b, BD21b, Bud24, Bui22, BLS20, BS24b, Bur20, Bur21, CRR20, CRSY22, CCU21, CC22a, CEJM21, CDM21, CL20, CDdM22, CS21, CD23a, CHY20, Che21, CZ22c, CT24, CJ20, CCK21b, CCP22, CCS22, CCKM22, CKKR23, CLNY22, CLS23, CN23a, CKLN23, CN23b, CW20, CCO23, CMT21, CIM22, CT20, CK21, CESS21]. matrices [CP22, CK23b, CKM20, Cos20a, Cos21, Cos22, CGMP23, CdSV22, CS20, CCL21, DGS20, DGS23, DGKP24, DY23a, Dew21, DJ23, DRG24, DMQD20, DGS21b, DF20c, DF22, DG22, EGJP21, EJS24, ES23, EHS22, EGV20, FCLP22, FR21, Fik20, FP21, FFSW23, FLT20, FG20, FG22, GS21a, GZ23, GA21, GHL^{+}24, GR21, GW22, Gau24, Gil20, GS21b, GK20, GM20, Goe21, GT21, GvRR21a, GvRR^{+}21b, GD20, GPR20, GMSG22, GZ20b, GP20, GS23b, HP23a, HW21a, HS23a, HS23b, HS22, HLZ17, Hou20, Hu21, HH23b, HL20, JLX22, JZZ24, JMPS24, JR24, KS23a, KT21, KKP21, KK20b, KK22a, KŠ22b, Kli20a, Kli20b, KPP21, Kol21, KFS21, KN20, Kus23, LR23, LV22b, Le22, LS23, LOŠ22, LLW20, LF20, LH20, LLW21, LW21b, LTWW23, Lim21, Lim22a, LZZ23, Loe21, Loe23, Los21b, Lub22, LMR21, Lyu20]. matrices [MPR22, MPR24, MA22, Mao20, MJ24, MMV21, MMV22, MRZ22, MR21a, MSZ21, MORŠ21, MS20a, MT21b, MC21a, MN24, MRV20, Men22, Mig21, Mil24, MT23, Moh22, Moh20, MSX22, MSZ23, MP24, Nar23, NS21, Ner20, NXS23, NV23b, OQ22, OW23, ORŠ23, OP21, PSW23, PŠ23a, dSP22, dSP23b, PQ22, PSAPNV20, PW21, PW19, PW20, Pri22, Ran22, RRvW24, RS21a, RKBA21, RSS21, RX21, Rub21, Rub22, SS21a, ST22, ŠH21a, Sha22, SM23, SS21c, Shi21c, Shi22, Sim20, SMM24, Sra21, SS20, SSvW20, Tan23, Tei22, TP23, Tom23, TBH20, VS20, Van24a, Var21, VMB22, Vit21, VB20, Wan21, WW22, Wei22, Xie21, YZ20, Yan21, Yaz21, ZS22a, ZM21, ZZ20, ZW20, ZFZ20, ZC23, ZBCI21, dC21, dHH23, dlCP20, dlCS21, – DMM^{+}22, vRvSTT21]. matricial [FKM20, FKKM22, Kim22]. Matrix [ABC^{+}22, BW24, CKKR23, CKLL20, Cra21, CMR21, GW20, Jon21, Kis22, KP23b, San23, ZK21, AT21, AOA23, ADGK21, AK20, ABG20, ACM22, ADHM23, AGV20, AdlC20, AS22, AJNS24, BGGA22, Bai24, BBG21, BS20e, BJP21, BR20a, Bar20, BR20b, BN23, BT20a, BFMS21, BDN24, BBS20, BS23a, B– DK^{+}24, BFPS21, BM22, BŠ23c, CP20, Car21, CZ22b, CK23a, CLW22, CAJ21, CLS23, CN21b, CMP21, CP23, CESS21, Chr24b, Cob22b, CER^{+}23, CMM22, Cos20b, DB23, DID22, DMS20, DZCY22, DdR20, DLL20, DP21, Djo23, DS24, DAG21, DL23, DN20, DN23, Dra22, DGS^{+}21a, DM24, EJ23, Elm23, Est23, FM22a, FMR20, FISW21, FGDY20, FM20b, FKKM22, FG20, FG23b, FJ24, FKK^{+}21b, GA22, GM22, GTW23, GW24, GKS21, GK21, GZ20a, GRR22, GvRR^{+}23, GMP22, Han21, HW20]. matrix [HLS21a, JJL20a, KL23, KS24, KCPMNS21, Kim20, KIS21, Kop20, Kov21, KP21, KLSP24, Kye23, LNN^{+}20, fLLV20, Li22, Li24, LDSS20, LHZ21, MMW21, Mel23, MC21b, ME21, MT23, MST24, MP21, MP23b, Mor23, MP20b, NGK23, NPkT21, NN24, Nic24, NV23a, O’L23, OPA23, dSP20a, PS23b, QCCS21, Rob20, RB22, RMG24, Ruž21, SK21a, SL20a, ST22, Sat21, Sch23, SNK22, Skr21, Sta21, Ste22a, SW23, TL21, DD23b, TXF24, Uch21, Uch23, Uru22, VS23, VD22, VB20, WCKQ20, WZL21, WWG22, WMCQ23, Wan23, Wer22, XY20a, XZ21, XWT20, XBSF23, Yan22, YZY24, ZG21, dVLG23, tHvdM21, vdMvSW24, GGGD23, KS23b, LLWW23]. matrix-sequences [CER^{+}23]. matrix-tree [GRR22, LLWW23]. matrix-weighted [Han21]. matrixwise [DFH^{+}20]. max [AdLN23, Jon21, KZPH22, MPS22, MP20b, MP24, SlW22, TV21, WFSS20]. max-min [MP20b]. max-plus [Jon21, MP24, SlW22, WFSS20]. Maxima [LMX23, ZL22, ZG22, JLW22, PSZ21]. Maximal [Ere21, GM23, APNP^{+}21, CZG21, CT23, LTYZ23, PSAPNV20]. Maximally [NZ24]. Maximally-fine [NZ24]. maximize [Li20b]. Maximizing [HLS21b, HLS22]. Maximum [ABW22, EG20, LLW23, LHH22, Yos21, AIK^{+}24, ADvdH21, AHLvdH21, ADvdH23, AJNS24, BCM21, Bud24, CLL20, CLT22, CKM20, FY23, GLT23, LW21c, SG20, WG20a, WYF^{+}20, WS22, YD21, ZTX24]. maxmin [MPS24]. maxmin- [MPS24]. May [Ano20-62, Ano20-61, Ano21-50, Ano21-52, Ano22-52, Ano22-53, Ano23-65, Ano23-66, Ano24-33, Ano24-37]. Mazur [BS21a, Ban22]. McCarthy [BL20]. McEnteggert [CZ22b]. McKay [BBCD21, QJW20]. Mean [LV22a, CMM20, ES23, FH23, FJ24, Ghi20, GLP21, GL22a, GL23, LL21b, LMW21]. means [BGOO21, CKLL20, CKL24, CIM22, DF20c, DF22, GLT21, Gye23, HK20, LL21a, LS20a, MS24, Nie24, PV21a, Uch21, Uch23]. Measure [BHHP21, EAHS21, HJBKR24]. Measure-theoretic [BHHP21]. measurements [KKM22b]. measures [AHNW22, FG20]. Measuring [AEGL21]. median [MC21a]. mesh [EGL21, SZ21, SZ22a]. Mesner [GF20, Mar21]. method [BWW21, CK23a, Dax23, EN22, GP21, GD20, Har22, KHU23, LG21b, OPA23, Pra20, SLTW22, SWW22, WCS20, XX20, XBSF23, ZL23a]. methods [BDFF23, DAG21, Gal21, KKM22a, RZ24]. Metric [BCR20, Bai24, BGLR22, CXZ22, DF20b, NPHT20, Oli21]. Metrics [Sra21, Min22, TP23]. Milner [WXY24]. Min [TV21, MP20b]. Min-max [TV21]. Minimal [EGC21, GS21b, ADMZ21, BP20, DPR20, DdF23, FM22a, GL21, GZZ21, HP22, IdSSV21, MRV20, MO20, NV23a]. Minimal-norm [GS21b]. minimality [Koz21]. minimally [ZG22]. minimax [FP20, SS21c]. minimization [fLLV20, LWZL23, LL24b]. minimizer [HLN24]. minimizing [UZ21]. Minimum [BBC^{+}22b, MT23, AFK^{+}23, BZZ22, BMM24, DJ23, FM22b, FdCP20, HLH23, KL23, LLX23, LLA23, MAG21, MMN22, ZWW23, ZWL20a]. miniversal [FKK^{+}21b]. Minkowski [LF20, ZJCA20]. minor [ADH^{+}24, LLW23, WCF21, ZL22]. minors [BDNO20, FMR20]. Mirsky [HR20]. Mixed [AGNN20, LLL22, ABR^{+}22, Ans21, BL23, BBCD21, GL21, HBHZ22, KB22, Kad23, KSY21, WFL20, XBSF23, YGW20, YWF21, ZW20]. mixed-unitary [GL21]. Möbius [Kli20a]. Model [WXL24, XJLL23, Ais24, BDSS21, CZ20b, ĆD23b, GQZ20, LM22b]. models [BCR^{+}23, HH23a]. modifications [MPRS21]. modular [CN23a]. module [AM23a, BBNS22, Mur22, XY20b]. modules [Ago24, BHKW22, CW23a, EGL21]. Moduli [AG24, ABHR22]. modulo [CLS23, JMP21a]. modulus [BMM24, MR21a]. moment [Cia22, EAHS21, FKM20, FKKM22, Kov20, Zal22]. Momenta [dlPG21]. momentary [BEFSC22]. Moments [BCR20, YBW24]. monodromy [Ayd23]. monoid [Got20]. monoids [BS23a, Wie22]. monomial [CDdM22, SDS21]. monomials [KLR22]. monotone [ADH^{+}24, CDR23, JLM20a, JLM20b]. Monotonicity [BJ20a, CZ22a, PTD22, BW24, CESS21]. Moody [FKM22]. Moore [FG23a, PŠ23a]. Morita [FPS22]. most [ADvdH21, AHLvdH21, BFF^{+}23b, DGS^{+}21a, GM23, HT23, Kub22, LBW23, NdLO21, Ran23, WLL23, dLMO20]. Motzkin [ZL23a]. MRD [GL20a, MMZ20]. MRD-codes [MMZ20]. MRI [Sar21]. MSE [Skr21]. multi [AKPS19, DO22, HWD23, KAPS20]. multi- [HWD23]. multi-angle [DO22]. multi-dimensional [AKPS19, KAPS20]. Multidimensional [CW20, Bay24, LT24]. multigrid [BDFF23]. multilevel [HT22]. multilinear [BFMS21, FPR20, HLYZ23, PSRS20, Skr20, Vit21]. multiparameter [VD22]. multipartite [SD21a, SD21d, TLC20]. Multiple [VZ20, CV21, CI21, Hal20, LPS23, LF20, Mao20]. multiplication [HS20]. Multiplicative [CI21, OQ22]. multiplicities [AAFM23, AGV20, GS23a, Iop20, KL21b, LL21e, SD21b, ZZW22]. multiplicity [AvDF20, CH21, DJ23, DH22, DH23, TWW20, TW21, TW22, TJ21, WWJ20, WFGT20, WZW^{+}21, WZT20, WWY23, XWT20, YW23, ZW22]. multiplied [MT23]. multiplier [Rai23]. multipliers [Chr24b, Skr20]. multisets [HHS20a]. multistructures [Pat20]. multivalued [ACMM24]. multivariable [CLY22]. multivariate [AS23, CI21, Cob22b, CFR23, KN20, MP22, PV21a, Zha23]. multiwavelets [PM23]. Murdock [Fik20, FP21]. Murray [HKS23]. mutual [AGK^{+}21]. mutually [FG22].

N [KIS21]. N-tessellable [KIS21]. Narayana [YY21]. Natural [Sar21, FGGG24]. near [MSX22]. near-Hadamard [MSX22]. Nearest [AGV20, DB23, LHK20, SK21a, Yu20]. Nearly [CGP21, FRS21, FK21]. necessary [BCR^{+}23, Kub23]. negative [Bas21, BCG24, Moh22, ZC23]. negativity [GZ23, JS22]. network [AHNW22, DN21, DTA23]. networks [AD20a, AD20b, BBW20, CEM21, Est20, GN20, HNS21, KL21c]. Neumann [CHHS20, HKS23, PPW20]. neural [HNS21]. Newton [ES20, KKM22a, WCS20]. Newton-like [ES20]. NIEP [JMP21b]. Nil [Šte22b]. Nil-clean [Šte22b]. nilpotence [DGL23a]. nilpotent [ABCO22, BRL23, B– DK^{+}24, BFPS21, BM20, BE20, BDV21, CL20, CdGS12, CdGS20, CMR20, DGL23a, FMS21, GLAdS23, GJV20, KK20a, LM22a, MORŠ21, dSP20a, SSvW20, TBH20, WJLW24]. nilpotents [Mor23]. NMF [Pol20]. No [LW21d, ADGK21, MZZ24, Obo24, TCC21, Yos21, ZWF20, ZL22, ZHG21]. nodes [BDGY21, KN20]. noise [KKM22b, SLTW22]. noise-blind [KKM22b]. noisy [CW21, HH23a]. Non [BZZ22, Eld22, JN22, LS20b, MDH23, AFS^{+}23, BGS22, BS21a, BL23, Bri22, CDdM22, CL21, CdSV22, FM22a, FGGG24, GZ23, GNP22, GK21, GM20, GZ22, GLZ21, JS22, KKSV21, KKS21, Lee20, LG21a, LMX23, MZZ24, Pan23, PR22a, Phu24, PV21b, SK21a, SS22, TL21]. Non-abelian [MDH23, BL23]. non-associative [Bri22]. non-backtracking [GK21, KKS21, MZZ24]. non-bipartite [GLZ21, LG21a, LMX23]. non-commutative [PR22a, SS22]. non-cospectral [GZ22]. non-cyclic [CL21]. non-degenerate [CDdM22, GNP22]. Non-group [Eld22]. non-Hermitian [GM20]. non-homogeneous [FGGG24]. non-invertible [FM22a]. non-isospectral [PV21b]. Non-minimum [BZZ22]. non-negativity [GZ23, JS22]. non-orthogonal [TL21]. non-prime [SK21a]. Non-singular [LS20b]. non-smooth [BS21a]. non-symmetric [BGS22, Lee20]. non-trivial [CdSV22]. non-uniform [Phu24]. Non-unital [JN22]. non-unitary [AFS^{+}23]. non-zero [KKSV21, Pan23]. nonassociative [KUZ22]. nonbipartite [BP21a]. noncommutative [Pre20]. Nonconcavity [AC20a]. noncontrollable [QQM24]. Nondegenerate [KLLS22, CFSiT20]. Nonderogatory [BM20]. nonhypoenergetic [ADGKM21]. noninteger [BBCK24]. Nonlinear [Cos20b, DMQD20, Lin23, Mar24, Tom23]. Nonnegative [JJS20b, AOA23, AAMR20, Arl20, Bui22, BŠ23c, CCU21, CP22, CMM22, DYB20, GR21, HH23b, JCD24, KZPH22, Kol21, LH20, Loe23, Miy24, NGK23, RX21, SM23, vRvSTT21]. Nonsingular [SZSMB20, BJP21, Est23]. nonsingularity [Che21]. nonstandard [RAK21]. nonsymmetric [Gal21]. Nontrivial [HPSZ21]. nonzero [CV21, WWY23]. Nordhaus [CLM20]. Norm [AHK23, CLY22, FLT20, ANBK22, BM22, Bud24, CDPR23, CKOS22, CER^{+}23, DM24, FH23, GMR21, GS21b, JLM20a, JLM20b, KPP21, KLSP24, LWW22, Mis22, MSZ23, NNSS20, Roy24, SZ23b, Sin21, ZXZW23]. norm-one [CDPR23]. Normal [BS20e, CLLK21, DHR20, RW22, DY23a, IGW17, Kad23, Kov21, KFS21, LPS23, LL21c, MP23b, SZ23a, Wan21, Wan23, dlCS21]. normality [Du23]. normalized [AMZ21, JT24, SD21b, SC23, TWW20, TW21]. normalizer [BK22]. normed [IOT21, KNS20, Wój22a]. norms [BDSS21, BLR21, CI21, DLL20, FISW21, GG21, LR23, PSRS20, WW22, ZS22a]. Note [Ano21-49, Ano24-41, Bib22, Csi22, YLZH21, CMP21, DLL20, FS23, FM20b, HR20, Han22, MK23, MP23a, Ran22, SZ23a, SLL22, ZZW22, dHH23, Ano24-40]. Notes [BS20a]. November [Ano20-71, Ano20-57, Ano21-54, Ano21-56, Ano22-58, Ano22-72, Ano23-68, Ano23-64]. NP [Shi21b]. NP-complete [Shi21b]. Null [JMP21a, Wer22, HMPT20, dMdSS23]. null-filiform [dMdSS23]. Nullities [EGV20, OdFVM21]. nullity [ADvdH21, AHLvdH21, ADvdH23, CL22, CLL20, CLT22, LW21d, WG20a, ZWT21]. nullspace [ACG^{+}21]. Nullstellensatz [GSV24]. number [AAFM23, AFK^{+}23, AADT23, BBG21, Bar20, Bas21, BP20, BR21, BBC^{+}22b, CRSY22, CWYZ23, CDG21, DJ23, DP21, DW21, FCLP22, FM22b, Fio20, Gan23, Goe21, HHD20, HLX20, HXLL21, HLH23, HH23b, HWD23, HLN24, HW21b, Kli20b, fLLV20, LQ21, LZ20, LLX23, MS20b, MP21, Nic24, PK21, PZ20, Pra20, Sch23, ŠH21b, TFT20, Van24b, WWJ20, WYF^{+}20, WYQ21, WHM23, XWT20, XLYS20, YBW24, YYS^{+}21, ZXL20]. number-constrained [fLLV20]. numbers [ACLS20, ACOM21, CK23a, CV21, DBV23, ES20, Gal21, Guo22, KNiO23, WWY23]. Numerator [Bur20, Bur21]. Numerical [CKLN23, DK21, GWW21, GW22, PS23b, AHK23, BDK21, BM23, BdPS21, BJ23, BP21b, BP23b, Bhu24, CCH^{+}21, CHSW22, CDM20, CDM21, COB22a, CN21a, CN22, Cos20a, Cos22, CLY22, DA23, Dru24, GW20, Gau24, KS20a, KP23b, MN24, MXZ20, Naj20, OM20, PK22, PW21, QSH22, RAK21, RW21a, SKS22, ZW21, Zam23, ZS22b, ZC23, vdMvSW24]. Numerically [ANAS20, MN24]. Numerorum [dC21]. nut [DS22].

obstruction [HLS21a]. obtained [Goe21, GKLZ20]. October [Ano20-60, Ano20-72, Ano21-55, Ano21-57, Ano22-63, Ano22-70, Ano23-69, Ano23-72]. Odd [BBNS22, LG21a, MLZ24, ZL23c]. Odd-quadratic [BBNS22]. ODEs [BCOvdD22, BCOvdD23]. Off [MRZ20, BL24, DL22, TJ21]. Off-diagonal [MRZ20, BL24, TJ21]. off-grid [DL22]. on- [DL22]. One [CKOS23, Alf20a, BR20b, BFPS21, CJ20, CAJ21, CDPR23, CMT21, DH21, FKMM21, GJV20, GMSG22, JMP21b, KAGH24, KKM22b, KMNP20, LWW22, LTYZ23, LWH21, Ran23, SW23, ZS22b]. one-dimensional [BFPS21]. one-dominating-vertex [LTYZ23]. One-parameter [CKOS23]. one-way [KMNP20]. onto [CVV21]. open [CKOS23, GAM20]. operation [DLL23]. operations [AGKS23, CMM20, Ram23]. Operator [Uch21, AHV20, BH24, BDK21, BCG24, BDN24, Che22, CLY22, DM24, EMŠV22, Fra22, JN22, JLM20a, JLM20b, KZ24, LS20a, MA24, MS22, MS24, Nie24, PV21a, SS23a, Shi20a, SZ20, ZTX24]. Operators [AGK^{+}21, BKR^{+}20, ĆD20, ĆD23b, TT21, AM23b, AHK23, AdF20, ACM20, AG15, AG20, BDSS21, BP23a, BKPS23, Bhu24, BFPS21, BKS^{+}21, Bot21, BKLT23, CFSiT20, COB22a, CER^{+}23, CLY22, DP23, DH21, Dru24, EJJL20, FPR20, Fek20, FG23a, GGGD23, GWW21, GS22, GS23b, Gye23, HP23b, JJS20a, JKX20, JLM20a, JLM20b, KK20a, KS23c, Mag22, MS21, MP20a, Min22, MXZ20, Naj20, O’L23, OM20, Pan21, PPŻ21, Pan22, PPŻ23, QSH22, Ram23, Rao21, RRF23, SH20, SZ23b, SS24b, TWD20, TD22, XY20b, ZW21, hB21]. Optimal [JLX22, KPP21, NMS20, ANAS20, BSC23, FRS21, LM22b, MSZ23, SC20a, Ste21b, WW20]. Optimization [GSS24, GLS22, KKM22a, Onn21]. Optimized [HLW23, RW21b]. Optimizing [PS22]. optimum [TFT20]. Orbit [LCH23, CL21, HMS21]. Orbit-injective [LCH23]. orbits [BL20, GLT21, KL21a, Poo20, SZ21, SZ22a]. Order [EE23, BDNO20, BW21, BCOvdD22, BCOvdD23, CT23, CLL20, CLT22, Coh20, EG20, FHMP21, FHMP22, FdCS20, GF20, HX24, LS22, LTYZ23, LJSZ23, LH21, hLqJ21, MA22, MC21b, Mis20, OQ22, Ram23, RMG24, Ruž21, SH20, Sch23, SZH22, SMM24, Uru22, WG20a, WXL24, XJLL23, YK21, ZWT21, ZBdAF20]. order- [BDNO20]. ordered [GZ20c]. Ordering [JLW22, YOV21]. orders [BW22, DHH^{+}20, FG23a]. orientation [YWF21]. Oriented [QWW21, Sta20b, AMZ21, Du23, GL20b, HBHZ24, Sta23b, ZXZW23]. Orthant [PSZ21]. Orthogonal [LOŠ22, AP20, Ais24, BCP24, CG23, CFR23, CS20, FR21, FG23b, FK21, GIJM22, HWK21, KLS23, LM24, MA22, MC21a, MT23, Mor22, Pet21, Sta21, TL21, VS21]. Orthogonality [Pan20, RSS21, AGK^{+}21, CPTP21, Ham22, SMP24, Wój22a]. orthogonally [HJBKR24]. Orthonormal [Mag22]. oscillatory [ZM21]. other [Kin21, RDFM23]. output [DK21, FM22a, WES21, ZK21]. ovals [BMS21]. over-relaxation [BHD^{+}21]. overdetermined [SV21]. overlapping [EGC21]. ovoidal [Adr22].

packing [HWD23]. packings [Kin21]. Pages [Ano20-63, Ano20-65, Ano20-50, Ano20-52, Ano20-55, Ano20-66, Ano20-68, Ano20-67, Ano20-69, Ano20-56, Ano20-59, Ano20-51, Ano20-58, Ano20-53, Ano20-70, Ano20-54, Ano20-62, Ano20-61, Ano20-71, Ano20-57, Ano20-60, Ano20-72, Ano20-49, Ano20-64, Ano21-53, Ano21-61, Ano21-62, Ano21-67, Ano21-71, Ano21-69, Ano21-68, Ano21-70, Ano21-51, Ano21-58, Ano21-63, Ano21-59, Ano21-64, Ano21-66, Ano21-50, Ano21-52, Ano21-54, Ano21-56, Ano21-55, Ano21-57, Ano21-65, Ano21-60, Ano22-55, Ano22-62, Ano22-57, Ano22-49, Ano22-69, Ano22-66, Ano22-54, Ano22-51, Ano22-67, Ano22-68, Ano22-65, Ano22-61, Ano22-59, Ano22-64, Ano22-50, Ano22-56, Ano22-52, Ano22-53, Ano22-58, Ano22-72, Ano22-63, Ano22-70, Ano22-60, Ano22-71, Ano23-52, Ano23-57, Ano23-53, Ano23-50, Ano23-61, Ano23-62, Ano23-63, Ano23-71, Ano23-70, Ano23-60, Ano23-51, Ano23-56, Ano23-54, Ano23-55, Ano23-49, Ano23-59, Ano23-65, Ano23-66]. Pages [Ano23-68, Ano23-64, Ano23-69, Ano23-72, Ano23-58, Ano23-67, Ano24-27, Ano24-35, Ano24-38, Ano24-34, Ano24-39, Ano24-29, Ano24-30, Ano24-31, Ano24-36, Ano24-28, Ano24-32, Ano24-33, Ano24-37]. pair [ADGK21, TJ21]. pairing [KW20]. pairings [BCP22]. Pairs [BFPS21, GMP22, BKS15, BKSP23, KLSP24, Lee22, Mag22, dSP24, PM23, SZ20, TT21, TŽ23]. Paley [SS23a, Sat21]. Palindromes [HS20]. Palindromic [Pet21, BN23]. para [BN23]. para-Hermitian [BN23]. parabolic [BCNS22]. paradox [Cia20a]. parallel [Arl20, Han22, KZ23, KLSP24, TWD20, TD22, Zal22]. Parallelization [IB22]. parameter [ADH^{+}24, CKOS23, EN22, Fik20, FP21, KP22]. parameters [AAFT22, CZG21]. parametric [LP22]. parametrization [GTW23]. Pareto [BS20c]. Parlett [WW20]. Parseval [KAPS20, AKPS19]. part [BBNS22, Buc20, SMM24]. Partial [vdMvSW24, AR21, DHH^{+}20, FG23a, Li22, YWF21]. particle [HH23a]. particular [Sta23a]. partition [AD24]. partitioned [BL24, Kad20, Wei22]. partitions [AHZ22, AS20a, DF20a, FG22, LW21a, LDW21, Mif21, RFP22, ZvD21]. parts [Alt22]. Pascal [AKVS23, Bay24, Cob22b]. Pascal-like [AKVS23]. pass [LY20]. Passive [Lew21]. path [BJ20a, BS23a, Mło22, dSP20a, Ste21a]. path-connected [dSP20a]. paths [AGNN20, KSY21, SD21c, YY21, YZY24]. pattern [BES21, BBC^{+}22b, GHL^{+}24, JWZ22]. patterns [BCOvdD22, BCOvdD23, Cav21, CS20, Das22, RX21, SM23, Shi20b]. PCA [NNSS20]. Peled [CLwW22]. Pell [ACOM21]. pencil [Bar20, DYT22, DS24, Fra22, MP21, NN24, vdMvSW24]. pencils [ADGK21, AK20, AHV20, BDRS20, BR20a, BR20b, DN23, FKK^{+}21b, KP23b, NV23a, Sch24, DD23b, ZMP23]. pendant [CB21, WWJ20, XWT20]. pendent [BP20, Cia20a]. Penrose [FG23a, PŠ23a]. pentadiagonal [Los21b]. pentagon [GLZ21]. perceptrons [DD23a]. Perfect [AHM20, ASG22, CW22, GL20b, KS22c, BC24, Tar20, WZW^{+}21, ZHW21]. Periodic [Zha23, APLN^{+}23, GP20, KS23c, Kub23, Szc21]. Periodicity [IMT20, KSY21, Kub22, SMA20, KCPMNS21, Pri22]. periods [CKKR23]. Peripherally [BKT23]. Permanent [GS20b, ABW22]. permanental [GS24]. permanents [Bar21b, CW20, dC21]. permit [JWZ22]. permutation [DF20a, Djo23, MT23, VB20]. permutations [OSC20]. permutative [MA22]. permuted [ES23]. perplectic [AdlCPS20, dlCS21]. Perron [ACD22, CA22, Cia22, ES23, KZPH22, Lin23, Miy24, Rum22, ZC23]. perspective [BR23a, BS23a]. perspectives [San23]. Persymmetric [VS20, JCD24]. Perturbation [FKK^{+}21b, IGW17, SZ23a, VCR21, AG21, AK20, Bar20, BLMT21, CI21, KZ23, LHZ21, MMV21, SL21, SMM24, TJ21, WW20, hB21]. Perturbations [Gil20, AMO24, BDRS20, BR20a, BR20b, DH21, DS24, GA21, OW23, PS22]. perturbed [MA24, MP21]. Pfaffian [Bär21a]. Phase [BFO23, IOT21, StW22, ZRQK22, EL20, WMCQ23]. Phase-isometries [IOT21]. Phase-Rank [ZRQK22]. phases [CDD21, WCKQ20, WMCQ23]. physical [FM22a]. PI [FK23, GKK^{+}20]. PI-eigenfunctions [GKK^{+}20]. piecewise [ABCH23]. pivot [BW24, PTD22]. placement [SV21]. planar [Ere21, GN20, UZ21]. plane [Hom20, KP23b, NYW22]. planes [BS20f, CLZ22]. platform [KCML20]. Play [HNS21]. Plug [HNS21]. Plug-and-Play [HNS21]. plus [Jon21, MP24, SlW22, WFSS20]. Podle [Soł20]. point [AJNS24, BS20e, BDFF23, BM22, CVV21, MPS22, RW21a, Win21, Zem22]. points [AG24, CCK21a, Chr24a, LPS23, Lim22b, Mań22, ŠH21b, Shi22]. Poisson [BE20, KK23a, KK23b]. polar [Bhu24, CGP21, CG23, Pat20]. Polarization [BC23]. Pole [SV21]. Polya [GS20b]. polygonal [CHSW22]. Polyhedral [May21]. Polynomial [Cav21, CLS23, CO24, KS20a, dMdSS23, Bol22, CB21, Cos20a, Cos22, DdR20, DGdSS23, DS20, GLS22, GvRR^{+}23, dAGK21, GZ20c, HCZ20, HY24, LTWW23, Mel23, MMA22, Nar23, NV23b, PS22, SK21a, Sud22, Uru22, VS21, VS23, Zhe21]. Polynomials [CP22, Adr22, ADHM23, BN23, BHBO21, BBCK24, BGOO21, BCP24, Bur20, Bur21, CP20, CLW22, CFR23, DA23, DB23, DN20, DN23, DvDW23, Fio20, FM20b, FG23b, GA22, GTW23, GW24, GSV24, GFD21, HHS20a, HHS21, JL22, Kad20, Lee20, LHK20, Loe23, LMR21, MPR22, MP22, MMW21, Mur22, NV23a, Oli21, Pet21, PZ20, RB22, SW21, SW23, TV21, Vit21, WZL21, XJLL23, YY21, YLZH21, dMdSS23]. polystochastic [CW20]. polytope [CC22a, Yu20, vED22]. polytopes [CCK21a, DJ22]. poset [CMR20, SS24a]. posets [CCKM22]. posinormal [BKLT23]. positional [PV22]. Positive [BFMM23, NN24, AKN23, ANBK22, AL21, Bar21b, BCE21, BSC23, BHMU20, BKT23, BL24, BKR^{+}20, BS20g, BLS20, CRR20, CZ22c, CDPR23, CIM22, DY24, Fek20, FOR22, FM20b, GW24, GT21, GD20, Gye23, HJBKR24, Har22, IB22, JJLS23, JS22, Kol21, KAGH24, LV22a, Le22, LWW22, Lim21, LWH21, MMV21, Min22, MS24, NXS23, Ram23, Ran23, SM23, Sim21, Sło20, Sło21a, Sra21, Tan23, TW22, Tom23, ZFZ20]. positive-definite [CIM22]. Positively [LR21]. Positivity [GPR20, PWW23, AC21, BJ23, BGOO21, Das22, FJS17, FJS21, JWZ22, MMW22, tHvdM21]. Positivstellensatz [SS22]. post [BE20]. post-Lie [BE20]. Potent [GK22a]. potential [CIM22]. potentially [Cav21]. Power [Dom20, HJ24, CvDB23, CLS23, CIM22, DA23, DF20c, HVX23, LL21a, MS21, Ost23, RRvW24, SSvW20, Zha23]. Powers [FFSW23, BBCK24, DD23a, FJS17, FJS21, FISW21, GW20, KCPMNS21, LL24a]. preconditioner [dAdCG^{+}21]. preconditioning [OPA23]. prediction [BBW20]. Preface [AHKT21, DKM^{+}21]. prescribed [APNP^{+}21, AGV20, BPZ21, CCU21, CDG21, LQ21, ZXL20]. presentations [BR21]. preservation [SMP24]. preserve [CP22]. preserver [Gye23, Tom23]. Preservers [DHH^{+}20, PR22b, Poo20, GK22a, GS24, Ham22, LMW21, QSH22]. Preserving [GA21, AGK^{+}21, BMM24, CL24, Cos20a, Cos20b, Cos21, Cos24, GTW23, GJV20, GJKM21, GMP22, GS23b, HP23b, HH23b, KFS21, KLSP24, LL24a, LTWW23, Loe23, MPRS21, Pan20, Pan22, Ruž21, Sch24, SMM24]. prevalent [MST24]. prime [BFO23, CLS23, HS23a, Kli20b, LS20b, LP21, SK21a]. primitive [HW21a]. primitivity [AKL^{+}21]. principal [BW24, BJ20a, BLR21, CTS23, FMR20, GQZ20, LHH22, PTD22, Zha24b]. principle [CZ20a, FP20, ZX20]. principles [LWZL23, LL24b]. probabilistic [CMS21, LM22b]. probabilities [PSZ21]. problem [AdLN23, AAMR20, AC22, BGS22, Beh21, CD23a, CO24, EAHS21, FM22a, FHLS22, FQ20, FKKM22, GA22, GR22, GS20b, Gye23, Hal20, Har22, JS20, JCD24, KS24, KKM22a, Kim22, Kis22, KS20b, KP22, Kov20, LDZ21, LOŠ21a, LOŠ23, MvWC23, MS20a, NYW22, dSP20b, dSP24, PWW23, Pol20, QCCS21, Row20, SS21c, SZ22a, Ste21a, Tom23, WES21, Wan24, XX20, Zal22, ZWF20, ZX20, dBS20, dC21]. problems [APLN^{+}23, BLMT21, Boj21, BDFF23, CRR20, CEM21, CKLL20, Des24, DMQD20, EN22, GK20, fLLV20, MMW21, MG21, MC21b, PSAPNV20, PLT23, PS22, SL20a, SLTW22, VD22, WCS20, YGW20, ZMP23, ZKLS20]. procedure [MS20b]. processes [EGJP21, GZ20a]. Procrustes [BGS22, CRR20, Min22]. product [AEGV21, AdlC20, BB22, BJ20a, BKS^{+}21, BKR^{+}20, CW23b, CS20, DYB20, Fek20, FKMM21, GT23, GFW24, HY20, HP23b, HK20, Kak20, KLL22, KMNP20, LH20, MQW20, Mil24, PS20, dSP23b, Ran22, Ste21b, TT24, Yos21]. production [EHS22]. Products [GT21, AL23b, AOCM20, BC21, BDHS22, BDHS23b, BKLT23, BNS22, DGET21, FJS17, FJS21, GJV20, Gow20, HW20, HLZ17, Hou20, Jai21, KT21, Kye23, QSH22]. programming [Ans21, ZMP23]. programs [LPV21]. progression [BLS20]. project [MMP21]. Projection [CVV21, GOZ21, GNR20, LHZ21]. projections [Ais24, AGKS23, ACMM24, BKRS20, BDI20, Fra20, HY24, HPSZ21, LPV21, PLT23, Poo20, RZ24, TXF24, XY20b, ZTX24]. projective [BS20f, CLZ22, DGL23b, Mań22, NYW22]. projectors [BR22]. Proof [CLL20, CJ20, Gau24, HD22, Loe23, LZ21b, Mur22, SD21a, SD21d, SD21c, XZ23, BD21a, CZ22c, GAMA20, KW20, SW21, SL20b, WZ23a]. Proofs [KK20b, AS23]. Proper [KZ23]. Properties [DO22, QSH22, AJM24, ACM22, ACOM21, BC22, BV24, BDGY21, CLRT22, CJMP22, CZ22c, CF24, CESS21, DFS23, Dra23, FHLS22, FOR22, GA21, GK21, GPR20, HS23b, JJS20a, KLL22, LZ21a, Men22, Miy24, MS24, MPRS21, Tar20, TD22, WSW20, YZY24, ZC23, MZZ24]. Property [CN20a, AG21, Bai24, BS21a, Ban22, BP21a, BW22, Cra21, CMR21, CS20, Dra22, FG23b, GOUY23, LOS20, MC21b, MRT20, Pik23, Ruž21, Sat21]. proximal [HNS21]. Pseudo [BMS21, DP23, AdlCPS20, ACLS20, Bol21, DG20, HS20, PŠ23a, XBSF23, dHH23]. pseudo-embeddings [DG20]. pseudo-inverse [Bol21]. pseudo-inversion [PŠ23a]. pseudo-involution [HS20]. pseudo-Jacobi [XBSF23]. Pseudo-ovals [BMS21]. pseudo-spectral [ACLS20]. pseudo-unitary [AdlCPS20]. pseudoskeleton [Ham23]. Pseudospectra [Ost23, RKBA21]. Publisher [Ano24-40, Ano24-41].

Q [NdLO21]. Q-integral [NdLO21]. QLP [WX20]. quadrangle [DG20]. Quadratic [CN24, Mło22, dBS20, BBNS22, BRL23, BLMT21, BCS24, CHHS20, GA22, Ger20, GCOP24, Har22, HLS21a, Hua22, IdSSV21, Mel23, MZ22, NYW22, dSP22, dSP23b, dSP24, RB22, Uch21, ZMP23]. quadrature [BF21, Sch23]. quadrics [BMS21]. Quantitative [FGPW24, OW23, hB21]. Quantitatively [AL21]. quantities [CMM20]. quantization [AP23, FISW21]. Quantum [CGJvB22, LL21a, AHM20, AGKS23, AFS^{+}23, CJMP22, CDS21, CFKO21, CKOS23, FP23, Gan23, God21, JS22, KS22a, KIS21, KMNP20, KSY21, Kub23, Li20a, LIW20, LCH23, MPR24, OSC20, SMA20, Shi21a, Soł20, Zha22].

Quasi [AT24, ADHM23, CW23a, CR21, GP20, Men22, PS23b, Rob20, Yan21, Yan22]. quasi-Cauchy [Yan21]. Quasi-ergodic [CR21]. Quasi-hereditary [AT24, Yan22]. quasi-linear [PS23b]. quasi-periodic [GP20]. quasi-Toeplitz [Men22, Rob20]. Quasi-triangularization [ADHM23]. Quasi-Whittaker [CW23a]. quasiabelian [WFY24]. quasinilpotent [HJ24]. quaternion [BBC^{+}22a, DL23, Mig21]. Quaternionic [CDM21, AG24, BG20, CDM20, DP23, DGL23b, GSV24, GvRR21a, MMA22, NS22, RAK21]. quaternions [ACLS20, LDZ21]. question [ACG^{+}21]. questions [GAM20]. quiver [ABHR22]. quivers [GLNW24]. quotient [PS22]. quotients [MS20a, Zak21].

Racah [CFR23]. radical [FMS21]. radii [AHK23, BS24b, CMS21, KZ24, LZ21b, OM20, SD21c, SC23, XW21, YD21, ZLG20]. radius [AOA23, ASS21, AIK^{+}24, AC20a, BGGA22, BDK21, BHHP21, BJ23, BP21b, BKPS23, BP23b, Bhu24, Bui22, CLZ20, CWZ22, CLM22, CDG21, CDG22a, Cos20a, CLY22, DA23, DYB20, DPR20, Dru24, DW21, Ere21, FY23, FLRS20, FW22, FS23, GM23, GLHNW20, GZ20b, GLMM23, HLH23, HBHZ24, HLS21b, HLS22, JGW24, JZZ24, JT24, KLLW21, LW20b, LC20, LQ21, LW21b, LZ20, LZZ23, LLX23, LBW23, LW24, LLA23, MK23, MT21b, MLH22, MR21b, MXZ20, Naj20, O.21, O.22, OPPS22, PK21, QSH22, SKS22, SWW21, Sta20b, SD20, WSW20, WG20b, WCF21, WS22, WZ23a, WZC23, XHW23, YYS^{+}21, ZW21, Zam23, ZXL20, ZZ21, ZG22, ZvD23, ZWW23, ZHG21, ZHW21, ZL23b, ZKLS20, ZWL20b, ZBWL20, ZvD21]. Rado [Adr22]. rainbow [GLMM23]. Randic [AA21, HBHZ24]. Random [GNP22, LPV21, PLT23, BAPF23, BW22, CLS23, Coo20, FR21, HW20, HBHZ22, HBHZ24, Kop20, LLW20, LW20a, LLW21, NS21, OW23, RRvW24, RW22]. Randomized [WX20, ZL23a, BWW21, LY20, LG21b, MN24, OPA23, SLTW22]. range [BM23, BKLT23, CHSW22, CDM20, CDM21, COB22a, CN21a, CN22, CF22, DY23b, EM23, FGPW24, GW20, HHS20a, Ham22, KP23b, RW21a, vdMvSW24]. range-kernel [DY23b]. ranges [BdPS21, CCH^{+}21, CR23, CKLN23, GWW21, GW22, Gau24, PW21, RAK21, ZS22b, ZC23]. Rank [BR20b, BES21, BCR20, HY20, SG20, ZRQK22, AMR20, AY23, BBS23, BDRS20, BR20a, Bar20, BZZ22, BdPS21, BF24b, BT20b, Bud24, BGLR22, CXZ22, CG22, CJ20, CAJ21, CMT21, Com21, CST21, CFR23, DB23, DH21, DS24, DG22, EGC21, FGM21, GJV20, GL21, GF20, Got20, GLS22, GLT23, GS24, HP23a, HHD20, Kak20, KCPMNS21, Kha24, KPP21, KKM22b, KST22, LS22, LJSZ23, LDZ21, LW21c, LHZ21, MMN22, MC21a, MN24, MR21b, MSZ23, MO20, NPHT20, Oli21, PPŻ21, PPŻ23, PW21, Rod21, Rub22, SK21a, SZH22, Ste21b, TL21, WLT22, XZWT20, ZXZW23, dBS20]. rank-1 [KPP21, MSZ23]. Rank-Metric [BCR20, NPHT20]. Rank-one [BR20b, DH21, GJV20, KKM22b]. rankability [CLS20]. ranks [AKN23, BBS20, FdCP20, Rub21, Shi20b]. rate [BWW21, Ros22]. ratio [Hae21, Li20b, LHH22, Zha24b]. Rational [Dew21, Rob20, AL23b, ADMZ21, DMQD20, KW20, LPS23, NV23b, PQ22, Rub21, Var21, AV23]. Rationality [BNS22]. ratios [Nic20]. Rayleigh [PS22]. rays [Yu20]. reachability [Sch22]. Real [GW24, Moh22, RT21, Wei22, AMO24, AL21, AdlC20, Bib22, BS24b, DJ23, EL20, HVX23, IKLP20, Kol21, LNN^{+}20]. realising [KŠ22b]. Reality [Bla20]. Reality-based [Bla20]. realizability [JJS20b, JS20, JMPS21]. realizable [AMS21, MM21]. realization [AL23b, FM22a, Rub21]. realizations [AV23]. reciprocal [BP21a, FJ24, MT21b]. reciprocity [KNiO23]. Reconstruction [ZW20, CO24, Sar21]. Recovering [StW22]. recovery [CW21, CZ20b, HMPT20, KS20b, KKM22b, LM22b]. rectangular [Kop20, MMV21]. recurrence [PWW23]. recursive [Var21]. reduced [Ayd23, BMM24]. reducibility [MORŠ21]. reducing [CR23]. Reduction [WES21, BR23a, CB21, KSSW20, LH21, MS20b, SS22, WXL24, XJLL23]. Rees [CDG22b]. Refined [BDK21, BR22]. reflection [RS21a]. reflexive [BK22]. regarding [ACG^{+}21]. region [CAJ21, KK20b, KŠ22b, NN24]. regions [KK22a]. regression [Ais24, CI21, GQZ20]. Regular [BFF^{+}23b, AG21, AFSS20, AdF20, BDRS20, BR20a, Bar20, CT24, CCO23, DF20a, DYT22, DS24, FS20, GA22, GK22b, GSV24, HO20, IKPR20, KGYY23, KAGH24, Kub22, Kub23, MP21, Nic24, QJMW23, YK21, Zha22, Zha24a]. Regularity [MP24]. Regularization [Kov20, HH23a]. regularized [DAG21]. Reinhardt [MT22]. related [AHNW22, BCOvdD23, BDI20, CZ22a, CMM20, Est20, FP20, FH23, JJL20b, Kli20b, LMW21, Li22, Mar21, MMW21, MS24, PV21a, Shi21a, Shi22, Sra21, Van24a]. Relation [ZXZW23, CIM22, GMR22, WJLW24, WWY23, XTLK23]. Relations [Sta23b, WFL20, ACSS22, Arl20, CMM22, GJKM21, Lim21, Nag23, SL21]. relationship [TFT20]. Relative [BLMT21, Shi21a]. relatively [Kus23]. relaxation [BHD^{+}21]. Rellich [BN23]. remark [Bar21b, Shi20b]. Remarks [Tei22]. removal [JST21]. Rényi [DF20c, Shi21a]. repeated [JMP21b]. repelling [Est20]. representation [Ago24, Bla20, CN22, PV22, SC20a, SW21]. Representations [AG15, AG20, CFR23, ABHR22, CM23a, CDD21, JL22, KMNP20, MRZ22, NPR22, O’L23, PQ22, RT21, Sch23, Shi21a, Wie22]. require [GHL^{+}24]. residue [KW20, NPkT21]. residues [Fra22]. resilient [KCML20]. Resistance [BEF20, Dev22, Est23]. Resolvent [Fra22, Zem22, AHV20]. resolvents [Fra20]. resonant [OSC20]. resource [JS22]. respect [AdlCPS20, CEM21, Ime23, KKS21, KLSP24, RAK21, Ram23, Sta21]. Restricted [NXS23, BD21b, CHSW22, MMA22, PGST23, Poo20, Sat21]. result [LLSV21]. Results [FJS21, BS20c, BEFSC22, DYLL23, Fra20, Iop20, LJSZ23, Nie24, OW23, Sch23, WHL24, FJS17]. retraction [GSS24]. retrieval [BFO23, EL20, StW22]. revealing [SG20, Ste21b]. Reverse [Pre20, BCP24]. Reversibility [GL22b]. Reversible [BG20]. revisited [CCO23, GH23, LIW20, Van24b]. Revisiting [LS20c]. revival [CCD^{+}22, GZ22, Mon23]. reweighted [CW21]. Riccati [BF24b, CF24, Ruž21]. Ricci [Sic21]. Richardson [KK20a]. ridge [GQZ20]. Riemann [Nag23]. Riemannian [ÁVKP20a, DAG21, KKM22a, NXS23, TP23]. Riesz [Fra20, GS20a]. right [AL23b, BMPS23, DHH^{+}20, FKK21a, Hal20, LS23]. right-angled [FKK21a]. right-hand [Hal20]. right-star [DHH^{+}20]. Rigidity [KNS20]. rigorous [Miy24]. Ring [HKS23, AT21, AM23a, CPS20, CGMP23, NPR22, SSvW20]. rings [AT24, BS20f, Bol22, BHKW22, CM23b, Ham22, HLZ17, Hou20, LP21, NPkT21, Nic24, NZ24, Pre20, Wer22]. Rio [AHKT21]. Riordan [Sło21a, AK23, Bur20, Bur21, CCP22, CCS22, CCKM22, Cob22b, Coh20, He20, KNiO23, LMR21, MMW22, Pet21, SS24a, Sło20, Sło21b, SS23c, YY21, YY23, ZZ20, ZZJ23]. Ritz [RW21a]. robust [KCML20]. robustness [KKM22b]. Roe [MT21a]. role [AFSS20, DN20]. Roos [De 20]. Root [DN20, NV23b, CA22, LHK20, NV23a, SZ21, SZ22a, TV21]. rooted [Cia22, DR20]. Roots [BHMU20, DGL23a, ES23, GvRR^{+}21b, Jon21, MZ22, dSP20a, PZ20, TV21, vRvSTT21]. Rota [MDH23]. rotational [NNSS20]. Rotfel’d [MJ24]. Row [RC23, ES23]. row-permuted [ES23]. rows [KS23a]. Ruelle [Mor20]. rule [NWW20]. Ruling [JMP21b].

Sachs [CC22b]. saddle [BDFF23]. Salem [RT23]. same [B– DK^{+}24, CDG22a, MR21a]. Sampling [FR21, CEF^{+}22, MB20]. satisfying [BKS^{+}21, DA23, Ime23]. Saturated [cRLY20]. Scaffolds [Mar21, LTTW22]. scalar [BKS^{+}21]. scale [BW21, BF24b, CHHS20]. scaling [FRS21, LT24, MS22]. Scattered [CLZ22, BCM21]. Schatten [Bot21, CI21, LHZ21]. Schatten- [LHZ21]. scheme [AHM20, KCML20, LIW20, SS20]. schemes [ABCH23, BMS21, LTTW22, XTLK23]. Schmidt [CLRT22, Min22, Zam23]. Schröder [YZY24]. Schrödinger [CW23a, CW23b, KS23c]. Schur [AV23, BJ20b, BS22, Chr24b, CMM22, Dev22, FKM20, Gow20, Skr20, TT21]. Schwarz [XX20]. Secant [KMZ20]. second [ABG20, BW21, BCOvdD22, BCOvdD23, CLM20, LWB20, LH21, MC21b, WQP23, XLS20, YK21]. second-order [BW21, MC21b]. secret [KCML20]. sector [Dru24, FLT20]. sectorial [AS20b, Ala21, MJ24, WMCQ23]. Sedentary [God21]. segments [LZ20]. Seidel [ACG^{+}21, BWW21, GKLZ20]. select [CRSY22, KHU23]. selection [Shi21b]. self [AC21, CCK21b, DH21, ES20, IMT20, MMP21, PPŻ21, Pan22, PPŻ23, SZ22a, SZ20]. self-adjoint [DH21, PPŻ21, Pan22, PPŻ23, SZ20]. self-adjointness [AC21]. self-conjugate [ES20]. self-dual [SZ22a]. self-inverse [CCK21b]. self-loops [IMT20]. self-similarities [MMP21]. selfadjoint [AMO24, Arl20, CFSiT20, CMM22, GvRR^{+}21b, PSW23, SH20]. selfadjoint-ideal [PSW23]. Semi [CCK21b, AdF20, BKPS23, Elm23, MRZ22, Men22, MXZ20, WMCQ23]. semi-Hilbertian [BKPS23, MXZ20]. semi-infinite [Men22]. semi-invariants [Elm23]. Semi-involutory [CCK21b]. semi-regular [AdF20]. semi-sectorial [WMCQ23]. semi-simple [MRZ22]. Semiclosed [ACMM24]. semidefinite [AKN23, ANBK22, BGS22, BKR^{+}20, CKM20, Fek20, FM20b, GW24, GNR20, HMPT20, MMN22, NXS23, NN24, PLT23, Tan23]. semifields [LS22]. semigroups [AG15, AG20, BS20d, CFKO21, PSW23]. semimonotone [CD23a]. Seminorm [MXZ20, Wój20]. seminorms [CG20, Ime23]. semipositive [HS23b]. semiring [DY23a]. semisimple [KK20a]. Semrl [Roy24]. sending [SS23b]. sense [ACM20]. sensing [Tsa23]. sensitivity [Mis20]. sensors [KHU23]. separable [GLS22, NGK23]. Separating [CL20, LM24, Elm23]. September [Ano20-49, Ano20-64, Ano21-65, Ano21-60, Ano22-60, Ano22-71, Ano23-58, Ano23-67]. Sequence [AK23, PSRS20]. Sequences [AOA23, CER^{+}23, FKM20, He20, PWW23, PW19, PW20, VS21, VS23]. Sergeichuk [SZ22a]. series [Bur21, Dom20, Zha23]. Sesqui [KGYY23]. Sesqui-regular [KGYY23]. set [BPZ21, CRR20, Cia20b, FPR20, HJ24, LHK20, LOŠ23, ZTX24]. Sets [HHS20a, CP20, CLZ22, CAEH^{+}22, ES20, GSV24, GZ20b, HHS21, HJBKR24, HW21a, LNP21, MM21, MMP21, MRT20, Pat20, RW21a, RW22, YLCB21]. setting [FP20, JST21]. several [DF22, LLH20, Naj20, QSS22, SS23a, SW23, ZM21]. SGD [SDTR23]. shadows [CTS23]. Shannon [Sra21]. shape [CWBB23]. sharing [KCML20]. sharp [MLH22, MKE^{+}20, WG20a]. sharpened [LS20c]. Sheffer [LMR21]. shell [CP23]. shell-extremal [CP23]. shift [COB22a, CN21b, CKLN23, CKLL20, EJJL20]. shifts [HJ24, JJLS23, Pik23]. short [CWZ22, SW21, SL20b]. shortest [Ste21a]. Shorting [Arl20]. sided [BB22, KLL22, WW20]. sides [Hal20]. Sidon [ZT23]. sigma [FISW21]. Sign [BD21b, CS20, Das22, RX21, SM23, AG21, BN23, BCOvdD22, BCOvdD23, BBC^{+}22b, BS24b, GHL^{+}24, JWZ22, KK22b, OQ22, Shi20b]. sign-real [BS24b]. Sign-restricted [BD21b]. signatures [BCP22]. Signed [ADvdH21, ADvdH23, HSS^{+}21, KP23a, RS21b, AAH^{+}22, AFG^{+}21, AHLvdH21, CH21, CG22, CvDB23, CCK21b, HT23, LW20a, LW21d, MMN22, MT23, OJQ20, SZ21, Sta23b, SLL22, VB20, WH20, WYQ21, WLL23, WHL24, Wan24, WLT22]. Signless [CWZ22, GH24, ZWL20a, AdL21, CLZ20, HLS21b, KLR22, LWB20, LQL21, LLWW23, LHH22, MK23, MT21b, Obo22, SSS21, VTBC20, XW21, ZXL20, ZHG21, ZWL20b]. similar [DBV23]. similarities [MMP21]. similarity [AdlCPS20, AR21, CN21b, JKX20, TBH20]. Simple [GG21, dlCS21, Eld22, IIM^{+}21, KLLS22, MRZ22]. simplex [PSZ21]. simplices [FS20, Dev22]. simplicial [KMZ20, Kni20, KL21c]. Simplifying [Zhe20]. simultaneous [Ost23, RZ24, Her20]. Single [LY20, CLY22, HH23a]. Single-pass [LY20]. Singular [Djo21, CV21, CJ20, CAJ21, CK23b, Cos21, DN23, FISW21, Fra22, KL23, KK22b, KP22, KN20, LS20b, LHK20, LH21, Loe21, MQW20, Nay23, Van24b, VCR21, ZS22b]. singularity [AHV20, GMP22]. Sinkhorn [CN20b]. SISO [WES21]. Six [CdGS12, CdGS20]. Six-dimensional [CdGS12, CdGS20]. size [BMR23, B– DK^{+}24, CT23, CDG22a, FY23, FPR20, FW22, HLS22, JLW22, LTYZ23, LW24, MP23a, MLH22, Nar23, Rob20, SWW21, ZG22]. skeleton [KLR22, dHH23]. skew [BHBO21, BDHS23a, Com21, Du23, DGS21b, EGV20, GS21a, HBHZ24, KT21, LR23, MP22, PS20, Pri22, QSH22, QWW21, SS21a, Sta20b, Sta23b, TFT20, Uru22, YY23, dlCP20]. skew-adjacency [Du23]. skew-eigenvalues [TFT20]. skew-Hamiltonian [SS21a, dlCP20]. skew-Hermitian [LR23]. skew-involution [PS20]. skew-symmetric [Com21, DGS21b, EGV20, GS21a, KT21, Pri22, Uru22]. SKT [BF24a]. slice [AdF20, AY23, GSV24]. small [AMO24, Alt22, BMR23, HW21b, PSRS20]. smallest [Csi22, KGYY23, KN20, LW21a, Loe21, LLL22, Obo22, Ran23, SWW21, WWT20, ZLG20]. Smirnov [AOCM20]. Smith [BS20e, MP23b, Wan21, Wan23]. Smooth [Lim22b, BS21a, Lub22, MP20a]. Smoothed [NGK23]. smoothers [HLW23]. SNIEP [MM21, MP23a]. solution [MvWC23, PS23b]. Solutions [GvRR^{+}23, BS21b, Djo23, DL23, EJ23, KST22, WvdVvdH24]. solvability [MP20b]. solvable [CNO23, LNN^{+}20, LS20b]. solve [VD22]. Solving [EN22, MPS24, ABCH23, BF21, HLW23, fLLV20, SDTR23]. Some [AGH20, ABR^{+}22, Chr24a, Fra20, GK21, Gow20, HVX23, Iop20, KUZ22, Kus23, LHFL20, LH20, Miy24, Naj20, Sło20, Sło21a, WG20b, AMO24, AAMR20, APNP^{+}21, BSS23, BD21a, BP23a, BdPS21, BO23, BDI20, Bri22, CNO23, CEJM21, Ces20, COB22a, GLLZ20, GAM20, GPR20, HJ24, Ime23, JJL20b, KZPH22, Li22, Los21b, Mar24, MS24, OP21, PZ20, Shi20a, TLC20, TWW20, WSW20, ZJCA20, ZC23]. space [AHV20, BS21a, Ban22, BKPS23, BKS^{+}21, CFSiT20, CGP21, DN23, FG23a, FP22, HMPT20, LR23, Mań22, OM20, dSP20a, SS23a, SlW22, ZW21, hB21]. spaces [ANAS20, ABHR22, BFMS21, BDI20, BKR^{+}20, CG23, CG20, ĆD20, ĆD23b, DP23, Djo21, FPR20, Fek20, Fra22, GKS21, IKLP20, IOT21, JN22, KNS20, LL24a, MP20a, MORŠ21, MV24, MXZ20, O’L23, ORŠ23, PSRS20, Ran22, Rao21, VMB22, WXY24, Wój22a, ZT23]. spanned [BR24]. spanning [BEF20, HJBKR24, HWD23]. sparse [AK20, ABW22, BR22, CW21, HLW23, HMPT20, Kli20b, KS20b, SLTW22]. sparseness [GL20a]. Sparsity [BFF^{+}23a]. SPD [TP23]. Specht [Che22]. special [DP23, GPR20, Mil24, Mor22, TLC20, AHKT21]. specified [AK20]. Spectra [BBC^{+}22a, CCPKP20, DF20a, DH21, JJL20a, AAMR20, AMS21, BS20c, BR23a, CN24, CFKO21, Coo20, DDFM^{+}21, DP23, FPS22, GK20, HO20, HS22, HBHZ22, JMP21b, KIS21, LW20b, RDFM23, SMA21, SB20, SS21b, WZFW23, WFY24, Zha21, ZS22b]. spectrahedra [CEF^{+}22, May21, MT22]. Spectral [AAH^{+}22, AFS^{+}23, BV24, Bou21, CJMP22, CER^{+}23, DFS23, DKL^{+}22, Des24, DTA23, FOR22, Gan23, GK22b, GM20, GLMM23, HWD23, KY21, LS20a, LBW23, LW24, O.21, WHM23, WZC23, Wan24, YYS^{+}21, ZY24, ZvD21, AOA23, ADGK21, ASS21, AFSS20, AIK^{+}24, AdST22, ACLS20, AC20a, AEGL21, APLN^{+}23, BGGA22, BC22, BHHP21, BDGY21, BR22, BCG24, BKPS23, BEFSC22, Bui22, BS24b, Ces20, CLZ20, CWZ22, CLM22, CWYZ23, CWBB23, CT24, CMS21, CZ20b, CDG21, CDG22a, DYB20, DPR20, DW21, Ere21, FY23, FLRS20, FW22, FS23, GS21a, GA21, GMR21, GM23, GLHNW20, GK21, GZ20b, GLZ21, HLH23, HBHZ24, HLS21b, HLS22, JGW24, JZZ24, JT24, KT21, KLLW21, KP23a, KPP21, LC20, LMW21, LQ21, LW21b, LY23, LJSZ23, LZ20, LOS20, LG21a, LZZ23, LLX23, LS24b, LZ21b, LLA23, MK23, MSZ21, MT21b, MLZ24, MLH22, MR21b, O.22]. spectral [OPPS22, PK21, QJMW23, QQM24, Ram23, ST20, SWW21, Sim20, Sim21, Sta20b, Ste21a, SD20, SD21c, SC23, WSW20, WG20b, WCF21, WS22, WZ23a, XW21, XHW23, YBW24, YD21, ZXL20, ZWF20, ZW20, ZLG20, ZZ21, ZG22, ZvD23, ZWW23, ZHG21, ZHW21, ZL23b, ZKLS20, ZWL20b, ZWL20a, ZBWL20, hB21]. spectrally [ANAS20]. Spectrum [KK22c, AGNN20, Ban21, BCFT21, CHT20, CGP22, Coo21, Cos24, DR20, Fio20, FB21, GH24, HP22, LQL21, LLSV21, MPS22, Mar24, QWW21, SSP22, Sta23a, Sta23b, XY20c, YW22, ZK21, Zem22, ZLLZ22, Zhe20]. spheres [CVV21, LR23, Soł20]. Spherical [HW21b, Alf20a, BNS22, Est23]. spillover [ADGK21]. spin [AD20a, AD20b, FHMP21, FHMP22]. splines [BGOO21]. split [SSS21]. splitting [Ayd23]. spread [BEHK22, BCE^{+}23, Bib22, CL21, LMG20, LLW23, MSZ21]. spring [UZ21, VS20]. Square [MSX22, vRvSTT21, Bar20, CP23, CGMP23, Ghi20, VS20]. square-integer [VS20]. squares [Boj21, CW21, CKTP20, Hal20, Kor21, KS23b, LV22b, Lyu20, SLTW22, SC20a]. SR [GSS24]. Stability [BCOvdD22, CT24, GOUY23, SW23, AMO24, AEGL21, BCR^{+}23, BBCK24, CW21, CLRT22, Cav21, KP21, LLSV21, Zha24b]. stabilizing [GG21]. Stable [Com21, ADvdH21, BBCK24, Cav21, CHY20, Kus23, TV21]. staggered [KIS21]. Standard [KLR22]. star [ARS24, CDM20, CDG22b, DLL23, DHH^{+}20, SS21c, GKK^{+}20, KKSV21]. star-center [CDM20]. star-clique [DLL23]. starlike [Sta23a]. State [NPR22, AHM20, ASG22, CW22, GL20b, KS22c]. states [CJMP22, CN20a, HX24, KMNP20, KS22c, SS22]. static [GS21b, ZK21]. status [JST21]. Steiner [AS22]. step [BDV21, LM22a, WCS20]. steps [CN20b]. Stiefel [GSS24]. Stieltjes [Kov20, YGW20]. Stirling [KNiO23]. Stochastic [KŠ22b, AKL^{+}21, BC21, CZ20b, Djo23, GK20, HT22, KK20b, KK22a, MvWC23, Pra20]. Storage [CLLK21]. straight [HLS24]. strength [HXLL21, RC23]. strengthenings [ELW24]. strict [DD23b]. Strictly [AJM24]. string [YGW20]. Strong [ARS24, AGK^{+}21, BHD^{+}21, CS20, FHLS22, GS21c, GS23b, LOS20, Wój22b]. stronger [Nay23]. Strongly [AEGV21, CCO23, GK22b]. strongly-regular [GK22b]. Structural [AJM24]. Structure [CPTP21, Ham22, XZ21, Bar22, BS21b, Bri22, CCU21, EJ23, FKMM21, GTW23, GFW24, JJS20b, KT21, KFS21, Lew21, LP21, Ran22, SMM24]. structure-preserving [GTW23, KFS21, SMM24]. Structured [AK20, DL22, SK21a, ADGK21, BEFSC22, Cos21, FGM21, GA21, HJBKR24, Kov21, KFS21, LHK20, dSP20b, PS22, RS21a]. structures [Ayd23, B– DK^{+}24, BE20, DLL23, GA22, KK23a, KK23b, Pat20, Tar20, VV22]. study [hLqJ21, Sim20]. subalgebra [PGST23]. subalgebras [BFLT23, MRZ20, MRZ22, Mig21]. subclass [SNK22]. subcubic [WH20]. subgraph [BD21a, FY23]. subgraphs [CvDB23, LW24]. subgroup [BL23]. subgroups [DHR20, HWK21, VZ20]. subject [DPR20, OW23, SDS21]. submajorization [BMS20]. submatrices [BJ20a, BLR21, CKM20, HW21a, RW21a]. submonoids [Got20]. subnormal [Kim22]. subproblems [NN24]. subrings [AM23a]. subsampling [ZL23a]. subset [Shi21b]. subsets [BMR23, Cos24, GMP22, Wer22]. subshift [CA22]. Subspace [AS20a, AD20a, AD20b, Ime23, Kin21, ZT23]. subspaces [BCM21, CGP21, GGGD23, GNP22, KK20a, LP22, MAG21, Pol20, RZ24, Rub22, SS21a, Van24b, WW20, Wój22b, XJLL23, Yos21]. substochastic [CC22a, JJL20b]. subtrees [LBW23]. Sufficient [ZWL20b, BCR^{+}23, ZBWL20, ZL23c]. Suleimanova [GK20]. Sum [LNP21, AT21, BGLR22, Csi22, DGL23a, ES23, GLT21, GPRT20, Han22, HS23a, HD22, KS20a, KZ23, LV22b, LS23, Onn21, dSP22, dSP24, TWD20]. Sum-full [LNP21]. sum-of-squares [LV22b]. sum-rank [BGLR22]. summable [TD22]. Sums [Jai21, KS23b, AS20a, Arl20, BM20, CWW24, FK23, FJS17, FJS21, GNP22, KL21b, Kli20a, SZ23b, dlCP20]. sun [AdST22, SL20b]. Superalgebras [IdSSV21, BBNS22, CM23a, CNO23, HLS21a, Iop20, KK23a, KLLS22]. supercoalgebras [KUZ22]. supereigenvector [SlW22]. superpairs [AO22]. supervisions [NMS20]. supported [AKPS19, KAPS20]. Supporting [dAdCG^{+}21, GLNW24]. Supports [MRV20]. surely [Roc20]. Surjective [CAEH^{+}22, GS20b]. surprising [Dax23]. Suslin [JR24]. Suzuki [CWW24]. Swift [FP23]. Switching [EJS24, BBCD21, GKLZ20]. Sylvester [AdF20, BF21, Djo21, Lim22a, Pra20, Sas21]. Symbol [BDFF23, BDvW24]. symbols [BEFSC22]. Symmetric [BŠ23c, CP21, Kim20, O’L23, Uch23, Uru22, AFK^{+}23, AD20a, AD20b, BGS22, BMR23, Bib22, BGOO21, BR21, CJMP22, CN21b, CL24, Com21, CKM20, DBV23, DJ23, DGS21b, EGV20, EJJL20, FMR20, FG20, GS21a, Gal21, GK20, HH23b, HLYZ21, JZZ24, JMP21b, KT21, KKM22a, KNS20, Kor21, KLL22, KLLS22, LW21a, Lee20, LOŠ22, MC21a, Moh22, MO20, Nag23, PŠ23a, PV21a, PS22, Pri22, Rod21, SZH22, Ste22a, SW21, Van24a, Wei22, Win21]. symmetries [AP20, LDW21]. Symmetrizable [MS20a, Sim20]. Symmetry [GLZ20, SS22, FKMM21, GA22]. Symplectic [Ayd23, AdlC20, BFLT23, BJ20b, EV20, GSS24, GT21, IGW17, Jai21, JLX22, Mis20, dSP24, ŠH21a, SZ23a, SMM24, Zem22, dlCP20, dlCS21]. system [ABCH23, BCOvdD22, BCOvdD23, FGM21, Ros22]. Systematic [Ner20]. systems [AOCM21, BH24, BPZ21, BW21, BDN24, Bri22, CMS21, Com21, DYT22, De 20, DK21, GP21, GJ23, HLW23, HMPT20, IB22, JN22, Kak20, KHU23, LPS23, Lew21, LH21, LG21b, Mar24, MMW21, MPS24, Rei23, RW22, SDTR23, ŠH21b, SV21, WXL24, XJLL23, ZK21, Zem22, ZS22b]. syzygies [BT20b]. Szego [Fik20, FP21].

T [hLqJ21]. T-eigenvalues [hLqJ21]. T. [ACM20]. tableaux [KK20a]. tables [PP20]. taking [AT24]. talented [BS23a]. tame [dBS20]. tau [EGJP21]. technique [KL21b, Sim20]. techniques [CLY22]. Tensor [AY23, HK20, AOCM20, Bar22, BZZ22, Buc20, GFW24, KKM22a, Kye23, LS22, LDZ21, Mar21, MQW20, SDS21, XX20, ZBdAF20]. tensors [BC24, CL24, DYB20, DBV23, GLZ20, HLYZ21, KZPH22, hLqJ21, Miy24, MO20, Rod21, SZH22, YLCB21]. term [FdCP20, PWW23]. terms [AAFM23, CLL20, CLT22, DW21, FLRS20, HHD20, LW21c, NWW20, NYW22, WWT20, WWJ20, WG20a, XWT20, ZWL20b]. ternary [AC20b]. Terwilliger [FM20a, LIW20]. tessellable [KIS21]. Testing [Ans21]. tetradiagonal [BFMM23]. tetrahedrally [De 20]. tetrahedron [Mor22]. th [CW23a, CW23b, GZZ21, GvRR^{+}21b, dSP20a]. their [AKN23, AM23b, AGNN20, Ala21, ADMZ21, APLN^{+}23, BDSS21, BC21, BKLT23, BCR20, CLRT22, CWBB23, CIM22, Coh20, CDG22a, DJ22, DN20, FLRS20, Fid22, FN23, HS23a, HS22, HVX23, IKLP20, JJS20a, KS23c, LZ21a, LL21e, Miy24, Moh22, OdFVM21, QWW21, SS21a, SB20, StW22, TFT20, WWY23, Zha21, DGdSS23]. theorem [AC20a, AFS^{+}23, BJ20b, BKPS23, CLNY22, CN21a, CC22b, ELW24, GS20a, HR20, Kli20b, Mur22, Nay23, PR22a, QJW20, Roy24, Shi22, Tei22, WW20, WXY24, XY20b, CZ22a, KKOO22, KKT22, LLWW23]. theorems [Adr22, ACSS22, AS23, GSV24, GRR22, KP23a, MT21a, Mar24, UZ21]. theoretic [BHHP21, Mar21]. Theoretical [Men22, KS22a]. Theory [MP22, NV23b, AD24, ABC^{+}22, BLMT21, CLY22, Djo21, DN20, DTA23, FT23, FKK^{+}21b, GCOP24, GZ20b, JS22, KS22a, KZPH22, LNN^{+}20, Lin23, MMV21, PŠ23a, RT21, SNK22, SMM24, dAdCG^{+}21, hB21]. There [MZZ24]. Theta [CN23b, AC22, Nag23]. third [GF20, hLqJ21, Vol23, ZBdAF20]. third-order [hLqJ21, ZBdAF20]. Thompson [CZ22b]. Thorin [GS20a]. those [Mel23, WFL20]. Three [DBV23, Alt22, CL20, MRT20, Nar23, NdLO21, PWW23, Row20, RS21b, WLL23, WZFW23, dBS20]. three-term [PWW23]. threshold [AFSS20, FM22b, GZZ21, Tur21, VTBC20]. thresholding [DL22]. tight [DO22, FS20]. tilting [EGL21]. time [BW21, CF24, CMS21, DK21, LV22a, Lew21, WXL24, XJLL23, Zha22]. time-delay [WXL24, XJLL23]. time-limited [BW21, DK21]. Tipping [Tho22]. Toeplitz [PW20, BDSS21, BEFSC22, CCGV^{+}20, CKKR23, EGV20, Los21b, Lub22, Men22, O’L23, OW23, PW19, Pri22, Rob20, RS21a, Shi21c]. token [BV24, DDFM^{+}21, RDFM23]. tolerance [Lim21]. tool [Mar21]. tools [CLY22]. topics [BDI20, FP20]. topological [GMR22]. topology [Win21]. toric [DJ22, KMZ20]. Total [FJS17, FJS21, BW22, CGMP23, MMW22]. Totally [GZ23, CCU21, CZ22c, GR21, MMV21, Sło20, Sło21a]. tournament [BCF^{+}20]. tournaments [BBS23, BBCL22]. Trace [KS23c, SL21, BO23, EMŠV22, FYS^{+}23, HT22, Li22, LWZL23, LL24b, MM21, MS20a, RRF23, UZ21, Vit21, ZXZW23]. trace-zero [MM21]. traces [Li22]. Tracial [Skr20, CN20a]. trains [ZBdAF20]. transfer [AHM20, ASG22, CW22, GL20b, KS22c]. transform [BW24, MB20, MT23, MST24, PTD22, SS24b]. Transformation [BS20d, MC21b]. transformations [AdlCPS20, AR21, BK22, DGL23b, GTW23, GOUY23, KFS21, Pan20, Pan22, Yam21, dlPG21]. transient [KCPMNS21]. transitions [BBW20]. translates [SS23b]. translation [CLZ22, LTTW22]. transportation [CCK21a]. transpose [CdSV22]. Transposed [KK23a, KK23b]. Transposition [KK22c, KK24]. TRASE [Sar21]. Tree [JWZ22, AS22, AJNS24, BJP21, DJ23, ET20, GR22, GRR22, HWD23, ZZW22, LLWW23]. Trees [AvDF20, GM22, LW21b, AADT23, ACD22, AS24, BEF20, CWBB23, Cia22, DFS23, DR20, DH22, FB21, JKS23, JGZ21, LQL21, LZ20, MK23, OdFVM21, Pik23, Ran23, SGGZ21, Sta23a, WZT20, WWY23, XY20c, XLYS20, YW22, YW23, Yaz21]. triangle [Csi22]. triangle-free [Csi22]. triangles [GL23, Pet21, ZHG21]. Triangular [Yan22, BS20b, BS20f, CLW22, CMT21, Dew21, DdR20, DRG24, FGDY20, GOZ21, JLX22, KS24, Lim22a, Loe21, Mis22, Nar23, RKBA21, TBH20, Uru22, WZL21, ZG21]. triangularization [ADHM23]. Triangulated [BC21]. tridiagonal [Hu21]. Tridiagonalization [GP21]. trifactorization [BŠ23c]. trigonometric [Guo22]. trinomial [MMZ20]. triple [RS21a]. triple-structured [RS21a]. triples [AG24, GLP21]. tripotents [AT21]. trivial [Alf20b, ABSV21, CdSV22]. TRO [EM23]. TRO-conditional [EM23]. Tropical [GN20, DZCY22, DY23a, KCPMNS21, NWW20, NYW22, Xie21, Yu20]. Trotter [CWW24]. truncated [FKKM22, O’L23, VCR21, Zal22]. truncation [BW21, DK21, Mis22]. trust [NN24]. Tucker [ZBdAF20]. tuples [BFMS21, Che22, Fek20, GS22]. Turań [KP23a, CWYZ23, Des24, ELW24, Wan24, ZWF20]. twin [BR23a, Bas21, Mon23]. twins [Cia20a]. Twirling [GL21]. Twisted [BCP22, GLNW24]. Two [AHLvdH21, BB22, DGdSS23, ELW24, KLL22, LM22a, LJSZ23, VD22, ZL23c, AT21, ADvdH21, ADvdH23, BFF^{+}23b, BFF^{+}23a, BR24, Bla20, B– DK^{+}24, Buc20, CCCS21, CFR23, EN22, GLT21, HT23, HHS20b, KS20a, Kor21, KP22, LHFL20, LM24, MS21, NdLO21, dSP22, dSP23b, RZ24, SZ21, WW20, WZW^{+}21, WCS20, XX20, XY20b, YD21, ZLG20, Zha24a, dLMO20]. two-band [CCCS21]. Two-connected [AHLvdH21]. Two-dimensional [DGdSS23, Bla20, LM24]. two-isometries [MS21]. two-level [XX20]. two-parameter [EN22, KP22]. Two-sided [BB22, KLL22, WW20]. Two-step [LM22a, WCS20]. type [ACSS22, AOCM21, CWW24, CLM20, CA22, CCL21, DN21, EE23, Fra20, GOZ21, GR22, GSV24, GS20a, HHS20b, HVX23, JL20, KNiO23, KP23a, KK23b, Le22, LLH20, LLW21, MT21a, MJ24, MMV22, MS24, Nag23, Sim21, SZ22a, Sos23, YZ20, YLCB21, KKOO22]. types [Alt22]. Typical [BBS20].

Ulam [BS21a, Ban22]. ultrametric [Yu20]. unbalanced [Sta20a]. unbounded [FPR20]. undecidable [Ros22]. underlying [WFL20]. understanding [CLS24]. undirected [ACDG21]. unequal [HT23]. unexpected [DFH^{+}20]. unicity [Wój22b]. unicyclic [BP20, BDVR21, DH23, LW20b, LQL21, QLCM20, YYS^{+}21]. unified [AKVS23, AWvW24, CZ22c, Lin23, VS21, WZFW23]. Uniform [MV24, ZLG20, BMR23, BMS20, CHT20, DvDW23, FYS^{+}23, LL21c, LYZ23, MT21a, Phu24, SSP22, Sch22, SWW21, SWW22, WSW20, XW21, XHW23, ZZ21, Zhe20, Zhe21]. uniformity [GRR22]. unifying [Min22]. unilateral [JJLS23]. unimodular [BBCL22, PSRS20]. union [HO20, Zal22]. unipotents [Pre20]. Unique [Chr24b, MLZ24]. uniqueness [ĆD23b, HHS20a, Li20b, WvdVvdH24]. unit [Alf20a, BBC^{+}22a, FG20, FG23b, HHD20, JLX22, LW21c, LWH21, SK21b, Wój22b, XZWT20]. unital [BKT23, BK22, CDPR23, CAEH^{+}22, JN22]. unitaries [FHMP21, FHMP22, FOP22, PR22b]. Unitarily [ANBK22, DLL20]. Unitary [CN21b, AdlCPS20, AFS^{+}23, BCS24, BL20, CXZ22, CAEH^{+}22, DID22, FR21, GL21, GvRR21a, Hua22, KIS21, KFS21, Mil24, NS21, PR22b, PV21b, Poo20, VB20]. units [DHR20]. unity [DGL23a]. Universal [HO20, JMPS21, Yas23, CGP21, CGP22, JS20, Ago24]. Universality [FOP22]. unknown [SL20a]. unmeasured [ADGK21]. Updating [ADGK21]. Upper [BS20b, CCGV^{+}20, JZZ24, MT21b, Ran23, AOA23, CLW22, DdR20, DRG24, MLH22, MT23, MP21, Nar23, Rum22, Uru22, WG20a, WG20b, WZL21]. use [Ais24]. Using [HJBKR24, KCPMNS21, AHV20, BBW20, BR23b, CLLK21, EN22, GS21b, MB20, Mur22].

value [CEM21, CV21, Cia22, KL23, KN20, LLW20, LHK20, Loe21, MQW20, Sha22, StW22, VCR21]. valued [MvWC23, RT23]. values [ACD22, AG15, AG20, CJ20, CAJ21, Cos21, JT24, KK22b, Li20a, LLW21, Nay23, ZS22a]. Vandermonde [BDGY21, KN20, MMV21, Yan21]. Vanishing [CK24, EGK^{+}24]. variable [GP21, SS23a]. variables [Ais24, GJ23, QSS22, SW23]. variant [ZX20]. Variational [Shi21a]. varieties [BS24a, CVV21, IdSSV21, KMZ20, Koz21]. variety [Elm23]. Vector [KMNP20, RT23, BKS^{+}21, BMM24, CFSiT20, CG20, ĆD20, ĆD23b, DN23, LL24a, MV24, Rum22, WXY24]. Vector-valued [RT23]. vectors [ACG^{+}21, AGNPSE23, CK23b, EL20, FISW21, HW21b, Miy24, NV23b, PV22, QJMW23]. Verdière [MS20b]. version [FGPW24, LS20c]. versions [CZ22a, dVLG23]. Vertex [BD21a, WYF^{+}20, AS24, GLNW24, GMR22, JST21, LTYZ23, PSZ21, RRF23, SZSMB20, WWT20, ZLLZ22, Zhu21]. Vertex-connectivity [WYF^{+}20]. vertex-degree-based [GMR22]. vertex-disjoint [ZLLZ22]. vertex-weighted [SZSMB20]. vertical [MG21]. Vertices [Wój22b, BBG21, BP20, FYS^{+}23, GM22, GZ22, Goe21, KSSW20, KS22c, Mon23, MPRS21, NdLO21, WWJ20, XWT20, dLMO20]. via [AdF20, Ans21, BDN24, Bhu24, BT20b, CDD21, CHHS20, CK23b, DYB20, FKMM21, FKK21a, GLS22, HW21a, KSSW20, LLWW23, LHZ21, MQW20, MPRS21, PP20, SDS21, ZT23, ZBdAF20]. view [Chr24a]. viewpoint [LIW20]. violations [Li20a]. volume [CEF^{+}22, CKM20, SG20].

walk [AFS^{+}23, CMP21, KIS21, MP23b, MPRS21, Wan21, WWG22, Wan23, WFSS20]. walks [BHHP21, BW22, CGJvB22, FP23, God21, IMT20, KKS21, KSY21, Kub22, KS22c, Kub23, SMA20, Zha22]. Ward [LMR21]. Waring [BT20b, KS24, Kis22]. Wasserstein [Bai24, CKL24, DLVV21, HK20, Min22]. wavelets [CCCS21]. way [KMNP20]. Weak [JJL20b, BBNS22, GOZ21, GS21c, MRT20, Roy24]. weak-type [GOZ21]. Weakly [AAF^{+}21, LR23, KS23a, Miy24, TD22]. weight [AHZ22, CEM21, CM23b]. weight-equitable [AHZ22]. Weighted [BGOO21, AHK23, CMM20, COB22a, CN21b, DdF23, EJJL20, FJ24, Gau24, GMSG22, Han21, HJ24, IIM^{+}21, JJLS23, LW21b, LS24b, Pik23, PV21a, SZSMB20, WS22, Zam23, Zhu21]. weights [GM22]. well [KZ23]. well-known [KZ23]. Wenzel [CJ20, CKOS23, GLLZ20]. Weyl [Ces20, FOP22, FKKM22, LP22, MZ22]. Weyr [BFPS21]. wheel [BBG21, Goe21]. whether [Ros22]. which [Alf20a, tHvdM21]. Whittaker [CW23a]. whole [KP23b]. whose [DJ23, GM23, GK20, KLS21, KCML20, LWB20, Lyu20, MR21a, Mel23, SS21c, Sta20b]. widths [BDHS23a]. Wiener [JKS23, SS23a]. Williamson [SZ23a, IGW17]. without [AHKT21, CWZ22, ĆD20, ĆD23b, JMP21a, LWYL20]. witnesses [CJMP22]. Witt [KK23b]. Wong [SL20b]. worst [CV21]. worst-case [CV21]. Wronskian [MPR22]. Wu [CRSY22].

Yamaguti [ABCO22]. Yamamoto [Nay23]. Yau [GLNW24]. Young [HVX23]. Young-type [HVX23].

Zagreb [FLRS20]. Zelisko [BS21b]. Zero [CW23b, GT23, GSV24, AEGV21, BB22, BES21, BMM24, CPS20, KKSV21, KL21b, KLL22, LNP21, MM21, Pan23, Ran23, Vit21]. Zero-product [GT23]. zero-sum-free [LNP21]. zeros [KST22, MMA22]. Zeta [BCR20, Her20, LLH20, Zak21, IIM^{+}21, Kad20, KKS21, Mor20, SC20b, Zhu21]. Zetaequivalent [Her20]. Zhou [SL20b]. Zygmund [RT23].

References

Arizmendi:2021:EGR

[AA21] Gerardo Arizmendi and Octavio Arizmendi. Energy of a graph and Randic index. Linear Algebra and its Applications, 609(??):332–338, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030447X.

Akbari:2022:ELG

[AAAH22] Saieed Akbari, Abdullah Alazemi, Milica An– delić, and Mohammad Ali Hosseinzadeh. On the energy of line graphs. Linear Algebra and its Applications, 636(??):143–153, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004286.

Akbari:2023:CTL

[AADT23] S. Akbari, M. Alaeiyan, M. Darougheh, and V. Trevisan. Classification of trees by Laplacian eigenvalue distribution and edge covering number. Linear Algebra and its Applications, 677(??):221–236, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003075.

Adm:2021:WHD

[AAF^{+}21] Mohammad Adm, Khawla Almuhtaseb, Shaun Fallat, Karen Meagher, Shahla Nasserasr, Mahsa N. Shirazi, and A. S. Razafimahatratra. Weakly Hadamard diagonalizable graphs.

Linear Algebra and its Applications, 610(??):86–119, February 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379520304663.

Abiad:2023:BDN

[AAFM23] A. Abiad, S. Akbari, M. H. Fakharan, and A. Mehdizadeh. A bound for the p-domination number of a graph in terms of its eigenvalue multiplicities. Linear Algebra and its Applications, 658(??):319– 330, February 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379522004086.

Ahanjideh:2022:LED

[AAFT22] M. Ahanjideh, S. Akbari, M. H. Fakharan, and V. Trevisan. Laplacian eigenvalue distribution and graph parameters. Linear Algebra and its Applications, 632(??):1–14, January 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003451.

Ahmad:2022:ASC

[AAH^{+}22] Uzma Ahmad, Saieed Akbari, Saira Hameed, Mohammad Ali Nematollahi, and Faiza Saeed. Addendum to “Spectral characterizations of signed cycles” [Linear Algebra Appl. 553 (2018) 307–327]. Linear Algebra and its Applications, 651(??):83–89, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002245.

Abiad:2022:CG

[AAHV22] Aida Abiad, Carlos A. Alfaro, Kristin Heysse, and Marcos C. Vargas. Codeterminantal graphs. Linear Algebra and its Applications, 650(??):1–25, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002142.

Andrade:2020:SSC

[AAMR20] Enide Andrade, Luis Arrieta, Cristina Manzaneda, and María Robbiano. On the spectra of some g-circulant matrices and applications to nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 590(??):1–21, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305439.

Araujo:2022:MTI

[ABC^{+}22] João Araújo, Wolfram Bentz, Peter J. Cameron, Michael Kinyon, and Janusz Konieczny. Matrix theory for independence algebras. Linear Algebra and its Applications, 642(??):221–250, June 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000647.

Armijo:2023:CIS

[ABCH23] Nicolas F. Armijo, Yunier Bello-Cruz, and Gabriel Haeser. On the convergence of iterative schemes for solving a piecewise linear system of equations. Linear Algebra and its Applications, 665(??):291–314, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000423.

Abdelwahab:2022:CNL

[ABCO22] Hani Abdelwahab, Elisabete Barreiro, Antonio J. Calderón, and Amir Fernández Ouaridi. The classification of nilpotent Lie–Yamaguti algebras. Linear Algebra and its Applications, 654(??):339–378, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379522003238.

Alhevaz:2020:SLE

[ABG20] Abdollah Alhevaz, Maryam Baghipur, and Hilal Ahmad Ganie. On the second largest eigenvalue of the generalized distance matrix of graphs. Linear Algebra and its Applications, 603(??):226–241, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302664.

Allem:2023:DFA

[ABH^{+}23] L. Emilio Allem, Rodrigo O. Braga, Carlos Hoppen, Elismar R. Oliveira, Lucas S. Sibemberg, and Vilmar Trevisan. Diminimal families of arbitrary diameter. Linear Algebra and its Applications, 676(??):318–351, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002781.

Armenta:2022:DFM

[ABHR22] Marco Armenta, Thomas Brüstle, Souheila Hassoun, and Markus Reineke. Double framed moduli spaces of quiver representations. Linear Algebra and its Applications, 650(??):98–131, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002129.

Arevalo:2021:GFD

[ABR21] N. E. Arévalo, R. O. Braga, and V. M. Rodrigues. Graphs with few distinct eigenvalues and extremal energy. Linear Algebra and its Applications, 620(??):147–167, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100094X.

Andrade:2022:SFI

[ABR^{+}22] Enide Andrade, Andréa Soares Bonifácio, María Robbiano, Jonnathan Rodríguez, and Katherine Tapia. Some families of integral mixed graphs. Linear Algebra and its Applications, 641(??):48–66, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000374.

Alfaro:2021:GFT

[ABSV21] Carlos A. Alfaro, Michael D. Barrus, John Sinkovic, and Ralihe R. Villagrán. Graphs with few trivial characteristic ideals. Linear Algebra and its Applications, 615(??):155–174, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305917.

Araujo:2022:MDP

[ABW22] Igor Araujo, József Balogh, and Yuzhou Wang. Maximum determinant and permanent of sparse 0–1 matrices. Linear Algebra and its Applications, 645(??):194–228, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001215.

Altenberg:2020:NSR

[AC20a] Lee Altenberg and Joel E. Cohen. Nonconcavity of the spectral radius in Levinger’s theorem. Linear Algebra and its Applications, 606 (??):201–218, December 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303517.

Angelini:2020:ITF

[AC20b] Elena Angelini and Luca Chiantini. On the identifiability of ternary forms. Linear Algebra and its Applications, 599(??):36–65, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301774.

Amorim:2021:CPS

[AC21] Érik Amorim and Eric A. Carlen. Complete positivity and self-adjointness. Linear Algebra and its Applications, 611(??):389–439, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305218.

Antipov:2022:LTA

[AC22] Mikhaylo Antipov and Danila Cherkashin. Lovász theta approach to eventown problem. Linear Algebra and its Applications, 655(??):302– 313, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437952200252X.

Andrade:2022:PVC

[ACD22] Enide Andrade, Lorenzo Ciardo, and Geir Dahl. Perron values and classes of trees. Linear Algebra and its Applications, 639(??):135–158, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000064.

Averkov:2021:LCU

[ACDG21] G. Averkov, A. Chavez, J. A. De Loera, and B. Gillespie. The lattice of cycles of an undirected graph. Linear Algebra and its Applications, 611(??):213–236, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305103.

Akbari:2021:QHR

[ACG^{+}21] S. Akbari, S. M. Cioabă, S. Goudarzi, A. Niaparast, and A. Tajdini. On a question of Haemers regarding vectors in the nullspace of Seidel matrices. Linear Algebra and its Applications, 615(??):194–206, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000057.

Alpay:2020:PSF

[ACLS20] Daniel Alpay, Fabrizio Colombo, Izchak Lewkowicz, and Irene Sabadini. On pseudo-spectral factorization over the complex numbers and quaternions. Linear Algebra and its Applications, 606(??):90–126, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303499.

Arias:2020:COS

[ACM20] M. Laura Arias, Gustavo Corach, and Alejandra Maestripieri. On complementable operators in the sense of T. Ando. Linear Algebra and its Applications, 594(??):158–176, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300781.

Amparore:2022:FFM

[ACM22] Elvio G. Amparore, Gianfranco Ciardo, and Andrew S. Miner. The footprint form of a matrix: Definition, properties, and an application. Linear Algebra and its Applications, 651(??):209–229, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002385.

Arias:2024:SMP

[ACMM24] M. Laura Arias, Maximiliano Contino, Alejandra Maestripieri, and Stefania Marcantognini. Semiclosed multivalued projections. Linear Algebra and its Applications, 683(??):180–200, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004433.

Andrade:2021:CLM

[ACOM21] Enide Andrade, Dante Carrasco-Olivera, and Cristina Manzaneda. On circulant like matrices properties involving Horadam, Fibonacci, Jacobsthal and Pell numbers. Linear Algebra and its Applications, 617(??):100–120, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000355.

Alpay:2022:BLT

[ACSS22] Daniel Alpay, Fabrizio Colombo, Irene Sabadini, and Baruch Schneider. Beurling–Lax type theorems and Cuntz relations. Linear Algebra and its Applications, 633(??):152–212, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003736.

Albertini:2020:CSC

[AD20a] Francesca Albertini and Domenico D’Alessandro. Corrigendum to “Subspace controllability of bipartite symmetric spin networks” [Linear Algebra Appl. 585 (2020) 1–23]. Linear Algebra and its Applications, 603(??):508–510, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303141.  See

[AD20b].

Albertini:2020:SCB

[AD20b] Francesca Albertini and Domenico D’Alessandro. Subspace controllability of bipartite symmetric spin networks. Linear Algebra and its Applications, 585(??):1–23, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304239. See cor-

rigendum [AD20a].

Andrade:2024:CFT

[AD24] Enide Andrade and Geir Dahl. Combinatorial Fiedler theory and graph partition. Linear Algebra and its Applications, 687(??):229–251, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000430.

Altavilla:2020:ESS

[AdF20] A. Altavilla and C. de Fabritiis. Equivalence of slice semi-regular functions via Sylvester operators. Linear Algebra and its Applications, 607(??):151–189, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303827.

Adhikari:2021:USM

[ADGK21] Bibhas Adhikari, Biswa Nath Datta, Tinku Ganai, and Michael Karow. Updating structured matrix pencils with no spillover effect on unmeasured spectral data and deflating pair. Linear Algebra and its Applications, 624(??):318–348, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001816.

Akbari:2021:HND

[ADGKM21] S. Akbari, K. C. Das, S. Khalashi Ghezelahmad, and F. Koorepazan-Moftakhar. Hypoenergetic and nonhypoenergetic digraphs. Linear Algebra and its Applications, 618(??):129–143, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000458.

Arav:2024:MMM

[ADH^{+}24] Marina Arav, Louis Deaett, H. Tracy Hall, Hein van der Holst, and Derek Young. A matching-minor monotone parameter for bipartite graphs. Linear Algebra and its Applications, 680(??):254–273, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003877.

Anguas:2023:QTM

[ADHM23] L. M. Anguas, F. M. Dopico, R. Hollister, and D. S. Mackey. Quasi-triangularization of matrix polynomials over arbitrary fields. Linear Algebra and its Applications, 665(??):61–106, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000290.

Abiad:2023:CMG

[ADHS23] Aida Abiad, Jozefien D’haeseleer, Willem H. Haemers, and Robin Simoens. Cospectral mates for generalized Johnson and Grassmann graphs. Linear Algebra and its Applications, 678(??):1–15, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003208.

Alencar:2021:GAS

[AdL21] Jorge Alencar and Leonardo de Lima. On graphs with adjacency and signless Laplacian matrices eigenvectors entries in {−1, +1}. Linear Algebra and its Applications, 614(??):301–315, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302378.

Awa:2020:ERS

[AdlC20] Dominic Awa and Ralph John de la Cruz. Every real symplectic matrix is a product of real symplectic involutions. Linear Algebra and its Applications, 589(??):85–95, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305105.

Afable:2020:DRP

[AdlCPS20] Edgar Aivan Afable, Ralph John de la Cruz, Agnes T. Paras, and Mary Elizabeth Segui. Diagonalizability with respect to perplectic and pseudounitary similarity transformations. Linear Algebra and its Applications, 591(??):61–71, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300148.

Alencar:2023:GEM

[AdLN23] Jorge Alencar, Leonardo de Lima, and Vladimir Nikiforov. On graphs with eigenvectors in {-1,0,1} and the max k-cut problem. Linear Algebra and its Applications, 663(??):222–240, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000058.

Amparan:2021:MBI

[ADMZ21] A. Amparan, F. M. Dopico, S. Marcaida, and I. Zaballa. On minimal bases and indices of rational matrices and their linearizations. Linear Algebra and its Applications, 623(??):14–67, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000331.

Adriaensen:2022:EKR

[Adr22] Sam Adriaensen. Erdős–Ko– Rado theorems for ovoidal circle geometries and polynomials over finite fields. Linear Algebra and its Applications, 643(??):1–38, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000568.

Allem:2022:SCS

[AdST22] L. Emilio Allem, Lucas G. M. da Silveira, and Vilmar Trevisan. On the spectral characterization of the p-sun and the (p, q)-double sun. Linear Algebra and its Applications, 636(??):1–24, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004109.

Arav:2021:SGS

[ADvdH21] Marina Arav, F. Scott Dahlgren, and Hein van der Holst. Signed graphs with stable maximum nullity at most two. Linear Algebra and its Applications, 620(??):124–146, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000768.

Arav:2023:SGM

[ADvdH23] Marina Arav, F. Scott Dahlgren, and Hein van der Holst. Signed graphs with maximum nullity two. Linear Algebra and its Applications, 675(??):29–47, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002410.

Andreotti:2021:MSS

[AEGL21] Eleonora Andreotti, Dominik Edelmann, Nicola Guglielmi, and Christian Lubich. Measuring the stability of spectral clustering. Linear Algebra and its Applications, 610(??):673–697, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304924.

Alaminos:2021:SZP

[AEGV21] J. Alaminos, J. Extremera, M. L. C. Godoy, and A. R. Villena. Strongly zero product determined Banach algebras. Linear Algebra and its Applications, 630 (??):326–354, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003359.

Akbari:2021:MES

[AFG^{+}21] S. Akbari, F. A. M. França, E. Ghasemian, M. Javarsineh, and L. S. de Lima. The main eigenvalues of signed graphs. Linear Algebra and its Applications, 614(??):270–280, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302238.

Abiad:2023:BSM

[AFK^{+}23] Aida Abiad, Shaun M. Fallat, Mark Kempton, Rupert H. Levene, Polona Oblak, Helena Šmigoc, Michael Tait, and Kevin N. Vander Meulen. Bordering of symmetric matrices and an application to the minimum number of distinct eigenvalues for the join of graphs. Linear Algebra and its Applications, 679(??):104–126, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003506.

Asahara:2023:SMT

[AFS^{+}23] Keisuke Asahara, Daiju Funakawa, Etsuo Segawa, Akito Suzuki, and Noriaki Teranishi. Spectral mapping theorem of an abstract nonunitary quantum walk. Linear Algebra and its Applications, 676(??):1–24, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002628.

Aguilar:2020:RAR

[AFSS20] Cesar O. Aguilar, Matthew Ficarra, Natalie Schurman, and Brittany Sullivan. The role of the anti-regular graph in the spectral analysis of threshold graphs. Linear Algebra and its Applications, 588(??):210–223, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305129.

Ay:2015:RSA

[AG15] Serdar Ay and Aurelian Gheondea. Representations of ∗-semigroups associated to invariant kernels with values adjointable operators. Linear Algebra and its Applications, 486(??):361–388, December 1, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379515004772. See cor-

rigendum [AG20].

Ay:2020:CRS

[AG20] Serdar Ay and Aurelian Gheondea. Corrigendum to “Representations of ∗- semigroups associated to invariant kernels with values adjointable operators” [Linear Algebra Appl. 486 (2015) 361–388]. Linear Algebra and its Applications, 589(??):242–246, March 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300136.  See

[AG15].

Adm:2021:CPI

[AG21] Mohammad Adm and Jürgen Garloff. Characterization, perturbation, and interval property of certain sign regular matrices. Linear Algebra and its Applications, 612(??):146–161, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379520305607.

Almeida:2024:MTP

[AG24] Igor Almeida and Nikolay Gusevskii. Moduli of triples of points in quaternionic hyperbolic geometry. Linear Algebra and its Applications, 686(??):33–63, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000144.

Akbari:2020:SLB

[AGH20] Saieed Akbari, Amir Hossein Ghodrati, and Mohammad Ali Hosseinzadeh. Some lower bounds for the energy of graphs. Linear Algebra and its Applications, 591(??):205–214, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030001X.

Arambasic:2021:OPM

[AGK^{+}21] Ljiljana Arambašić, Alexander Guterman, Bojan Kuzma, Rajna Rajić, and Svetlana Zhilina. Operators preserving mutual strong Birkhoff–James orthogonality on B(H). Linear Algebra and its Applications, 624(??):27–43, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379521001592.

Araiza:2023:APQ

[AGKS23] Roy Araiza, Colton Griffin, Aneesh Khilnani, and Thomas Sinclair. Approximating projections by quantum operations. Linear Algebra and its Applications, 663(??):178–199, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000083.

Akbari:2020:MPC

[AGNN20] S. Akbari, A. Ghafari, M. Nahvi, and M. A. Nematollahi. Mixed paths and cycles determined by their spectrum. Linear Algebra and its Applications, 586(??):325–346, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519303970.

Alonso-Gonzalez:2023:FCD

[AGNPSE23] Clementa Alonso-González, Miguel Ángel Navarro-Pérez, and Xaro Soler-Escrivà. Flag codes: Distance vectors and cardinality bounds. Linear Algebra and its Applications, 656(??):27–62, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S002437952200341X.

Agore:2024:FBR

[Ago24] A. L. Agore. Functors between representation categories. Universal modules. Linear Algebra and its Applications, 688(??):104–119, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000533.

Armentia:2020:NMP

[AGV20] Gorka Armentia, Juan-Miguel Gracia, and Francisco-Enrique Velasco. Nearest matrix with a prescribed eigenvalue of bounded multiplicities. Linear Algebra and its Applications, 592(??):188–209, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300367.

Alrimawi:2023:NII

[AHK23] Fadi Alrimawi, Omar Hirzallah, and Fuad Kittaneh. Norm inequalities involving the weighted numerical radii of operators. Linear Algebra and its Applications, 657(??):127–146, January 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379522003834.

Abreu:2021:PSI

[AHKT21] Nair Abreu, Christoph Helmberg, Gitta Kutyniok, and Vilmar Trevisan. Preface to the Special Issue of ILAS 2019: Linear Algebra without Borders, The 22nd ILAS Conference, Rio de Janeiro, Brazil, July 08–12, 2019. Linear Algebra and its Applications, 614(??):1–4, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304055.

Arav:2021:TCS

[AHLvdH21] Marina Arav, Frank J. Hall, Zhongshan Li, and Hein van der Holst. Two-connected signed graphs with maximum nullity at most two. Linear Algebra and its Applications, 611(??):82–93, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305541.

Ahmadi:2020:PQS

[AHM20] Bahman Ahmadi, M. H. Shirdareh Haghighi, and Ahmad Mokhtar. Perfect quantum state transfer on the Johnson scheme. Linear Algebra and its Applications, 584(??):326–342, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304045.

Arrigo:2022:DKR

[AHNW22] Francesca Arrigo, Desmond J. Higham, Vanni Noferini, and Ryan Wood. Dynamic Katz and related network measures. Linear Algebra and its Applications, 655 (??):159–185, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003068.

Albrecht:2020:IOP

[AHV20] Amie Albrecht, Phil Howlett, and Geetika Verma. Inversion of operator pencils on Banach space using Jordan chains when the generalized resolvent has an isolated essential singularity. Linear Algebra and its Applications, 595(??):33–62, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301026.

Abiad:2022:CCW

[AHZ22] Aida Abiad, Christopher Hojny, and Sjanne Zeijlemaker. Characterizing and computing weight-equitable partitions of graphs. Linear Algebra and its Applications, 645(??):30–51, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000982.

Ai:2024:SRG

[AIK^{+}24] Jiangdong Ai, Seonghyuk Im, Jaehoon Kim, Hyunwoo Lee, Suil O, and Liwen Zhang. On the spectral radius of graphs with given maximum degree and girth. Linear Algebra and its Applications, 691(??):182–195, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001162.

Aishima:2024:CEU

[Ais24] Kensuke Aishima. Consistent estimation with the use of orthogonal projections for a linear regression model with errors in the variables. Linear Algebra and its Applications, 684(??):101–126, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004652.

Abreu:2021:ILE

[AJM21] Nair Abreu, Claudia Marcela Justel, and Lilian Markenzon. Integer Laplacian eigenvalues of chordal graphs. Linear Algebra and its Applications, 614(??):68–81, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305440.

Abreu:2024:SCG

[AJM24] Nair Abreu, Claudia Marcela Justel, and Lilian Markenzon. Strictly chordal graphs: Structural properties and integer Laplacian eigenvalues. Linear Algebra and its Applications, 682(??):351–362, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004263.

Azimi:2024:MFP

[AJNS24] Ali Azimi, Rakesh Jana, Mukesh Kumar Nagar, and Sivaramakrishnan Sivasubramanian. The maximum four point condition matrix of a tree. Linear Algebra and its Applications, 691(??):133–150, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001137.

Ahmad:2020:SPA

[AK20] Sk. Safique Ahmad and Prince Kanhya. Structured perturbation analysis of sparse matrix pencils with s-specified eigenpairs. Linear Algebra and its Applications, 602(??):93–119, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302317.

Alp:2023:SCA

[AK23] Yasemin Alp and E. Gokcen Kocer. Sequence characterization of almost-Riordan arrays. Linear Algebra and its Applications, 664(??):1–23, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000186.

Ahiable:2021:EBC

[AKL^{+}21] Jennifer Ahiable, David W. Kribs, Jeremy Levick, Rajesh Pereira, and Mizanur Rahaman. Entanglement breaking channels, stochastic matrices, and primitivity. Linear Algebra and its Applications, 629(??):219–231, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003104.

Abbasi:2023:ACP

[AKN23] Paria Abbasi, Andreas Klingler, and Tim Netzer. Approximate completely positive semidefinite factorizations and their ranks. Linear Algebra and its Applications, 677(??):323–336, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300304X.

Atreas:2019:DMD

[AKPS19] N. Atreas, N. Karantzas, M. Papadakis, and T. Stavropoulos. On the design of multidimensional compactly supported Parseval framelets with directional characteristics. Linear Algebra and its Applications, 582(??):1–36, December 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519303155. See er-

ratum [KAPS20].

Akkus:2023:UAG

[AKVS23] Ilker Akkus, Gonca Kizilaslan, and Luis Verde-Star. A unified approach to generalized Pascal-like matrices: q-analysis. Linear Algebra and its Applications, 673(??):138–159, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300188X.

Alpay:2021:QHP

[AL21] Daniel Alpay and Izchak Lewkowicz. Quantitatively hyper-positive real functions. Linear Algebra and its Applications, 623(??):316–334, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305395.

Akbari:2023:ILB

[AL23a] Saieed Akbari and Hongying Lin. Improved lower bound on the energy of line graphs. Linear Algebra and its Applications, 674(??):442–452, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002069.

Alpay:2023:RLR

[AL23b] Daniel Alpay and Izchak Lewkowicz. On realization of left and right products of rational functions. Linear Algebra and its Applications, 666(??):42–61, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000630.

Alakhrass:2021:SMT

[Ala21] Mohammad Alakhrass. On sectorial matrices and their inequalities. Linear Algebra and its Applications, 617(??):179–189, May 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000550.

Alfakih:2020:USE

[Alf20a] A. Y. Alfakih. On unit spherical Euclidean distance matrices which differ in one entry. Linear Algebra and its Applications, 594(??):29–50, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300793.

Alfaro:2020:GTD

[Alf20b] Carlos A. Alfaro. On graphs with 2 trivial distance ideals. Linear Algebra and its Applications, 597(??):69–85, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301452.

Alizadeh:2020:CEH

[Ali20] Rahim Alizadeh. Corrigendum to “An extension of Haagerup–Harpe inequality” [Linear Algebra Appl. 577 (2019) 69–77]. Linear Algebra and its Applications, 600(??):161–164, September 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520302202.  See

[HAD19].

Andrade:2020:EGE

[ALMZ^{+}20] Enide Andrade, Eber Lenes, Exequiel Mallea-Zepeda, María Robbiano, and Jonnathan Rodríguez Z. Extremal graphs for Estrada indices. Linear Algebra and its Applications, 588(??):54–73, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304690.

Altafi:2022:JTS

[Alt22] Nasrin Altafi. Jordan types with small parts for Artinian Gorenstein algebras of codimension three. Linear Algebra and its Applications, 646(??):54–83, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001148.

Agore:2021:ACJ

[AM21] A. L. Agore and G. Militaru. Algebraic constructions for Jacobi–Jordan algebras. Linear Algebra and its Applications, 630 (??):158–178, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002949.

Abyzov:2023:ADS

[AM23a] Adel Abyzov and Alexander Maklakov. Automorphisms of dense subrings of the endomorphism ring of a free module. Linear Algebra and its Applications, 679(??):220–230, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300366X.

Abyzov:2023:LAL

[AM23b] Adel Abyzov and Alexander Maklakov. Locally algebraic linear operators and their centralizers. Linear Algebra and its Applications, 662(??):1–17, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004621.

Akgul:2024:LSS

[AMO24] S. Dogruer Akgul, A. Minenkova, and V. Olshevsky. Lipschitz stability of some canonical Jordan bases of real H -selfadjoint matrices under small perturbations. Linear Algebra and its Applications, 686(??):64–82, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000181.

Agudelo:2020:CDG

[AMR20] Natalia Agudelo, Juan Monsalve, and Juan Rada. On the characterization of digraphs with given rank. Linear Algebra and its Applications, 587(??):215–227, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304781.

Arrieta:2021:SRD

[AMS21] Luis E. Arrieta, Alfredo D. Millano, and Ricardo L. Soto. On spectra realizable and diagonalizably realizable. Linear Algebra and its Applications, 612(??):273–288, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305267.

Abiad:2021:CNL

[AMZ21] Aida Abiad, Raffaella Mulas, and Dong Zhang. Coloring the normalized Laplacian for oriented hypergraphs. Linear Algebra and its Applications, 629(??):192–207, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379521002834.

Arabyani-Neyshaburi:2020:NSO

[ANAS20] Fahimeh Arabyani-Neyshaburi, Ali Akbar Arefijamaal, and Ghadir Sadeghi. Numerically and spectrally optimal dual frames in Hilbert spaces. Linear Algebra and its Applications, 604(??):52–71, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302962.

Al-Natoor:2022:UIN

[ANBK22] Ahmad Al-Natoor, Sakina Benzamia, and Fuad Kittaneh. Unitarily invariant norm inequalities for positive semidefinite matrices. Linear Algebra and its Applications, 633(??):303–315, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003773.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Linear Algebra and its Applications, 584(??):i, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304379.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Linear Algebra and its Applications, 584(??):ii, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304380.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Linear Algebra and its Applications, 585(??):i, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304574.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Linear Algebra and its Applications, 585(??):ii, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304586.

Anonymous:2020:EBe

[Ano20e] Anonymous. Editorial Board. Linear Algebra and its Applications, 586(??):i, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304811.

Anonymous:2020:EBf

[Ano20f] Anonymous. Editorial Board. Linear Algebra and its Applications, 586(??):ii, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304823.

Anonymous:2020:EBg

[Ano20g] Anonymous. Editorial Board. Linear Algebra and its Applications, 587(??):i, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305270.

Anonymous:2020:EBh

[Ano20h] Anonymous. Editorial Board. Linear Algebra and its Applications, 587(??):ii, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305282.

Anonymous:2020:EBi

[Ano20i] Anonymous. Editorial Board. Linear Algebra and its Applications, 588(??):i, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305518.

Anonymous:2020:EBj

[Ano20j] Anonymous. Editorial Board. Linear Algebra and its Applications, 588(??):ii, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930552X.

Anonymous:2020:EBk

[Ano20k] Anonymous. Editorial Board. Linear Algebra and its Applications, 589(??):i, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030029X.

Anonymous:2020:EBl

[Ano20l] Anonymous. Editorial Board. Linear Algebra and its Applications, 589(??):ii, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300306.

Anonymous:2020:EBm

[Ano20m] Anonymous. Editorial Board. Linear Algebra and its Applications, 590(??):i, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300495.

Anonymous:2020:EBn

[Ano20n] Anonymous. Editorial Board. Linear Algebra and its Applications, 590(??):ii, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300501.

Anonymous:2020:EBo

[Ano20o] Anonymous. Editorial Board. Linear Algebra and its Applications, 591(??):i, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300707.

Anonymous:2020:EBp

[Ano20p] Anonymous. Editorial Board. Linear Algebra and its Applications, 591(??):ii, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300719.

Anonymous:2020:EBq

[Ano20q] Anonymous. Editorial Board. Linear Algebra and its Applications, 592(??):i, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300872.

Anonymous:2020:EBr

[Ano20r] Anonymous. Editorial Board. Linear Algebra and its Applications, 592(??):ii, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300884.

Anonymous:2020:EBs

[Ano20s] Anonymous. Editorial Board. Linear Algebra and its Applications, 593(??):i, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301105.

Anonymous:2020:EBt

[Ano20t] Anonymous. Editorial Board. Linear Algebra and its Applications, 593(??):ii, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301117.

Anonymous:2020:EBu

[Ano20u] Anonymous. Editorial Board. Linear Algebra and its Applications, 594(??):i, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301245.

Anonymous:2020:EBv

[Ano20v] Anonymous. Editorial Board. Linear Algebra and its Applications, 594(??):ii, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301257.

Anonymous:2020:EBw

[Ano20w] Anonymous. Editorial Board. Linear Algebra and its Applications, 595(??):i, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301373.

Anonymous:2020:EBx

[Ano20x] Anonymous. Editorial Board. Linear Algebra and its Applications, 595(??):ii, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301385.

Anonymous:2020:EBy

[Ano20y] Anonymous. Editorial Board. Linear Algebra and its Applications, 596(??):i, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301622.

Anonymous:2020:EBz

[Ano20z] Anonymous. Editorial Board. Linear Algebra and its Applications, 596(??):ii, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301634.

Anonymous:2020:EBba

[Ano20-27] Anonymous. Editorial Board. Linear Algebra and its Applications, 597(??):i, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301968.

Anonymous:2020:EBbb

[Ano20-28] Anonymous. Editorial Board. Linear Algebra and its Applications, 597(??):ii, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030197X.

Anonymous:2020:EBbc

[Ano20-29] Anonymous. Editorial Board. Linear Algebra and its Applications, 598(??):i, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030207X.

Anonymous:2020:EBbd

[Ano20-30] Anonymous. Editorial Board. Linear Algebra and its Applications, 598(??):ii, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302081.

Anonymous:2020:EBbe

[Ano20-31] Anonymous. Editorial Board. Linear Algebra and its Applications, 599(??):i, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302263.

Anonymous:2020:EBbf

[Ano20-32] Anonymous. Editorial Board. Linear Algebra and its Applications, 599(??):ii, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302275.

Anonymous:2020:EBbg

[Ano20-33] Anonymous. Editorial Board. Linear Algebra and its Applications, 600(??):i, September 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302457.

Anonymous:2020:EBbh

[Ano20-34] Anonymous. Editorial Board. Linear Algebra and its Applications, 600(??):ii, September 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302469.

Anonymous:2020:EBbi

[Ano20-35] Anonymous. Editorial Board. Linear Algebra and its Applications, 601(??):i, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302743.

Anonymous:2020:EBbj

[Ano20-36] Anonymous. Editorial Board. Linear Algebra and its Applications, 601(??):ii, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302755.

Anonymous:2020:EBbk

[Ano20-37] Anonymous. Editorial Board. Linear Algebra and its Applications, 602(??):i, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302883.

Anonymous:2020:EBbl

[Ano20-38] Anonymous. Editorial Board. Linear Algebra and its Applications, 602(??):ii, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302895.

Anonymous:2020:EBbm

[Ano20-39] Anonymous. Editorial Board. Linear Algebra and its Applications, 603(??):i, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303578.

Anonymous:2020:EBbn

[Ano20-40] Anonymous. Editorial Board. Linear Algebra and its Applications, 603(??):ii, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030358X.

Anonymous:2020:EBbo

[Ano20-41] Anonymous. Editorial Board. Linear Algebra and its Applications, 604(??):i, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303712.

Anonymous:2020:EBbp

[Ano20-42] Anonymous. Editorial Board. Linear Algebra and its Applications, 604(??):ii, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303724.

Anonymous:2020:EBbq

[Ano20-43] Anonymous. Editorial Board. Linear Algebra and its Applications, 605(??):i, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304018.

Anonymous:2020:EBbr

[Ano20-44] Anonymous. Editorial Board. Linear Algebra and its Applications, 605(??):ii, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030402X.

Anonymous:2020:EBbs

[Ano20-45] Anonymous. Editorial Board. Linear Algebra and its Applications, 606(??):i, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304237.

Anonymous:2020:EBbt

[Ano20-46] Anonymous. Editorial Board. Linear Algebra and its Applications, 606(??):ii, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304249.

Anonymous:2020:EBbu

[Ano20-47] Anonymous. Editorial Board. Linear Algebra and its Applications, 607(??):i, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304584.

Anonymous:2020:EBbv

[Ano20-48] Anonymous. Editorial Board. Linear Algebra and its Applications, 607(??):ii, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304596.

Anonymous:2020:PSa

[Ano20-49] Anonymous. Pages 1–164 (1 September 2020). Linear Algebra and its Applications, 600 (??):1–164, September 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PAc

[Ano20-50] Anonymous. Pages 1–186 (1 August 2020). Linear Algebra and its Applications, 598(??): 1–186, August 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PJd

[Ano20-51] Anonymous. Pages 1–192 (15 June 2020). Linear Algebra and its Applications, 595(??): 1–192, June 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PAd

[Ano20-52] Anonymous. Pages 1–204 (15 August 2020). Linear Algebra and its Applications, 599(??): 1–204, August 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PJf

[Ano20-53] Anonymous. Pages 1–206 (15 July 2020). Linear Algebra and its Applications, 597(??): 1–206, July 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PMb

[Ano20-54] Anonymous. Pages 1–246 (15 March 2020). Linear Algebra and its Applications, 589(??): 1–246, March 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PDa

[Ano20-55] Anonymous. Pages 1–248 (1 December 2020). Linear Algebra and its Applications, 606 (??):1–248, December 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PJb

[Ano20-56] Anonymous. Pages 1–262 (15 January 2020). Linear Algebra and its Applications, 585(??): 1–262, January 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PNb

[Ano20-57] Anonymous. Pages 1–262 (15 November 2020). Linear Algebra and its Applications, 605(??):1–262, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PJe

[Ano20-58] Anonymous. Pages 1–280 (1 July 2020). Linear Algebra and its Applications, 596(??): 1–280, July 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PJc

[Ano20-59] Anonymous. Pages 1–300 (1 June 2020). Linear Algebra and its Applications, 594(??): 1–300, June 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:POa

[Ano20-60] Anonymous. Pages 1–300 (1 October 2020). Linear Algebra and its Applications, 602(??): 1–300, October 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PMd

[Ano20-61] Anonymous. Pages 1–304 (15 May 2020). Linear Algebra and its Applications, 593(??): 1–304, May 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PMc

[Ano20-62] Anonymous. Pages 1–306 (1 May 2020). Linear Algebra and its Applications, 592(??): 1–306, May 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PAa

[Ano20-63] Anonymous. Pages 1–332 (1 April 2020). Linear Algebra and its Applications, 590(??): 1–332, April 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PSb

[Ano20-64] Anonymous. Pages 1–338 (15 September 2020). Linear Algebra and its Applications, 601(??):1–338, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PAb

[Ano20-65] Anonymous. Pages 1–352 (15 April 2020). Linear Algebra and its Applications, 591(??): 1–352, April 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PDb

[Ano20-66] Anonymous. Pages 1–378 (15 December 2020). Linear Algebra and its Applications, 607(??):1–378, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PFb

[Ano20-67] Anonymous. Pages 1–392 (15 February 2020). Linear Algebra and its Applications, 587 (??):1–392, February 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PFa

[Ano20-68] Anonymous. Pages 1–410 (1 February 2020). Linear Algebra and its Applications, 586 (??):1–410, February 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PJa

[Ano20-69] Anonymous. Pages 1–484 (1 January 2020). Linear Algebra and its Applications, 584(??): 1–484, January 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PMa

[Ano20-70] Anonymous. Pages 1–496 (1 March 2020). Linear Algebra and its Applications, 588(??): 1–496, March 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:PNa

[Ano20-71] Anonymous. Pages 1–508 (1 November 2020). Linear Algebra and its Applications, 604 (??):1–508, November 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2020:POb

[Ano20-72] Anonymous. Pages 1–510 (15 October 2020). Linear Algebra and its Applications, 603(??): 1–510, October 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Linear Algebra and its Applications, 608(??):i, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304845.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Linear Algebra and its Applications, 608(??):ii, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304857.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Linear Algebra and its Applications, 609(??):i, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305036.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Linear Algebra and its Applications, 609(??):ii, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305048.

Anonymous:2021:EBe

[Ano21e] Anonymous. Editorial Board. Linear Algebra and its Applications, 610(??):i, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305486.

Anonymous:2021:EBf

[Ano21f] Anonymous. Editorial Board. Linear Algebra and its Applications, 610(??):ii, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305498.

Anonymous:2021:EBg

[Ano21g] Anonymous. Editorial Board. Linear Algebra and its Applications, 611(??):i, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305796.

Anonymous:2021:EBh

[Ano21h] Anonymous. Editorial Board. Linear Algebra and its Applications, 611(??):ii, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305802.

Anonymous:2021:EBi

[Ano21i] Anonymous. Editorial Board. Linear Algebra and its Applications, 612(??):i, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305942.

Anonymous:2021:EBj

[Ano21j] Anonymous. Editorial Board. Linear Algebra and its Applications, 612(??):ii, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305954.

Anonymous:2021:EBk

[Ano21k] Anonymous. Editorial Board. Linear Algebra and its Applications, 613(??):i, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000112.

Anonymous:2021:EBl

[Ano21l] Anonymous. Editorial Board. Linear Algebra and its Applications, 613(??):ii, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000124.

Anonymous:2021:EBm

[Ano21m] Anonymous. Editorial Board. Linear Algebra and its Applications, 614(??):i, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000276.

Anonymous:2021:EBn

[Ano21n] Anonymous. Editorial Board. Linear Algebra and its Applications, 614(??):ii, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000288.

Anonymous:2021:EBo

[Ano21o] Anonymous. Editorial Board. Linear Algebra and its Applications, 615(??):i, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000483.

Anonymous:2021:EBp

[Ano21p] Anonymous. Editorial Board. Linear Algebra and its Applications, 615(??):ii, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000495.

Anonymous:2021:EBq

[Ano21q] Anonymous. Editorial Board. Linear Algebra and its Applications, 616(??):i, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100063X.

Anonymous:2021:EBr

[Ano21r] Anonymous. Editorial Board. Linear Algebra and its Applications, 616(??):ii, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000641.

Anonymous:2021:EBs

[Ano21s] Anonymous. Editorial Board. Linear Algebra and its Applications, 617(??):i, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000872.

Anonymous:2021:EBt

[Ano21t] Anonymous. Editorial Board. Linear Algebra and its Applications, 617(??):ii, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000884.

Anonymous:2021:EBu

[Ano21u] Anonymous. Editorial Board. Linear Algebra and its Applications, 618(??):i, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001245.

Anonymous:2021:EBv

[Ano21v] Anonymous. Editorial Board. Linear Algebra and its Applications, 618(??):ii, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001257.

Anonymous:2021:EBw

[Ano21w] Anonymous. Editorial Board. Linear Algebra and its Applications, 619(??):i, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001415.

Anonymous:2021:EBx

[Ano21x] Anonymous. Editorial Board. Linear Algebra and its Applications, 619(??):ii, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001427.

Anonymous:2021:EBy

[Ano21y] Anonymous. Editorial Board. Linear Algebra and its Applications, 620(??):i, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001543.

Anonymous:2021:EBz

[Ano21z] Anonymous. Editorial Board. Linear Algebra and its Applications, 620(??):ii, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001555.

Anonymous:2021:EBba

[Ano21-27] Anonymous. Editorial Board. Linear Algebra and its Applications, 621(??):i, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001725.

Anonymous:2021:EBbb

[Ano21-28] Anonymous. Editorial Board. Linear Algebra and its Applications, 621(??):ii, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001737.

Anonymous:2021:EBbc

[Ano21-29] Anonymous. Editorial Board. Linear Algebra and its Applications, 622(??):i, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001907.

Anonymous:2021:EBbd

[Ano21-30] Anonymous. Editorial Board. Linear Algebra and its Applications, 622(??):ii, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001919.

Anonymous:2021:EBbe

[Ano21-31] Anonymous. Editorial Board. Linear Algebra and its Applications, 623(??):i, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001981.

Anonymous:2021:EBbf

[Ano21-32] Anonymous. Editorial Board. Linear Algebra and its Applications, 623(??):ii, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001993.

Anonymous:2021:EBbg

[Ano21-33] Anonymous. Editorial Board. Linear Algebra and its Applications, 624(??):i, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002184.

Anonymous:2021:EBbh

[Ano21-34] Anonymous. Editorial Board. Linear Algebra and its Applications, 624(??):ii, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002196.

Anonymous:2021:EBbi

[Ano21-35] Anonymous. Editorial Board. Linear Algebra and its Applications, 625(??):i, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002305.

Anonymous:2021:EBbj

[Ano21-36] Anonymous. Editorial Board. Linear Algebra and its Applications, 625(??):ii, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002317.

Anonymous:2021:EBbk

[Ano21-37] Anonymous. Editorial Board. Linear Algebra and its Applications, 626(??):i, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002548.

Anonymous:2021:EBbl

[Ano21-38] Anonymous. Editorial Board. Linear Algebra and its Applications, 626(??):ii, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100255X.

Anonymous:2021:EBbm

[Ano21-39] Anonymous. Editorial Board. Linear Algebra and its Applications, 627(??):i, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002718.

Anonymous:2021:EBbn

[Ano21-40] Anonymous. Editorial Board. Linear Algebra and its Applications, 627(??):ii, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100272X.

Anonymous:2021:EBbo

[Ano21-41] Anonymous. Editorial Board. Linear Algebra and its Applications, 628(??):i, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003013.

Anonymous:2021:EBbp

[Ano21-42] Anonymous. Editorial Board. Linear Algebra and its Applications, 628(??):ii, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003025.

Anonymous:2021:EBbq

[Ano21-43] Anonymous. Editorial Board. Linear Algebra and its Applications, 629(??):i, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100330X.

Anonymous:2021:EBbr

[Ano21-44] Anonymous. Editorial Board. Linear Algebra and its Applications, 629(??):ii, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003311.

Anonymous:2021:EBbs

[Ano21-45] Anonymous. Editorial Board. Linear Algebra and its Applications, 630(??):i, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003529.

Anonymous:2021:EBbt

[Ano21-46] Anonymous. Editorial Board. Linear Algebra and its Applications, 630(??):ii, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003530.

Anonymous:2021:EBbu

[Ano21-47] Anonymous. Editorial Board. Linear Algebra and its Applications, 631(??):i, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100361X.

Anonymous:2021:EBbv

[Ano21-48] Anonymous. Editorial Board. Linear Algebra and its Applications, 631(??):ii, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003621.

Anonymous:2021:NEC

[Ano21-49]  Anonymous.  Note from

the Editors-in-Chief. Linear Algebra and its Applications, 617(??):iii, May 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379521000914.

Anonymous:2021:PMc

[Ano21-50] Anonymous. Pages 1–200 (1 May 2021). Linear Algebra and its Applications, 616(??): 1–200, May 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PJc

[Ano21-51] Anonymous. Pages 1–204 (1 June 2021). Linear Algebra and its Applications, 618(??): 1–204, June 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PMd

[Ano21-52] Anonymous. Pages 1–220 (15 May 2021). Linear Algebra and its Applications, 617(??): 1–220, May 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PAa

[Ano21-53] Anonymous. Pages 1–228 (15 April 2021). Linear Algebra and its Applications, 615(??): 1–228, April 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PNa

[Ano21-54] Anonymous. Pages 1–230 (1 November 2021). Linear Algebra and its Applications, 628 (??):1–230, November 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:POa

[Ano21-55] Anonymous. Pages 1–234 (1 October 2021). Linear Algebra and its Applications, 626(??): 1–234, October 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PNb

[Ano21-56] Anonymous. Pages 1–258 (15 November 2021). Linear Algebra and its Applications, 629(??):1–258, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:POb

[Ano21-57] Anonymous. Pages 1–308 (15 October 2021). Linear Algebra and its Applications, 627(??): 1–308, October 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PJd

[Ano21-58] Anonymous. Pages 1–340 (15 June 2021). Linear Algebra and its Applications, 619(??): 1–340, June 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PJf

[Ano21-59] Anonymous. Pages 1–344 (15 July 2021). Linear Algebra and its Applications, 621(??):

1–344, July 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PSb

[Ano21-60] Anonymous. Pages 1–348 (15 September 2021). Linear Algebra and its Applications, 625(??):1–348, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PAb

[Ano21-61] Anonymous. Pages 1–354 (1 August 2021). Linear Algebra and its Applications, 622(??): 1–354, August 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PDa

[Ano21-62] Anonymous. Pages 1–354 (1 December 2021). Linear Algebra and its Applications, 630 (??):1–354, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PJe

[Ano21-63] Anonymous. Pages 1–368 (1 July 2021). Linear Algebra and its Applications, 620(??): 1–368, July 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PMa

[Ano21-64] Anonymous. Pages 1–374 (1 March 2021). Linear Algebra and its Applications, 612(??): 1–374, March 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PSa

[Ano21-65] Anonymous. Pages 1–386 (1 September 2021). Linear Algebra and its Applications, 624 (??):1–386, September 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PMb

[Ano21-66] Anonymous. Pages 1–396 (15 March 2021). Linear Algebra and its Applications, 613(??): 1–396, March 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PDb

[Ano21-67] Anonymous. Pages 1–406 (15 December 2021). Linear Algebra and its Applications, 631(??):1–406, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PJa

[Ano21-68] Anonymous. Pages 1–410 (1 January 2021). Linear Algebra and its Applications, 608(??): 1–410, January 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PFb

[Ano21-69] Anonymous. Pages 1–440 (15 February 2021). Linear Algebra and its Applications, 611 (??):1–440, February 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PJb

[Ano21-70] Anonymous. Pages 1–442 (15 January 2021). Linear Algebra and its Applications, 609(??): 1–442, January 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2021:PFa

[Ano21-71] Anonymous. Pages 1–858 (1 February 2021). Linear Algebra and its Applications, 610 (??):1–858, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Linear Algebra and its Applications, 632(??):i, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003918.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Linear Algebra and its Applications, 632(??):ii, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100392X.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Linear Algebra and its Applications, 633(??):i, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100402X.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Linear Algebra and its Applications, 633(??):ii, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004031.

Anonymous:2022:EBe

[Ano22e] Anonymous. Editorial Board. Linear Algebra and its Applications, 634(??):i, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004158.

Anonymous:2022:EBf

[Ano22f] Anonymous. Editorial Board. Linear Algebra and its Applications, 634(??):ii, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100416X.

Anonymous:2022:EBg

[Ano22g] Anonymous. Editorial Board. Linear Algebra and its Applications, 635(??):i, February 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004456.

Anonymous:2022:EBh

[Ano22h] Anonymous. Editorial Board. Linear Algebra and its Applications, 635(??):ii, February 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004468.

Anonymous:2022:EBi

[Ano22i] Anonymous. Editorial Board. Linear Algebra and its Applications, 636(??):i, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004560.

Anonymous:2022:EBj

[Ano22j] Anonymous. Editorial Board. Linear Algebra and its Applications, 636(??):ii, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004572.

Anonymous:2022:EBk

[Ano22k] Anonymous. Editorial Board. Linear Algebra and its Applications, 637(??):i, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000143.

Anonymous:2022:EBl

[Ano22l] Anonymous. Editorial Board. Linear Algebra and its Applications, 637(??):ii, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000155.

Anonymous:2022:EBm

[Ano22m] Anonymous. Editorial Board. Linear Algebra and its Applications, 638(??):i, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000313.

Anonymous:2022:EBn

[Ano22n] Anonymous. Editorial Board. Linear Algebra and its Applications, 638(??):ii, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000325.

Anonymous:2022:EBo

[Ano22o] Anonymous. Editorial Board. Linear Algebra and its Applications, 639(??):i, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000404.

Anonymous:2022:EBp

[Ano22p] Anonymous. Editorial Board. Linear Algebra and its Applications, 639(??):ii, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000416.

Anonymous:2022:EBq

[Ano22q] Anonymous. Editorial Board. Linear Algebra and its Applications, 640(??):i, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000684.

Anonymous:2022:EBr

[Ano22r] Anonymous. Editorial Board. Linear Algebra and its Applications, 640(??):ii, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000696.

Anonymous:2022:EBs

[Ano22s] Anonymous. Editorial Board. Linear Algebra and its Applications, 641(??):i, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000933.

Anonymous:2022:EBt

[Ano22t] Anonymous. Editorial Board. Linear Algebra and its Applications, 641(??):ii, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000945.

Anonymous:2022:EBu

[Ano22u] Anonymous. Editorial Board. Linear Algebra and its Applications, 642(??):i, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001070.

Anonymous:2022:EBv

[Ano22v] Anonymous. Editorial Board. Linear Algebra and its Applications, 642(??):ii, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001082.

Anonymous:2022:EBw

[Ano22w] Anonymous. Editorial Board. Linear Algebra and its Applications, 643(??):i, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001252.

Anonymous:2022:EBx

[Ano22x] Anonymous. Editorial Board. Linear Algebra and its Applications, 643(??):ii, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001264.

Anonymous:2022:EBy

[Ano22y] Anonymous. Editorial Board. Linear Algebra and its Applications, 644(??):i, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001422.

Anonymous:2022:EBz

[Ano22z] Anonymous. Editorial Board. Linear Algebra and its Applications, 644(??):ii, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001434.

Anonymous:2022:EBba

[Ano22-27] Anonymous. Editorial Board. Linear Algebra and its Applications, 645(??):i, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001562.

Anonymous:2022:EBbb

[Ano22-28] Anonymous. Editorial Board. Linear Algebra and its Applications, 645(??):ii, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001574.

Anonymous:2022:EBbc

[Ano22-29] Anonymous. Editorial Board. Linear Algebra and its Applications, 646(??):i, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001732.

Anonymous:2022:EBbd

[Ano22-30] Anonymous. Editorial Board. Linear Algebra and its Applications, 646(??):ii, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001744.

Anonymous:2022:EBbe

[Ano22-31] Anonymous. Editorial Board. Linear Algebra and its Applications, 647(??):i, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001847.

Anonymous:2022:EBbf

[Ano22-32] Anonymous. Editorial Board. Linear Algebra and its Applications, 647(??):ii, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001859.

Anonymous:2022:EBbg

[Ano22-33] Anonymous. Editorial Board. Linear Algebra and its Applications, 648(??):i, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002063.

Anonymous:2022:EBbh

[Ano22-34] Anonymous. Editorial Board. Linear Algebra and its Applications, 648(??):ii, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002075.

Anonymous:2022:EBbi

[Ano22-35] Anonymous. Editorial Board. Linear Algebra and its Applications, 649(??):i, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002294.

Anonymous:2022:EBbj

[Ano22-36] Anonymous. Editorial Board. Linear Algebra and its Applications, 649(??):ii, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002300.

Anonymous:2022:EBbk

[Ano22-37] Anonymous. Editorial Board. Linear Algebra and its Applications, 650(??):i, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002555.

Anonymous:2022:EBbl

[Ano22-38] Anonymous. Editorial Board. Linear Algebra and its Applications, 650(??):ii, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002567.

Anonymous:2022:EBbm

[Ano22-39] Anonymous. Editorial Board. Linear Algebra and its Applications, 651(??):i, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002713.

Anonymous:2022:EBbn

[Ano22-40] Anonymous. Editorial Board. Linear Algebra and its Applications, 651(??):ii, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002725.

Anonymous:2022:EBbo

[Ano22-41] Anonymous. Editorial Board. Linear Algebra and its Applications, 652(??):i, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002993.

Anonymous:2022:EBbp

[Ano22-42] Anonymous. Editorial Board. Linear Algebra and its Applications, 652(??):ii, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003007.

Anonymous:2022:EBbq

[Ano22-43] Anonymous. Editorial Board. Linear Algebra and its Applications, 653(??):i, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003330.

Anonymous:2022:EBbr

[Ano22-44] Anonymous. Editorial Board. Linear Algebra and its Applications, 653(??):ii, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003342.

Anonymous:2022:EBbs

[Ano22-45] Anonymous. Editorial Board. Linear Algebra and its Applications, 654(??):i, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003615.

Anonymous:2022:EBbt

[Ano22-46] Anonymous. Editorial Board. Linear Algebra and its Applications, 654(??):ii, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003627.

Anonymous:2022:EBbu

[Ano22-47] Anonymous. Editorial Board. Linear Algebra and its Applications, 655(??):i, December 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003767.

Anonymous:2022:EBbv

[Ano22-48] Anonymous. Editorial Board. Linear Algebra and its Applications, 655(??):ii, December 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003779.

Anonymous:2022:PAd

[Ano22-49] Anonymous. Pages 1–152 (15 August 2022). Linear Algebra and its Applications, 647(??): 1–152, August 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PMa

[Ano22-50] Anonymous. Pages 1–154 (1 March 2022). Linear Algebra and its Applications, 636(??): 1–154, March 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PFb

[Ano22-51] Anonymous. Pages 1–202 (15 February 2022). Linear Algebra and its Applications, 635 (??):1–202, February 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PMc

[Ano22-52] Anonymous. Pages 1–208 (1 May 2022). Linear Algebra and its Applications, 640(??): 1–208, May 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PMd

[Ano22-53] Anonymous. Pages 1–208 (15 May 2022). Linear Algebra and its Applications, 641(??): 1–208, May 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PFa

[Ano22-54] Anonymous. Pages 1–210 (1 February 2022). Linear Algebra and its Applications, 634 (??):1–210, February 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PAa

[Ano22-55] Anonymous. Pages 1–218 (1 April 2022). Linear Algebra and its Applications, 638(??):

1–218, April 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PMb

[Ano22-56] Anonymous. Pages 1–240 (15 March 2022). Linear Algebra and its Applications, 637(??): 1–240, March 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PAc

[Ano22-57] Anonymous. Pages 1–252 (1 August 2022). Linear Algebra and its Applications, 646(??): 1–252, August 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PNa

[Ano22-58] Anonymous. Pages 1–252 (1 November 2022). Linear Algebra and its Applications, 652 (??):1–252, November 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PJe

[Ano22-59] Anonymous. Pages 1–270 (1 July 2022). Linear Algebra and its Applications, 644(??): 1–270, July 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PSa

[Ano22-60] Anonymous. Pages 1–270 (1 September 2022). Linear Algebra and its Applications, 648 (??):1–270, September 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PJd

[Ano22-61] Anonymous. Pages 1–280 (15 June 2022). Linear Algebra and its Applications, 643(??): 1–280, June 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PAb

[Ano22-62] Anonymous. Pages 1–282 (15 April 2022). Linear Algebra and its Applications, 639(??): 1–282, April 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:POa

[Ano22-63] Anonymous. Pages 1–298 (1 October 2022). Linear Algebra and its Applications, 650(??): 1–298, October 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PJf

[Ano22-64] Anonymous. Pages 1–306 (15 July 2022). Linear Algebra and its Applications, 645(??): 1–306, July 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PJc

[Ano22-65] Anonymous. Pages 1–310 (1 June 2022). Linear Algebra and its Applications, 642(??): 1–310, June 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PDb

[Ano22-66] Anonymous. Pages 1–314 (15 December 2022). Linear Algebra and its Applications, 655(??):1–314, December 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PJa

[Ano22-67] Anonymous. Pages 1–334 (1 January 2022). Linear Algebra and its Applications, 632(??): 1–334, January 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PJb

[Ano22-68] Anonymous. Pages 1–366 (15 January 2022). Linear Algebra and its Applications, 633(??): 1–366, January 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PDa

[Ano22-69] Anonymous. Pages 1–390 (1 December 2022). Linear Algebra and its Applications, 654 (??):1–390, December 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:POb

[Ano22-70] Anonymous. Pages 1–426 (15 October 2022). Linear Algebra and its Applications, 651(??): 1–426, October 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PSb

[Ano22-71] Anonymous. Pages 1–490 (15 September 2022). Linear Algebra and its Applications, 649(??):1–490, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2022:PNb

[Ano22-72] Anonymous. Pages 1–500 (15 November 2022). Linear Algebra and its Applications, 653(??):1–500, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial Board. Linear Algebra and its Applications, 656(??):i, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003974.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial Board. Linear Algebra and its Applications, 656(??):ii, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003986.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial Board. Linear Algebra and its Applications, 657(??):i, January 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004116.

Anonymous:2023:EBd

[Ano23d] Anonymous. Editorial Board. Linear Algebra and its Applications, 657(??):ii, January 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004128.

Anonymous:2023:EBe

[Ano23e] Anonymous. Editorial Board. Linear Algebra and its Applications, 658(??):i, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004244.

Anonymous:2023:EBf

[Ano23f] Anonymous. Editorial Board. Linear Algebra and its Applications, 658(??):ii, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004256.

Anonymous:2023:EBg

[Ano23g] Anonymous. Editorial Board. Linear Algebra and its Applications, 659(??):i, February 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004463.

Anonymous:2023:EBh

[Ano23h] Anonymous. Editorial Board. Linear Algebra and its Applications, 659(??):ii, February 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004475.

Anonymous:2023:EBi

[Ano23i] Anonymous. Editorial Board. Linear Algebra and its Applications, 660(??):i, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200458X.

Anonymous:2023:EBj

[Ano23j] Anonymous. Editorial Board. Linear Algebra and its Applications, 660(??):ii, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004591.

Anonymous:2023:EBk

[Ano23k] Anonymous. Editorial Board. Linear Algebra and its Applications, 661(??):i, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000149.

Anonymous:2023:EBl

[Ano23l] Anonymous. Editorial Board. Linear Algebra and its Applications, 661(??):ii, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000150.

Anonymous:2023:EBm

[Ano23m] Anonymous. Editorial Board. Linear Algebra and its Applications, 662(??):i, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000381.

Anonymous:2023:EBn

[Ano23n] Anonymous. Editorial Board. Linear Algebra and its Applications, 662(??):ii, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000393.

Anonymous:2023:EBo

[Ano23o] Anonymous. Editorial Board. Linear Algebra and its Applications, 663(??):i, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000514.

Anonymous:2023:EBp

[Ano23p] Anonymous. Editorial Board. Linear Algebra and its Applications, 663(??):ii, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000526.

Anonymous:2023:EBq

[Ano23q] Anonymous. Editorial Board. Linear Algebra and its Applications, 664(??):i, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300068X.

Anonymous:2023:EBr

[Ano23r] Anonymous. Editorial Board. Linear Algebra and its Applications, 664(??):ii, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000691.

Anonymous:2023:EBs

[Ano23s] Anonymous. Editorial Board. Linear Algebra and its Applications, 665(??):i, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000873.

Anonymous:2023:EBt

[Ano23t] Anonymous. Editorial Board. Linear Algebra and its Applications, 665(??):ii, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000885.

Anonymous:2023:EBu

[Ano23u] Anonymous. Editorial Board. Linear Algebra and its Applications, 666(??):i, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001106.

Anonymous:2023:EBv

[Ano23v] Anonymous. Editorial Board. Linear Algebra and its Applications, 666(??):ii, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001118.

Anonymous:2023:EBw

[Ano23w] Anonymous. Editorial Board. Linear Algebra and its Applications, 667(??):i, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001222.

Anonymous:2023:EBx

[Ano23x] Anonymous. Editorial Board. Linear Algebra and its Applications, 667(??):ii, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001234.

Anonymous:2023:EBy

[Ano23y] Anonymous. Editorial Board. Linear Algebra and its Applications, 668(??):i, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001374.

Anonymous:2023:EBz

[Ano23z] Anonymous. Editorial Board. Linear Algebra and its Applications, 668(??):ii, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001386.

Anonymous:2023:EBba

[Ano23-27] Anonymous. Editorial Board. Linear Algebra and its Applications, 669(??):i, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001593.

Anonymous:2023:EBbb

[Ano23-28] Anonymous. Editorial Board. Linear Algebra and its Applications, 669(??):ii, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300160X.

Anonymous:2023:EBbc

[Ano23-29] Anonymous. Editorial Board. Linear Algebra and its Applications, 670(??):i, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001751.

Anonymous:2023:EBbd

[Ano23-30] Anonymous. Editorial Board. Linear Algebra and its Applications, 670(??):ii, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001763.

Anonymous:2023:EBbe

[Ano23-31] Anonymous. Editorial Board. Linear Algebra and its Applications, 671(??):i, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001921.

Anonymous:2023:EBbf

[Ano23-32] Anonymous. Editorial Board. Linear Algebra and its Applications, 671(??):ii, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001933.

Anonymous:2023:EBbg

[Ano23-33] Anonymous. Editorial Board. Linear Algebra and its Applications, 672(??):i, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002227.

Anonymous:2023:EBbh

[Ano23-34] Anonymous. Editorial Board. Linear Algebra and its Applications, 672(??):ii, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002239.

Anonymous:2023:EBbi

[Ano23-35] Anonymous. Editorial Board. Linear Algebra and its Applications, 673(??):i, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002306.

Anonymous:2023:EBbj

[Ano23-36] Anonymous. Editorial Board. Linear Algebra and its Applications, 673(??):ii, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002318.

Anonymous:2023:EBbk

[Ano23-37] Anonymous. Editorial Board. Linear Algebra and its Applications, 674(??):i, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002562.

Anonymous:2023:EBbl

[Ano23-38] Anonymous. Editorial Board. Linear Algebra and its Applications, 674(??):ii, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002574.

Anonymous:2023:EBbm

[Ano23-39] Anonymous. Editorial Board. Linear Algebra and its Applications, 675(??):i, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002860.

Anonymous:2023:EBbn

[Ano23-40] Anonymous. Editorial Board. Linear Algebra and its Applications, 675(??):ii, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002872.

Anonymous:2023:EBbo

[Ano23-41] Anonymous. Editorial Board. Linear Algebra and its Applications, 676(??):i, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003105.

Anonymous:2023:EBbp

[Ano23-42] Anonymous. Editorial Board. Linear Algebra and its Applications, 676(??):ii, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003117.

Anonymous:2023:EBbq

[Ano23-43] Anonymous. Editorial Board. Linear Algebra and its Applications, 677(??):i, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003427.

Anonymous:2023:EBbr

[Ano23-44] Anonymous. Editorial Board. Linear Algebra and its Applications, 677(??):ii, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003439.

Anonymous:2023:EBbs

[Ano23-45] Anonymous. Editorial Board. Linear Algebra and its Applications, 678(??):i, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003580.

Anonymous:2023:EBbt

[Ano23-46] Anonymous. Editorial Board. Linear Algebra and its Applications, 678(??):ii, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003592.

Anonymous:2023:EBbu

[Ano23-47] Anonymous. Editorial Board. Linear Algebra and its Applications, 679(??):i, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003920.

Anonymous:2023:EBbv

[Ano23-48] Anonymous. Editorial Board. Linear Algebra and its Applications, 679(??):ii, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003932.

Anonymous:2023:PMa

[Ano23-49] Anonymous. Pages 1–118 (1 March 2023). Linear Algebra and its Applications, 660(??): 1–118, March 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PAd

[Ano23-50] Anonymous. Pages 1–134 (15 August 2023). Linear Algebra and its Applications, 671(??): 1–134, August 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PJc

[Ano23-51] Anonymous. Pages 1–144 (1 June 2023). Linear Algebra and its Applications, 666(??): 1–144, June 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PAa

[Ano23-52] Anonymous. Pages 1–162 (1 April 2023). Linear Algebra and its Applications, 662(??): 1–162, April 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PAc

[Ano23-53] Anonymous. Pages 1–168 (1 August 2023). Linear Algebra and its Applications, 670(??): 1–168, August 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PJe

[Ano23-54] Anonymous. Pages 1–172 (1 July 2023). Linear Algebra and its Applications, 668(??): 1–172, July 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PJf

[Ano23-55] Anonymous. Pages 1–174 (15 July 2023). Linear Algebra and its Applications, 669(??): 1–174, July 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PJd

[Ano23-56] Anonymous. Pages 1–192 (15 June 2023). Linear Algebra and its Applications, 667(??): 1–192, June 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PAb

[Ano23-57] Anonymous. Pages 1–240 (15 April 2023). Linear Algebra and its Applications, 663(??): 1–240, April 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PSa

[Ano23-58] Anonymous. Pages 1–270 (1 September 2023). Linear Algebra and its Applications, 672 (??):1–270, September 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PMb

[Ano23-59] Anonymous. Pages 1–270 (15 March 2023). Linear Algebra and its Applications, 661(??): 1–270, March 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PJb

[Ano23-60] Anonymous. Pages 1–286 (15 January 2023). Linear Algebra and its Applications, 657(??):

1–286, January 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PDa

[Ano23-61] Anonymous. Pages 1–316 (1 December 2023). Linear Algebra and its Applications, 678 (??):1–316, December 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PDb

[Ano23-62] Anonymous. Pages 1–326 (15 December 2023). Linear Algebra and its Applications, 679(??):1–326, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PFa

[Ano23-63] Anonymous. Pages 1–350 (1 February 2023). Linear Algebra and its Applications, 658 (??):1–350, February 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PNb

[Ano23-64] Anonymous. Pages 1–352 (15 November 2023). Linear Algebra and its Applications, 677(??):1–352, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PMc

[Ano23-65] Anonymous. Pages 1–368 (1 May 2023). Linear Algebra and its Applications, 664(??): 1–368, May 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PMd

[Ano23-66] Anonymous. Pages 1–420 (15 May 2023). Linear Algebra and its Applications, 665(??): 1–420, May 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PSb

[Ano23-67] Anonymous. Pages 1–420 (15 September 2023). Linear Algebra and its Applications, 673(??):1–420, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PNa

[Ano23-68] Anonymous. Pages 1–458 (1 November 2023). Linear Algebra and its Applications, 676 (??):1–458, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:POa

[Ano23-69] Anonymous. Pages 1–518 (1 October 2023). Linear Algebra and its Applications, 674(??): 1–518, October 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PJa

[Ano23-70] Anonymous. Pages 1–558 (1 January 2023). Linear Algebra and its Applications, 656(??): 1–558, January 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:PFb

[Ano23-71] Anonymous. Pages 1–72 (15 February 2023). Linear Algebra and its Applications, 659 (??):1–72, February 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2023:POb

[Ano23-72] Anonymous. Pages 1-350 (15 October 2023). Linear Algebra and its Applications, 675(??): 1–350, October 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial Board. Linear Algebra and its Applications, 680(??):i, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004111.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial Board. Linear Algebra and its Applications, 680(??):ii, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004123.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial Board. Linear Algebra and its Applications, 681(??):i, January 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004330.

Anonymous:2024:EBd

[Ano24d] Anonymous. Editorial Board. Linear Algebra and its Applications, 681(??):ii, January 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004342.

Anonymous:2024:EBe

[Ano24e] Anonymous. Editorial Board. Linear Algebra and its Applications, 682(??):i, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004500.

Anonymous:2024:EBf

[Ano24f] Anonymous. Editorial Board. Linear Algebra and its Applications, 682(??):ii, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004512.

Anonymous:2024:EBg

[Ano24g] Anonymous. Editorial Board. Linear Algebra and its Applications, 683(??):i, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004706.

Anonymous:2024:EBh

[Ano24h] Anonymous. Editorial Board. Linear Algebra and its Applications, 683(??):ii, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004718.

Anonymous:2024:EBi

[Ano24i] Anonymous. Editorial Board. Linear Algebra and its Applications, 684(??):i, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000089.

Anonymous:2024:EBj

[Ano24j] Anonymous. Editorial Board. Linear Algebra and its Applications, 684(??):ii, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000090.

Anonymous:2024:EBk

[Ano24k] Anonymous. Editorial Board. Linear Algebra and its Applications, 685(??):i, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000247.

Anonymous:2024:EBl

[Ano24l] Anonymous. Editorial Board. Linear Algebra and its Applications, 685(??):ii, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000259.

Anonymous:2024:EBm

[Ano24m] Anonymous. Editorial Board. Linear Algebra and its Applications, 686(??):i, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952400048X.

Anonymous:2024:EBn

[Ano24n] Anonymous. Editorial Board. Linear Algebra and its Applications, 686(??):ii, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000491.

Anonymous:2024:EBo

[Ano24o] Anonymous. Editorial Board. Linear Algebra and its Applications, 687(??):i, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000697.

Anonymous:2024:EBp

[Ano24p] Anonymous. Editorial Board. Linear Algebra and its Applications, 687(??):ii, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000703.

Anonymous:2024:EBq

[Ano24q] Anonymous. Editorial Board. Linear Algebra and its Applications, 688(??):i, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001071.

Anonymous:2024:EBr

[Ano24r] Anonymous. Editorial Board. Linear Algebra and its Applications, 688(??):ii, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001083.

Anonymous:2024:EBs

[Ano24s] Anonymous. Editorial Board. Linear Algebra and its Applications, 689(??):i, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001204.

Anonymous:2024:EBt

[Ano24t] Anonymous. Editorial Board. Linear Algebra and its Applications, 689(??):ii, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001216.

Anonymous:2024:EBu

[Ano24u] Anonymous. Editorial Board. Linear Algebra and its Applications, 690(??):i, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001423.

Anonymous:2024:EBv

[Ano24v] Anonymous. Editorial Board. Linear Algebra and its Applications, 690(??):ii, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001435.

Anonymous:2024:EBw

[Ano24w] Anonymous. Editorial Board. Linear Algebra and its Applications, 691(??):i, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001599.

Anonymous:2024:EBx

[Ano24x] Anonymous. Editorial Board. Linear Algebra and its Applications, 691(??):ii, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001605.

Anonymous:2024:EBy

[Ano24y] Anonymous. Editorial Board. Linear Algebra and its Applications, 692(??):i, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001782.

Anonymous:2024:EBz

[Ano24z] Anonymous. Editorial Board. Linear Algebra and its Applications, 692(??):ii, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001794.

Anonymous:2024:PAa

[Ano24-27] Anonymous. Pages 1–134 (1 April 2024). Linear Algebra and its Applications, 686(??): 1–134, April 1, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PMa

[Ano24-28] Anonymous. Pages 1–160 (1 March 2024). Linear Algebra and its Applications, 684(??): 1–160, March 1, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PJb

[Ano24-29] Anonymous. Pages 1–186 (15 January 2024). Linear Algebra and its Applications, 681(??): 1–186, January 15, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PJc

[Ano24-30] Anonymous. Pages 1–198 (1 June 2024). Linear Algebra and its Applications, 690(??): 1–198, June 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PJd

[Ano24-31] Anonymous. Pages 1–232 (15 June 2024). Linear Algebra and its Applications, 691(??): 1–232, June 15, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PMb

[Ano24-32] Anonymous. Pages 1–246 (15 March 2024). Linear Algebra and its Applications, 685(??): 1–246, March 15, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PMc

[Ano24-33] Anonymous. Pages 1–256 (1 May 2024). Linear Algebra and its Applications, 688(??): 1–256, May 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PFb

[Ano24-34] Anonymous. Pages 1–264 (15 February 2024). Linear Algebra and its Applications, 683 (??):1–264, February 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PAb

[Ano24-35] Anonymous. Pages 1–280 (15 April 2024). Linear Algebra and its Applications, 687(??): 1–280, April 15, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PJe

[Ano24-36] Anonymous. Pages 1–292 (1 July 2024). Linear Algebra and its Applications, 692(??): 1–292, July 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PMd

[Ano24-37] Anonymous. Pages 1–294 (15 May 2024). Linear Algebra and its Applications, 689(??): 1–294, May 15, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PFa

[Ano24-38] Anonymous. Pages 1–362 (1 February 2024). Linear Algebra and its Applications, 682 (??):1–362, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PJa

[Ano24-39] Anonymous. Pages 1–390 (1 January 2024). Linear Algebra and its Applications, 680(??): 1–390, January 1, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2024:PNa

[Ano24-40] Anonymous. Publisher’s Note. Linear Algebra and its Applications, 688(??):1–20, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000454.

Anonymous:2024:PNb

[Ano24-41] Anonymous. Publisher’s note. Linear Algebra and its Applications, 688(??):59–77, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000442.

Anstreicher:2021:TCM

[Ans21] Kurt M. Anstreicher. Testing copositivity via mixed– integer linear programming. Linear Algebra and its Applications, 609(??):218–230, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304171.

Aranda-Orna:2022:FCG

[AO22] Diego Aranda-Orna. On the Faulkner construction for generalized Jordan superpairs. Linear Algebra and its Applications, 646(??):1–28, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200115X.

Adam:2023:SLU

[AOA23] Maria Adam, Iro Oikonomou, and Aikaterini Aretaki. Sequences of lower and upper bounds for the spectral radius of a nonnegative matrix. Linear Algebra and its Applications, 667(??):165–191, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000836.

Aranda-Orna:2020:GTP

[AOCM20] Diego Aranda-Orna and Alejandra S. Córdova-Martínez. Gradings on tensor products of composition algebras and on the Smirnov algebra. Linear Algebra and its Applications, 584(??):1–36, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519303726.

Aranda-Orna:2021:FGK

[AOCM21] Diego Aranda-Orna and Alejandra S. Córdova-Martínez. Fine gradings on Kantor systems of Hurwitz type. Linear Algebra and its Applications, 613(??):201–240, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030570X.

Aznay:2023:GKD

[AOZ23] Zakariae Aznay, Abdelmalek Ouahab, and Hassan Zariouh. Generalization of Kato’s decomposition. Linear Algebra and its Applications, 672(??):108–121, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001477.

Agapito:2020:OMS

[AP20] Theeex D. Agapito and Agnes T. Paras. ψ_{S}- orthogonal matrices and ψ_{S}- symmetries. Linear Algebra and its Applications, 584(??):185–196, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304008.

Ashbrock:2023:DDF

[AP23] Jonathan Ashbrock and Alexander M. Powell. Dynamical dual frames with an application to quantization. Linear Algebra and its Applications, 658(??):151–185, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003469.

Arela-Perez:2023:NIE

[APLN^{+}23] S. Arela-Pérez, Charlie Lozano, Hans Nina, H. Pickmann-Soto, and Jonnathan Rodríguez. The new inverse eigenvalue problems for periodic and generalized periodic Jacobi matrices from their extremal spectral data. Linear Algebra and its Applications, 659(??):55–72, February 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004207.

Arela-Perez:2021:CLD

[APNP^{+}21] S. Arela-Pérez, Hans Nina, Jésica Pantáz, H. Pickmann-Soto, and Elvis Valero. Construction of Lefkovitch and doubly Lefkovitch matrices with maximal eigenvalues and some diagonal elements prescribed. Linear Algebra and its Applications, 626(??):152–170, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001944.

Arora:2021:EPL

[AR21] Akansha Arora and Samrith Ram. Enumerating partial linear transformations in a similarity class. Linear Algebra and its Applications, 625(??):196–211, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002093.

Arlinskii:2020:SPA

[Arl20] Yu. M. Arlinskiĭ. Shorting, parallel addition and form sums of nonnegative selfadjoint linear relations. Linear Algebra and its Applications, 599(??):156–200, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301890.

Anuelic:2024:SSC

[ARS24] Milica An– delić, Peter Rowlinson, and Zoran Stanić. Strong star complements in graphs. Linear Algebra and its Applications, 688(??):179–194, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000594.

Akman:2020:SPC

[AS20a] Fusun Akman and Papa A. Sissokho. Subspace partitions of F_{q}^{n} containing direct sums II: General case. Linear Algebra and its Applications, 598(??):1–28, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301695.

Alakhrass:2020:LFS

[AS20b] Mohammad Alakhrass and Mohammad Sababheh. Lieb functions and sectorial matrices. Linear Algebra and its Applications, 586(??):308– 324, February 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379519304689.

Azimi:2022:SDM

[AS22] Ali Azimi and Sivaramakrishnan Sivasubramanian. The 2-Steiner distance matrix of a tree. Linear Algebra and its Applications, 655 (??):65–86, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003251.

Ayyer:2023:CPM

[AS23] Arvind Ayyer and Naren Sundaravaradan. Combinatorial proofs of multivariate Cayley– Hamilton theorems. Linear Algebra and its Applications, 661(??):247–269, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004414.

Arizmendi:2024:CVE

[AS24] Octavio Arizmendi and Saylé Sigarreta. The change of vertex energy when joining trees. Linear Algebra and its Applications, 687(??):117–131, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000284.

Arezoomand:2022:PST

[ASG22] Majid Arezoomand, Farideh Shafiei, and Modjtaba Ghorbani. Perfect state transfer on Cayley graphs over the dicyclic group. Linear Algebra and its Applications, 639(??):116–134, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004602.

Afshari:2021:LBL

[ASS21] B. Afshari, M. T. Saadati, and R. Saadati. Lower bounds for the Laplacian spectral radius of graphs. Linear Algebra and its Applications, 631(??):136–142, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003232.

Alfaro:2020:DIG

[AT20] Carlos A. Alfaro and Libby Taylor. Distance ideals of graphs. Linear Algebra and its Applications, 584(??):127–144, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519303957.

Abyzov:2021:WEM

[AT21] A. N. Abyzov and D. T. Tapkin. When is every matrix over a ring the sum of two tripotents? Linear Algebra and its Applications, 630(??):316–325, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003402.

Adachi:2024:QHR

[AT24] Takahide Adachi and Mayu Tsukamoto. Quasi-hereditary rings are closed under taking block extensions. Linear Algebra and its Applications, 692(??):24–36, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001307.

Alpay:2023:SAD

[AV23] Daniel Alpay and Dan Volok. Schur analysis and discrete analytic functions: Rational functions and co-isometric realizations. Linear Algebra and its Applications, 675 (??):179–202, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002458.

Akbari:2020:TLL

[AvDF20] S. Akbari, E. R. van Dam, and M. H. Fakharan. Trees with a large Laplacian eigenvalue multiplicity. Linear Algebra and its Applications, 586(??):262–273, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304458.

Alvarez-Vizoso:2020:LED

[ÁVKP20a] Javier Álvarez-Vizoso, Michael Kirby, and Chris Peterson. Local eigenvalue decomposition for embedded Riemannian manifolds. Linear Algebra and its Applications, 604(??):21–51, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302974.

Alvarez-Vizoso:2020:MCL

[ÁVKP20b] Javier Álvarez-Vizoso, Michael Kirby, and Chris Peterson. Manifold curvature learning from hypersurface integral invariants. Linear Algebra and its Applications, 602(??):179–205, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302585.

Aliniaeifard:2024:DCU

[AWvW24] Farid Aliniaeifard, Victor Wang, and Stephanie van Willigenburg. Deletioncontraction for a unified Laplacian and applications. Linear Algebra and its Applications, 691(??):50–95, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000958.

Amanov:2023:TSR

[AY23] Alimzhan Amanov and Damir Yeliussizov. Tensor slice rank and Cayley’s first hyperdeterminant. Linear Algebra and its Applications, 656(??):224–246, January 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003524.

Aydin:2023:SSH

[Ayd23]  Cengiz Aydin.  Symplec-

tic splitting of Hamiltonian structures and reduced monodromy matrices. Linear Algebra and its Applications, 674(??):426–441, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379523002264.

Bai:2024:WMM

[Bai24] Zhong-Zhi Bai. The Wasserstein metric matrix and its computational property. Linear Algebra and its Applications, 681(??):150–186, January 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004020.

Banerjee:2021:SH

[Ban21] Anirban Banerjee. On the spectrum of hypergraphs. Linear Algebra and its Applications, 614(??):82–110, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300161.

Banakh:2022:EDB

[Ban22]  Taras Banakh.  Every 2-

dimensional Banach space has the Mazur–Ulam property. Linear Algebra and its Applications, 632(??):268–280, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003657.

Baz:2023:CMG

[BAPF23] Juan Baz, Pedro Alonso, and Raúl Pérez-Fernández. Correlation matrices of Gaussian Markov random fields over cycle graphs. Linear Algebra and its Applications, 663(??):32–61, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004682.

Baragana:2020:NDE

[Bar20] Itziar Baragaña. The number of distinct eigenvalues of a regular pencil and of a square matrix after rank perturbation. Linear Algebra and its Applications, 588(??):101–121, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304987.

Bar:2021:FLA

[Bär21a] Christian Bär. The Faddeev-LeVerrier algorithm and the Pfaffian. Linear Algebra and its Applications, 630(??):39–55, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002883.

Barvinok:2021:RAP

[Bar21b] Alexander Barvinok. A remark on approximating permanents of positive definite matrices. Linear Algebra and its Applications, 608(??):399–406, January 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304547.

Bari:2022:ST

[Bar22] Kashif K. Bari. On the structure tensor of sl_{n}. Linear Algebra and its Applications, 653(??):266–286, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002890.

Basso:2021:FTF

[Bas21] Giuliano Basso. False-twin-free graphs with a fixed number of negative eigenvalues. Linear Algebra and its Applications, 618(??):144–149, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000562.

Bayat:2024:PFM

[Bay24] Morteza Bayat. The Pascal functional matrices with multidimensional binomial coefficients. Linear Algebra and its Applications, 688(??):195–214, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S002437952400065X.

Bajuk:2022:TSZ

[BB22] Žan Bajuk and Matej Brešar. Two-sided zero product determined algebras. Linear Algebra and its Applications, 643(??):125–136, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000659.

Belardo:2022:SQU

[BBC^{+}22a] Francesco Belardo, Maurizio Brunetti, Nolan J. Coble, Nathan Reff, and Howard Skogman. Spectra of quaternion unit gain graphs. Linear Algebra and its Applications, 632(??):15–49, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003426.

Breen:2022:MND

[BBC^{+}22b] J. Breen, C. Brouwer, M. Catral, M. Cavers, P. van den Driessche, and K. N. Vander Meulen. Minimum number of distinct eigenvalues allowed by a sign pattern. Linear Algebra and its Applications, 654(??):311–338, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003202.

Belardo:2021:GMS

[BBCD21] Francesco Belardo, Maurizio Brunetti, Matteo Cavaleri, and Alfredo Donno. Godsil– McKay switching for mixed and gain graphs over the circle group. Linear Algebra and its Applications, 614(??):256–269, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302196.

Bialas:2024:HSN

[BBCK24] Stanisław Białas, Leokadia Białas-Cież, and Michał Kudra. On the Hurwitz stability of noninteger Hadamard powers of stable polynomials. Linear Algebra and its Applications, 683(??):111–124, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004421.

Belkouche:2022:UT

[BBCL22] Wiam Belkouche, Abderrahim Boussaïri, Abdelhak Chaïchaâ, and Soufiane Lakhlifi. On unimodular tournaments. Linear Algebra and its Applications, 632(??):50–60, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003475.

Balaji:2021:IFD

[BBG21] R. Balaji, R. B. Bapat, and Shivani Goel. An inverse formula for the distance matrix of a wheel graph with an even number of vertices. Linear Algebra and its Applications, 610(??):274–292, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304742.

Barreiro:2022:OQL

[BBNS22] Elisabete Barreiro, Saïd Benayadi, Rosa M. Navarro, and José M. Sánchez. Oddquadratic Lie superalgebras with a weak filiform module as an odd part. Linear Algebra and its Applications, 649(??):22–46, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001641.

Bernstein:2020:TGR

[BBS20] Daniel Irving Bernstein, Grigoriy Blekherman, and Rainer Sinn. Typical and generic ranks in matrix completion. Linear Algebra and its Applications, 585(??):71–104, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519303763.

Balachandran:2023:EHR

[BBS23] Niranjan Balachandran, Srimanta Bhattacharya, and Brahadeesh Sankarnarayanan. An ensemble of high rank matrices arising from tournaments. Linear Algebra and its Applications, 658(??):310–318, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004049.

Balls-Barker:2020:LPN

[BBW20] Bryn Balls-Barker and Benjamin Webb. Link prediction in networks using effective transitions. Linear Algebra and its Applications, 599(??):79–104, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301889.

Belabbas:2021:TLG

[BC21] Mohamed-Ali Belabbas and Xudong Chen. Triangulated Laman graphs, local stochastic matrices, and limits of their products. Linear Algebra and its Applications, 619(??):176–209, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000756.

Barbarino:2022:CSP

[BC22] Giovanni Barbarino and Antonio Cicone. Conjectures on spectral properties of ALIF algorithm. Linear Algebra and its Applications, 647(??):127–152, August 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200163X.

Bender:2023:PI

[BC23] Chase Bender and Debraj Chakrabarti. Polarization identities. Linear Algebra and its Applications, 665(??):211–252, May 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000472.

Byrne:2024:CPB

[BC24] Eimear Byrne and Giuseppe Cotardo. Constructions of perfect bases for classes of 3-tensors. Linear Algebra and its Applications, 683(??):1–30, February 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379523004391.

Benhida:2021:CPD

[BCE21] Chafiq Benhida, Raúl E. Curto, and George R. Exner. Conditional positive definiteness as a bridge between k-hyponormality and n-contractivity. Linear Algebra and its Applications, 625(??):146–170, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002068.

Barrett:2023:LSD

[BCE^{+}23] Wayne Barrett, Thomas R. Cameron, Emily Evans, H. Tracy Hall, and Mark Kempton. On the Laplacian spread of digraphs. Linear Algebra and its Applications, 664(??):126–146, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000228.

Brown:2020:ETD

[BCF^{+}20] David E. Brown, Eric Culver, Bryce Frederickson, Sidney Tate, and Brent J. Thomas. Entropy of tournament digraphs. Linear Algebra and its Applications, 584(??):371–393, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930415X.

Bravo:2021:CSD

[BCFT21] Diego Bravo, Florencia Cubría, Marcelo Fiori, and Vilmar Trevisan. Complementarity spectrum of digraphs. Linear Algebra and its Applications, 627(??):24–40, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002354.

Bera:2024:SDC

[BCG24] Santu Bera, Sameer Chavan, and Soumitra Ghara. A spectral dichotomy for commuting m-isometries with negative core operator. Linear Algebra and its Applications, 688(??):244–255, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000624.

Bartoli:2021:CAM

[BCM21] Daniele Bartoli, Bence Csajbók, and Maria Montanucci. On a conjecture about maximum scattered subspaces of F_{q^{6}} × F_{q^{6}}. Linear Algebra and its Applications, 631 (??):111–135, December 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003207.

Barrera:2022:DPG

[BCNS22] Waldemar Barrera, Angel Cano, Juan Pablo Navarrete, and José Seade. Discrete parabolic groups in PSL(3, C). Linear Algebra and its Applications, 653 (??):430–500, November 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003111.

Berliner:2022:SSP

[BCOvdD22] Adam H. Berliner, Minerva Catral, D. D. Olesky, and P. van den Driessche. Stability of sign patterns from a system of second order ODEs. Linear Algebra and its Applications, 632(??):61–78, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100327X.

Berliner:2023:IMS

[BCOvdD23] Adam H. Berliner, Minerva Catral, D. D. Olesky, and P. van den Driessche. Inertias of matrices and sign patterns related to a system of second order ODEs. Linear Algebra and its Applications, 661(??):116–136, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004372.

Borodzik:2022:TBP

[BCP22] Maciej Borodzik, Anthony Conway, and Wojciech Politarczyk. Twisted Blanchfield pairings and twisted signatures I: Algebraic background. Linear Algebra and its Applications, 655 (??):236–290, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003305.

Bracciali:2024:CRM

[BCP24] Cleonice F. Bracciali, Glalco S. Costa, and Teresa E. Pérez. Centrosymmetric and reverse matrices in bivariate orthogonal polynomials. Linear Algebra and its Applications, 692(??):212–240, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001332.

Byrne:2020:RMC

[BCR20] Eimear Byrne, Giuseppe Cotardo, and Alberto Ravagnani. Rank-metric codes, generalized binomial moments and their zeta functions. Linear Algebra and its Applications, 604(??):92–128, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302937.

Bachelier:2023:NSC

[BCR^{+}23] Olivier Bachelier, Thomas Cluzeau, Alexandre Rigaud, Francisco José Silva Àlvarez, and Nima Yeganefar. A necessary and sufficient condition of asymptotic stability for a class of Fornasini– Marchesini models. Linear Algebra and its Applications, 658(??):206–232, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003913.

Bessades:2024:UAG

[BCS24] D. C. L. Bessades, W. D. S. Costa, and M. L. O. Santos. On unitary algebras with graded involution of quadratic growth. Linear Algebra and its Applications, 689(??):260–293, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000831.

Ballini:2021:VDS

[BD21a] Francesco Ballini and Nikita Deniskin. Vertex distinction with subgraph centrality: a proof of Estrada’s conjecture and some generalizations. Linear Algebra and its Applications, 612(??):334–344, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305310.

Brualdi:2021:SRM

[BD21b] Richard A. Brualdi and Geir Dahl. Sign-restricted matrices of 0’s, 1’s, and −1’s. Linear Algebra and its Applications, 615(??):77–103, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100001X.

Bolten:2023:SBC

[BDFF23] Matthias Bolten, Marco Donatelli, Paola Ferrari, and Isabella Furci. Symbol based convergence analysis in multigrid methods for saddle point problems. Linear Algebra and its Applications, 671(??):67–108, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001507.

Batenkov:2021:SPV

[BDGY21] Dmitry Batenkov, Benedikt Diederichs, Gil Goldman, and Yosef Yomdin. The spectral properties of Vandermonde matrices with clustered nodes. Linear Algebra and its Applications, 609(??):37–72, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304134.

Bien:2022:DMD

[BDHS22] Mai Hoang Bien, Truong Huu Dung, Nguyen Thi Thai Ha, and Tran Nam Son. Decompositions of matrices over division algebras into products of commutators. Linear Algebra and its Applications, 646(??):119–131, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001331. See cor-

rigendum [BDHS23b].

Bien:2023:IWS

[BDHS23a] M. H. Bien, T. H. Dung, N. T. T. Ha, and T. N. Son. Involution widths of skew linear groups generated by involutions. Linear Algebra and its Applications, 679(??):305–326, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003646.

Bien:2023:CDM

[BDHS23b] Mai Hoang Bien, Truong Huu Dung, Nguyen Thi Thai Ha, and Tran Nam Son. Corrigendum to “Decompositions of matrices over division algebras into products of commutators” [Linear Algebra Appl. 646 (2022) 119–131]. Linear Algebra and its Applications, 666(??):37–41, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000435.  See

[BDHS22].

Botelho:2020:HPS

[BDI20] Fernanda Botelho, Priyadarshi Dey, and Dijana Ilisević. Hermitian projections on some Banach spaces and related topics. Linear Algebra and its Applications, 598(??):92–104, August 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301701.

Bani-Domi:2021:RGN

[BDK21] Watheq Bani-Domi and Fuad Kittaneh. Refined and generalized numerical radius inequalities for 2 × 2 operator matrices. Linear Algebra and its Applications, 624(??):364–386, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001506.

Berger:2024:CLM

[BDN24] Martin Berger, Tom Drescher, and Tim Netzer. Classifying linear matrix inequalities via abstract operator systems. Linear Algebra and its Applications, 682(??):28–49, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300407X.

Bag:2020:OMC

[BDNO20] Priyabrata Bag, Santanu Dey, Masaru Nagisa, and Hiroyuki Osaka. The order-n minors of certain (n + k) × n matrices. Linear Algebra and its Applications, 603(??):368–389, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303074.

Bebiano:2021:KCH

[BdPS21] Natália Bebiano, Joáo da Providéncia, and Ilya M. Spitkovsky. On Kippenhahn curves and higher-rank numerical ranges of some matrices. Linear Algebra and its Applications, 629(??):246–257, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003074.

Baragana:2020:BRP

[BDRS20] Itziar Baragaña, Marija Dodig, Alicia Roca, and Marko Stošić. Bounded rank perturbations of regular pencils over arbitrary fields. Linear Algebra and its Applications, 601(??):180–188, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302536.

Bala:2021:IMT

[BDSS21] Neeru Bala, Kousik Dhara, Jaydeb Sarkar, and Aryaman Sensarma. Idempotent, model, and Toeplitz operators attaining their norms. Linear Algebra and its Applications, 622(??):150–165, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001361.

Burde:2021:AID

[BDV21] Dietrich Burde, Karel Dekimpe, and Bert Verbeke. Almost inner derivations of 2-step nilpotent Lie algebras of genus 2. Linear Algebra and its Applications, 608(??):185–202, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303967.

Braga:2021:IUG

[BDVR21] Rodrigo O. Braga, Renata R. Del-Vecchio, and Virgínia M. Rodrigues. Integral unicyclic graphs. Linear Algebra and its Applications, 614(??):281–300, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302330.

Boboc:2024:CAS

[BDvW24] C. Boboc, S. Dăscălescu, and L. van Wyk. Cyclic algebras, symbol algebras and gradings on matrices. Linear Algebra and its Applications, 688(??):157–178, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000612.

Burde:2020:CPL

[BE20] Dietrich Burde and Christof Ender. Commutative post-Lie algebra structures on nilpotent Lie algebras and Poisson algebras. Linear Algebra and its Applications, 584(??):107–126, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379519303933.

Barrett:2020:RDS

[BEF20] Wayne Barrett, Emily J. Evans, and Amanda E. Francis. Resistance distance and spanning 2-forest matrices of linear 2-trees. Linear Algebra and its Applications, 606(??):41–67, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303220.

Bolten:2022:TMS

[BEFSC22] Matthias Bolten, Sven-Erik Ekström, Isabella Furci, and Stefano Serra-Capizzano. Toeplitz momentary symbols: definition, results, and limitations in the spectral analysis of structured matrices. Linear Algebra and its Applications, 651(??):51–82, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002397.

Behera:2021:GIE

[Beh21] Kiran Kumar Behera. A generalized inverse eigenvalue problem and m-functions. Linear Algebra and its Applications, 622(??):46–65, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001294.

Barrett:2022:NCA

[BEHK22] Wayne Barrett, Emily Evans, H. Tracy Hall, and Mark Kempton. New conjectures on algebraic connectivity and the Laplacian spread of graphs. Linear Algebra and its Applications, 648 (??):104–132, September 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001793.

Bart:2021:RDU

[BES21] H. Bart, T. Ehrhardt, and B. Silbermann. Rank decomposition under zero pattern constraints and L-free directed graphs. Linear Algebra and its Applications, 621(??):135–180, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001075.

Bezushchak:2022:ADA

[Bez22] Oksana Bezushchak. Automorphisms and derivations of algebras of infinite matrices. Linear Algebra and its Applications, 650(??):42–59, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522002130.

Bertram:2021:QFS

[BF21] Christian Bertram and Heike Faßbender. A quadrature framework for solving Lyapunov and Sylvester equations. Linear Algebra and its Applications, 622(??):66–103, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001336.

Beaufort:2024:LCS

[BF24a] Louis-Brahim Beaufort and Anna Fino. Locally conformal SKT almost abelian Lie algebras. Linear Algebra and its Applications, 684(??):1–22, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300352X.

Bertram:2024:FLR

[BF24b] Christian Bertram and Heike Faßbender. On a family of low-rank algorithms for largescale algebraic Riccati equations. Linear Algebra and its Applications, 687(??):38–67, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952400034X.

Barrett:2023:SGA

[BFF^{+}23a] Wayne Barrett, Shaun Fallat, Veronika Furst, Franklin Kenter, Shahla Nasserasr, Brendan Rooney, Michael Tait, and Hein van der Holst. Sparsity of graphs that allow two distinct eigenvalues. Linear Algebra and its Applications, 674(??):377–395, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002161.

Barrett:2023:RGD

[BFF^{+}23b] Wayne Barrett, Shaun Fallat, Veronika Furst, Shahla Nasserasr, Brendan Rooney, and Michael Tait. Regular graphs of degree at most four that allow two distinct eigenvalues. Linear Algebra and its Applications, 679(??):127–164, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003488.

Bazzoni:2023:CSL

[BFLT23] Giovanni Bazzoni, Marco Freibert, Adela Latorre, and Nicoletta Tardini. Complex symplectic Lie algebras with large Abelian subalgebras. Linear Algebra and its Applications, 677 (??):254–305, November 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003051.

Branquinho:2023:PBF

[BFMM23] Amílcar Branquinho, Ana Foulquié-Moreno, and Manuel Mañas. Positive bidiagonal factorization of tetradiagonal Hessenberg matrices. Linear Algebra and its Applications, 677(??):132–160, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003002.

Belitskii:2021:CMS

[BFMS21] Genrich R. Belitskii, Vyacheslav Futorny, Mikhail Muzychuk, and Vladimir V. Sergeichuk. Congruence of matrix spaces, matrix tuples, and multilinear maps. Linear Algebra and its Applications, 609(??):317–331, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304407.

Bartusel:2023:PRA

[BFO23] David Bartusel, Hartmut Führ, and Vignon Oussa. Phase retrieval for affine groups over prime fields. Linear Algebra and its Applications, 677(??):161–193, November 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379523002999.

Bondarenko:2021:PCN

[BFPS21] Vitalij M. Bondarenko, Vyacheslav Futorny, Anatolii P. Petravchuk, and Vladimir V. Sergeichuk. Pairs of commuting nilpotent operators with one-dimensional intersection of kernels and matrices commuting with a Weyr matrix. Linear Algebra and its Applications, 612(??):188–205, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305231.

Bhunia:2020:RQH

[BG20] Sushil Bhunia and Krishnendu Gongopadhyay. Reversible quaternionic hyperbolic isometries. Linear Algebra and its Applications, 591(??):268–283, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305646.

Baghipur:2022:SRG

[BGGA22] Maryam Baghipur, Hilal A. Ganie, Modjtaba Ghorbani, and Enide Andrade. On the spectral radius of the generalized adjacency matrix of a digraph. Linear Algebra and its Applications, 653(??):287–313, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002944.

Byrne:2022:ASR

[BGLR22] Eimear Byrne, Heide Gluesing-Luerssen, and Alberto Ravagnani. Anticodes in the sum-rank metric. Linear Algebra and its Applications, 643(??):80–98, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000556.

Bottcher:2021:WMB

[BGOO21] Albrecht Böttcher, Stephan Ramon Garcia, Mohamed Omar, and Christopher O’Neill. Weighted means of B-splines, positivity of divided differences, and complete homogeneous symmetric polynomials. Linear Algebra and its Applications, 608(??):68–83, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303918.

Baghel:2022:NSS

[BGS22] Mohit Kumar Baghel, Nicolas Gillis, and Punit Sharma. On the non-symmetric semidefinite Procrustes problem. Linear Algebra and its Applications, 648(??):133–159, September 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437952200146X.

Balasubramanian:2024:EEB

[BH24] Sriram Balasubramanian and Neha Hotwani. C^{⋆}-extreme entanglement breaking maps on operator systems. Linear Algebra and its Applications, 685(??):182–213, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000041.

Beaupre:2021:IGP

[BHBO21] Aria Beaupré, Emily Hoopes-Boyd, and Grace O’Brien. On the images of generalized polynomials evaluated on matrices over an algebraically closed skew field. Linear Algebra and its Applications, 610(??):827–836, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304985.

Borgard:2021:ADK

[BHD^{+}21] Riley Borgard, Steven N.

Harding, Haley Duba, Chloe Makdad, Jay Mayfield, Randal Tuggle, and Eric S. Weber. Accelerating the distributed Kaczmarz algorithm by strong over-relaxation. Linear Algebra and its Applications, 611(??):334–355, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305188.

Barreras:2021:MTB

[BHHP21] Francisco Barreras, Mikhail Hayhoe, Hamed Hassani, and Victor M. Preciado. Measuretheoretic bounds on the spectral radius of graphs from walks. Linear Algebra and its Applications, 625(??):126–145, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001877.

Byrne:2022:DFM

[BHKW22] Eimear Byrne, Anna-Lena Horlemann, Karan Khathuria, and Violetta Weger. Density of free modules over finite chain rings. Linear Algebra and its Applications, 651(??):1–25, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S002437952200235X.

Bhat:2020:RCP

[BHMU20] B. V. Rajarama Bhat, Robin Hillier, Nirupama Mallick, and Vijaya Kumar U. Roots of completely positive maps. Linear Algebra and its Applications, 587(??):143–165, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304677.

Bhunia:2024:IBN

[Bhu24]  Pintu Bhunia.  Improved

bounds for the numerical radius via polar decomposition of operators. Linear Algebra and its Applications, 683(??):31–45, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300441X.

Biborski:2022:NSR

[Bib22]  Iwo Biborski.  Note on

the spread of real symmetric matrices with entries in fixed interval. Linear Algebra and its Applications, 632(??):246–257, January 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003566.

Bechtold:2020:MPP

[BJ20a] B. K. Bechtold and C. R. Johnson. Monotonicity, path product matrices, and principal submatrices. Linear Algebra and its Applications, 593(??):269–275, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300604.

Bhatia:2020:SHT

[BJ20b] Rajendra Bhatia and Tanvi Jain. A Schur–Horn theorem for symplectic eigenvalues. Linear Algebra and its Applications, 599(??):133–139, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301877.

Bhatia:2023:NRP

[BJ23] Rajendra Bhatia and Tanvi Jain. The numerical radius and positivity of block matrices. Linear Algebra and its Applications, 656(??):463–482, January 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003676.

Bapat:2021:BDM

[BJP21] R. B. Bapat, Rakesh Jana, and S. Pati. The bipartite distance matrix of a nonsingular tree. Linear Algebra and its Applications, 631(??):254–281, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003347.

Bracic:2022:NRC

[BK22] Janko Bračič and Marko Kandić. On the normalizer of the reflexive cover of a unital algebra of linear transformations. Linear Algebra and its Applications, 653(??):207–230, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002907.

Bogdanic:2024:JSN

[B– DK^{+}24] Duško Bogdanić, Alen – Durić, Sara Koljančić, Polona Oblak, and Klemen Šivic. Jordan structures of nilpotent matrices in the centralizer of a nilpotent matrix with two Jordan blocks of the same size. Linear Algebra and its Applications, 690(??):59–90, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000910.

Bourdon:2023:CRP

[BKLT23] Paul Bourdon, Carlos Kubrusly, Trieu Le, and Derek Thompson. Closed-range posinormal operators and their products. Linear Algebra and its Applications, 671(??):38–58, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300143X.

Bhunia:2023:ATS

[BKPS23] Pintu Bhunia, Fuad Kittaneh, Kallol Paul, and Anirban Sen. Anderson’s theorem and A-spectral radius bounds for semi-Hilbertian space operators. Linear Algebra and its Applications, 657(??):147–162, January 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003846.

Bovdi:2020:OPS

[BKR^{+}20] Victor A. Bovdi, Tetiana Klymchuk, Tetiana Rybalkina, Mohamed A. Salim, and Vladimir V. Sergeichuk. Operators on positive semidefinite inner product spaces. Linear Algebra and its Applications, 596(??):82–105, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S002437952030118X.

Bargetz:2020:API

[BKRS20] Christian Bargetz, Jona Klemenc, Simeon Reich, and Natalia Skorokhod. On angles, projections and iterations. Linear Algebra and its Applications, 603(??):41–56, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302615.

Burdak:2015:CPC

[BKS15] Zbigniew Burdak, Marek Kosiek, and Marek Slociński. Compatible pairs of commuting isometries. Linear Algebra and its Applications, 479(??):216–259, August 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379515002384. See cor-

rigendum [BKSP23].

Borges:2021:CLO

[BKS^{+}21] Victor Senoguchi Borges, Iryna Kashuba, Vladimir V. Sergeichuk, Eduardo Ventilari Sodré, and André Zaidan. Classification of linear operators satisfying (Au, v) = (u, A^{r}v) or (Au, A^{r}v) = (u, v) on a vector space with indefinite scalar product. Linear Algebra and its Applications, 611(??):118–134, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305589.

Burdak:2023:CCP

[BKSP23] Zbigniew Burdak, Marek Kosiek, Marek Słociński, and Patryk Pagacz. Corrigendum to “Compatible pairs of commuting isometries” [Linear Algebra Appl. 479 (2015) 216–259]. Linear Algebra and its Applications, 675(??):106–117, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002367.  See

[BKS15].

Bhat:2023:PAU

[BKT23] B. V. Rajarama Bhat, Samir Kar, and Bharat Talwar. Peripherally automorphic unital completely positive maps. Linear Algebra and its Applications, 678(??):191–205, December 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379523003257.

Bourin:2020:CMI

[BL20] Jean-Christophe Bourin and Eun-Young Lee. Clarkson– McCarthy inequalities with unitary and isometry orbits. Linear Algebra and its Applications, 601(??):170– 179, September 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379520302135.

Behajaina:2022:CGG

[BL22] Angelot Behajaina and François Legrand. On Cayley graphs over generalized dicyclic groups. Linear Algebra and its Applications, 642(??):264–284, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000775.

Behajaina:2023:IMC

[BL23] Angelot Behajaina and François Legrand. On integral mixed Cayley graphs over nonabelian finite groups admitting an abelian subgroup of index 2. Linear Algebra and its Applications, 675(??):256–273, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002446.

Bourin:2024:DDB

[BL24] Jean-Christophe Bourin and Eun-Young Lee. Diagonal and off-diagonal blocks of positive definite partitioned matrices. Linear Algebra and its Applications, 684(??):87–100, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004627.

Blau:2020:RBA

[Bla20]  Harvey I. Blau.  Reality-

based algebras with a twodimensional representation. Linear Algebra and its Applications, 590(??):97–123, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305488.

Benner:2021:RPT

[BLMT21] Peter Benner, Xin Liang, Suzana Miodragović, and Ninoslav Truhar. Relative perturbation theory for quadratic Hermitian eigenvalue problems. Linear Algebra and its Applications, 618(??):97–128, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000422.

Bunger:2021:NPS

[BLR21] F. Bünger, M. Lange, and S. M. Rump. On norms of principal submatrices. Linear Algebra and its Applications, 620(??):27–36, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100080X.

Bukovsek:2020:CPF

[BLS20] Damjana Kokol Bukovsek, Thomas Laffey, and Helena Smigoc. Completely positive factorizations associated with Euclidean distance matrices corresponding to an arithmetic progression. Linear Algebra and its Applications, 597(??):113–132, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301506.

Breaz:2020:NMS

[BM20] Simion Breaz and Sergiu Megieşan. Nonderogatory matrices as sums of idempotent and nilpotent matrices. Linear Algebra and its Applications, 605(??):239–248, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030344X.

Bouthat:2022:CPN

[BM22] Ludovick Bouthat and Javad Mashreghi. The critical point and the p-norm of the Hilbert L-matrix. Linear Algebra and its Applications, 634(??):1–14, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003761.

Basha:2023:ENR

[BM23] Aishah Ibraheam Basha and Judi J. McDonald. An extension of the numerical range over finite fields. Linear Algebra and its Applications, 673(??):14–27, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001714.

Bourhim:2024:MPM

[BMM24] Abdellatif Bourhim, Mohamed Mabrouk, and Mostafa Mbekhta. Maps preserving matrices of local reduced minimum modulus zero at a fixed vector. Linear Algebra and its Applications, 680(??):65–82, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003671.

Bursics:2023:ARA

[BMPS23] Balázs Bursics, Dávid Matolcsi, Péter Pál Pach, and Jakab Schrettner. Avoiding right angles and certain Hamming distances. Linear Algebra and its Applications, 677(??):71–87, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300294X.

Behajaina:2023:IDS

[BMR23] Angelot Behajaina, Roghayeh Maleki, and Andriaherimanana Sarobidy Razafimahatratra. On the intersection density of the symmetric group acting on uniform subsets of small size. Linear Algebra and its Applications, 664(??):61–103, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000101.

Bikchentaev:2020:AIU

[BMS20] A. Bikchentaev, E. McDonald, and F. Sukochev. Ando’s inequality for uniform submajorization. Linear Algebra and its Applications, 605(??):206–226, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303384.

Bamberg:2021:POE

[BMS21] John Bamberg, Giusy Monzillo, and Alessandro Siciliano. Pseudo-ovals of elliptic quadrics as Delsarte designs of association schemes. Linear Algebra and its Applications, 624(??):281–317, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001786.

Barbarino:2023:REH

[BN23] Giovanni Barbarino and Vanni Noferini. On the Rellich eigendecomposition of para-Hermitian matrices and the sign characteristics of ∗-palindromic matrix polynomials. Linear Algebra and its Applications, 672(??): 1–27, September 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379523001635.

Boyvalenkov:2022:RIP

[BNS22] Peter Boyvalenkov, Hiroshi Nozaki, and Navid Safaei. Rationality of the inner products of spherical s-distance tdesigns for t ≥ 2s − 2, s ≥ 3. Linear Algebra and its Applications, 646(??):107–118, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001367.

Bekjan:2023:JCS

[BO23] Turdebek N. Bekjan and Kordan N. Ospanov. On jointly concavity of some trace functions. Linear Algebra and its Applications, 664(??):147–164, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000241.

Bojanczyk:2021:AIL

[Boj21] Adam W. Bojanczyk. Algorithms for indefinite linear least squares problems. Linear Algebra and its Applications, 623(??):104–127, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304274.

Boley:2021:FCD

[Bol21] Daniel Boley. On fast computation of directed graph Laplacian pseudo-inverse. Linear Algebra and its Applications, 623(??):128–148, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030495X.

Bolotnikov:2022:CMP

[Bol22] Vladimir Bolotnikov. Cyclic matrices and polynomial interpolation over division rings. Linear Algebra and its Applications, 646(??):132–174, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001380.

Bottazzi:2021:BAD

[Bot21] Tamara Bottazzi. Best approximation by diagonal operators in Schatten ideals. Linear Algebra and its Applications, 620(??):1–26, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000847.

Bourget:2021:SDF

[Bou21] A. Bourget. Spectral distribution of families of Hankel matrices. Linear Algebra and its Applications, 624(??):103–120, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001658.

Bhat:2020:UGG

[BP20] Mushtaq A. Bhat and S. Pirzada. On unicyclic graphs with given number of pendent vertices and minimal energy. Linear Algebra and its Applications, 597(??):86–93, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030152X.

Barik:2021:CNG

[BP21a] Sasmita Barik and Sukanta Pati. Classes of nonbipartite graphs with reciprocal eigenvalue property. Linear Algebra and its Applications, 612(??):177–187, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030522X.

Bhunia:2021:DIC

[BP21b] Pintu Bhunia and Kallol Paul. Development of inequalities and characterization of equality conditions for the numerical radius. Linear Algebra and its Applications, 630(??):306–315, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003116. See cor-

rigendum [BP23b].

Banerjee:2023:SGO

[BP23a] Anirban Banerjee and Samiron Parui. On some general operators of hypergraphs. Linear Algebra and its Applications, 667(??):97–132, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000782.

Bhunia:2023:CDI

[BP23b] Pintu Bhunia and Kallol Paul. Corrigendum to “Development of inequalities and characterization of equality conditions for the numerical radius” [Linear Algebra Appl. 630 (2021) 306–315]. Linear Algebra and its Applications, 679(??):1–3, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003245.  See

[BP21b].

Bisai:2023:CCE

[BPS23] Bappa Bisai, Sourav Pal, and Prajakta Sahasrabuddhe. On q-commuting co-extensions and q-commutant lifting. Linear Algebra and its Applications, 658(??):186–205, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004037.

Baragana:2021:GSC

[BPZ21] Itziar Baragaña, F. Puerta, and I. Zaballa. On the geometry of the set of control matrices of linear systems with prescribed controllability indices. Linear Algebra and its Applications, 624(??):214–246, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001774.

Baragana:2020:FRP

[BR20a] Itziar Baragaña and Alicia Roca. Fixed rank perturbations of regular matrix pencils. Linear Algebra and its Applications, 589(??):201–221, March 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305361.

Baragana:2020:ROP

[BR20b] Itziar Baragaña and Alicia Roca. Rank-one perturbations of matrix pencils. Linear Algebra and its Applications, 606(??):170–191, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303530.

Brassil:2021:NSP

[BR21] Matthew Brassil and Zinovy Reichstein. On the number of symmetric presentations of a determinantal hypersurface. Linear Algebra and its Applications, 617(??):39–48, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379521000215.

Benzi:2022:RDB

[BR22] Michele Benzi and Michele Rinelli. Refined decay bounds on the entries of spectral projectors associated with sparse Hermitian matrices. Linear Algebra and its Applications, 647(??):1–30, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001501.

Barik:2023:LSC

[BR23a] Sasmita Barik and Sane Umesh Reddy. Laplacian spectra of cographs: a twin reduction perspective. Linear Algebra and its Applications, 675(??):203–224, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002471.

Breaz:2023:DMU

[BR23b] Simion Breaz and Cristian Rafiliu. Decompositions of matrices by using commutators. Linear Algebra and its Applications, 662(??):39–48, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004530.

Biswas:2024:DGI

[BR24] Rounak Biswas and Falguni Roy. Drazin and group invertibility in algebras spanned by two idempotents. Linear Algebra and its Applications, 689(??):155–175, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000764.

Briec:2022:DLS

[Bri22]  Walter Briec.  Determi-

nants and limit systems in some idempotent and nonassociative algebraic structure. Linear Algebra and its Applications, 651(??):162–208, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002415.

Benito:2023:ECN

[BRL23] Pilar Benito and Jorge Roldán-López. Equivalent constructions of nilpotent quadratic Lie algebras. Linear Algebra and its Applications, 657(??):1–31, January 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003809.

Brualdi:2021:EC

[Bru21] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 608(??):407–409, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303785.

Brualdi:2022:EC

[Bru22] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 632(??):332–334, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003700.

Brualdi:2024:EC

[Bru24] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 684(??):158–159, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004639.

Baake:2020:NME

[BS20a] Michael Baake and Jeremy Sumner. Notes on Markov embedding. Linear Algebra and its Applications, 594(??):262–299, June 1, 2020. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300823.

Baeth:2020:UTM

[BS20b] Nicholas R. Baeth and Rylan Sampson. Upper triangular matrices over information algebras. Linear Algebra and its Applications, 587(??):334–357, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304902.

Baillon:2020:NRP

[BS20c] Jean-Bernard Baillon and Alberto Seeger. New results on Pareto spectra. Linear Algebra and its Applications, 588(??):338–363, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305051.

Bakhvalov:2020:TBD

[BS20d] Pavel A. Bakhvalov and Mikhail D. Surnachev. Transformation to a block-diagonal form of matrices generating bounded semigroups. Linear Algebra and its Applications, 603(??):275–288, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302858.

Bapat:2020:SNF

[BS20e] R. B. Bapat and Sivaramakrishnan Sivasubramanian. Smith Normal Form of a distance matrix inspired by the four-point condition. Linear Algebra and its Applications, 603(??):301–312, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302792.

Bartnicka:2020:APP

[BS20f] Edyta Bartnicka and Metod Saniga. Affine and projective planes linked with projective lines over certain rings of lower triangular matrices. Linear Algebra and its Applications, 586(??):377–409, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930446X.

Buckley:2020:NEE

[BS20g] Anita Buckley and Klemen Sivic. New examples of extremal positive linear maps. Linear Algebra and its Applications, 598(??):110–144, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379520301671.

Banakh:2021:ENS

[BS21a] Taras Banakh and Javier Cabello Sánchez. Every nonsmooth 2-dimensional Banach space has the Mazur– Ulam property. Linear Algebra and its Applications, 625(??):1–19, September 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001841.

Bovdi:2021:GSL

[BS21b] V. A. Bovdi and V. P. Shchedryk. Generating solutions of a linear equation and structure of elements of the Zelisko group. Linear Algebra and its Applications, 625(??):55–67, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100183X.

Bhatia:2022:ESM

[BS22] Rajendra Bhatia and Rajesh Sharma. Extensions of the Schur majorisation inequalities. Linear Algebra and its Applications, 651 (??):385–389, October 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522002488.

Bock:2023:AMP

[BS23a] Wolfgang Bock and Alfilgen N. Sebandal. An adjacency matrix perspective of talented monoids and Leavitt path algebras. Linear Algebra and its Applications, 678(??):295–316, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003300.

Braun:2023:HFI

[BS23b] Phillip Braun and Hristo Sendov. On the Hadamard– Fischer inequality, the inclusionexclusion formula, and bipartite graphs. Linear Algebra and its Applications, 668(??):64–92, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000927.

Bukovsek:2023:SNM

[BŠ23c] Damjana Kokol Bukovšek and Helena Šmigoc. Symmetric nonnegative matrix trifactorization. Linear Algebra and its Applications, 665(??):36–60, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Bolognini:2024:IV

[BS24a] Davide Bolognini and Paolo Sentinelli. Immanant varieties. Linear Algebra and its Applications, 682 (??):164–190, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300424X.

Bunger:2024:SRS

[BS24b] Florian Bünger and Alberto Seeger. On sign-real spectral radii and sign-real expansive matrices. Linear Algebra and its Applications, 680(??):293–324, January 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003841.

Bera:2023:COP

[BSC23] Anindita Bera, Gniewomir Sarbicki, and Dariusz Chruściński. A class of optimal positive maps in M_{n}. Linear Algebra and its Applications, 668(??):131–148, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300099X.

Bajpai:2023:ASH

[BSS23] Jitendra Bajpai, Sandip Singh, and Shashank Vikram Singh. Arithmeticity of some hypergeometric groups. Linear Algebra and its Applications, 661(??):137–148, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004438.

Beckerleg:2020:DCA

[BT20a] Melanie Beckerleg and Andrew Thompson. A divide-and-conquer algorithm for binary matrix completion. Linear Algebra and its Applications, 601(??):113–133, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302111.

Boij:2020:BWR

[BT20b] Mats Boij and Zach Teitler. A bound for the Waring rank of the determinant via syzygies. Linear Algebra and its Applications, 587 (??):195–214, February 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930477X.

Buczynska:2020:HCT

[Buc20] Weronika Buczyńska. The Hackbusch conjecture on tensor formats — Part two. Linear Algebra and its Applications, 584(??):221–232, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304033.

Budzinskiy:2024:DLR

[Bud24] Stanislav Budzinskiy. On the distance to low-rank matrices in the maximum norm. Linear Algebra and its Applications, 688(??):44–58, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000570.

Bui:2021:GBM

[Bui21] Vuong Bui. Growth of bilinear maps. Linear Algebra and its Applications, 624(??):198–213, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001671.

Bui:2022:JSR

[Bui22] Vuong Bui. On the joint spectral radius of nonnegative matrices. Linear Algebra and its Applications, 654(??):89–101, December 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003135.

Burlachenko:2020:NPR

[Bur20] E. Burlachenko. Numerator polynomials of Riordan matrices. Linear Algebra and its Applications, 603(??):20–40, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302718.

Burlachenko:2021:NPR

[Bur21] E. Burlachenko. Numerator polynomials of Riordan matrices and generalized Lagrange series. Linear Algebra and its Applications, 620(??):322–343, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001105.

Barik:2024:SPT

[BV24] Sasmita Barik and Piyush Verma. Spectral properties of token graphs. Linear Algebra and its Applications, 687(??):181–206, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

Benner:2021:FTL

[BW21] Peter Benner and Steffen W. R. Werner. Frequency- and time-limited balanced truncation for large-scale secondorder systems. Linear Algebra and its Applications, 623(??):68–103, August 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303153.

Britnell:2022:IRW

[BW22] John R. Britnell and Mark Wildon. Involutive random walks on total orders and the anti-diagonal eigenvalue property. Linear Algebra and its Applications, 641(??):1–47, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200026X.

Beard:2024:MMC

[BW24] Kenneth Beard and Aaron Welters. Matrix monotonicity and concavity of the principal pivot transform. Linear Algebra and its Applications, 682(??):323–350, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379523004305.

Bai:2021:CRR

[BWW21] Zhong-Zhi Bai, Lu Wang, and Wen-Ting Wu. On convergence rate of the randomized Gauss–Seidel method. Linear Algebra and its Applications, 611(??):237–252, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305115.

Bartoli:2022:NMT

[BZZ22] Daniele Bartoli, Giovanni Zini, and Ferdinando Zullo. Non-minimum tensor rank Gabidulin codes. Linear Algebra and its Applications, 650(??):248–266, October 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002348.

Cheriyath:2022:PRE

[CA22] Haritha Cheriyath and Nikita Agarwal. On the Perron root and eigenvectors associated with a subshift of finite type. Linear Algebra and its Applications, 633(??):42–70, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003682.

Cueto-Avellaneda:2022:SIB

[CAEH^{+}22] María Cueto-Avellaneda, Yuta Enami, Daisuke Hirota, Takeshi Miura, and Antonio M. Peralta. Surjective isometries between unitary sets of unital JB^{⋆}-algebras. Linear Algebra and its Applications, 643(??):39–79, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000465.

Cheng:2021:CBR

[CAJ21] Che-Man Cheng, Daniel Oluwadamilare Akintoye, and Ruiqiang Jiao. Commutator bounds and region of singular values of the commutator with a rank one matrix. Linear Algebra and its Applications, 613(??):347–376, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305383.

Carlsson:2021:LCI

[Car21] Marcus Carlsson. Lipschitz continuity for isotropic matrix functions. Linear Algebra and its Applications, 624(??):259–266, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379521001634.

Cavers:2021:PSP

[Cav21]  Michael Cavers.  Polyno-

mial stability and potentially stable patterns. Linear Algebra and its Applications, 613(??):87–114, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305693.

Chen:2021:RFC

[CB21] Lixiang Chen and Changjiang Bu. A reduction formula for the characteristic polynomial of hypergraph with pendant edges. Linear Algebra and its Applications, 611(??):171–186, February 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305085.

Cao:2022:FPD

[CC22a] Lei Cao and Zhi Chen. Faces of the polytope of doubly substochastic matrices. Linear Algebra and its Applications, 637(??):59–81, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004122.

Clark:2022:AHS

[CC22b] Gregory J. Clark and Joshua N. Cooper. Applications of the Harary–Sachs theorem for hypergraphs. Linear Algebra and its Applications, 649(??):354–374, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001999.

Charina:2021:BTB

[CCCS21] Maria Charina, Costanza Conti, Mariantonia Cotronei, and Tomas Sauer. Bivariate two-band wavelets demystified. Linear Algebra and its Applications, 608(??):13–36, January 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303864.

Chan:2022:FFR

[CCD^{+}22] Ada Chan, Gabriel Coutinho, Whitney Drazen, Or Eisenberg, Chris Godsil, Mark Kempton, Gabor Lippner, Christino Tamon, and Hanmeng Zhan. Fundamentals of fractional revival in graphs. Linear Algebra and its Applications, 655 (??):129–158, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003287.

Chan:2020:UHT

[CCGV^{+}20] Eunice Y. S. Chan, Robert M. Corless, Laureano Gonzalez-Vega, J. Rafael Sendra, Juana Sendra, and Steven E. Thornton. Upper Hessenberg and Toeplitz Bohemians. Linear Algebra and its Applications, 601(??):72–100, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301713.

Camenga:2021:GNR

[CCH^{+}21] Kristin A. Camenga, Brandon Collins, Gage Hoefer, Jonny Quezada, Patrick X. Rault, James Willson, and Rebekah B. Johnson Yates. On the geometry of numerical ranges over finite fields. Linear Algebra and its Applications, 628(??):182–201, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002664.

Chen:2021:EPC

[CCK21a] Zhi Chen, Lei Cao, and Selcuk Koyuncu. The extreme points of centrosymmetric transportation polytopes. Linear Algebra and its Applications, 608(??):214–235, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304080.

Cheon:2021:SIM

[CCK21b] Gi-Sang Cheon, Bryan Curtis, and Hana Kim. Semiinvolutory matrices and signed self-inverse. Linear Algebra and its Applications, 622(??):294–315, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100149X.

Cheon:2022:RPA

[CCKM22] Gi-Sang Cheon, Bryan Curtis, Gukwon Kwon, and Arnauld Mesinga Mwafise. Riordan posets and associated incidence matrices. Linear Algebra and its Applications, 632(??):308–331, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003670.

Cvetkovic:2021:CTM

[CCL21] Dragana Lj. Cvetković, Ljiljana Cvetković, and Chaoqian Li. CKV-type matrices with applications. Linear Algebra and its Applications, 608(??):158–184, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304079.

Chiru:2023:SRM

[CCO23] Iulia-Elena Chiru, Septimiu Crivei, and Gabriela Olteanu. Strongly regular matrices revisited. Linear Algebra and its Applications, 658(??):233–249, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004062.

Cheon:2022:DER

[CCP22] Gi-Sang Cheon, Marshall M. Cohen, and Nikolaos Pantelidis. Decompositions and eigenvectors of Riordan matrices. Linear Algebra and its Applications, 642(??):118–138, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000738.

Caputo:2020:SCC

[CCPKP20] J.-G. Caputo, G. Cruz-Pacheco, A. Knippel, and P. Panayotaros. Spectra of chains connected to complete graphs. Linear Algebra and its Applications, 605(??):29–62, November 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303347.

Cheon:2022:RKM

[CCS22] Gi-Sang Cheon, Bryan Curtis, and Bryan Shader. Riordan– Krylov matrices over an algebra. Linear Algebra and its Applications, 636(??):93–114, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004213.

Canto:2021:ITN

[CCU21] Begoña Cantó, Rafael Cantó, and Ana M. Urbano. Irreducible totally nonnegative matrices with a prescribed Jordan structure. Linear Algebra and its Applications, 609(??):129–151, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030416X.

Curgus:2020:OEF

[ĆD20] Branko Ćurgus and Aad Dijksma. Operators without eigenvalues in finitedimensional vector spaces. Linear Algebra and its Applications, 605(??):63–117, November 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437952030330X.

Chauhan:2023:ASM

[CD23a] Bharat Pratap Chauhan and Dipti Dubey. On almost semimonotone matrices and the linear complementarity problem. Linear Algebra and its Applications, 661(??):35–50, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004359.

Curgus:2023:OEF

[ĆD23b] Branko Ćurgus and Aad Dijksma. Operators without eigenvalues in finitedimensional vector spaces: Essential uniqueness of the model. Linear Algebra and its Applications, 679 (??):86–98, December 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003312.

Cavaleri:2021:GLG

[CDD21] Matteo Cavaleri, Daniele D’Angeli, and Alfredo Donno. Gain-line graphs via Gphases and group representations. Linear Algebra and its Applications, 613(??):241–270, March 15, 2021. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305346.

Centrone:2022:GMI

[CDdM22] Lucio Centrone, Diogo Diniz, and Thiago Castilho de Mello. Graded monomial identities and almost non-degenerate gradings on matrices. Linear Algebra and its Applications, 653(??):246–265, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200283X.

Conde:2021:SRB

[CDG21] Cristian M. Conde, Ezequiel Dratman, and Luciano N. Grippo. On the spectral radius of block graphs with prescribed independence number α. Linear Algebra and its Applications, 614(??):111–124, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300628.

Conde:2022:SRB

[CDG22a] Cristian M. Conde, Ezequiel Dratman, and Luciano N. Grippo. On the spectral radius of block graphs having all their blocks of the same size. Linear Algebra and its Applications, 634(??):137–148, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004067.

Costantini:2022:RAI

[CDG22b] A. Costantini, B. Drabkin, and L. Guerrieri. Rees algebras of ideals of star configurations. Linear Algebra and its Applications, 645(??):91–122, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001045.

Cicalo:2012:SDN

[CdGS12] Serena Cicalò, Willem A. de Graaf, and Csaba Schneider. Six-dimensional nilpotent Lie algebras. Linear Algebra and its Applications, 436(1):163–189, January 1, 2012. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379511004861.  See

[CdGS20].

Cicalo:2020:CSD

[CdGS20] Serena Cicalò, Willem A. de Graaf, and Csaba Schneider. Corrigendum to “Sixdimensional nilpotent Lie algebras” [Linear Algebra Appl. 436 (2012) 163–189]. Linear Algebra and its Applications, 604(??):507–508, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303189.  See

[CdGS12].

Carvalho:2020:SCQ

[CDM20] Luís Carvalho, Cristina Diogo, and Sérgio Mendes. The star-center of the quaternionic numerical range. Linear Algebra and its Applications, 603(??):166–185, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302706.

Carvalho:2021:QNR

[CDM21] Luís Carvalho, Cristina Diogo, and Sérgio Mendes. Quaternionic numerical range of complex matrices. Linear Algebra and its Applications, 620(??):168–181, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000975.

Chiribella:2023:CEP

[CDPR23] Giulio Chiribella, Kenneth R. Davidson, Vern I. Paulsen, and Mizanur Rahaman. Counterexamples to the extendibility of positive unital norm-one maps. Linear Algebra and its Applications, 663(??):102–115, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000034.

Crismale:2023:MA

[CDR23] Vitonofrio Crismale, Simone Del Vecchio, and Stefano Rossi. On the monotone C^{⋆}-algebra. Linear Algebra and its Applications, 659(??):33–41, February 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004189.

Chavez-Dominguez:2021:CQG

[CDS21] Javier Alejandro Chávez-Domínguez and Andrew T. Swift. Connectivity for quantum graphs. Linear Algebra and its Applications, 608(??):37–53, January 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303931.

Cruz:2022:ICA

[CdSV22] J. P. Cruz, R. B. dos Santos, and A. C. Vieira. Identities and cocharacters of the algebra of 3 × 3 matrices with non-trivial grading and transpose graded involution. Linear Algebra and its Applications, 652(??):173–195, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002634.

Cardoso:2022:AEH

[CDVPT22] Kauê Cardoso, Renata Del-Vecchio, Lucas Portugal, and Vilmar Trevisan. Adjacency energy of hypergraphs. Linear Algebra and its Applications, 648(??):181–204, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001690.

Chalkis:2022:ESS

[CEF^{+}22] Apostolos Chalkis, Ioannis Z. Emiris, Vissarion Fisikopoulos, Panagiotis Repouskos, and Elias Tsigaridas. Efficient sampling in spectrahedra and volume approximation. Linear Algebra and its Applications, 648(??):205–232, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001471.

Carmona:2021:GIS

[CEJM21] A. Carmona, A. M. Encinas, M. J. Jiménez, and M. Mitjana. The group inverse of some circulant matrices. Linear Algebra and its Applications, 614(??):415–436, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305279.

Carmona:2021:ERW

[CEM21] A. Carmona, A. M. Encinas, and M. Mitjana. Eigenvalues with respect to a weight for general boundary value problems on networks. Linear Algebra and its Applications, 614(??):208–243, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301816.

Coco:2023:SNE

[CER^{+}23] Armando Coco, Sven-Erik Ekström, Giovanni Russo, Stefano Serra-Capizzano, and Santina Chiara Stissi. Spectral and norm estimates for matrix-sequences arising from a finite difference approximation of elliptic operators. Linear Algebra and its Applications, 667(??):10–43, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000824.

Cesi:2020:SGS

[Ces20] Filippo Cesi. On the spectral gap of some Cayley graphs on the Weyl group W (B_{n}). Linear Algebra and its Applications, 586(??):274–295, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304641.

Choudhury:2021:MCI

[CESS21] Projesh Nath Choudhury, N. Eagambaram, K. C. Sivakumar, and P. Sushmitha. On the matrix class Q_{0} and inverse monotonicity properties of bordered matrices. Linear Algebra and its Applications, 612(??):206–222, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305243.

Cowen:2022:CBR

[CF22] Carl C. Cowen and Christopher Felder. Convexity of the Berezin range. Linear Algebra and its Applications, 647(??):47–63, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001483.

Chiang:2024:IPC

[CF24] Chun-Yueh Chiang and Hung-Yuan Fan. Inheritance properties of the conjugate discrete-time algebraic Riccati equation. Linear Algebra and its Applications, 683(??):71–97, February 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004251.

Chan:2020:CHD

[CFK^{+}20] Ada Chan, Shaun Fallat, Steve Kirkland, Jephian C.-H. Lin, Shahla Nasserasr, and Sarah Plosker. Complex Hadamard diagonalisable graphs. Linear Algebra and its Applications, 605(??):158–179, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303414.

Chruscinski:2021:CSG

[CFKO21] Dariusz Chruściński, Ryohei Fujii, Gen Kimura, and Hiromichi Ohno. Constraints for the spectra of generators of quantum dynamical semigroups. Linear Algebra and its Applications, 630(??):293–305, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003098.

Crampe:2023:RRT

[CFR23] Nicolas Crampé, Luc Frappat, and Eric Ragoucy. Representations of the rank two Racah algebra and orthogonal multivariate polynomials. Linear Algebra and its Applications, 664(??):165–215, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300023X.

Caalim:2020:ISO

[CFSiT20] Jonathan V. Caalim, Vyacheslav Futorny, Vladimir V. Sergeichuk, and Yu ichi Tanaka. Isometric and selfadjoint operators on a vector space with nondegenerate diagonalizable form. Linear Algebra and its Applications, 587(??):92–110, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304732.

Chmielinski:2020:CDS

[CG20] Jacek Chmieliński and Moshe Goldberg. Continuity and discontinuity of seminorms on infinite-dimensional vector spaces. II. Linear Algebra and its Applications, 594(??):249–261, June 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300951.

Chen:2022:RSG

[CG22] Qian-Qian Chen and Ji-Ming Guo. The rank of a signed graph. Linear Algebra and its Applications, 651(??):407–425, October 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002476.

Cardinali:2023:OPS

[CG23] Ilaria Cardinali and Luca Giuzzi. On orthogonal polar spaces. Linear Algebra and its Applications, 674(??):493–518, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002380.

Casado:2021:CEA

[CGG^{+}21] Yolanda Cabrera Casado, Maria Inez Cardoso Gonçalves, Daniel Gonçalves, Dolores Martín Barquero, and Cándido Martín González. Chains in evolution algebras. Linear Algebra and its Applications, 622(??):104–149, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001300.

Coutinho:2022:QWD

[CGJvB22] Gabriel Coutinho, Chris Godsil, Emanuel Juliano, and Christopher M. van Bommel. Quantum walks do not like bridges. Linear Algebra and its Applications, 652(??):155–172, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002671.

Costara:2023:ATD

[CGMP23] C. Costara, A. E. Guterman, A. M. Maksaev, and V. V. Promyslov. Automorphisms of the total digraph for the ring of square matrices over a field. Linear Algebra and its Applications, 666(??):129–143, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000848.

Cardinali:2021:NAS

[CGP21] I. Cardinali, L. Giuzzi, and A. Pasini. Nearly all subspaces of a classical polar space arise from its universal embedding. Linear Algebra and its Applications, 627(??):287–307, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002482.

Cardoso:2022:JAF

[CGP22] Domingos M. Cardoso, Helena Gomes, and Sofia J. Pinheiro. The H-join of arbitrary families of graphs — the universal adjacency spectrum. Linear Algebra and its Applications, 648(??):160–180, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001665.

Chen:2021:EMC

[CH21] Yu Chen and Yaoping Hou. Eigenvalue multiplicity in cubic signed graphs. Linear Algebra and its Applications, 630(??):95–111, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002937.

Chen:2021:NCM

[Che21] Zhangchi Chen. On nonsingularity of circulant matrices. Linear Algebra and its Applications, 612(??):162–176, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305644.

Chen:2022:SIL

[Che22] Li Chen. Specht’s invariant and localization of operator tuples. Linear Algebra and its Applications, 633(??):290–302, January 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003840.

Choi:2020:FCN

[CHHS20] Hayoung Choi, Jinglian He, Hang Hu, and Yuanming Shi. Fast computation of von Neumann entropy for largescale graphs via quadratic approximations. Linear Algebra and its Applications, 585(??):127–146, January 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304203.

Christensen:2024:SPV

[Chr24a] Erik Christensen. Some points of view on Grothendieck’s inequalities. Linear Algebra and its Applications, 691(??):196–215, June 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000995.

Christensen:2024:UMF

[Chr24b]  Erik Christensen.  Unique

matrix factorizations associated to bilinear forms and Schur multipliers. Linear Algebra and its Applications, 688(??):215–231, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000648.

Cameron:2022:DPR

[CHSW22] Thomas R. Cameron, H. Tracy Hall, Ben Small, and Alexander Wiedemann. On digraphs with polygonal restricted numerical range. Linear Algebra and its Applications, 642(??):285–310, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000829.

Cardoso:2020:SCU

[CHT20] Kauê Cardoso, Carlos Hoppen, and Vilmar Trevisan. The spectrum of a class of uniform hypergraphs. Linear Algebra and its Applications, 590(??):243–257, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305634.

Chen:2020:SIM

[CHY20] Zhibing Chen, Zejun Huang, and Jingru Yan. The stable index of 0–1 matrices. Linear Algebra and its Applications, 600(??):148–160, September 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302044.

Chi:2021:MPB

[CI21] Jocelyn T. Chi and Ilse C. F. Ipsen. Multiplicative perturbation bounds for multivariate multiple linear regression in Schatten p-norms. Linear Algebra and its Applications, 624(??):87–102, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001518.

Ciardo:2020:BPP

[Cia20a] Lorenzo Ciardo. The Braess’ paradox for pendent twins. Linear Algebra and its Applications, 590(??):304–316, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305610.

Ciardo:2020:FCG

[Cia20b] Lorenzo Ciardo. A Fiedler center for graphs generalizing the characteristic set. Linear Algebra and its Applications, 584(??):197–220, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379519303982.

Ciardo:2022:PVM

[Cia22] Lorenzo Ciardo. Perron value and moment of rooted trees. Linear Algebra and its Applications, 635(??):69–94, February 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003980.

Chouaieb:2022:GCP

[CIM22] Nadia Chouaieb, Bruno Iannazzo, and Maher Moakher. Geometries on the cone of positive-definite matrices derived from the power potential and their relation to the power means. Linear Algebra and its Applications, 638(??):80–109, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004407.

Cheng:2020:PWC

[CJ20] Che-Man Cheng and Ruiqiang Jiao. Proof of Wenzel’s conjecture concerning singular values of the commutator of rank one matrices. Linear Algebra and its Applications, 592(??):165–174, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379520300380.

Champagne:2022:SPS

[CJMP22] Gabriel Champagne, Nathaniel Johnston, Mitchell MacDonald, and Logan Pipes. Spectral properties of symmetric quantum states and symmetric entanglement witnesses. Linear Algebra and its Applications, 649(??):273–300, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001914.

Choudhury:2021:IHM

[CK21] Projesh Nath Choudhury and M. Rajesh Kannan. Interval hulls of N-matrices and almost P -matrices. Linear Algebra and its Applications, 617(??):27–38, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000161.

Cervantes:2023:CNC

[CK23a] Jonathan Cervantes and Mike Krebs. Chromatic numbers of Cayley graphs of abelian groups: a matrix method. Linear Algebra and its Applications, 676 (??):277–295, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002768.

Cole:2023:CMC

[CK23b] Sam Cole and Steve Kirkland. Clusters in Markov chains via singular vectors of Laplacian matrices. Linear Algebra and its Applications, 660(??):1–39, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004219.

Cheraghpour:2024:VI

[CK24] Hassan Cheraghpour and Bojan Kuzma. Vanishing immanants. Linear Algebra and its Applications, 692(??):1–23, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000892.

Cheon:2023:MPC

[CKKR23] Gi-Sang Cheon, Bumtle Kang, Suh-Ryung Kim, and Homoon Ryu. Matrix periods and competition periods of Boolean Toeplitz matrices. Linear Algebra and its Applications, 672(??):228–250, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001647.

Choi:2024:LCW

[CKL24] Hayoung Choi, Sejong Kim, and Yongdo Lim. Linearity of Cartan and Wasserstein means. Linear Algebra and its Applications, 681(??):66–88, January 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004007.

Choi:2020:MEP

[CKLL20] Hayoung Choi, Sejong Kim, Hosoo Lee, and Yongdo Lim. Matrix extremal problems and shift invariant means. Linear Algebra and its Applications, 587(??):166–194, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304768.

Chien:2023:NRC

[CKLN23] Mao-Ting Chien, Steve Kirkland, Chi-Kwong Li, and Hiroshi Nakazato. Numerical ranges of cyclic shift matrices. Linear Algebra and its Applications, 678 (??):268–294, December 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003294.

Cortinovis:2020:MVS

[CKM20] Alice Cortinovis, Daniel Kressner, and Stefano Massei. On maximum volume submatrices and cross approximation for symmetric semidefinite and diagonally dominant matrices. Linear Algebra and its Applications, 593(??):251–268, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300768.

Chruscinski:2022:BFN

[CKOS22] Dariusz Chruściński, Gen Kimura, Hiromichi Ohno, and Tanmay Singal. Bounding the Frobenius norm of a qdeformed commutator. Linear Algebra and its Applications, 646(??):95–106, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001227.

Chruscinski:2023:OPG

[CKOS23] Dariusz Chruściński, Gen Kimura, Hiromichi Ohno, and Tanmay Singal. Oneparameter generalization of the Böttcher–Wenzel inequality and its application to open quantum dynamics. Linear Algebra and its Applications, 656(??):158–166, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003470.

Chang:2020:EBC

[CKTP20] Xiao-Wen Chang, Peng Kang, and David Titley-Peloquin. Error bounds for computed least squares estimators. Linear Algebra and its Applications, 586(??):28–42, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304483.

Cavalcante:2020:SIT

[CL20] Felipe Barbosa Cavalcante and Artem Lopatin. Separating invariants of three nilpotent 3 × 3 matrices. Linear Algebra and its Applications, 607(??):9–28, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303669.

Chen:2021:CSC

[CL21] Shangdi Chen and Junying Liang. Construction of spread codes based on Abelian noncyclic orbit codes. Linear Algebra and its Applications, 608(??):54–67, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520303943.

Chang:2022:GN

[CL22] Sarula Chang and Jianxi Li. Graphs G with nullity n(G) − g(G) − 1. Linear Algebra and its Applications, 642(??):251–263, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000842.

Chooi:2024:APM

[CL24] Wai Leong Chooi and Jinting Lau. Adjacency preserving maps on symmetric tensors. Linear Algebra and its Applications, 690(??):27–58, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000818.

Cheng:2020:PCN

[CLL20] Bo Cheng, Muhuo Liu, and Bolian Liu. Proof of a conjecture on the nullity of a connected graph in terms of order and maximum degree. Linear Algebra and its Applications, 587(??):291–301, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304872.

Cho:2021:SEA

[CLLK21] Gook Hwa Cho, Hyang-Sook Lee, Seongan Lim, and Yoonjeong Kim. Storage efficient algorithm for Hermite Normal Form using LLL. Linear Algebra and its Applications, 613(??):183–200, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305899.

Chen:2020:NGT

[CLM20] Yuanyuan Chen, Dan Li, and Jixiang Meng. Nordhaus– Gaddum type inequalities of the second A_{α}-eigenvalue of a graph. Linear Algebra and its Applications, 602(??):57–72, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302329.

Chen:2022:SRH

[CLM22] Yuanyuan Chen, Dan Li, and Jixiang Meng. On the A α- spectral radius of Halin graphs. Linear Algebra and its Applications, 645(??):153–164, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000866.

Cheong:2022:JLT

[CLNY22] Gilyoung Cheong, Jungin Lee, Hayan Nam, and Myungjun Yu. Jordan–Landau theorem for matrices over finite fields. Linear Algebra and its Applications, 655 (??):100–128, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003275.

Carson:2022:BGS

[CLRT22] Erin Carson, Kathryn Lund, Miroslav Rozložník, and Stephen Thomas. Block Gram– Schmidt algorithms and their stability properties. Linear Algebra and its Applications, 638(??):150–195, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004523.

Cameron:2020:GLR

[CLS20] Thomas R. Cameron, Amy N. Langville, and Heather C. Smith. On the graph Laplacian and the rankability of data. Linear Algebra and its Applications, 588(??):81–100, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305026.

Cheong:2023:PEM

[CLS23] Gilyoung Cheong, Yunqi Liang, and Michael Strand. Polynomial equations for matrices over integers modulo a prime power and the cokernel of a random matrix. Linear Algebra and its Applications, 677(??):1–30, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002975.

Carson:2024:TUC

[CLS24] Erin Carson, Jörg Liesen, and Zdeněk Strakoš. Towards understanding CG and GM-RES through examples. Linear Algebra and its Applications, 692(??):241–291, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001381.

Cheng:2022:NCG

[CLT22] Bo Cheng, Muhuo Liu, and Bit-Shun Tam. On the nullity of a connected graph in terms of order and maximum degree. Linear Algebra and its Applications, 632(??):193–232, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003487.

Chen:2022:IPU

[CLW22] Qian Chen, Yingyu Luo, and Yu Wang. The images of polynomials on 3 × 3 upper triangular matrix algebras. Linear Algebra and its Applications, 648(??):254–269, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001975.

Cheng:2022:CBF

[CLwW22] Yen-Jen Cheng, Chia-An Liu, and Chih wen Weng. Counterexamples of the Bhattacharya– Friedland–Peled conjecture. Linear Algebra and its Applications, 641(??):200–207, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000635.

Curto:2022:NNR

[CLY22] Raúl E. Curto, Sang Hoon Lee, and Jasang Yoon. Norm and numerical radius of single operators through tools and techniques from multivariable operator theory. Linear Algebra and its Applications, 649(??):301–325, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001963.

Chen:2020:SLS

[CLZ20] Ming-Zhu Chen, A-Ming Liu, and Xiao-Dong Zhang. The signless Laplacian spectral radius of graphs with forbidding linear forests. Linear Algebra and its Applications, 591(??):25–43, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305592.

Casarino:2022:SLS

[CLZ22] Valentina Casarino, Giovanni Longobardi, and Corrado Zanella. Scattered linear sets in a finite projective line and translation planes. Linear Algebra and its Applications, 650(??):286–298, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002361.

Calixto:2023:FDR

[CM23a] Lucas Calixto and Tiago Macedo. Finite-dimensional representations of map superalgebras. Linear Algebra and its Applications, 676 (??):104–130, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002719.

Cavicchioni:2023:WDC

[CM23b] Giulia Cavicchioni and Alessio Meneghetti. The weight distribution of codes over finite chain rings. Linear Algebra and its Applications, 675(??):90–105, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002409.

Chabbabi:2020:CJI

[CMM20] Fadil Chabbabi, Mostafa Mbekhta, and Lajos Molnár. Characterizations of Jordan*- isomorphisms of C^{∗}-algebras by weighted geometric mean related operations and quantities. Linear Algebra and its Applications, 588(??):364–390, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305002.

Contino:2022:MFS

[CMM22] Maximiliano Contino, Alejandra Maestripieri, and Stefania Marcantognini. A matrix formula for Schur complements of nonnegative selfadjoint linear relations. Linear Algebra and its Applications, 654(??):143–176, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379522003214.

Choi:2021:NIF

[CMP21] Jinwon Choi, Sunyo Moon, and Seungkook Park. A note on the invariant factors of the walk matrix of a graph. Linear Algebra and its Applications, 631(??):362–378, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003396.

Coll:2020:INL

[CMR20] Vincent E. Coll, Nicholas W. Mayers, and Nicholas V. Russoniello. The index of nilpotent Lie poset algebras. Linear Algebra and its Applications, 605(??):118–129, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303396.

Cramer:2021:MACb

[CMR21] Zachary Cramer, Laurent W. Marcoux, and Heydar Radjavi. Matrix algebras with a certain compression property I. Linear Algebra and its Applications, 621(??):50–85, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001026.

Chitour:2021:GBD

[CMS21] Yacine Chitour, Guilherme Mazanti, and Mario Sigalotti. On the gap between deterministic and probabilistic joint spectral radii for discretetime linear systems. Linear Algebra and its Applications, 613(??):24–45, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030567X.

Chooi:2021:CMR

[CMT21] W. L. Chooi, M. H. A. Mutalib, and L. Y. Tan. Commuting maps on rank one triangular matrices. Linear Algebra and its Applications, 626(??):34–55, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002226.

Chen:2020:PAA

[CN20a] Yuzhang Chen and Chi-Keung Ng. Property T C^{⋆}-algebras with amenable tracial states. Linear Algebra and its Applications, 599(??):66–78, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301749.

Cohen:2020:SLF

[CN20b] Alex Cohen and Melvyn B. Nathanson. Sinkhorn limits in finitely many steps. Linear Algebra and its Applications, 589(??):1–8, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305233.

Chien:2021:ATI

[CN21a] Mao-Ting Chien and Hiroshi Nakazato. Abel theorem and inverse numerical range. Linear Algebra and its Applications, 626(??):20–33, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002159.

Chien:2021:USW

[CN21b] Mao-Ting Chien and Hiroshi Nakazato. Unitary similarity of a weighted shift matrix to a symmetric matrix. Linear Algebra and its Applications, 611(??):356–364, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030519X.

Chien:2022:INR

[CN22] Mao-Ting Chien and Hiroshi Nakazato. Inverse numerical range and Abel–Jacobi map of Hermitian determinantal representation. Linear Algebra and its Applications, 633(??):227–243, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003803.

Chien:2023:EMI

[CN23a] Mao-Ting Chien and Hiroshi Nakazato. Elliptic modular invariants of 4-by-4 matrices. Linear Algebra and its Applications, 662(??):49–66, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004657.

Chien:2023:TDA

[CN23b] Mao-Ting Chien and Hiroshi Nakazato. Theta divisor and Abel map for 4by-4 matrices. Linear Algebra and its Applications, 676(??):296–317, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300277X.

Choudhury:2024:QEC

[CN24] Projesh Nath Choudhury and Raju Nandi. Quadratic embedding constants of graphs: Bounds and distance spectra.

Linear Algebra and its Applications, 680(??):108–125, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003695.

Camacho:2023:CES

[CNO23] L. M. Camacho, R. M. Navarro, and B. A. Omirov. Central extensions of some solvable Leibniz superalgebras. Linear Algebra and its Applications, 656(??):63– 91, January 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437952200338X.

Cooper:2024:PRP

[CO24] Joshua Cooper and Utku Okur. Polynomial reconstruction problem for hypergraphs. Linear Algebra and its Applications, 692(??):37–61, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001253.

Chakraborty:2022:NRS

[COB22a] Bikshan Chakraborty, Sarita Ojha, and Riddhick Birbonshi. On the numerical range of some weighted shift operators. Linear Algebra and its Applications, 640(??):179–190, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000258.

Cobo:2022:PMM

[Cob22b]  Helena Cobo.  The Pas-

cal matrix in the multivariate Riordan group. Linear Algebra and its Applications, 635(??):1–23, February 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004080.

Cohen:2020:EFO

[Coh20] Marshall M. Cohen. Elements of finite order in the Riordan group and their eigenvectors. Linear Algebra and its Applications, 602(??):264–280, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302524.

Comaschi:2021:SLS

[Com21] Gaia Comaschi. Stable linear systems of skew-symmetric forms of generic rank ≤ 4. Linear Algebra and its Applications, 624(??):121–152, September 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437952100166X.

Cooper:2020:ASR

[Coo20] Joshua Cooper. Adjacency spectra of random and complete hypergraphs. Linear Algebra and its Applications, 596(??):184–202, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301348.

Cooper:2021:CBL

[Coo21] Joshua N. Cooper. Constraints on Brouwer’s Laplacian spectrum conjecture. Linear Algebra and its Applications, 615(??):11–27, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520306017.

Costara:2020:LMP

[Cos20a] Constantin Costara. Linear maps preserving the polynomial numerical radius of matrices. Linear Algebra and its Applications, 595(??):63–71, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301075.

Costara:2020:NIP

[Cos20b] Constantin Costara. Nonlinear invertibility preserving maps on matrix algebras. Linear Algebra and its Applications, 602(??):216–222, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030241X.

Costara:2021:LMP

[Cos21]  Constantin Costara.  Lin-

ear maps preserving structured singular values of matrices. Linear Algebra and its Applications, 620(??):76–91, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000999.

Costara:2022:LMC

[Cos22]  Constantin Costara.  Lin-

ear maps compressing the polynomial numerical hulls of matrices. Linear Algebra and its Applications, 637(??):49–58, March 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100433X.

Costara:2024:LMP

[Cos24]  Constantin Costara.  Lin-

ear maps preserving inclusion of fixed subsets into the spectrum. Linear Algebra and its Applications, 682(??):285–294, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004287.

Cameron:2020:HSM

[CP20] Thomas R. Cameron and Panayiotis J. Psarrakos. On Householder sets for matrix polynomials. Linear Algebra and its Applications, 585(??):105–126, January 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304264.

Cohen:2021:SCC

[CP21] Nir Cohen and Edgar Pereira. Symmetric completions of cycles and bipartite graphs. Linear Algebra and its Applications, 614(??):164–175, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301579.

Clark:2022:PPN

[CP22] Benjamin J. Clark and Pietro Paparella. Polynomials that preserve nonnegative matrices. Linear Algebra and its Applications, 637(??):110– 118, March 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379521004365.

Chorianopoulos:2023:SSE

[CP23] Christos Chorianopoulos and Panayiotis Psarrakos. On the shell and the shell-extremal eigenvalues of a square matrix. Linear Algebra and its Applications, 665(??):354–381, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000587.

Chattopadhyay:2020:LEZ

[CPS20] Sriparna Chattopadhyay, Kamal Lochan Patra, and Binod Kumar Sahoo. Laplacian eigenvalues of the zero divisor graph of the ring Z_{n}. Linear Algebra and its Applications, 584(??):267–286, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519303556.

Chang:2021:SLO

[CPTP21] Xiao-Wen Chang, Christopher C. Paige, and David Titley-Peloquin. Structure in loss of orthogonality. Linear Algebra and its Applications, 623(??):149–167, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303475.

Colonius:2021:QEL

[CR21] Fritz Colonius and Martin Rasmussen. Quasi-ergodic limits for finite absorbing Markov chains. Linear Algebra and its Applications, 609(??):253–288, January 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030433X.

Chavan:2023:BIR

[CR23] Sameer Chavan and Md. Ramiz Reza. Bi-isometries reducing the hyper-ranges of the coordinates. Linear Algebra and its Applications, 668(??):51–63, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000964.

Cramer:2021:MACa

[Cra21] Zachary Cramer. Matrix algebras with a certain compression property II. Linear Algebra and its Applications, 619(??):210–263, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379521001014.

Lin:2020:SCN

[cRLY20] Yen chi Roger Lin and Wei-Hsuan Yu. Saturated configuration and new large construction of equiangular lines. Linear Algebra and its Applications, 588(??):272–281, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305099.

Calderon:2020:LEP

[CRR20] Pablo Calderón, Noelia B. Rios, and Mariano A. Ruiz. Local extrema for Procrustes problems in the set of positive definite matrices. Linear Algebra and its Applications, 602(??):252–263, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302597.

Camenga:2022:GWN

[CRSY22] Kristin A. Camenga, Patrick X. Rault, Ilya M. Spitkovsky, and Rebekah B. Johnson Yates. The Gau–Wu number for 4×4 and select arrowhead matrices. Linear Algebra and its Applications, 644(??):192–218, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379522000763.

Curtis:2020:SPO

[CS20] Bryan A. Curtis and Bryan L. Shader. Sign patterns of orthogonal matrices and the strong inner product property. Linear Algebra and its Applications, 592(??):228–259, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300434.

Chatterjee:2021:IMI

[CS21] Manami Chatterjee and K. C. Sivakumar. Intervals of H-matrices and inverse Mmatrices. Linear Algebra and its Applications, 614 (??):24–43, April 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379519305087.

Csikvari:2022:NSS

[Csi22]  Péter Csikvári.  Note on

the sum of the smallest and largest eigenvalues of a triangle-free graph. Linear Algebra and its Applications, 650(??):92–97, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002208.

Correia:2021:FRC

[CST21] David Munhá Correia, Benny Sudakov, and István Tomon. Flattening rank and its combinatorial applications. Linear Algebra and its Applications, 625(??):113–125, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002044.

Choudhury:2020:ADC

[CT20] Projesh Nath Choudhury and Michael J. Tsatsomeros. Algorithmic detection and construction of N-matrices. Linear Algebra and its Applications, 602(??):46–56, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302226.

Chang:2023:GFO

[CT23] Ting-Chung Chang and Bit-Shun Tam. Graphs of fixed order and size with maximal A_{α}-index. Linear Algebra and its Applications, 673(??):69–100, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001866.

Chen:2024:SLA

[CT24] Tyler Chen and Thomas Trogdon. Stability of the Lanczos algorithm on matrices with regular spectral distributions. Linear Algebra and its Applications, 682(??):191–237, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004202.

Clark:2023:CPE

[CTS23] Gregory J. Clark, Felipe Thomaz, and Andrew T. Stephen. Comparing the principal eigenvector of a hypergraph and its shadows. Linear Algebra and its Applications, 673(??):46–68, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001830.

Chen:2021:WCC

[CV21] Xiao Shan Chen and Seak-Weng Vong. On worst-case condition numbers of a multiple nonzero finite generalized singular value. Linear Algebra and its Applications, 616(??):1–18, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379521000069.

Chen:2023:AEP

[CvDB23] Lixiang Chen, Edwin R. van Dam, and Changjiang Bu. All eigenvalues of the power hypergraph and signed subgraphs of a graph. Linear Algebra and its Applications, 676(??):205–210, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002720.

Caseiro:2021:PPI

[CVV21] R. Caseiro, M. A. Facas Vicente, and José Vitória. Projection of a point onto the intersection of spheres in linear varieties. Linear Algebra and its Applications, 610(??):40–51, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304353.

Child:2020:MPP

[CW20] Billy Child and Ian M. Wanless. Multidimensional permanents of polystochastic matrices. Linear Algebra and its Applications, 586(??):89–102, February 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379519304422.

Cai:2021:CSA

[CW21] Yun Cai and Ying Wang. Convergence and stability analysis of iteratively reweighted least squares for noisy block sparse recovery. Linear Algebra and its Applications, 631(??):181–202, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003219.

Cao:2022:PES

[CW22] Xiwang Cao and Jinlong Wan. Perfect edge state transfer on abelian Cayley graphs. Linear Algebra and its Applications, 653(??):44–65, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002804.

Chen:2023:QWM

[CW23a] Zhengxin Chen and Yu Wang. Quasi-Whittaker modules for the n-th Schrödinger algebra. Linear Algebra and its Applications, 677 (??):51–70, November 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002963.

Chen:2023:ZPD

[CW23b] Zhengxin Chen and Yu Wang. Zero product determined n-th Schrödinger algebra. Linear Algebra and its Applications, 679(??):165–193, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003464.

Chen:2023:DHS

[CWBB23] Zhen Chen, Jianfeng Wang, Maurizio Brunetti, and Francesco Belardo. On the divisibility of H-shape trees and their spectral determination. Linear Algebra and its Applications, 675(??):312–337, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002537.

Chehade:2024:STE

[CWW24] Sarah Chehade, Shuzhou Wang, and Zhenhua Wang. Suzuki type estimates for exponentiated sums and generalized Lie–Trotter formulas in JB-algebras. Linear Algebra and its Applications, 680(??):156–169, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003786.

Chen:2023:BTN

[CWYZ23] Ming-Zhu Chen, Ning Wang, Long-Tu Yuan, and Xiao-Dong Zhang. The bipartite Turán number and spectral extremum for linear forests. Linear Algebra and its Applications, 676(??):150–173, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002707.

Chen:2022:SLS

[CWZ22] Wenwen Chen, Bing Wang, and Mingqing Zhai. Signless Laplacian spectral radius of graphs without short cycles or long cycles. Linear Algebra and its Applications, 645(??):123–136, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001124.

Chen:2022:CRM

[CXZ22] Shangdi Chen, Qin Xu, and Junmei Zhang. Constructions of rank metric codes under actions of the unitary groups. Linear Algebra and its Applications, 645(??):293–306, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522001392.

Cimpric:2020:LGP

[CZ20a] Jaka Cimpric and Aljaz Zalar. A local-global principle for linear dependence in enveloping algebras of Lie algebras. Linear Algebra and its Applications, 596(??):230–255, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301324.

Cole:2020:ERH

[CZ20b] Sam Cole and Yizhe Zhu. Exact recovery in the hypergraph stochastic block model: a spectral algorithm. Linear Algebra and its Applications, 593(??):45–73, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300562.

Carlen:2022:MVE

[CZ22a] Eric A. Carlen and Haonan Zhang. Monotonicity versions of Epstein’s Concavity Theorem and related inequalities. Linear Algebra and its Applications, 654 (??):289–310, December 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379522003196.

Castillo:2022:FTM

[CZ22b] Kenier Castillo and Ion Zaballa. On a formula of Thompson and McEnteggert for the adjugate matrix. Linear Algebra and its Applications, 634(??):37–56, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003864.

Chen:2022:UPI

[CZ22c] Xi Chen and Sai-Nan Zheng. A unified proof of interlacing properties of eigenvalues of totally positive matrices. Linear Algebra and its Applications, 632(??):241–245, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100358X.

Cai:2021:MEL

[CZG21] Yiting Cai, Bo Zhou, and Mengmeng Gao. On maximal energy of line graphs with given parameters. Linear Algebra and its Applications, 619(??):12–49, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000690.

Dadar:2023:PSP

[DA23] Elham Dadar and Rahim Alizadeh. On polynomials satisfying power inequality for numerical radius. Linear Algebra and its Applications, 665(??):1–11, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000319.

deAlmeida:2021:STB

[dAdCG^{+}21] Moisés Ceni de Almeida, Julia Sekiguchi da Cruz, Paulo Goldfeld, Luiz Mariano Carvalho, and Michael Souza. Supporting theory for a block approximate inverse preconditioner. Linear Algebra and its Applications, 614(??):325–342, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303086.

Dong:2021:RGD

[DAG21] Shuyu Dong, P.-A. Absil, and K. A. Gallivan. Riemannian gradient descent methods for graph-regularized matrix completion. Linear Algebra and its Applications, 623(??):193–235, August 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Guimaraes:2021:GPI

[dAGK21] Alan de Araújo Guimarães and Plamen Koshlukov. Zgraded polynomial identities of the Grassmann algebra. Linear Algebra and its Applications, 617(??):190–214, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000537.

Das:2022:SPA

[Das22]  Sunil Das.  Sign pat-

terns that allow algebraic positivity. Linear Algebra and its Applications, 653 (??):151–182, November 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002841.

Dax:2023:KAS

[Dax23]  Achiya Dax.  Kaczmarz’s

anomaly: a surprising feature of Kaczmarz’s method. Linear Algebra and its Applications, 662(??):136–162, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004633.

Das:2023:NRD

[DB23] Biswajit Das and Shreemayee Bora. Nearest rank deficient matrix polynomials. Linear Algebra and its Applications, 674(??):304–350, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002021.

deBondt:2020:QTG

[dBS20] Michiel de Bondt and Xiaosong Sun. Quadratic tame generators problem of rank three. Linear Algebra and its Applications, 587(??):1–22, February 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304549.

Dewaele:2023:TDS

[DBV23] Nick Dewaele, Paul Breiding, and Nick Vannieuwenhoven. Three decompositions of symmetric tensors have similar condition numbers. Linear Algebra and its Applications, 664(??):253–263, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000265.

deCamargo:2021:DMD

[dC21] André Pierro de Camargo. Dirichlet matrices: Determinants, permanents and the Factorisatio Numerorum problem. Linear Algebra and its Applications, 628 (??):115–129, November 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002640.

Damm:2023:HPK

[DD23a] Tobias Damm and Nicolas Dietrich. Hadamard powers and kernel perceptrons. Linear Algebra and its Applications, 672(??):93–107, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001556.

Teran:2023:BMP

[DD23b] Fernando De Terán and Froilán M. Dopico. On bundles of matrix pencils under strict equivalence. Linear Algebra and its Applications, 658(??):1–31, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003949.

Du:2023:CMW

[DdF23] Zhibin Du and Carlos M. da Fonseca. The characterization of the minimal weighted acyclic graphs. Linear Algebra and its Applications, 672(??):75–92, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001544.

Dalfo:2021:LST

[DDFM^{+}21] C. Dalfó, F. Duque, R. Fabila-Monroy, M. A. Fiol, C. Huemer, A. L. Trujillo-Negrete, and F. J. Zaragoza Martínez. On the Laplacian spectra of token graphs. Linear Algebra and its Applications, 625(??):322–348, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100207X.

DiVincenzo:2020:FPI

[DdR20] Onofrio Mario Di Vincenzo, Marcos Antônio da Silva Pinto, and Viviane Ribeiro Tomaz da Silva. On the factorability of polynomial identities of upper block triangular matrix algebras graded by cyclic groups. Linear Algebra and its Applications, 601(??):311–337, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302391.

DeBruyn:2020:TCL

[De 20] Bart De Bruyn. On tetrahedrally closed line systems and a generalization of the Haemers–Roos inequality. Linear Algebra and its Applications, 600 (??):130–147, September 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302123.

Desai:2024:STP

[Des24] Dheer Noal Desai. Spectral Turán problems for intersecting even cycles. Linear Algebra and its Applications, 683(??):46–70, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300438X.

Devriendt:2022:ERM

[Dev22] Karel Devriendt. Effective resistance is more than distance: Laplacians, Simplices and the Schur complement. Linear Algebra and its Applications, 639(??):24–49, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522000039.

Dewilde:2021:RFB

[Dew21]  Patrick Dewilde.  Ratio-

nal forms for block lowertriangular matrices. Linear Algebra and its Applications, 623(??):168–192, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030450X.

Dalfo:2020:SER

[DF20a] C. Dalfó and M. A. Fiol. Spectra and eigenspaces from regular partitions of Cayley (di)graphs of permutation groups. Linear Algebra and its Applications, 597(??):94–112, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301439.

Dumitru:2020:GHM

[DF20b] Raluca Dumitru and Jose A. Franco. Generalized Hellinger metric and Audenaert’s inbetweenness. Linear Algebra and its Applications, 585(??):191–198, January 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304306.

Dumitru:2020:RPM

[DF20c] Raluca Dumitru and Jose A. Franco. The Rényi power means of matrices. Linear Algebra and its Applications, 607(??):45–57, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303682.

Dumitru:2022:GBM

[DF22] Raluca Dumitru and Jose A. Franco. Geodesic in-betweenness for means of several matrices. Linear Algebra and its Applications, 636(??):77–92, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004237.

Dumnicki:2020:MAU

[DFH^{+}20] Marcin Dumnicki, Lucja Farnik, Brian Harbourne, Grzegorz Malara, Justyna Szpond, and Halszka Tutaj-Gasińska. A matrixwise approach to unexpected hypersurfaces. Linear Algebra and its Applications, 592(??):113–133, May 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300331.

Damnjanovic:2023:SPB

[DFS23] Ivan Damnjanović, Slobodan Filipovski, and Dragan Stevanović. Spectral properties of balanced trees and dendrimers. Linear Algebra and its Applications, 657 (??):163–196, January 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003858.

DeBruyn:2020:HPE

[DG20] Bart De Bruyn and Mou Gao. The homogeneous pseudoembeddings and hyperovals of the generalized quadrangle H(3, 4). Linear Algebra and its Applications, 593(??):90–115, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300586.

Dwivedi:2022:RHM

[DG22] Omesh Dhar Dwivedi and Darij Grinberg. On the rank of Hankel matrices over finite fields. Linear Algebra and its Applications, 641(??):156–181, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200057X.

Diniz:2023:TDJ

[DGdSS23] Diogo Diniz, Dimas José Gonçalves, Viviane Ribeiro Tomaz da Silva, and Manuela Souza. Two-dimensional Jordan algebras: Their classification and polynomial identities. Linear Algebra and its Applications, 664(??):104–125, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000216.

Daza-Garcia:2021:CPA

[DGET21] Alberto Daza-García, Alberto Elduque, and Liming Tang. Cross products, automorphisms, and gradings. Linear Algebra and its Applications, 610(??):227–256, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304717.

Danielyan:2024:IM

[DGKP24] S. Danielyan, A. Guterman, E. Kreines, and F. Pakovich. Integrability of matrices. Linear Algebra and its Applications, 684(??):37–62, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004597.

Danchev:2023:DES

[DGL23a] Peter Danchev, Esther García, and Miguel Gómez Lozano. Decompositions of endomorphisms into a sum of roots of the unity and nilpotent endomorphisms of fixed nilpotence. Linear Algebra and its Applications, 676(??):44–55, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002653.

Dutta:2023:CDQ

[DGL23b] Sandipan Dutta, Krishnendu Gongopadhyay, and Tejbir Lohan. Classification and decomposition of quaternionic projective transformations. Linear Algebra and its Applications, 676(??):66–87, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300263X.

Dahl:2020:MM

[DGS20] Geir Dahl, Alexander Guterman, and Pavel Shteyner. Majorization for (0, 1)-matrices. Linear Algebra and its Applications, 585(??):147–163, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379519304276.

Dratman:2021:DDM

[DGS^{+}21a] Ezequiel Dratman, Luciano N. Grippo, Martín D. Safe, Celso M. da Silva, and Renata R. Del-Vecchio. The determinant of the distance matrix of graphs with blocks at most bicyclic. Linear Algebra and its Applications, 614(??):437–454, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305334.

Duffner:2021:CIS

[DGS21b] M. A. Duffner, A. E. Guterman, and I. A. Spiridonov. Converting immanants on skew-symmetric matrices. Linear Algebra and its Applications, 618(??):76–96, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000446.

Dahl:2023:MM

[DGS23] Geir Dahl, Alexander Guterman, and Pavel Shteyner. Majorization for (0, ±1)-matrices. Linear Algebra and its Applications, 663(??):200–221, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379523000095.

Dobosevych:2021:SRO

[DH21] Oles Dobosevych and Rostyslav Hryniv. Spectra of rank-one perturbations of self-adjoint operators. Linear Algebra and its Applications, 609(??):339–364, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304493.

Du:2022:MET

[DH22] Zhibin Du and Yinhao Huang. The multiplicity of eigenvalues of trees. Linear Algebra and its Applications, 654(??):56–68, December 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002956.

Du:2023:MEU

[DH23] Zhibin Du and Yinhao Huang. The multiplicity of eigenvalues of unicyclic graphs. Linear Algebra and its Applications, 670(??):19–41, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001192.

Dolinar:2020:PLS

[DHH^{+}20] Gregor Dolinar, Sait Halicioglu, Abdullah Harmanci, Bojan Kuzma, Janko Marovt, and Burcu Ungor. Preservers of the left-star and rightstar partial orders. Linear Algebra and its Applications, 587(??):70–91, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304720.

deHoog:2023:NEB

[dHH23] Frank de Hoog and Markus Hegland. A note on error bounds for pseudo skeleton approximations of matrices. Linear Algebra and its Applications, 669(??):102–117, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001155.

Dugas:2020:NSG

[DHR20] Manfred Dugas, Daniel Herden, and Jack Rebrovich. Normal subgroups of the group of units of incidence algebras. Linear Algebra and its Applications, 586(??):64–88, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304513.

DeVos:2022:CDD

[DID22] Alexis De Vos, Martin Idel, and Stijn De Baerdemacker. Conjectured DXZ decompositions of a unitary matrix. Linear Algebra and its Applications, 652(??):196–212, November 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379522002762.

Domokos:2022:LDF

[DJ22] M. Domokos and D. Joó. Low dimensional flow polytopes and their toric ideals. Linear Algebra and its Applications, 654(??):210–249, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200307X.

Ding:2023:MNM

[DJ23] Wenxuan Ding and Charles R. Johnson. The minimum number of multiplicity 1 eigenvalues among real symmetric matrices whose graph is a linear tree. Linear Algebra and its Applications, 665(??):315–338, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000484.

Djordjevic:2021:SSE

[Djo21] Bogdan D. Djordjević. Singular Sylvester equation in Banach spaces and its applications: Fredholm theory approach. Linear Algebra and its Applications, 622(??):189–214, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001452.

Djordjevic:2023:DSP

[Djo23] Bogdan D. Djordjević. Doubly stochastic and permutation solutions to AXA = XAX when A is a permutation matrix. Linear Algebra and its Applications, 661(??):79–105, March 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004402.

Duff:2021:NCN

[DK21] Igor Pontes Duff and Patrick Kürschner. Numerical computation and new output bounds for time-limited balanced truncation of discretetime systems. Linear Algebra and its Applications, 623(??):367–397, August 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520304511.

Desai:2022:SEG

[DKL^{+}22] Dheer Noal Desai, Liying Kang, Yongtao Li, Zhenyu Ni, Michael Tait, and Jing Wang. Spectral extremal graphs for intersecting cliques. Linear Algebra and its Applications, 644(??):234–258, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001161.

Dopico:2021:P

[DKM^{+}21] Froilán Dopico, Daniel Kressner, Nicola Mastronardi, Volker Mehrmann, and Raf Vandebril. Preface. Linear Algebra and its Applications, 623(??):1–13, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000239.

Donato:2022:SIH

[DL22] Joseph S. Donato and Howard W. Levinson. Structured iterative hard thresholding with on- and off-grid applications. Linear Algebra and its Applications, 638(??):46–79, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S002437952100450X.

Dong:2023:SQM

[DL23] Liqiang Dong and Jicheng Li. The solutions of the quaternion matrix equation AX^{ε} + BX^{δ} = 0 . Linear Algebra and its Applications, 678(??):227–267, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003269.

delaCruz:2020:SSH

[dlCP20] Ralph John de la Cruz and Agnes T. Paras. The sums of symplectic, Hamiltonian, and skew-Hamiltonian matrices. Linear Algebra and its Applications, 603(??):84–90, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302809.

delaCruz:2021:SFP

[dlCS21] Ralph John de la Cruz and Philip Saltenberger. Simple forms for perplectic and symplectic normal matrices. Linear Algebra and its Applications, 620(??):201–227, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001002.

Ding:2020:NUI

[DLL20] Wenxuan Ding, Chi-Kwong Li, and Yuqiao Li. A note on unitarily invariant matrix norms. Linear Algebra and its Applications, 607(??):341–346, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304328.

Diaz-Lopez:2023:CGA

[DLL23] Alexander Diaz-Lopez and Joel Louwsma. Critical groups of arithmetical structures under a generalized star-clique operation. Linear Algebra and its Applications, 656(??):324–344, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003548.

deLima:2020:IGM

[dLMO20] L. S. de Lima, A. Mohammadian, and C. S. Oliveira. On integral graphs with at most two vertices of degree larger than two. Linear Algebra and its Applications, 584(??):164–184, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304021.

delaPena:2021:MCT

[dlPG21] José Antonio de la Peña and Jesús Arturo Jiménez González. Momenta of Coxeter transformations. Linear Algebra and its Applications, 625(??):171–195, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002056.

Dinh:2021:BWD

[DLVV21] Trung Hoa Dinh, Cong Trinh Le, Bich Khue Vo, and Trung Dung Vuong. The α–z-Bures Wasserstein divergence. Linear Algebra and its Applications, 624 (??):267–280, September 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001646.

Duncan:2024:HMA

[DM24] John Duncan and Colin M. McGregor. Hermitians in matrix algebras with operator norm and associated Lie algebras. Linear Algebra and its Applications, 682(??):238–256, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004238.

deMello:2023:PII

[dMdSS23] Thiago Castilho de Mello and Manuela da Silva Souza. Polynomial identities and images of polynomials on null-filiform Leibniz algebras. Linear Algebra and its Applications, 679(??):246–260, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003531.

Dopico:2020:LLR

[DMQD20] Froilán M. Dopico, Silvia Marcaida, María C. Quintana, and Paul Van Dooren. Local linearizations of rational matrices with application to rational approximations of nonlinear eigenvalue problems. Linear Algebra and its Applications, 604(??):441–475, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030327X.

Dellacherie:2020:IMN

[DMS20] Claude Dellacherie, Servet Martínez, and Jaime San Martín. Inverse M-matrix, a new characterization. Linear Algebra and its Applications, 595(??):182–191, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379520300963.

Dopico:2020:RPT

[DN20] Froilán M. Dopico and Vanni Noferini. Root polynomials and their role in the theory of matrix polynomials. Linear Algebra and its Applications, 584(??):37–78, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519303891.

DeVille:2021:CTF

[DN21] Lee DeVille and Eddie Nijholt. Circulant type formulas for the eigenvalues of linear network maps. Linear Algebra and its Applications, 610(??):379–439, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304456.

Dopico:2023:VSP

[DN23] Froilán M. Dopico and Vanni Noferini. The DL(P) vector space of pencils for singular matrix polynomials. Linear Algebra and its Applications, 677(??):88–131, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002938.

Datta:2022:PCM

[DO22] Somantika Datta and Jesse Oldroyd. Properties of certain multi-angle tight frames in R^{d}. Linear Algebra and its Applications, 650(??):75–91, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002154.

Dominici:2020:PSE

[Dom20] Diego Dominici. Power series expansion of a Hankel determinant. Linear Algebra and its Applications, 601(??):17–54, September 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302020.

Diniz:2021:NGM

[DP21] Diogo Diniz and Daniel Pellegrino. On the number of gradings on matrix algebras. Linear Algebra and its Applications, 624(??):14–26, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001622.

Dhara:2023:PSS

[DP23] Kousik Dhara and Santhosh Kumar Pamula. Pseudo S-spectra of special operators in quaternionic Hilbert spaces. Linear Algebra and its Applications, 656(??):345–367, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003536.

Diaz:2020:MSR

[DPR20] Roberto C. Díaz, Germain Pastén, and Oscar Rojo. On the minimal D_{α}-spectral radius of graphs subject to fixed connectivity. Linear Algebra and its Applications, 584(??):353–370, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304161.

DeFord:2020:SFR

[DR20] Daryl DeFord and Daniel N. Rockmore. On the spectrum of finite, rooted homogeneous trees. Linear Algebra and its Applications, 598(??):165–185, August 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301750.

Draisma:2022:ICP

[Dra22]  Jan Draisma.  The irre-

ducible control property in matrix groups. Linear Algebra and its Applications, 634(??):15–29, February 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003852.

Drazin:2023:ELP

[Dra23] Michael P. Drazin. EP-like properties of (b, c)-inverses. Linear Algebra and its Applications, 669(??):136–146, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000794.

Diniz:2024:GII

[DRG24] Diogo Diniz, Alex Ramos, and José Lucas Galdino. Graded identities with involution for the algebra of upper triangular matrices. Linear Algebra and its Applications, 688(??):120–156, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000569.

Drury:2024:NRI

[Dru24] Stephen Drury. A numerical radius inequality for sector operators. Linear Algebra and its Applications, 687(??):108–116, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Dodig:2020:CPC

[DS20] Marija Dodig and Marko Stosić. Combinatorics of polynomial chains. Linear Algebra and its Applications, 589(??):130–157, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305427.

Damnjanovic:2022:CNG

[DS22] Ivan Damnjanović and Dragan Stevanović. On circulant nut graphs. Linear Algebra and its Applications, 633(??):127–151, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003712.

Dodig:2024:BRP

[DS24] Marija Dodig and Marko Stošić. Bounded rank perturbations of a regular matrix pencil. Linear Algebra and its Applications, 689(??):126–154, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000806.

Pazzis:2020:SAR

[dSP20a] Clément de Seguins Pazzis. The space of all p-th roots of a nilpotent complex matrix is path-connected. Linear Algebra and its Applications, 596(??):106–116, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301002.

Pazzis:2020:SGP

[dSP20b] Clément de Seguins Pazzis. The structured Gerstenhaber problem (III). Linear Algebra and its Applications, 601(??):134–169, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302159.

Pazzis:2022:STQ

[dSP22] Clément de Seguins Pazzis. The sum of two quadratic matrices: Exceptional cases. Linear Algebra and its Applications, 653(??):357–394, November 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379522003044.

Pazzis:2023:LFC

[dSP23a] Clément de Seguins Pazzis. Locally finite cycles of linear mappings in countable dimension. Linear Algebra and its Applications, 669(??):57–101, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000988.

Pazzis:2023:PTI

[dSP23b] Clément de Seguins Pazzis. The product of two invertible quadratic matrices: Exceptional cases. Linear Algebra and its Applications, 662(??):67–109, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004670.

Pazzis:2024:QSP

[dSP24] Clément de Seguins Pazzis. The quadratic sum problem for symplectic pairs. Linear Algebra and its Applications, 683(??):151–179, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004548.

Duttweiler:2023:SBN

[DTA23] Luke Duttweiler, Sally W. Thurston, and Anthony Almudevar. Spectral Bayesian network theory. Linear Algebra and its Applications, 674(??):282–303, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300215X.

Du:2023:ANA

[Du23] Zhibin Du. The asymptotic normality of adjacency coefficients of bipartite graphs and skew-adjacency coefficients of oriented graphs. Linear Algebra and its Applications, 656(??):541–558, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200355X.

Duan:2023:CPU

[DvDW23] Cunxiang Duan, Edwin R. van Dam, and Ligong Wang. The characteristic polynomials of uniform double hyperstars and uniform hypertriangles. Linear Algebra and its Applications, 678(??):16–32, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003166.

deVries:2023:LVM

[dVLG23] Christopher de Vries, Nico Lombardi, and Eugenia Saorín Gómez. On linearized versions of matrix inequalities. Linear Algebra and its Applications, 674(??):21–45, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002033.

Duan:2021:BSR

[DW21] Cunxiang Duan and Ligong Wang. Bounds on the spectral radius of general hypergraphs in terms of clique number. Linear Algebra and its Applications, 610(??):120–134, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304675.

Deng:2023:CCS

[DY23a] Weina Deng and Baomin Yu. Certain congruences on the semiring of normal tropical matrices. Linear Algebra and its Applications, 668(??):110–130, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001040.

Drivaliaris:2023:GCR

[DY23b] Dimos Drivaliaris and Nikos Yannakakis. A geometric characterization of rangekernel complementarity. Linear Algebra and its Applications, 667(??):88–96, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000976.

Duan:2024:GGP

[DY24] Fang Duan and Qi Yang. On graphs with girth g and positive inertia index of ⌈g⌉/2 − 1 and ⌈g⌉/2. Linear Algebra and its Applications, 683(??):98–110, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004457.

Deng:2020:BSR

[DYB20] Chunli Deng, Hongmei Yao, and Changjiang Bu. The bounds on spectral radius of nonnegative tensors via general product of tensors. Linear Algebra and its Applications, 602(??):206–215, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302573.

Du:2023:FRA

[DYLL23] Zenan Du, Lihua You, Hechao Liu, and Fenjin Liu. Further results on almost controllable graphs. Linear Algebra and its Applications, 677(??):31–50, November 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002951.

Dastgeer:2022:DDS

[DYT22] Zoubia Dastgeer, Awais Younus, and Cemil Tunç. Distinguishability of the descriptor systems with regular pencil. Linear Algebra and its Applications, 652(??):82–96, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002622.

Deng:2022:GAT

[DZCY22] Weina Deng, Xianzhong Zhao, Yanliang Cheng, and Baomin Yu. On the groups associated with a tropical n × n matrix. Linear Algebra and its Applications, 639(??):1–17, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004626.

El-Azhar:2021:IMP

[EAHS21] Hamza El-Azhar, Ayoub Harrat, and Jan Stochel. Idempotents and moment problem for discrete measure. Linear Algebra and its Applications, 628(??):202–227, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379521002767.

ElAbidineAbdelali:2023:OIE

[EE23] Zine El Abidine Abdelali and Youssef El Khatiri. Order isomorphisms on effect algebras of the C^{⋆}-algebras of type CX ⊗ BH. Linear Algebra and its Applications, 674(??):91–123, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002100.

Esmailian:2020:MOG

[EG20] Hossein Esmailian and Ebrahim Ghorbani. Maximum order of graphs with a given corank. Linear Algebra and its Applications, 588(??):122–133, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304975.

Epperly:2021:MRC

[EGC21] Ethan N. Epperly, Nithin Govindarajan, and Shivkumar Chandrasekaran. Minimal rank completions for overlapping blocks. Linear Algebra and its Applications, 627(??):185–198, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379521002469.

Ekstrom:2021:EET

[EGJP21] Sven-Erik Ekström, Carlo Garoni, Adam Jozefiak, and Jesse Perla. Eigenvalues and eigenvectors of tau matrices with applications to Markov processes and economics. Linear Algebra and its Applications, 627(??):41–71, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002408.

Erickson:2024:CVG

[EGK^{+}24] Craig Erickson, Luyining Gan, Jürgen Kritschgau, Jephian C. H. Lin, and Sam Spiro. Complementary vanishing graphs. Linear Algebra and its Applications, 692(??):185–211, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001265.

Erdmann:2021:CTM

[EGL21] Karin Erdmann, Sira Gratz, and Lisa Lamberti. Cluster tilting modules for mesh algebras. Linear Algebra and its Applications, 630 (??):112–157, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100286X.

Evans:2020:NCS

[EGV20] Ron Evans, John Greene, and Mark Van Veen. Nullities for a class of skew-symmetric Toeplitz band matrices. Linear Algebra and its Applications, 593(??):276–304, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300756.

Esteban:2022:NPM

[EHS22] Guillermo Esteban, Clemens Huemer, and Rodrigo I. Silveira. New production matrices for geometric graphs. Linear Algebra and its Applications, 633(??):244–280, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003785.

Edelman:2023:SSM

[EJ23] Alan Edelman and Sungwoo Jeong. On the structure of the solutions to the matrix equation G^{⋆}JG = J. Linear Algebra and its Applications, 657(??):241–273, January 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Exner:2020:CSW

[EJJL20] G. R. Exner, J. Y. Jin, I. B. Jung, and J. E. Lee. On mcomplex symmetric weighted shift operators on C^{n}. Linear Algebra and its Applications, 603(??):130–153, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302688.

Ellison:2024:SCM

[EJS24] David Ellison, Bertrand Jouve, and Lewi Stone. Switching checkerboards in (0, 1)- matrices. Linear Algebra and its Applications, 680(??):274–292, January 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003828.

Evans:2020:CPR

[EL20] Luke Evans and Chun-Kit Lai. Conjugate phase retrieval on CM by real vectors. Linear Algebra and its Applications, 587(??):45–69, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304707.

Elduque:2022:NGG

[Eld22] Alberto Elduque. Non-group gradings on simple Lie algebras. Linear Algebra and its Applications, 640(??):149–161, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000222.

Elmer:2023:SVM

[Elm23] Jonathan Elmer. The separating variety for matrix semi-invariants. Linear Algebra and its Applications, 674(??):466–492, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002379.

Elphick:2024:TCS

[ELW24] Clive Elphick, William Linz, and Pawel Wocjan. Two conjectured strengthenings of Turán’s theorem. Linear Algebra and its Applications, 684(??):23–36, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004603.

Estaremi:2023:RTC

[EM23] Y. Estaremi and M. Mathieu. On the range of TROconditional expectations. Linear Algebra and its Applications, 670(??):78–103, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001295.

Evert:2022:CCO

[EMŠV22] Eric Evert, Scott McCullough, Tea Štrekelj, and Anna Vershynina. Convexity of a certain operator trace functional. Linear Algebra and its Applications, 643(??):218–234, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000830.

Eisenmann:2022:STP

[EN22] Henrik Eisenmann and Yuji Nakatsukasa. Solving twoparameter eigenvalue problems using an alternating method. Linear Algebra and its Applications, 643 (??):137–160, June 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437952200074X.

Erey:2021:MDS

[Ere21] Aysel Erey. Maximal distance spectral radius of 4chromatic planar graphs. Linear Algebra and its Applications, 618(??):183–202, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000574.

Ellard:2020:FNL

[ES20] Richard Ellard and Helena Smigoc. Families of Newtonlike inequalities for sets of selfconjugate complex numbers. Linear Algebra and its Applications, 597(??):46–68, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301427.

Engel:2023:BRS

[ES23] Gernot Michael Engel and Sergeĭ Sergeev. Bounding the row sum arithmetic mean by Perron roots of rowpermuted matrices. Linear Algebra and its Applications, 673(??):220–232, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001970.

Estrada:2020:HRL

[Est20]  Ernesto Estrada.  ‘Hubs-

repelling’ Laplacian and related diffusion on graphs/ networks. Linear Algebra and its Applications, 596 (??):256–280, July 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379520301336.

Estrada:2023:ENS

[Est23] Ernesto Estrada. Every nonsingular spherical Euclidean distance matrix is a resistance distance matrix. Linear Algebra and its Applications, 656(??):198–209, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003500.

Erzan:2020:ECE

[ET20] Ayse Erzan and Asli Tuncer. Explicit construction of the eigenvectors and eigenvalues of the graph Laplacian on the Cayley tree. Linear Algebra and its Applications, 586(??):111–129, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437951930463X.

Ellers:2020:GSG

[EV20] Erich W. Ellers and Oliver Villa. Generation of the symplectic group by involutions. Linear Algebra and its Applications, 591(??):154–159, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300070.

Franca:2021:MSG

[FB21] Francisca Andrea Macedo França and André Ebling Brondani. On the main spectrum of generalized Bethe trees. Linear Algebra and its Applications, 628(??):56– 71, November 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521002597.

Fabila-Carrasco:2022:MNH

[FCLP22] John Stewart Fabila-Carrasco, Fernando Lledó, and Olaf Post. Matching number, Hamiltonian graphs and magnetic Laplacian matrices. Linear Algebra and its Applications, 642(??):86–100, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000490.

Fernandes:2020:MMT

[FdCP20] Rosário Fernandes, Henrique F. da Cruz, and Susana C. Palheira. Matrices in A(R, S) with minimum t-term ranks. Linear Algebra and its Applications, 586(??):239–261, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304446.

Fernandes:2020:CCB

[FdCS20] Rosário Fernandes, Henrique F. da Cruz, and Domingos Salomão. On a conjecture concerning the Bruhat order. Linear Algebra and its Applications, 600(??):82–95, September 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302032.

Feki:2020:TCO

[Fek20] Kais Feki. On tuples of commuting operators in positive semidefinite inner product spaces. Linear Algebra and its Applications, 603 (??):313–328, October 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303062.

Friday:2023:PDM

[FFSW23] Sarah Friday, Fumiko Futamura, Jordan Smith, and Aaron Waclawczyk. Powers of defective matrices from diagonalizable dilations. Linear Algebra and its Applications, 661(??):202–221, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004426.

Fuentes:2020:CBM

[FG20] Edinson Fuentes and Luis E. Garza. CMV block matrices for symmetric matrix measures on the unit circle. Linear Algebra and its Applications, 587(??):314–333, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304914.

Fulman:2022:CLP

[FG22] Jason Fulman and Robert Guralnick. Cohen Lenstra partitions and mutually annihilating matrices over a finite field. Linear Algebra and its Applications, 645(??):1–8, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001008.

Fongi:2023:MPI

[FG23a] Guillermina Fongi and M. Celeste Gonzalez. Moore– Penrose inverse and partial orders on Hilbert space operators. Linear Algebra and its Applications, 674(??):1–20, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379523002045.

Fuentes:2023:HPM

[FG23b] Edinson Fuentes and Luis E. Garza. On the Hahn’s property for matrix orthogonal polynomials on the unit circle. Linear Algebra and its Applications, 675(??):118–141, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002422.

Fidelis:2020:GIB

[FGDY20] Claudemir Fidelis, Dimas José Gonçalves, Diogo Diniz, and Felipe Yukihide Yasumura. Graded involutions on blocktriangular matrix algebras. Linear Algebra and its Applications, 585(??):24–44, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304124.

Fideles:2024:CNG

[FGGG24] Claudemir Fideles, Ana Beatriz Gomes, Alexandre Grishkov, and Alan Guimarães. A characterization of the natural grading of the Grassmann algebra and its nonhomogeneous Z_{2}-gradings. Linear Algebra and its Applications, 680(??):93–107, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003750.

Fazzi:2021:GSA

[FGM21] Antonio Fazzi, Nicola Guglielmi, and Ivan Markovsky. A gradient system approach for Hankel structured low-rank approximation. Linear Algebra and its Applications, 623(??):236–257, August 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305413.

Fang:2024:QVG

[FGPW24] Licheng Fang, Shuzheng Guo, Yaqun Peng, and Fengpeng Wang. Quantitative version of Gordon’s lemma for Hamiltonian with finite range. Linear Algebra and its Applications, 687(??):91–107, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000211.

Freewan:2023:NIR

[FH23] Shaima’a Freewan and Mostafa Hayajneh. On norm inequalities related to the geometric mean. Linear Algebra and its Applications, 670(??):104–120, August 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001465.

Fallat:2022:BLS

[FHLS22] Shaun M. Fallat, H. Tracy Hall, Jephian C. H. Lin, and Bryan L. Shader. The bifurcation lemma for strong properties in the inverse eigenvalue problem of a graph. Linear Algebra and its Applications, 648(??):70–87, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001604.

Farenick:2021:COE

[FHMP21] Douglas Farenick, Farrah Huntinghawk, Adili Masanika, and Sarah Plosker. Complete order equivalence of spin unitaries. Linear Algebra and its Applications, 610(??):1–28, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304559.

Farenick:2022:CCO

[FHMP22] Douglas Farenick, Farrah Huntinghawk, Adili Masanika, and Sarah Plosker. Corrigendum to “Complete order equivalence of spin unitaries” [Linear Algebra Appl. 610 (2021) 1–28]. Linear Algebra and its Applications, 635(??):201–202, February 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004298.

Fidelis:2022:GIT

[Fid22] Claudemir Fidelis. Graded identities of M_{n}(E) and their generalizations over infinite fields. Linear Algebra and its Applications, 643(??):258–280, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000787.

Fikioris:2020:EBK

[Fik20] George Fikioris. Eigenvalue bifurcations in Kac–Murdock– Szegő matrices with a complex parameter. Linear Algebra and its Applications, 607(??):118–150, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303797.

Fiol:2020:NCP

[Fio20] M. A. Fiol. A new class of polynomials from the spectrum of a graph, and its application to bound the kindependence number. Linear Algebra and its Applications, 605(??):1–20, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303323.

Faust:2021:NSV

[FISW21] Theodore Faust, Mark Iwen, Rayan Saab, and Rongrong Wang. On the ℓ^{∞}-norms of the singular vectors of arbitrary powers of a difference matrix with applications to sigma–delta quantization. Linear Algebra and its Applications, 626(??):79–151, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002238.

Furtado:2024:EWG

[FJ24] Susana Furtado and Charles R. Johnson. Efficiency of any weighted geometric mean of the columns of a reciprocal matrix. Linear Algebra and its Applications, 680(??):83–92, January 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003762.

Fallat:2017:TPS

[FJS17] Shaun Fallat, Charles R. Johnson, and Alan D. Sokal. Total positivity of sums, Hadamard products and Hadamard powers: Results and counterexamples. Linear Algebra and its Applications, 520(??):242–259, May 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379517300411. See cor-

rigendum [FJS21].

Fallat:2021:CTP

[FJS21] Shaun Fallat, Charles R. Johnson, and Alan D. Sokal. Corrigendum to “Total positivity of sums, Hadamard products and Hadamard powers: Results and counterexamples” [Linear Algebra Appl. 520 (2017) 242–259]. Linear Algebra and its Applications, 613(??):393–396, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305863.  See

[FJS17].

Fukshansky:2021:GNO

[FK21] Lenny Fukshansky and David Kogan. On the geometry of nearly orthogonal lattices. Linear Algebra and its Applications, 629 (??):112–137, November 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002809.

Fagundes:2023:SGP

[FK23] Pedro Fagundes and Plamen Koshlukov. On sums of gr-PI algebras. Linear Algebra and its Applications, 677(??):200–220, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003026.

Flores:2021:HCR

[FKK21a] Ramón Flores, Delaram Kahrobaei, and Thomas Koberda. Hamiltonicity via cohomology of right-angled Artin groups. Linear Algebra and its Applications, 631(??):94–110, December 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521003165.

Futorny:2021:PTM

[FKK^{+}21b] Vyacheslav Futorny, Tetiana Klymchuk, Olena Klymenko, Vladimir V. Sergeichuk, and Nadiya Shvai. Perturbation theory of matrix pencils through miniversal deformations. Linear Algebra and its Applications, 614(??):455–499, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305620.

Fritzsche:2022:WMB

[FKKM22] Bernd Fritzsche, Bernd Kirstein, Susanne Kley, and Conrad Mädler. The Weyl matrix balls corresponding to the matricial truncated Hamburger moment problem. Linear Algebra and its Applications, 651(??):230–311, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002191.

Fritzsche:2020:SAM

[FKM20] Bernd Fritzsche, Bernd Kirstein, and Conrad Mädler. Schur analysis of matricial Hausdorff moment sequences. Linear Algebra and its Applications, 590(??):133–209, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305415.

Fidelis:2022:GIK

[FKM22] Claudemir Fidelis, Plamen Koshlukov, and David Macêdo. Graded identities for Kac– Moody and Heisenberg algebras with the Cartan grading. Linear Algebra and its Applications, 638(??):196–218, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379521004511.

Filipiak:2021:AKP

[FKMM21] Katarzyna Filipiak, Daniel Klein, Augustyn Markiewicz, and Monika Mokrzycka. Approximation with a Kronecker product structure with one component as compound symmetry or autoregression via entropy loss function. Linear Algebra and its Applications, 610(??):625–646, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304900.

Li:2020:EAS

[fLLV20] Jiao fen Li, Wen Li, and Seak-Weng Vong. Efficient algorithms for solving condition number-constrained matrix minimization problems. Linear Algebra and its Applications, 607(??):190–230, December 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379520303803.

Feng:2020:BSR

[FLRS20] Lihua Feng, Lu Lu, Tamás Réti, and Dragan Stevanović. A bound on the spectral radius of graphs in terms of their Zagreb indices. Linear Algebra and its Applications, 597(??):33–45, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030149X.

Fu:2020:NIS

[FLT20] Xiaohui Fu, Pan-Shun Lau, and Tin-Yau Tam. Norm inequalities for sector block matrices. Linear Algebra and its Applications, 605 (??):249–262, November 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303463.

Fernandez:2020:TAD

[FM20a] Blas Fernández and Stefko Miklavic. On the Terwilliger algebra of distance-biregular graphs. Linear Algebra and its Applications, 597(??):18–32, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301440.

Friedland:2020:NHP

[FM20b] S. Friedland and A. Melman. A note on Hermitian positive semidefinite matrix polynomials. Linear Algebra and its Applications, 598(??):105–109, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301725.

Faccio:2022:LAM

[FM22a] Chiara Faccio and Fabio Marcuzzi. A linear algorithm for the minimal realization problem in physical coordinates with a non-invertible output matrix. Linear Algebra and its Applications, 644(??):149–171, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200088X.

Fallat:2022:MND

[FM22b] Shaun Fallat and Seyed Ahmad Mojallal. On the minimum number of distinct eigenvalues of a threshold graph. Linear Algebra and its Applications, 642(??):1–29, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000441.

Fallat:2020:APM

[FMR20] Shaun M. Fallat and Xavier Martínez-Rivera. On the almost-principal minors of a symmetric matrix. Linear Algebra and its Applications, 606(??):219–243, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520303633.

Filho:2021:LFC

[FMS21] G. Santos Filho, L. Murakami, and I. Shestakov. Locally finite coalgebras and the locally nilpotent radical I. Linear Algebra and its Applications, 621(??):235–253, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001208.

Figula:2023:CFD

[FN23] Ágota Figula and Péter T. Nagy. Classification of a family of 4-dimensional anticommutative algebras and their automorphisms. Linear Algebra and its Applications, 656(??):385–408, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200369X.

Farenick:2022:UWU

[FOP22] Douglas Farenick, Oluwatobi Ruth Ojo, and Sarah Plosker. Universality of Weyl unitaries. Linear Algebra and its Applications, 634(??):57–76, February 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379521003827.

Fidaleo:2022:SEP

[FOR22] Francesco Fidaleo, Federico Ottomano, and Stefano Rossi. Spectral and ergodic properties of completely positive maps and decoherence. Linear Algebra and its Applications, 633(??):104–126, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003724.

Fernandez-Polo:2020:MPR

[FP20] Francisco J. Fernández-Polo. The minimax principle and related topics in the Jordan setting. Linear Algebra and its Applications, 604(??):293–323, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303177.

Fikioris:2021:ECL

[FP21] George Fikioris and Christos Papapanos. Eigenvalue contour lines of Kac–Murdock– Szegő matrices with a complex parameter. Linear Algebra and its Applications, 629(??):87–111, November 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379521002810.

Fornaroli:2022:AII

[FP22] Érica Z. Fornaroli and Roger E. M. Pezzott. Antiisomorphisms and involutions on the idealization of the incidence space over the finitary incidence algebra. Linear Algebra and its Applications, 637(??):82–109, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004341.

Frigerio:2023:SCQ

[FP23] Massimo Frigerio and Matteo G. A. Paris. Swift chiral quantum walks. Linear Algebra and its Applications, 673(??):28–45, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001842.

Favaro:2020:SSU

[FPR20] V. V. Fávaro, D. M. Pellegrino, and P. Rueda. On the size of the set of unbounded multilinear operators between Banach spaces. Linear Algebra and its Applications, 606(??):144–158, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303529.

Farsi:2022:SDM

[FPS22] Carla Farsi, Emily Proctor, and Christopher Seaton. The spectra of digraphs with Morita equivalent C^{⋆}- algebras. Linear Algebra and its Applications, 655 (??):28–64, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200324X.

Fang:2020:IEB

[FQ20] Ximing Fang and Zhijun Qiao. Improved error bounds based on α(M) for the linear complementarity problem. Linear Algebra and its Applications, 589(??):186–200, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305166.

Fasi:2021:SER

[FR21] Massimiliano Fasi and Leonardo Robol. Sampling the eigenvalues of random orthogonal and unitary matrices. Linear Algebra and its Applications, 620(??):297–321, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379521000987.

Franchi:2020:SRE

[Fra20] Massimo Franchi. Some results on eigenvalues of finite type, resolvents and Riesz projections. Linear Algebra and its Applications, 588(??):238–271, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305142.

Franchi:2022:RLR

[Fra22]  Massimo Franchi.  Resol-

vent and logarithmic residues of a singular operator pencil in Hilbert spaces. Linear Algebra and its Applications, 639(??):243–281, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000210.

Fassbender:2021:NOS

[FRS21] Heike Faßbender, Miroslav Rozložník, and Sanja Singer. Nearly optimal scaling in the SR decomposition. Linear Algebra and its Applications, 613(??):295–319, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030536X.

Fickus:2020:HET

[FS20] Matthew Fickus and Courtney A. Schmitt. Harmonic equiangular tight frames comprised of regular simplices. Linear Algebra and its Applications, 586(??):130–169, February 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379519304537.

Filipovski:2023:NBS

[FS23] Slobodan Filipovski and Dragan Stevanović. A note on the bounds for the spectral radius of graphs. Linear Algebra and its Applications, 667(??):1–9, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000769.

Fernandez-Sanchez:2022:LLD

[FSRVSST22] J. Fernández-Sánchez, D. L. Rodríguez-Vidanes, J. B. Seoane-Sepúlveda, and W. Trutschnig. Lineability and K-linear discontinuous functions. Linear Algebra and its Applications, 645(??):52–67, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001021.

Friedman:2023:CDL

[FT23] Joel Friedman and Amir Tootooni. Coordination and discoordination in linear algebra, linear information theory, and coded caching. Linear Algebra and its Applications, 673(??):304–420, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001982.

Feng:2022:SRG

[FW22] Zhimin Feng and Wei Wei. On the A_{α}-spectral radius of graphs with given size and diameter. Linear Algebra and its Applications, 650(??):132–149, October 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002221.

Fang:2023:MSR

[FY23] Xiaona Fang and Lihua You. The maximum spectral radius of graphs of given size with forbidden subgraph. Linear Algebra and its Applications, 666(??):114–128, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000745.

Fan:2023:TEI

[FYS^{+}23] Yi-Zheng Fan, Ya Yang, Chuan-Ming She, Jian Zheng, Yi-Min Song, and Hong-Xia Yang. The trace and Estrada index of uniform hypergraphs with cut vertices. Linear Algebra and its Applications, 660(??):89–117, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004335.

Ganai:2021:PSP

[GA21] Tinku Ganai and Bibhas Adhikari. Preserving spectral properties of structured matrices under structured perturbations. Linear Algebra and its Applications, 629 (??):168–191, November 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002822.

Ganai:2022:EEP

[GA22] Tinku Ganai and Bibhas Adhikari. Eigenvalue embedding problem for quadratic regular matrix polynomials with symmetry structures. Linear Algebra and its Applications, 640(??):118–148, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522000271.

Galvis:2021:CNS

[Gal21]  Juan Galvis.  On condi-

tion numbers of symmetric and nonsymmetric domain decomposition methods. Linear Algebra and its Applications, 610(??):293–320, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304730.

Ghabries:2020:SOQ

[GAM20] Mohammad M. Ghabries, Hassane Abbas, and Bassam Mourad. On some open questions concerning determinantal inequalities. Linear Algebra and its Applications, 596(??):169–183, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301300.

Ghabries:2020:PCD

[GAMA20] Mohammad M. Ghabries, Hassane Abbas, Bassam Mourad, and Abdallah Assi. A proof of a conjectured determinantal inequality. Linear Algebra and its Applications, 605(??):21–28, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379520303360.

Ganesan:2023:SBQ

[Gan23]  Priyanga Ganesan.  Spec-

tral bounds for the quantum chromatic number of quantum graphs. Linear Algebra and its Applications, 674(??):351–376, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002197.

Gau:2024:PCN

[Gau24] Hwa-Long Gau. Proof of a conjecture on numerical ranges of weighted cyclic matrices. Linear Algebra and its Applications, 682 (??):295–308, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004378.

Gnilke:2024:IQF

[GCOP24] Oliver W. Gnilke, Padraig Ó Catháin, Oktay Olmez, and Guillermo Nuñez Ponasso. Invariants of quadratic forms and applications in design theory. Linear Algebra and its Applications, 682(??):1– 27, February 1, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379523004160.

Groetzner:2020:FMC

[GD20] Patrick Groetzner and Mirjam Dür. A factorization method for completely positive matrices. Linear Algebra and its Applications, 591(??):1–24, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305385.

Gargate:2023:NAL

[GdM23] Ivan Gonzales Gargate and Thiago Castilho de Mello. A new approach to the Lvov–Kaplansky conjecture through gradings. Linear Algebra and its Applications, 674(??):453–465, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002355.

Gerstein:2020:IQF

[Ger20] Larry J. Gerstein. Integral quadratic forms and graphs. Linear Algebra and its Applications, 585(??):60–70, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304215.

Gnang:2020:BMR

[GF20] Edinah K. Gnang and Yuval Filmus. On the Bhattacharya-Mesner rank of third order hypermatrices. Linear Algebra and its Applications, 588(??):391–418, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304999.

Guimaraes:2021:GIC

[GFD21] Alan Guimarães, Claudemir Fidelis, and Laise Dias. Z_{q}- graded identities and central polynomials of the Grassmann algebra. Linear Algebra and its Applications, 609(??):12–36, January 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303876.

Gu:2024:HAB

[GFW24] Jiaqi Gu, Shenghao Feng, and Yimin Wei. Hypergraph analysis based on a compatible tensor product structure. Linear Algebra and its Applications, 682(??):122–151, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004172.

Gaal:2021:SLG

[GG21] Marcell Gaál and Robert M. Guralnick. Simple Lie groups stabilizing G-invariant norms. Linear Algebra and its Applications, 609(??):308–316, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304341.

Gallardo-Gutierrez:2023:BDO

[GGGD23] Eva A. Gallardo-Gutiérrez and F. Javier González-Doña. Band-diagonal operators on Banach lattices: Matrix dynamics and invariant subspaces. Linear Algebra and its Applications, 674(??):155–191, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002008.

Gluck:2023:ECI

[GH23] Jochen Glück and Julian Hölz. Eventual cone invariance revisited. Linear Algebra and its Applications, 675(??):274–293, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002392.

Ghodrati:2024:SLS

[GH24] Amir Hossein Ghodrati and Mohammad Ali Hosseinzadeh. Signless Laplacian spectrum of a graph. Linear Algebra and its Applications, 682(??):257–267, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004214.

Ghiglioni:2020:NDA

[Ghi20] Eduardo M. Ghiglioni. New deterministic approaches to the least square mean. Linear Algebra and its Applications, 585(??):50–59, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304227.

Gao:2024:ISP

[GHL^{+}24] Yubin Gao, Frank J. Hall, Zhongshan Li, Victor Bailey, and Paul Kim. 4×4 Irreducible sign pattern matrices that require four distinct eigenvalues. Linear Algebra and its Applications, 680(??):1–27, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003701.

Greaves:2022:FFF

[GIJM22] Gary R. W. Greaves, Joseph W. Iverson, John Jasper, and Dustin G. Mixon. Frames over finite fields: Equiangular lines in orthogonal geometry. Linear Algebra and its Applications, 639(??):50–80, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000027.

Gil:2020:PDM

[Gil20] Michael Gil’. Perturbations of determinants of matrices. Linear Algebra and its Applications, 590(??):235–242, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300021.

Gollin:2023:MVE

[GJ23] J. Pascal Gollin and Attila Joó. Matching variables to equations in infinite linear equation systems. Linear Algebra and its Applications, 660(??):40–46, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004293.

Guterman:2021:LFP

[GJKM21] Alexander Guterman, Marianne Johnson, Mark Kambites, and Artem Maksaev. Linear functions preserving Green’s relations over fields. Linear Algebra and its Applications, 611 (??):310–333, February 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305164.

Ginsburg:2020:MPL

[GJV20] Victor Ginsburg, Hayden Julius, and Ricardo Velasquez. On maps preserving Lie products equal to a rank-one nilpotent. Linear Algebra and its Applications, 593(??):212–227, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300665.

Gnacik:2020:IPS

[GK20] Michał Gnacik and Tomasz Kania. Inverse problems for symmetric doubly stochastic matrices whose Suleimanova spectra are bounded below by 1/2. Linear Algebra and its Applications, 592(??):175–187, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300392.

Glover:2021:SSP

[GK21] Cory Glover and Mark Kempton. Some spectral properties of the non-backtracking matrix of a graph. Linear Algebra and its Applications, 618(??):37–57, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000410.

Garces:2022:PPI

[GK22a] Jorge J. Garcés and Mykola Khrypchenko. Potent preservers of incidence algebras. Linear Algebra and its Applications, 635(??):171–200, February 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379521004262.

Ghorbani:2022:SCS

[GK22b] Ebrahim Ghorbani and Masoumeh Koohestani. Spectral classes of strongly-regular and distance-regular graphs. Linear Algebra and its Applications, 641(??):182–199, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000623.

Goryainov:2020:PES

[GKK^{+}20] Sergey Goryainov, Vladislav Kabanov, Elena Konstantinova, Leonid Shalaginov, and Alexandr Valyuzhenich. PIeigenfunctions of the Star graphs. Linear Algebra and its Applications, 586(??):7– 27, February 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379519304525.

Goryainov:2020:IGO

[GKLZ20] Sergey Goryainov, Elena V. Konstantinova, Honghai Li, and Da Zhao. Integral graphs obtained by dual Seidel switching. Linear Algebra and its Applications, 604(??):476–489, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303335.

Girard:2021:CCM

[GKS21] Mark Girard, Seung-Hyeok Kye, and Erling Størmer. Convex cones in mapping spaces between matrix algebras. Linear Algebra and its Applications, 608(??):248–269, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304298.

Gluesing-Luerssen:2020:SCL

[GL20a] Heide Gluesing-Luerssen. On the sparseness of certain linear MRD codes. Linear Algebra and its Applications, 596(??):145–168, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301208.

Godsil:2020:PST

[GL20b] Chris Godsil and Sabrina Lato. Perfect state transfer on oriented graphs. Linear Algebra and its Applications, 604(??):278–292, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303165.

Girard:2021:TCM

[GL21] Mark Girard and Jeremy Levick. Twirling channels have minimal mixed-unitary rank. Linear Algebra and its Applications, 615(??):207–227, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520306005.

Ghiglioni:2022:HKM

[GL22a] Eduardo Ghiglioni and Yongdo Lim. Hyperbolicity of the Karcher mean. Linear Algebra and its Applications, 643(??):196–217, June 15, 2022. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000611.

Gongopadhyay:2022:RHI

[GL22b] Krishnendu Gongopadhyay and Tejbir Lohan. Reversibility of Hermitian isometries. Linear Algebra and its Applications, 639(??):159–176, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000106.

Ghiglioni:2023:KME

[GL23] Eduardo Ghiglioni and Yongdo Lim. The Karcher mean of equilateral triangles. Linear Algebra and its Applications, 656(??):421–445, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003731.

Garcia:2023:GIN

[GLAdS23] Esther García, Miguel Gómez Lozano, Rubén Muñoz Alcázar, and Guillermo Vera de Salas. Gradings induced by nilpotent elements. Linear Algebra and its Applications, 656(??):92–111, January 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522003421.

Ghorbani:2020:BAS

[GLHNW20] Modjtaba Ghorbani, Xueliang Li, Mardjan Hakimi-Nezhaad, and Junming Wang. Bounds on the ABC spectral radius and ABC energy of graphs. Linear Algebra and its Applications, 598(??):145–164, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301786.

Ge:2020:SCL

[GLLZ20] Jianquan Ge, Fagui Li, Zhiqin Lu, and Yi Zhou. On some conjectures by Lu and Wenzel. Linear Algebra and its Applications, 592(??):134–164, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300379.

Guo:2023:SRR

[GLMM23] Mingyang Guo, Hongliang Lu, Xinxin Ma, and Xiao Ma. Spectral radius and rainbow matchings of graphs. Linear Algebra and its Applications, 679(??):30–37, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003361.

Gaddis:2024:FVQ

[GLNW24] Jason Gaddis, Thomas Lamkin, Thy Nguyen, and Caleb Wright. Four-vertex quivers supporting twisted graded Calabi–Yau algebras. Linear Algebra and its Applications, 689(??):86–107, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000752.

Ghiglioni:2021:KML

[GLP21] Eduardo Ghiglioni, Yongdo Lim, and Miklós Pálfia. The Karcher mean of linearly independent triples. Linear Algebra and its Applications, 610(??):203–221, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304699.

Gribling:2022:BSR

[GLS22] Sander Gribling, Monique Laurent, and Andries Steenkamp. Bounding the separable rank via polynomial optimization. Linear Algebra and its Applications, 648(??):1–55, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001616.

Gan:2021:TGM

[GLT21] Luyining Gan, Xuhua Liu, and Tin-Yau Tam. On two geometric means and sum of adjoint orbits. Linear Algebra and its Applications, 631(??):156–173, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003268.

Gupta:2023:EIC

[GLT23] Somi Gupta, Giovanni Longobardi, and Rocco Trombetti. On the equivalence issue of a class of 2dimensional linear maximum rank distance codes. Linear Algebra and its Applications, 678(??):33–56, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300318X.

Gao:2020:SCT

[GLZ20] Rui Gao, Heguo Liu, and Feng Zhou. Symmetry classes of tensors associated with certain groups. Linear Algebra and its Applications, 602(??):240–251, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302548.

Guo:2021:SCE

[GLZ21] Hangtian Guo, Huiqiu Lin, and Yanhua Zhao. A spectral condition for the existence of a pentagon in non-bipartite graphs. Linear Algebra and its Applications, 627(??):140–149, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002378.

Gnang:2020:SAN

[GM20] Edinah K. Gnang and James M. Murphy. Spectral analysis of non-Hermitian matrices and directed graphs. Linear Algebra and its Applications, 604(??):72–91, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302998.

Ganesh:2022:TMW

[GM22] Swetha Ganesh and Sumit Mohanty. Trees with matrix weights: Laplacian matrix and characteristic-like vertices. Linear Algebra and its Applications, 646(??):195–237, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001379.

Gavrilyuk:2023:MDW

[GM23] Alexander L. Gavrilyuk and Akihiro Munemasa. Maximal digraphs whose Hermitian spectral radius is at most 2. Linear Algebra and its Applications, 658 (??):331–349, February 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004074.

Guterman:2022:PMP

[GMP22] A. E. Guterman, A. M. Maksaev, and V. V. Promyslov. Pairs of maps preserving singularity on subsets of matrix algebras. Linear Algebra and its Applications, 644(??):1–27, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000854.

Garcia:2021:LBS

[GMR21] Jazmín García, Juan Monsalve, and Juan Rada. Lower bounds for the spectral norm of digraphs. Linear Algebra and its Applications, 617(??):151–167, May 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000197.

Gutman:2022:RBV

[GMR22] Ivan Gutman, Juan Monsalve, and Juan Rada. A relation between a vertex-degree-based topological index and its energy. Linear Algebra and its Applications, 636(??):134–142, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004274.

Grudsky:2022:ELM

[GMSG22] Sergei M. Grudsky, Egor A. Maximenko, and Alejandro Soto-González. Eigenvalues of the Laplacian matrices of the cycles with one weighted edge. Linear Algebra and its Applications, 653(??):86–115, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002750.

Gaubert:2020:TPN

[GN20] Stéphane Gaubert and Adi Niv. Tropical planar networks. Linear Algebra and its Applications, 595(??):123–144, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300914.

Glasby:2022:RGD

[GNP22] Stephen P. Glasby, Alice C. Niemeyer, and Cheryl E. Praeger. Random generation of direct sums of finite nondegenerate subspaces. Linear Algebra and its Applications, 649(??):408–432, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002038.

Goulart:2020:AAP

[GNR20] Paul J. Goulart, Yuji Nakatsukasa, and Nikitas Rontsis. Accuracy of approximate projection to the semidefinite cone. Linear Algebra and its Applications, 594(??):177–192, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030080X.

Godsil:2021:SQW

[God21]  Chris Godsil.  Sedentary

quantum walks. Linear Algebra and its Applications, 614(??):356–375, April 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304067.

Goel:2021:DMH

[Goe21] Shivani Goel. On distance matrices of helm graphs obtained from wheel graphs with an even number of vertices. Linear Algebra and its Applications, 621(??):86–104, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001051.

Golubev:2020:GDE

[Gol20] Konstantin Golubev. Graphical designs and extremal combinatorics. Linear Algebra and its Applications, 604(??):490–506, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303359.

Gordienko:2024:GGA

[Gor24] A. S. Gordienko. Graded group actions and generalized H-actions compatible with gradings. Linear Algebra and its Applications, 682(??):96–121, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004196.

Gotti:2020:GCA

[Got20] Felix Gotti. Geometric and combinatorial aspects of submonoids of a finite-rank free commutative monoid. Linear Algebra and its Applications, 604(??):146–186, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303001.

Golla:2023:SPI

[GOUY23] Ramesh Golla, Hiroyuki Osaka, Yoichi Udagawa, and Takeaki Yamazaki. Stability of AN -property for the induced Aluthge transformations. Linear Algebra and its Applications, 678 (??):206–226, December 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300321X.

Gowda:2020:SMI

[Gow20] M. Seetharama Gowda. Some majorization inequalities induced by Schur products in Euclidean Jordan algebras. Linear Algebra and its Applications, 600(??):1–21, September 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301841.

Galazka:2021:WTI

[GOZ21] Tomasz Gała̧zka, Adam Osȩkowski, and Yahui Zuo. A weaktype (∞, ∞) inequality for the triangular projection. Linear Algebra and its Applications, 612(??):112–127, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305619.

Guo:2020:LBD

[GP20] Shuzheng Guo and Daxiong Piao. Lyapunov behavior and dynamical localization for quasi-periodic CMV matrices. Linear Algebra and its Applications, 606(??):68–89, December 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303542.

Giscard:2021:TSC

[GP21] Pierre-Louis Giscard and Stefano Pozza. Tridiagonalization of systems of coupled linear differential equations with variable coefficients by a Lanczos-like method. Linear Algebra and its Applications, 624(??):153–173, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001683.

Grover:2020:PPS

[GPR20] Priyanka Grover, Veer Singh Panwar, and A. Satyanarayana Reddy. Positivity properties of some special matrices. Linear Algebra and its Applications, 596(??):203–215, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301294.

Ganie:2020:FDB

[GPRT20] Hilal A. Ganie, S. Pirzada, Bilal A. Rather, and Vilmar Trevisan. Further developments on Brouwer’s conjecture for the sum of Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 588(??):1–18, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304963.

Guo:2020:GRP

[GQZ20] Wenxing Guo, Shanshan Qin, and Zhiwen Zhao. Generalized ridge and principal correlation estimator of the regression coefficient in growth curve model. Linear Algebra and its Applications, 591(??):115–133, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300100.

Gatephan:2021:DLM

[GR21] Peeraphat Gatephan and Kijti Rodtes. Divisibility of LCM matrices by totally nonnegative GCD matrices. Linear Algebra and its Applications, 619(??):146–159, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000793.

Gernandt:2022:CTI

[GR22] Hannes Gernandt and Jonathan Rohleder. A Calderón type inverse problem for tree graphs. Linear Algebra and its Applications, 646(??):29–42, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001197.

Grilliette:2022:IHI

[GRR22] Will Grilliette, Josephine Reynes, and Lucas J. Rusnak. Incidence hypergraphs: Injectivity, uniformity, and matrix-tree theorems. Linear Algebra and its Applications, 634(??):77–105, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003888.

Gowda:2020:RTT

[GS20a] M. Seetharama Gowda and Roman Sznajder. A Riesz– Thorin type interpolation theorem in Euclidean Jordan algebras. Linear Algebra and its Applications, 585(??):178–190, January 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304185.

Guterman:2020:PPP

[GS20b] A. E. Guterman and I. A. Spiridonov. Permanent Polya problem for additive surjective maps. Linear Algebra and its Applications, 599(??):140–155, August 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030183X.

Gaal:2021:SIS

[GS21a] Marcell Gaál and Brigitta Szilágyi. c-spectral isometries of skew-symmetric matrices: the eight dimensional case. Linear Algebra and its Applications, 609(??):108–113, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304109.

Gillis:2021:MNS

[GS21b] Nicolas Gillis and Punit Sharma. Minimal-norm static feedbacks using dissipative Hamiltonian matrices. Linear Algebra and its Applications, 623(??):258–281, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300677.

Guterman:2021:LCW

[GS21c] Alexander Guterman and Pavel Shteyner. Linear converters of weak, directional and strong majorizations. Linear Algebra and its Applications, 613(??):320–346, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305371.

Grover:2022:DFT

[GS22] Priyanka Grover and Sushil Singla. A distance formula for tuples of operators. Linear Algebra and its Applications, 650(??):267–285, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200218X.

Gavrilyuk:2023:MDE

[GS23a] Alexander L. Gavrilyuk and Sho Suda. On the multiplicities of digraph eigenvalues. Linear Algebra and its Applications, 673 (??):101–137, September 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001416.

Guterman:2023:LOP

[GS23b] Alexander Guterman and Pavel Shteyner. Linear operators preserving strong majorization of (0, 1)-matrices. Linear Algebra and its Applications, 658(??):116–150, February 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379522003937.

Guterman:2024:LPP

[GS24] A. E. Guterman and I. A. Spiridonov. On linear preservers of permanental rank. Linear Algebra and its Applications, 680(??):325–340, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003968.

Gao:2024:OSS

[GSS24] Bin Gao, Nguyen Thanh Son, and Tatjana Stykel. Optimization on the symplectic Stiefel manifold: SR decomposition-based retraction and applications. Linear Algebra and its Applications, 682(??):50–85, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004056.

Gori:2024:ZSN

[GSV24] Anna Gori, Giulia Sarfatti, and Fabio Vlacci. Zero sets and Nullstellensatz type theorems for slice regular quaternionic polynomials. Linear Algebra and its Applications, 685(??):162–181, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000053.

Granario:2021:PPD

[GT21] Daryl Q. Granario and Tin-Yau Tam. Products of positive definite symplectic matrices. Linear Algebra and its Applications, 626(??):188–203, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100224X.

Gardella:2023:ZPB

[GT23] Eusebio Gardella and Hannes Thiel. Zero-product balanced algebras. Linear Algebra and its Applications, 670(??):121–153, August 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001490.

Garvey:2023:PSP

[GTW23] Seamus D. Garvey, Françoise Tisseur, and Shujuan Wang. A parametrization of structurepreserving transformations for matrix polynomials. Linear Algebra and its Applications, 674(??):68–90, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002070.

Guo:2022:DTF

[Guo22] Xuejun Guo. Determinants of trigonometric functions and class numbers. Linear Algebra and its Applications, 653(??):33–43, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002786.

Groenewald:2021:ACF

[GvRR21a] G. J. Groenewald, D. B. Janse van Rensburg, and A. C. M. Ran. An alternative canonical form for quaternionic H-unitary matrices. Linear Algebra and its Applications, 623(??):282–298, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301932.

Groenewald:2021:RSM

[GvRR^{+}21b] G. J. Groenewald, D. B. Janse van Rensburg, A. C. M. Ran, F. Theron, and M. van Straaten. m-th roots of Hselfadjoint matrices. Linear Algebra and its Applications, 610(??):804–826, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304973.

Groenewald:2023:SME

[GvRR^{+}23] G. J. Groenewald, D. B. Janse van Rensburg, A. C. M. Ran, F. Theron, and M. van Straaten. Solutions of the matrix equation p(X) = A, with polynomial function p(λ) over field extensions of Q. Linear Algebra and its Applications, 665(??):107–138, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000344.

Gau:2020:MPC

[GW20] Hwa-Long Gau and Kuo-Zhong Wang. Matrix powers with circular numerical range. Linear Algebra and its Applications, 603(??):190–211, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520302834.

Gau:2022:NRH

[GW22] Hwa-Long Gau and Pei Yuan Wu. Numerical ranges of Hankel matrices. Linear Algebra and its Applications, 650(??):60–74, October 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002105.

Gift:2024:RFP

[GW24] Sarah Gift and Hugo J. Woerdeman. Real factorization of positive semidefinite matrix polynomials. Linear Algebra and its Applications, 683(??):125–150, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300455X.

Gau:2021:NRF

[GWW21] Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Numerical ranges of Foguel operators. Linear Algebra and its Applications, 610 (??):766–784, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304948.

Gyenti:2023:PPC

[Gye23] Bálint Gyenti. On a preserver problem concerning Kubo– Ando means of positive operators. Linear Algebra and its Applications, 665(??):404–419, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000563.

Golyandina:2020:BDC

[GZ20a] Nina Golyandina and Anatoly Zhigljavsky. Blind deconvolution of covariance matrix inverses for autoregressive processes. Linear Algebra and its Applications, 593(??):188–211, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300641.

Guglielmi:2020:ATS

[GZ20b] Nicola Guglielmi and Marino Zennaro. An antinorm theory for sets of matrices: Bounds and approximations to the lower spectral radius. Linear Algebra and its Applications, 607(??):89–117, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303761.

Guo:2020:GOB

[GZ20c] Wan-Ming Guo and Bao-Xuan Zhu. A generalized ordered Bell polynomial. Linear Algebra and its Applications, 588(??):458–470, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305130.

Godsil:2022:FRN

[GZ22] Chris Godsil and Xiaohong Zhang. Fractional revival on non-cospectral vertices. Linear Algebra and its Applications, 654(??):69–88, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003081.

Galvin:2023:TNN

[GZ23] David Galvin and Yufei Zhang. Totally non-negativity of a family of change-of-basis matrices. Linear Algebra and its Applications, 676(??):88–103, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002690.

Gong:2021:EGH

[GZZ21] Shi-Cai Gong, Peng Zou, and Xiao-Dong Zhang. Each (n, m)-graph having the i-th minimal Laplacian coefficient is a threshold graph. Linear Algebra and its Applications, 631(??):398–406, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003463.

Hamedani:2019:EHH

[HAD19] Ali Memar Hamedani, Rahim Alizadeh, and Silvestru Sever Dragomir. An extension of Haagerup–Harpe inequality. Linear Algebra and its Applications, 577(??):69– 77, September 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951930179X.

See corrigendum [Ali20].

Haemers:2021:HRB

[Hae21] Willem H. Haemers. Hoffman’s ratio bound. Linear Algebra and its Applications, 617(??):215–219, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000689.

Hallman:2020:EBE

[Hal20] Eric Hallman. Estimating the backward error for the leastsquares problem with multiple right-hand sides. Linear Algebra and its Applications, 605(??):227–238, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303438.

Hamhalter:2022:SPR

[Ham22] J. Hamhalter. Structure of preservers of range orthogonality on ⋆-rings and C^{⋆}-algebras. Linear Algebra and its Applications, 642(??):139–159, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000751.

Hamm:2023:GPD

[Ham23]  Keaton Hamm.  General-

ized pseudoskeleton decompositions. Linear Algebra and its Applications, 664(??):236–252, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000253.

Hansen:2021:EMW

[Han21] Jakob Hansen. Expansion in matrix-weighted graphs. Linear Algebra and its Applications, 630(??):252–273, December 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521003062.

Hansen:2022:NPS

[Han22]  Frank Hansen.  A note

on the parallel sum. Linear Algebra and its Applications, 636(??):69–76, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004195.

Hari:2022:QCC

[Har22] Vjeran Hari. On the quadratic convergence of the complex HZ method for the positive definite generalized eigenvalue problem. Linear Algebra and its Applications, 632(??):153–192, January 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003190.

han:2021:QSP

[hB21] Ayşe Güven Sarıhan and Oscar F. Bandtlow. Quantitative spectral perturbation theory for compact operators on a Hilbert space. Linear Algebra and its Applications, 610(??):169–202, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304122.

Hu:2022:SRM

[HBHZ22] Dan Hu, Hajo Broersma, Jiangyou Hou, and Shenggui Zhang. The spectra of random mixed graphs. Linear Algebra and its Applications, 653(??):320–338, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002968.

Hu:2024:SSR

[HBHZ24] Dan Hu, Hajo Broersma, Jiangyou Hou, and Shenggui Zhang. The skew spectral radius and skew Randić spectral radius of general random oriented graphs. Linear Algebra and its Applications, 685(??):125–137, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952400003X.

Hou:2020:HPA

[HCZ20] Yuan Hou, An Chang, and Lei Zhang. A homogeneous polynomial associated with general hypergraphs and its applications. Linear Algebra and its Applications, 591(??):72–86, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300057.

Huang:2022:PCC

[HD22] Xueyi Huang and Kinkar Chandra Das. Proof of a conjecture on communicability distance sum index of graphs. Linear Algebra and its Applications, 645(??):278–292, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001355.

He:2020:HRA

[He20] Tian-Xiao He. Half Riordan array sequences. Linear Algebra and its Applications, 604(??):236–264, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303104.

Herbrich:2020:ZED

[Her20] Peter Herbrich. Zeta-equivalent digraphs: Simultaneous cospectrality. Linear Algebra and its Applications, 602 (??):281–299, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302214.

Havelkova:2023:IHR

[HH23a] Eva Havelková and Iveta Hnětynková. Iterative hybrid regularization for extremely noisy full models in single particle analysis. Linear Algebra and its Applications, 656(??):131–157, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003457.

Hu:2023:LMN

[HH23b] Yanan Hu and Zejun Huang. Linear maps on nonnegative symmetric matrices preserving the independence number. Linear Algebra and its Applications, 670(??):154–167, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001441.

He:2020:RCU

[HHD20] Shengjie He, Rong-Xia Hao, and Fengming Dong. The rank of a complex unit gain graph in terms of the matching number. Linear Algebra and its Applications, 589(??):158–185, March 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437951930521X.

Halbeisen:2020:SMR

[HHS20a] Lorenz Halbeisen, Norbert Hungerbühler, and Salome Schumacher. Sets and multisets of range uniqueness for polynomials. Linear Algebra and its Applications, 589(??):39–61, March 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305221.

Hunnell:2020:ITF

[HHS20b] Mark Hunnell, John Hutchens, and Nathaniel Schwartz. On involutions of type O(q, k) over a field of characteristic two. Linear Algebra and its Applications, 593(??):228–250, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300653.

Halbeisen:2021:MSP

[HHS21] Lorenz Halbeisen, Norbert Hungerbühler, and Salome Schumacher. Magic sets for polynomials of degree n. Linear Algebra and its Applications, 609(??):413–441, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304481.

Hu:2024:PSS

[HJ24] ChaoLong Hu and YouQing Ji. Power set of some quasinilpotent weighted shifts on ℓ^{p}. Linear Algebra and its Applications, 686(??):111–133, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000193.

Hare:2024:UOS

[HJBKR24] Warren Hare, Gabriel Jarry-Bolduc, Sébastien Kerleau, and Clément W. Royer. Using orthogonally structured positive bases for constructing positive k-spanning sets with cosine measure guarantees. Linear Algebra and its Applications, 680(??):183–207, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003798.

Hwang:2020:TPH

[HK20] Jinmi Hwang and Sejong Kim. Tensor product and Hadamard product for the Wasserstein means. Linear Algebra and its Applications, 603(??):496–507, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303232.

Huang:2023:RDM

[HKS23] Jinghao Huang, Karimbergen Kudaybergenov, and Fedor Sukochev. Ring derivations of Murray–von Neumann algebras. Linear Algebra and its Applications, 672(??):28–52, September 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001453.

Huang:2020:IM

[HL20] Zejun Huang and Huiqiu Lin. On k-idempotent 0–1 matrices. Linear Algebra and its Applications, 597(??):187–197, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301488.

Hu:2023:DCM

[HLH23] Yarong Hu, Zhenzhen Lou, and Qiongxiang Huang. Disproof of a conjecture on the minimum spectral radius and the domination number. Linear Algebra and its Applications, 677 (??):237–253, November 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003038.

Hu:2024:MGG

[HLN24] Yarong Hu, Zhenzhen Lou, and Wenjie Ning. The Qminimizer graph with given independence number. Linear Algebra and its Applications, 685(??):1–23, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004664.

Liu:2021:SET

[hLqJ21] Wei hui Liu and Xiao qing Jin. A study on T-eigenvalues of third-order tensors. Linear Algebra and its Applications, 612(??):357–374, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305292.

Higham:2021:IMF

[HLS21a] Nicholas J. Higham, Matthew C. Lettington, and Karl Michael Schmidt. Integer matrix factorisations, superalgebras and the quadratic form obstruction. Linear Algebra and its Applications, 622(??):250–267, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001324.

Huang:2021:MSL

[HLS21b] Peng Huang, Jianxi Li, and Wai Chee Shiu. Maximizing the signless Laplacian spectral radius of k-connected graphs with given diameter. Linear Algebra and its Applications, 617(??):78–99, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000343.

Huang:2022:MSR

[HLS22] Peng Huang, Jianxi Li, and Wai Chee Shiu. Maximizing the A_{α}-spectral radius of graphs with given size and diameter. Linear Algebra and its Applications, 651(??):116–130, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002440.

Huang:2024:CSL

[HLS24] Zejun Huang, Chi-Kwong Li, and Nung-Sing Sze. Constructing a straight line intersecting four lines. Linear Algebra and its Applications, 683(??):201–210, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004445.

He:2023:OSA

[HLW23] Yunhui He, Jun Liu, and Xiang-Sheng Wang. Optimized sparse approximate inverse smoothers for solving Laplacian linear systems. Linear Algebra and its Applications, 656(??):304–323, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003561.

Hong:2020:CEG

[HLX20] Zhen-Mu Hong, Hong-Jian Lai, and Zheng-Jiang Xia. Connectivity and eigenvalues of graphs with given girth or clique number. Linear Algebra and its Applications, 607(??):319–340, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303888.

Huang:2021:DHD

[HLYZ21] Hua-Lin Huang, Huajun Lu, Yu Ye, and Chi Zhang. Diagonalizable higher degree forms and symmetric tensors. Linear Algebra and its Applications, 613(??):151–169, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305851.

Huang:2023:CMF

[HLYZ23] Hua-Lin Huang, Huajun Lu, Yu Ye, and Chi Zhang. Centers of multilinear forms and applications. Linear Algebra and its Applications, 673(??):160–176, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001891.

Hou:2017:EIM

[HLZ17] Xin Hou, Shangzhi Li, and Qingqing Zheng. Expressing infinite matrices over rings as products of involutions. Linear Algebra and its Applications, 532(??):257–265, November 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379517304111. See cor-

rigendum [Hou20].

Heuer:2020:BSR

[HMPT20] Janin Heuer, Frederic Matter, Marc E. Pfetsch, and Thorsten Theobald. Blocksparse recovery of semidefinite systems and generalized null space conditions. Linear Algebra and its Applications, 603(??):470–495, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303190.

Huang:2021:WFO

[HMS21] Jiahui Huang, David McKinnon, and Matthew Satriano. What fraction of an S_{n}- orbit can lie on a hyperplane?

Linear Algebra and its Applications, 613(??):1–23, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305656.

Hertrich:2021:CPN

[HNS21] Johannes Hertrich, Sebastian Neumayer, and Gabriele Steidl. Convolutional proximal neural networks and Plugand-Play algorithms. Linear Algebra and its Applications, 631(??):203–234, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003372.

Haemers:2020:USD

[HO20] Willem H. Haemers and Mohammad Reza Oboudi. Universal spectra of the disjoint union of regular graphs. Linear Algebra and its Applications, 606 (??):244–248, December 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303621.

Homma:2020:FPF

[Hom20] Masaaki Homma. Fragments of plane filling curves of degree q + 2 over the finite field of q elements, and of affine-plane filling curves of degree q + 1. Linear Algebra and its Applications, 589(??):9–27, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305191.

Hou:2020:CEI

[Hou20] Xin Hou. Corrigendum to “Expressing infinite matrices over rings as products of involutions” [Linear Algebra Appl. 532 (2017) 257–265]. Linear Algebra and its Applications, 587(??):387–391, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304896.  See

[HLZ17].

Howlader:2022:DSM

[HP22] Aditi Howlader and Pratima Panigrahi. On the distance spectrum of minimal cages and associated distance biregular graphs. Linear Algebra and its Applications, 636(??):115–133, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004201.

Haviv:2023:BRC

[HP23a] Ishay Haviv and Michal Parnas. The binary rank of circulant block matrices. Linear Algebra and its Applications, 656(??):277–303, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003652.

Hosseinzadeh:2023:MPA

[HP23b] Roja Hosseinzadeh and Tatjana Petek. Maps preserving the ascent and descent of product of operators. Linear Algebra and its Applications, 671(??):109–120, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001337.

Huang:2021:NLP

[HPSZ21] Yanhe Huang, George Petroulakis, Frank Sottile, and Igor Zelenko. Nontrivial linear projections on the Grassmannian Gr_{3}(C^{6}). Linear Algebra and its Applications, 610(??):488–505, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304778.

Hadwin:2020:NMT

[HR20] Don Hadwin and Heydar Radjavi. A note on Mirsky’s theorem. Linear Algebra and its Applications, 603 (??):186–189, October 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030269X.

He:2020:PPI

[HS20] Tian-Xiao He and Louis Shapiro. Palindromes and pseudo-involution multiplication. Linear Algebra and its Applications, 593(??):1–17, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300410.

Hofmann:2022:ECM

[HS22] Tobias Hofmann and Uwe Schwerdtfeger. Edge-connectivity matrices and their spectra. Linear Algebra and its Applications, 640(??):34–47, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000209.

Higashitani:2023:AMF

[HS23a] Akihiro Higashitani and Yuya Sugishita. Adjacency matrices over a finite prime field and their direct sum decompositions. Linear Algebra and its Applications, 677(??):306–322, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003063.

Hisabia:2023:AGP

[HS23b] Aritra Narayan Hisabia and Manideepa Saha. Algebraic and geometric properties of L_{+}^{n}-semipositive matrices and L_{+}^{n}-semipositive cones. Linear Algebra and its Applications, 672(??):210–227, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001829.

Hameed:2021:SDS

[HSS^{+}21] Shahul K. Hameed, T. V. Shijin, P. Soorya, K. A. Germina, and Thomas Zaslavsky. Signed distance in signed graphs. Linear Algebra and its Applications, 608(??):236–247, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303979.

Hallman:2022:MAS

[HT22] Eric Hallman and Devon Troester. A multilevel approach to stochastic trace estimation. Linear Algebra and its Applications, 638(??):125–149, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Haemers:2023:SGM

[HT23] Willem H. Haemers and Hatice Topcu. On signed graphs with at most two eigenvalues unequal to ±1. Linear Algebra and its Applications, 670(??):68–77, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001283.

Hu:2021:EEC

[Hu21] Zhiguang Hu. Eigenvalues and eigenvectors of a class of irreducible tridiagonal matrices. Linear Algebra and its Applications, 619(??):328–337, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001117.

Huang:2022:QUC

[Hua22] Jing Huang. On the quadratic unitary Cayley graphs. Linear Algebra and its Applications, 644(??):219–233, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000799.

Huy:2023:SGR

[HVX23] Duong Quoc Huy, Doan Thi Thuy Van, and Dinh Thi Xinh. Some generalizations of real power form for Young-type inequalities and their applications. Linear Algebra and its Applications, 656(??):368–384, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003585.

Henriksen:2020:CIR

[HW20] Amelia Henriksen and Rachel Ward. Concentration inequalities for random matrix products. Linear Algebra and its Applications, 594(??):81–94, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300574.

Herr:2021:ICH

[HW21a] John E. Herr and Troy M. Wiegand. Identifying complex Hadamard submatrices of the Fourier matrices via primitive sets. Linear Algebra and its Applications, 617(??):121–150, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000367.

Hughes:2021:SDS

[HW21b] Daniel Hughes and Shayne Waldron. Spherical (t, t)- designs with a small number of vectors. Linear Algebra and its Applications, 608(??):84–106, January 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303839.

Hu:2023:SCE

[HWD23] Yang Hu, Ligong Wang, and Cunxiang Duan. Spectral conditions for edge connectivity and spanning tree packing number in (multi-)graphs. Linear Algebra and its Applications, 664(??):324–348, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000289.

Hauffe-Waschbusch:2021:CSO

[HWK21] Adrian Hauffe-Waschbüsch and Aloys Krieg. Congruence subgroups and orthogonal groups. Linear Algebra and its Applications, 618(??):22–36, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000227.

Han:2024:CSH

[HX24] Lixing Han and Jianhong Xu. On classification of states in higher order Markov chains. Linear Algebra and its Applications, 685(??):24–45, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004676.

Hong:2021:FAS

[HXLL21] Zhen-Mu Hong, Zheng-Jiang Xia, Hong-Jian Lai, and Ruifang Liu. Fractional arboricity, strength and eigenvalues of graphs with fixed girth or clique number. Linear Algebra and its Applications, 611(??):135–147, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305000.

Horn:2020:RHP

[HY20] Roger A. Horn and Zai Yang. Rank of a Hadamard product. Linear Algebra and its Applications, 591(??):87–98, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300094.

Howell:2024:CPP

[HY24] Kate Howell and Rongwei Yang. The characteristic polynomial of projections. Linear Algebra and its Applications, 680(??):170–182, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003774.

Inghelbrecht:2022:PHP

[IB22] Gilles Inghelbrecht and Kurt Barbé. Parallelization of Hermitian positive definite systems of equations: a hierarchical Jacobi approach. Linear Algebra and its Applications, 639(??):81–115, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000052.

Ioppolo:2021:SGI

[IdSSV21] A. Ioppolo, R. B. dos Santos, M. L. O. Santos, and A. C. Vieira. Superalgebras with graded involution: Classifying minimal varieties of quadratic growth. Linear Algebra and its Applications, 621(??):105–134, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001087.

Idel:2017:PBW

[IGW17] Martin Idel, Sebastián Soto Gaona, and Michael M.

Wolf. Perturbation bounds for Williamson’s symplectic normal form. Linear Algebra and its Applications, 525(??):45–58, July 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379517301751. See note

[SZ23a].

Ide:2021:IEG

[IIM^{+}21] Yusuke Ide, Ayaka Ishikawa, Hideaki Morita, Iwao Sato, and Etsuo Segwa. The Ihara expression for the generalized weighted zeta function of a finite simple graph. Linear Algebra and its Applications, 627(??):227–241, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002433.

Ilisevic:2020:CRB

[IKLP20] Dijana Ilisević, Bojan Kuzma, Chi-Kwong Li, and Edward Poon. Complexifications of real Banach spaces and their isometries. Linear Algebra and its Applications, 589(??):222–241, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305208.

Iqbal:2020:DRG

[IKPR20] Quaid Iqbal, Jack H. Koolen, Jongyook Park, and Masood Ur Rehman. Distanceregular graphs with diameter 3 and eigenvalue a_{2} − c_{3}. Linear Algebra and its Applications, 587(??):271–290, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304616.

Imekraz:2023:CSS

[Ime23] Rafik Imekraz. Characterizations of seminorms satisfying some inequalities with respect to a finite-dimensional subspace. Linear Algebra and its Applications, 672 (??):251–270, September 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001817.

Ito:2020:PGW

[IMT20] Naoharu Ito, Toyoki Matsuyama, and Tatsuya Tsurii. Periodicity of Grover walks on complete graphs with self-loops. Linear Algebra and its Applications, 599(??):121–132, August 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520301853.

Ioppolo:2020:SRC

[Iop20] Antonio Ioppolo. Some results concerning the multiplicities of cocharacters of superalgebras with graded involution. Linear Algebra and its Applications, 594(??):51–70, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300811.

Ilisevic:2021:PIB

[IOT21] Dijana Ilišević, Matjaž Omladič, and Aleksej Turnšek. Phase-isometries between normed spaces. Linear Algebra and its Applications, 612(??):99–111, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305590.

Jain:2021:SPS

[Jai21] Tanvi Jain. Sums and products of symplectic eigenvalues. Linear Algebra and its Applications, 631 (??):67–82, December 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100313X.

Julio:2024:APN

[JCD24] Ana I. Julio, Valeria Collants, and Roberto C. Díaz. About the persymmetric nonnegative inverse eigenvalue problem in low dimension. Linear Algebra and its Applications, 690(??):142–161, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000983.

Ji:2024:ESR

[JGW24] Xue Ji, Ji-Ming Guo, and Zhiwen Wang. The extremal spectral radius of generalized block graphs. Linear Algebra and its Applications, 680(??):239–253, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003853.

Jin:2021:DEC

[JGZ21] Ya-Lei Jin, Rui Gu, and Xiao-Dong Zhang. The distance energy of clique trees. Linear Algebra and its Applications, 615(??):1–10, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030598X.

Jankowski:2020:SCH

[JJL20a] Eric Jankowski, Charles R. Johnson, and Derek Lim. Spectra of convex hulls of matrix groups. Linear Algebra and its Applications, 593(??):74–89, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300227.

Jeong:2020:WMD

[JJL20b] Juyoung Jeong, Yoon Mo Jung, and Yongdo Lim. Weak majorization, doubly substochastic maps, and some related inequalities in Euclidean Jordan algebras. Linear Algebra and its Applications, 597(??):133–154, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301567.

Jablonski:2023:CPD

[JJLS23] Zenon Jan Jabłoński, Il Bong Jung, Eun Young Lee, and Jan Stochel. Conditionally positive definite unilateral weighted shifts. Linear Algebra and its Applications, 671(??):1–37, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379523001489.

Jablonski:2020:IOT

[JJS20a] Zenon Jan Jabłoński, Il Bong Jung, and Jan Stochel. m-Isometric operators and their local properties. Linear Algebra and its Applications, 596(??):49–70, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301166.

Johnson:2020:NRJ

[JJS20b] Charles R. Johnson, Ana I. Julio, and Ricardo L. Soto. Nonnegative realizability with Jordan structure. Linear Algebra and its Applications, 587(??):302–313, February 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304926.

Jang:2023:KCW

[JKS23] Jihyeug Jang, Sooyeong Kim, and Minho Song. Kemeny’s constant and Wiener index on trees. Linear Algebra and its Applications, 674(??):230–243, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002124.

Ji:2020:HSC

[JKX20] Kui Ji, Hyun-Kyoung Kwon, and Jing Xu. N-hypercontractivity and similarity of Cowen– Douglas operators. Linear Algebra and its Applications, 592(??):20–47, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300215.

Jiang:2020:HTE

[JL20] Zhaolin Jiang and Mathieu Lin. A Harnack type eigenvalue inequality. Linear Algebra and its Applications, 585(??):45–49, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304252.

Jiang:2022:CPF

[JL22] Tianyi Jiang and Shoumin Liu. Characteristic polynomials and finitely dimensional representations of ∫↕(2, C). Linear Algebra and its Applications, 647(??):78–88, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001537.

Jocic:2020:CIG

[JLM20a] Danko R. Jocić, Milan Lazarević, and Stefan Milošević. Corrigendum to “Inequalities for generalized derivations of operator monotone functions in norm ideals of compact operators” [Linear Algebra Appl. 586 (2020) 43–63]. Linear Algebra and its Applications, 599(??):201–204, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302007.  See

[JLM20b].

Jocic:2020:IGD

[JLM20b] Danko R. Jocić, Milan Lazarević, and Stefan Milosević. Inequalities for generalized derivations of operator monotone functions in norm ideals of compact operators. Linear Algebra and its Applications, 586(??):43–63, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304434. See cor-

rigendum [JLM20a].

Jia:2022:OMI

[JLW22] Huiming Jia, Shuchao Li, and Shujing Wang. Ordering the maxima of L-index and Q-index: Graphs with given size and diameter. Linear Algebra and its Applications, 652(??):18–36, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002518.

Jin:2022:OUT

[JLX22] Pengzhan Jin, Zhangli Lin, and Bo Xiao. Optimal unit triangular factorization of symplectic matrices. Linear Algebra and its Applications, 650(??):236–247, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002257.

Jaume:2021:NDB

[JMP21a] Daniel A. Jaume, Gonzalo Molina, and Adrián Pastine. Null decomposition of bipartite graphs without cycles of length 0 modulo 4. Linear Algebra and its Applications, 614(??):176–196, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301580.

Johnson:2021:RCS

[JMP21b] C. R. Johnson, C. Marijuán, and M. Pisonero. Ruling out certain 5-spectra with one repeated eigenvalue for the symmetric NIEP. Linear Algebra and its Applications, 612(??):75–98, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305577.

Julio:2021:URL

[JMPS21] Ana I. Julio, Carlos Marijuán, Miriam Pisonero, and Ricardo L. Soto. Universal realizability in low dimension. Linear Algebra and its Applications, 619(??):107–136, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000719.

Johnson:2024:DEI

[JMPS24] C. R. Johnson, C. Marijuán, M. Pisonero, and I. M. Spitkovsky. Diagonal entries of inverses of diagonally dominant matrices. Linear Algebra and its Applications, 692(??):84–90, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001460.

Jia:2022:NUO

[JN22] Yu-Shu Jia and Chi-Keung Ng. Non-unital operator systems that are dual spaces. Linear Algebra and its Applications, 650(??):150–161, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002166.

Jones:2021:MRM

[Jon21] Daniel Jones. Matrix roots in the max-plus algebra. Linear Algebra and its Applications, 631(??):10–34, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003050.

Jose:2024:SME

[JR24] Selby Jose and Ravi A. Rao. Suslin matrices and the Euler characteristic. Linear Algebra and its Applications, 682(??):268–284, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004275.

Julio:2020:URP

[JS20] Ana I. Julio and Ricardo L. Soto. On the universal realizability problem. Linear Algebra and its Applications, 597(??):170–186, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301543.

Johnston:2022:CPC

[JS22] Nathaniel Johnston and Jamie Sikora. Completely positive completely positive maps (and a resource theory for non-negativity of quantum amplitudes). Linear Algebra and its Applications, 653(??):395–429, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002932.

Johnson:2021:CVS

[JST21] Charles R. Johnson, Carlos M. Saiago, and Kenji Toyonaga. Change in vertex status after removal of another vertex in the general setting. Linear Algebra and its Applications, 612(??):128–145, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305632.

Johnston:2024:EVS

[JT24] Jacob Johnston and Michael Tait. Extremal values for the spectral radius of the normalized distance Laplacian. Linear Algebra and its Applications, 691(??):37–49, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379524001046.

Johnson:2022:TSP

[JWZ22] Charles Johnson, M. Waugh, and Yulin Zhang. Tree sign pattern that permit eventual positivity. Linear Algebra and its Applications, 650(??):162–168, October 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002269.

Jin:2024:UBS

[JZZ24] Ya-Lei Jin, Jie Zhang, and Xiao-Dong Zhang. Upper bounds of spectral radius of symmetric matrices and graphs. Linear Algebra and its Applications, 682(??):152–163, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004226.

Kada:2020:CPZ

[Kad20] Osamu Kada. Characteristic polynomials and zeta functions of equitably partitioned graphs. Linear Algebra and its Applications, 588(??):471–488, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305063.

Kadyan:2023:HIE

[Kad23] Monu Kadyan. HS-integral and Eisenstein integral normal mixed Cayley graphs. Linear Algebra and its Applications, 669(??):1–23, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001064.

Koolen:2024:DRG

[KAGH24] Jack H. Koolen, Mamoon Abdullah, Brhane Gebremichel, and Sakander Hayat. Distanceregular graphs with exactly one positive q-distance eigenvalue. Linear Algebra and its Applications, 689(??):230–246, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952400082X.

Kakariadis:2020:AEP

[Kak20] Evgenios T. A. Kakariadis. Applications of entropy of product systems: Higher-rank graphs. Linear Algebra and its Applications, 594(??):124–157, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300598.

Karantzas:2020:EDM

[KAPS20] N. Karantzas, N. Atreas, M. Papadakis, and T. Stavropoulos. Erratum to “On the design of multi-dimensional compactly supported Parseval framelets with directional characteristics [Linear Algebra Appl. 582 (2019) 1–36]. Linear Algebra and its Applications, 584(??):483–484, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304197.  See

[AKPS19].

Kadyan:2022:HIE

[KB22] Monu Kadyan and Bikash Bhattacharjya. HS-integral and Eisenstein integral mixed Cayley graphs over abelian groups. Linear Algebra and its Applications, 645(??):68–90, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001033.

Ku-Cauich:2020:LCB

[KCML20] Juan Carlos Ku-Cauich and Guillermo Morales-Luna. A linear code based on resilient Boolean maps whose dual is a platform for a robust secret sharing scheme. Linear Algebra and its Applications, 596(??):216–229, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379520301312.

Kennedy-Cochran-Patrick:2021:NBP

[KCPMNS21] Arthur Kennedy-Cochran-Patrick, Glenn Merlet, Thomas Nowak, and Sergeĭ Sergeev. New bounds on the periodicity transient of the powers of a tropical matrix: Using cyclicity and factor rank. Linear Algebra and its Applications, 611(??):279–309, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305152.

Kovac:2021:NSM

[KFS21] Erna Begović Kovač, Heike Faßbender, and Philip Saltenberger. On normal and structured matrices under unitary structurepreserving transformations. Linear Algebra and its Applications, 608(??):322–342, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304419.

Koolen:2023:SRG

[KGYY23] Jack H. Koolen, Brhane Gebremichel, Jae Young Yang, and Qianqian Yang. Sesquiregular graphs with fixed smallest eigenvalue. Linear Algebra and its Applications, 668(??):1–10, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001003.

Khan:2024:CGG

[Kha24] Suliman Khan. On connected T-gain graphs with rank equal to girth. Linear Algebra and its Applications, 688(??):232–243, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000740.

Kouri:2023:GGM

[KHU23] Drew P. Kouri, Zuhao Hua, and Madeleine Udell. A greedy Galerkin method to efficiently select sensors for linear dynamical systems. Linear Algebra and its Applications, 679(??):275–304, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003336.

Kim:2020:SBM

[Kim20]  Ik-Pyo Kim.  Symmetric

Bernoulli matrix and its Cholesky decomposition. Linear Algebra and its Applications, 603(??):402–417, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S002437952030313X.

Kimsey:2022:MSC

[Kim22]  David P. Kimsey.  The

matricial subnormal completion problem. Linear Algebra and its Applications, 645(??):170–193, July 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001185.

King:2021:CSP

[Kin21] Emily J. King. Constructing subspace packings from other packings. Linear Algebra and its Applications, 625(??):68–80, September 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001865.

Konno:2021:SUM

[KIS21] Norio Konno, Yusuke Ide, and Iwao Sato. The spectra of the unitary matrix of an N-tessellable staggered quantum walk on a graph. Linear Algebra and its Applications, 611(??):187–212, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305097.

Kishore:2022:MWP

[Kis22]  Krishna Kishore.  Matrix

Waring problem. Linear Algebra and its Applications, 646(??):84–94, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001306.

Kaniecki:2020:ALR

[KK20a] Mariusz Kaniecki and Justyna Kosakowska. Applications of Littlewood–Richardson tableaux to computing generic extension of semisimple invariant subspaces of nilpotent linear operators. Linear Algebra and its Applications, 588(??):134–159, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304951.

Kim:2020:PCK

[KK20b] Bara Kim and Jeongsim Kim. Proofs of conjectures on the Karpelevich arcs in the region of eigenvalues of stochastic matrices. Linear Algebra and its Applications, 595(??):13–23, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301014.

Kim:2022:CAD

[KK22a] Bara Kim and Jeongsim Kim. Conjectures about determining the regions of eigenvalues of stochastic and doubly stochastic matrices. Linear Algebra and its Applications, 637(??):157–174, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004419.

Kim:2022:GMS

[KK22b] Sooyeong Kim and Steve Kirkland. Gram mates, sign changes in singular values, and isomorphism. Linear Algebra and its Applications, 644(??):108–148, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000891.

Konstantinova:2022:STG

[KK22c] Elena V. Konstantinova and Artem Kravchuk. Spectrum of the Transposition graph. Linear Algebra and its Applications, 654 (??):379–389, December 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003160.

Kaygorodov:2023:TPSa

[KK23a] Ivan Kaygorodov and Mykola Khrypchenko. Transposed Poisson structures on Block Lie algebras and superalgebras. Linear Algebra and its Applications, 656 (??):167–197, January 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003494.

Kaygorodov:2023:TPSb

[KK23b] Ivan Kaygorodov and Mykola Khrypchenko. Transposed Poisson structures on Witt type algebras. Linear Algebra and its Applications, 665(??):196–210, May 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000447.

Konstantinova:2024:DET

[KK24] Elena V. Konstantinova and Artem Kravchuk. Distinct eigenvalues of the Transposition graph. Linear Algebra and its Applications, 690(??):132–141, June 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952400096X.

Khouja:2022:RNO

[KKM22a] Rima Khouja, Houssam Khalil, and Bernard Mourrain. Riemannian Newton optimization methods for the symmetric tensor approximation problem. Linear Algebra and its Applications, 637(??):175–211, March 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004377.

Krahmer:2022:RNB

[KKM22b] Felix Krahmer, Christian Kümmerle, and Oleh Melnyk. On the robustness of noiseblind low-rank recovery from rank-one measurements. Linear Algebra and its Applications, 652(??):37–81, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002609.

Kim:2022:CTT

[KKOO22] Jaehoon Kim, Soyeon Kim, Suil O., and Semin Oh. A Cvetković-type Theorem for coloring of digraphs. Linear Algebra and its Applications, 634(??):30–36, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379521003876.

Kannan:2021:BEE

[KKP21] M. Rajesh Kannan, Navish Kumar, and Shivaramakrishna Pragada. Bounds for the extremal eigenvalues of gain Laplacian matrices. Linear Algebra and its Applications, 625(??):212–240, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002111.

Komatsu:2021:ZFR

[KKS21] Takashi Komatsu, Norio Konno, and Iwao Sato. A zeta function with respect to non-backtracking alternating walks for a digraph. Linear Algebra and its Applications, 620(??):344–367, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001178.

Kabanov:2021:SGE

[KKSV21] Vladislav V. Kabanov, Elena V. Konstantinova, Leonid Shalaginov, and Alexandr Valyuzhenich. The Star graph eigenfunctions with non-zero eigenvalues. Linear Algebra and its Applications, 610 (??):222–226, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304705.

Kocharovsky:2022:HMT

[KKT22] Vitaly V. Kocharovsky, Vladimir V. Kocharovsky, and Sergey V. Tarasov. The Hafnian Master Theorem. Linear Algebra and its Applications, 651(??):144–161, October 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002439.

Kauffman:2021:OAC

[KL21a] Louis H. Kauffman and Pedro Lopes. On the orbits associated with the Collatz conjecture. Linear Algebra and its Applications, 615(??):143–154, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000021.

Kenter:2021:ZFT

[KL21b] Franklin H. J. Kenter and Jephian C.-H. Lin. A zero forcing technique for bounding sums of eigenvalue multiplicities. Linear Algebra and its Applications, 629 (??):138–167, November 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379521002792.

Kook:2021:KIS

[KL21c] Woong Kook and Kang-Ju Lee. Kirchhoff index of simplicial networks. Linear Algebra and its Applications, 626(??):1–19, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002147.

Kaur:2023:NLB

[KL23] Avleen Kaur and S. H. Lui. New lower bounds on the minimum singular value of a matrix. Linear Algebra and its Applications, 666(??):62–95, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000617.

Kline:2020:BHM

[Kli20a] Jeffery Kline. Bordered Hermitian matrices and sums of the Möbius function. Linear Algebra and its Applications, 588(??):224–237, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305117.

Kline:2020:ESM

[Kli20b] Jeffery Kline. On the eigenstructure of sparse matrices related to the prime number theorem. Linear Algebra and its Applications, 584(??):409–430, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304057.

Kosan:2022:TSZ

[KLL22] M. Tamer Koşan, Tsiu-Kwen Lee, and Jheng-Huei Lin. Two-sided zero product properties on symmetric algebras. Linear Algebra and its Applications, 655(??):186–201, December 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003391.

Krutov:2022:NIS

[KLLS22] Andrey Krutov, Alexei Lebedev, Dimitry Leites, and Irina Shchepochkina. Nondegenerate invariant symmetric bilinear forms on simple Lie superalgebras in characteristic 2. Linear Algebra and its Applications, 649(??):1–21, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522001781.

Kang:2021:ESR

[KLLW21] Liying Kang, Lele Liu, Linyuan Lu, and Zhiyu Wang. The extremal p-spectral radius of Berge hypergraphs. Linear Algebra and its Applications, 610(??):608–624, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304894.

Kumar:2022:SMS

[KLR22] Chanchal Kumar, Gargi Lather, and Amit Roy. Standard monomials of 1-skeleton ideals of graphs and generalized signless Laplacians. Linear Algebra and its Applications, 637(??):24–48, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004328.

Kalauch:2021:MWG

[KLS21] A. Kalauch, S. Lavanya, and K. C. Sivakumar. Matrices whose group inverses are M-matrices. Linear Algebra and its Applications, 614(??):44–67, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379519305403.

Kuo:2023:OFO

[KLS23] Yueh-Cheng Kuo, Huey-Er Lin, and Shih-Feng Shieh. The orthogonal flows for orthogonal iteration. Linear Algebra and its Applications, 679(??):67–85, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003324.

Kuzma:2024:LMP

[KLSP24] Bojan Kuzma, Chi-Kwong Li, Sushil Singla, and Edward Poon. Linear maps preserving parallel matrix pairs with respect to the Ky–Fan k-norm. Linear Algebra and its Applications, 687(??):68–90, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000314.

Kribs:2020:VRG

[KMNP20] David W. Kribs, Comfort Mintah, Michael Nathanson, and Rajesh Pereira. Vector representations of graphs and distinguishing quantum product states with oneway LOCC. Linear Algebra and its Applications, 602(??):223–239, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030255X.

Koutis:2023:GCI

[KMP23] Ioannis Koutis, Gary Miller, and Richard Peng. A generalized Cheeger inequality. Linear Algebra and its Applications, 665(??):139–152, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000204.

Khadam:2020:SVT

[KMZ20] M. Azeem Khadam, Mateusz Michałek, and Piotr Zwiernik. Secant varieties of toric varieties arising from simplicial complexes. Linear Algebra and its Applications, 588(??):428–457, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305154.

Kunis:2020:SSV

[KN20] Stefan Kunis and Dominik Nagel. On the smallest singular value of multivariate Vandermonde matrices with clustered nodes. Linear Algebra and its Applications, 604(??):1–20, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

Knill:2020:ESC

[Kni20] Oliver Knill. The energy of a simplicial complex. Linear Algebra and its Applications, 600(??):96–129, September 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302019.

Kamiyoshi:2023:ETE

[KNiO23] Tomohiro Kamiyoshi, Makoto Nagura, and Shin ichi Otani. Exponential-type extended Riordan arrays and reciprocity law for generalized Stirling numbers. Linear Algebra and its Applications, 674(??):396–425, October 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002173.

Kitson:2020:RSF

[KNS20] D. Kitson, A. Nixon, and B. Schulze. Rigidity of symmetric frameworks in normed spaces. Linear Algebra and its Applications, 607(??):231–285, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520303773.

Kolegov:2021:RAG

[Kol21] N. A. Kolegov. On real algebras generated by positive and nonnegative matrices. Linear Algebra and its Applications, 611(??):46–65, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305450.

Koplewitz:2020:CRR

[Kop20]  Shaked Koplewitz.  The

corank of a rectangular random integer matrix. Linear Algebra and its Applications, 591(??):160–168, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305397.

Korhonen:2021:DES

[Kor21] Mikko Korhonen. Decomposition of exterior and symmetric squares in characteristic two. Linear Algebra and its Applications, 624(??):349–363, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001828.

Kovalyov:2020:RIS

[Kov20] Ivan Kovalyov. Regularization of the indefinite Stieltjes moment problem. Linear Algebra and its Applications, 594(??):1–28, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030077X.

Kovac:2021:FCN

[Kov21] Erna Begović Kovač. Finding the closest normal structured matrix. Linear Algebra and its Applications, 617(??):49–77, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000240.

Kozhasov:2021:MDV

[Koz21]  Khazhgali Kozhasov.  On

minimality of determinantal varieties. Linear Algebra and its Applications, 626(??):56–78, October 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002135.

Kushel:2021:GCD

[KP21] Olga Y. Kushel and Raffaella Pavani. Generalization of the concept of diagonal dominance with applications to matrix D-stability. Linear Algebra and its Applications, 630(??):204–224, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002950.

Kosir:2022:STP

[KP22] Tomaž Košir and Bor Plestenjak. On the singular twoparameter eigenvalue problem II. Linear Algebra and its Applications, 649 (??):433–451, September 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002002.

Kannan:2023:SST

[KP23a] M. Rajesh Kannan and Shivaramakrishna Pragada. Signed spectral Turań type theorems. Linear Algebra and its Applications, 663(??):62–79, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000022.

Koval:2023:MPN

[KP23b] Vadym Koval and Patryk Pagacz. Matrix pencils with the numerical range equal to the whole complex plane. Linear Algebra and its Applications, 657(??):274–286, January 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200386X.

Knirsch:2021:ORH

[KPP21] Hanna Knirsch, Markus Petz, and Gerlind Plonka. Optimal rank-1 Hankel approximation of matrices: Frobenius norm and spectral norm and Cadzow’s algorithm. Linear Algebra and its Applications, 629(??):1–39, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002627.

Karami:2020:PNH

[KS20a] Saeed Karami and Abbas Salemi. Polynomial numerical hulls of the direct sum of two Jordan blocks. Linear Algebra and its Applications, 585(??):209–226, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304331.

Konyagin:2020:EPI

[KS20b] Sergei Konyagin and Benny Sudakov. An extremal problem for integer sparse recovery. Linear Algebra and its Applications, 586(??):1–6, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930432X.

Kalachev:2022:LAL

[KS22a] Gleb Kalachev and Sergey Sadov. A linear-algebraic and lattice-theoretical look at the Cleaning Lemma of quantum coding theory. Linear Algebra and its Applications, 649(??):96–121, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001896.

Kirkland:2022:SMR

[KŠ22b] Stephen Kirkland and Helena Šmigoc. Stochastic matrices realising the boundary of the Karpelevič region. Linear Algebra and its Applications, 635(??):116–138, February 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004225.

Kubota:2022:PST

[KS22c] Sho Kubota and Etsuo Segawa. Perfect state transfer in Grover walks between states associated to vertices of a graph. Linear Algebra and its Applications, 646(??):238–251, August 1, 2022. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001513.

Kager:2023:CMR

[KS23a] Wouter Kager and Pieter Jacob Storm. On the class of matrices with rows that weakly decrease cyclicly from the diagonal. Linear Algebra and its Applications, 673(??):200–219, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001969.

Korobenko:2023:SSI

[KS23b] Lyudmila Korobenko and Eric Sawyer. Sums of squares II: Matrix functions. Linear Algebra and its Applications, 664(??):264–313, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004669.

Korotyaev:2023:TFM

[KS23c] Evgeny Korotyaev and Natalia Saburova. Trace formulas for magnetic Schrödinger operators on periodic graphs and their applications. Linear Algebra and its Applications, 676(??):395–440, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002914.

Kaushik:2024:WPT

[KS24] Rahul Kaushik and Anupam Singh. Waring problem for triangular matrix algebra. Linear Algebra and its Applications, 692(??):135–159, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001277.

Kempton:2020:CCV

[KSSW20] Mark Kempton, John Sinkovic, Dallas Smith, and Benjamin Webb. Characterizing cospectral vertices via isospectral reduction. Linear Algebra and its Applications, 594(??):226–248, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300926.

Kubjas:2022:ESL

[KST22] Kaie Kubjas, Luca Sodomaco, and Elias Tsigaridas. Exact solutions in low-rank approximation with zeros. Linear Algebra and its Applications, 641(??):67–97, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379522000362.

Kubota:2021:PQW

[KSY21] Sho Kubota, Hiroto Sekido, and Harunobu Yata. Periodicity of quantum walks defined by mixed paths and mixed cycles. Linear Algebra and its Applications, 630(??):15–38, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002871.

Kalantarova:2021:SSJ

[KT21] Nargiz Kalantarova and Levent Tunçel. On the spectral structure of Jordan-Kronecker products of symmetric and skew-symmetric matrices. Linear Algebra and its Applications, 608(??):343–362, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303955.

Kubota:2022:PGW

[Kub22] Sho Kubota. Periodicity of Grover walks on bipartite regular graphs with at most five distinct eigenvalues. Linear Algebra and its Applications, 654(??):125–142, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522003147.

Kubota:2023:CNC

[Kub23] Sho Kubota. Combinatorial necessary conditions for regular graphs to induce periodic quantum walks. Linear Algebra and its Applications, 673(??):259–279, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952300201X.

Kushel:2023:SBD

[Kus23]  Olga Y. Kushel.  Some

bounds for determinants of relatively D-stable matrices. Linear Algebra and its Applications, 656(??):9–26, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003433.

Kozybaev:2022:SEN

[KUZ22] Daniyar Kozybaev, Ualbai Umirbaev, and Viktor Zhelyabin. Some examples of nonassociative coalgebras and supercoalgebras. Linear Algebra and its Applications, 643(??):235–257, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000817.

Kass:2020:CPA

[KW20] Jesse Leo Kass and Kirsten Wickelgren. A classical proof that the algebraic homotopy class of a rational function is the residue pairing. Linear Algebra and its Applications, 595(??):157–181, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305622.

Key:2021:SIF

[KY21] Fatemeh Azari Key and Rongwei Yang. Spectral invariants for finite dimensional Lie algebras. Linear Algebra and its Applications, 611 (??):148–170, February 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305073.

Kye:2023:CTP

[Kye23] Seung-Hyeok Kye. Compositions and tensor products of linear maps between matrix algebras. Linear Algebra and its Applications, 658(??):283–309, February 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004025.

Kittaneh:2023:PIW

[KZ23] Fuad Kittaneh and Ali Zamani. Proper improvements of well-known perturbation estimations for the parallel sum. Linear Algebra and its Applications, 678(??):125–135, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003221.

Kittaneh:2024:OR

[KZ24] Fuad Kittaneh and Ali Zamani. On the ρ-operator radii. Linear Algebra and its Applications, 687(??):132–156, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000375.

Khaleghzade:2022:PFT

[KZPH22] Sedighe Khaleghzade, Mostafa Zangiabadi, Aljoša Peperko, and Masoud Hajarian. Perron– Frobenius theory for some classes of nonnegative tensors in the max algebra. Linear Algebra and its Applications, 641(??):115–142, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000544.

Liu:2023:SRC

[LBW23] Xiangxiang Liu, Hajo Broersma, and Ligong Wang. Spectral radius conditions for the existence of all subtrees of diameter at most four. Linear Algebra and its Applications, 663(??):80–101, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000046.

Li:2020:ESR

[LC20] Wei Li and An Chang. The effect on the spectral radius of rgraphs by grafting or contracting edges. Linear Algebra and its Applications, 597(??):1–17, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301464.

Liu:2023:OIC

[LCH23] Kai Liu, Chuangxun Cheng, and Deguang Han. Orbitinjective covariant quantum channels. Linear Algebra and its Applications, 668 (??):28–50, July 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379523001027.

Liu:2020:LEM

[LDSS20] Shuting Liu, Kinkar Chandra Das, Shaowei Sun, and Jinlong Shu. On the least eigenvalue of A_{α}-matrix of graphs. Linear Algebra and its Applications, 586(??):347–376, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304653.

Lund:2021:FDG

[LDW21] Darren Lund, Joseph Drapeau, and Benjamin Webb. Fourier decompositions of graphs with symmetries and equitable partitions. Linear Algebra and its Applications, 627(??):199–226, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002275.

Liang:2021:TRP

[LDZ21] Y. G. Liang, Sergio Da Silva, and Yang Zhang. The tensor rank problem over the quaternions. Linear Algebra and its Applications, 620(??):37–60, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000781.

Le:2022:HTI

[Le22]  Nam Q. Le.  Hadamard-

type inequalities for k-positive matrices. Linear Algebra and its Applications, 635 (??):159–170, February 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004249.

Lee:2020:GGD

[Lee20]  Jae-Ho Lee.  Grassmann

graphs, degenerate DAHA, and non-symmetric dual q-Hahn polynomials. Linear Algebra and its Applications, 588(??):160–195, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305038.

Lee:2022:CHP

[Lee22] Jae-Ho Lee. Circular Hessenberg pairs. Linear Algebra and its Applications, 655(??):202–235, December 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003408.

Lewkowicz:2021:PLD

[Lew21] Izchak Lewkowicz. Passive linear discrete-time systems: Characterization through structure. Linear Algebra and its Applications, 623(??):299– 315, August 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379521000586.

Li:2020:EBM

[LF20] Yongtao Li and Lihua Feng. Extensions of Brunn–Minkowski’s inequality to multiple matrices. Linear Algebra and its Applications, 603(??):91–100, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302810.

Lin:2021:SCO

[LG21a] Huiqiu Lin and Hangtian Guo. A spectral condition for odd cycles in nonbipartite graphs. Linear Algebra and its Applications, 631(??):83–93, December 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003177.

Liu:2021:GRB

[LG21b] Yong Liu and Chuan-Qing Gu. On greedy randomized block Kaczmarz method for consistent linear systems. Linear Algebra and its Applications, 616(??):178–200, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000434.

Li:2020:SNI

[LH20] Jing Li and Han Hai. Some new inequalities for the Hadamard product of nonnegative matrices. Linear Algebra and its Applications, 606(??):159–169, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303487.

Linh:2021:IRS

[LH21] Vu Hoang Linh and Phi Ha. Index reduction for second order singular systems of difference equations. Linear Algebra and its Applications, 608(??):107–132, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030392X.

Li:2020:SAT

[LHFL20] Yongtao Li, Yang Huang, Lihua Feng, and Weijun Liu. Some applications of two completely copositive maps. Linear Algebra and its Applications, 590(??):124–132, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305579.

Liu:2022:MPR

[LHH22] Lele Liu, Shengming Hu, and Changxiang He. Maximum principal ratio of the signless Laplacian of graphs. Linear Algebra and its Applications, 653(??):339–356, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200310X.

Limantseva:2020:NCR

[LHK20] Olga Limantseva, George Halikias, and Nicos Karcanias. Nearest common root of a set of polynomials: a structured singular value approach. Linear Algebra and its Applications, 584(??):233–256, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437951930388X.

Luo:2021:SLR

[LHZ21] Yuetian Luo, Rungang Han, and Anru R. Zhang. A Schatten-q low-rank matrix perturbation analysis via perturbation projection error bound. Linear Algebra and its Applications, 630 (??):225–240, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

Li:2020:EVQ

[Li20a] Ben Li. Exact values of quantum violations in lowdimensional Bell correlation inequalities. Linear Algebra and its Applications, 603 (??):289–300, October 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302846.

Li:2020:UFM

[Li20b] Kenan Li. On the uniqueness of functions that maximize the Crouzeix ratio. Linear Algebra and its Applications, 599(??):105–120, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301804.

Li:2022:ESM

[Li22] Yongtao Li. Extensions of some matrix inequalities related to trace and partial traces. Linear Algebra and its Applications, 639(??):205–224, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000088.

Li:2024:MEJ

[Li24] Chi-Kwong Li. A matrix equation and the Jordan canonical form. Linear Algebra and its Applications, 687(??):1–7, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000326.

Lim:2021:ITR

[Lim21] Yongdo Lim. Invariant tolerance relations on positive definite matrices. Linear Algebra and its Applications, 619(??):1–11, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000604.

Lim:2022:HSE

[Lim22a] Yongdo Lim. Homogeneous Sylvester equations of lower triangular matrices. Linear Algebra and its Applications, 653(??):183–192, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002865.

Lim:2022:SDF

[Lim22b] Yongdo Lim. Smooth dependence of fixed points of Hamiltonians. Linear Algebra and its Applications, 655(??):87–99, December 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003263.

Lins:2023:UAN

[Lin23] Brian Lins. A unified approach to nonlinear Perron– Frobenius theory. Linear Algebra and its Applications, 675(??):48–89, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002434.

Liang:2020:TAG

[LIW20] Xiaoye Liang, Tatsuro Ito, and Yuta Watanabe. The Terwilliger algebra of the Grassmann scheme J_{q}(N, D) revisited from the viewpoint of the quantum affine al-

gebra U_{q}(ˆ∫↕^{2}). Linear Algebra and its Applications, 596(??):117–144, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301191.

Li:2023:TSE

[LJSZ23] Xiuqing Li, Xian’an Jin, Chao Shi, and Ruiling Zheng. Two spectral extremal results for graphs with given order and rank. Linear Algebra and its Applications, 672(??):53–74, September 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001660.

Liu:2020:KCC

[LL20] Yuanyuan Liu and Fangfang Lyu. Kemeny’s constant for countable Markov chains. Linear Algebra and its Applications, 604 (??):425–440, November 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303244.

Lam:2021:QDP

[LL21a] Nguyen Lam and Phi Long Le. Quantum divergences with p-power means. Linear Algebra and its Applications, 609(??):289–307, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304316.

Lawson:2021:AKM

[LL21b] Jimmie Lawson and Yongdo Lim. Analyticity of the Karcher mean coefficient maps. Linear Algebra and its Applications, 627(??):162– 184, October 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521002445.

Li:2021:NLG

[LL21c] Wei Li and Lele Liu. The α-normal labeling for generalized directed uniform hypergraphs. Linear Algebra and its Applications, 611 (??):365–388, February 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305206.

Li:2021:CMF

[LL21d] Yun-Zhang Li and Ya-Nan Li. Constructing more Kframes. Linear Algebra and its Applications, 616(??):45–65, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000070.

Lou:2021:DEC

[LL21e] Zhenzhen Lou and Huiqiu Lin. Distance eigenvalues of a cograph and their multiplicities. Linear Algebra and its Applications, 608(??):1–12, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030389X.

Lau:2024:APM

[LL24a] Jinting Lau and Ming Huat Lim. Adjacency preserving maps between exterior powers of vector spaces. Linear Algebra and its Applications, 686(??):83–101, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952400017X.

Liang:2024:GTM

[LL24b] Xin Liang and Ren-Cang Li. On generalizing trace minimization principles, II. Linear Algebra and its Applications, 687(??):8–37, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000296.

Lou:2023:FMF

[LLA23] Jing Lou, Ruifang Liu, and Guoyan Ao. Fractional matching, factors and spectral radius in graphs involving minimum degree. Linear Algebra and its Applications, 677(??):337–351, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003154.

Li:2020:ZFS

[LLH20] Deqiong Li, Jing Li, and Yaoping Hou. Zeta functions of several corona-type graphs. Linear Algebra and its Applications, 591(??):134–153, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300112.

Lu:2022:MGS

[LLL22] Junying Lu, Weijun Liu, and Lu Lu. Mixed graphs with smallest eigenvalue greater than −3. Linear Algebra and its Applications, 635 (??):95–115, February 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004092.

Lucas:2021:SRS

[LLSV21] Lisa Hernandez Lucas, Ivan Landjev, Leo Storme, and Peter Vandendriessche. A stability result and a spectrum result on constant dimension codes. Linear Algebra and its Applications, 621(??):193–213, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001166.

Li:2020:AVE

[LLW20] Xueliang Li, Yiyang Li, and Zhiqian Wang. The asymptotic value of energy for matrices with degree-distance-based entries of random graphs. Linear Algebra and its Applications, 603(??):390–401, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303116.

Li:2021:AVF

[LLW21] Xueliang Li, Yiyang Li, and Zhiqian Wang. Asymptotic values of four Laplacian-type energies for matrices with degree-distance-based entries of random graphs. Linear Algebra and its Applications, 612(??):318–333, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305309.

Linz:2023:MSM

[LLW23] William Linz, Linyuan Lu, and Zhiyu Wang. Maximum spread of K_{2,t}-minor-free graphs. Linear Algebra and its Applications, 676(??):352–373, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379523002902.

Li:2023:MTT

[LLWW23] Shu Li, Lu Lu, Jianfeng Wang, and Weifang Wang. Matrix-Tree Theorem of digraphs via signless Laplacians. Linear Algebra and its Applications, 658(??):32–48, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003925.

Liu:2023:MSR

[LLX23] Chang Liu, Jianping Li, and Zheng Xie. The minimum spectral radius of graphs with a given domination number. Linear Algebra and its Applications, 673(??):233–258, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001702.

LaRosa:2022:TSN

[LM22a] Gianmarco La Rosa and Manuel Mancini. Twostep nilpotent Leibniz algebras. Linear Algebra and its Applications, 637(??):119– 137, March 15, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521004420.

Loukili:2022:NPM

[LM22b] S. Loukili and M. Maslouhi. A new probabilistic model for optimal frames in erasure’s recovery. Linear Algebra and its Applications, 633(??):23–41, January 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100375X.

Lopatin:2024:SIT

[LM24] Artem Lopatin and Pedro Antonio Muniz Martins. Separating invariants for twodimensional orthogonal groups over finite fields. Linear Algebra and its Applications, 692(??):71–83, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001344.

Lin:2020:SG

[LMG20] Zhen Lin, Lianying Miao, and Shu-Guang Guo. The A_{α}- spread of a graph. Linear Algebra and its Applications, 606(??):1–22, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303451.

Luzon:2021:WDC

[LMR21] Ana Luzón, Manuel A. Morón, and José L. Ramírez. On Ward’s differential calculus, Riordan matrices and Sheffer polynomials. Linear Algebra and its Applications, 610(??):440–473, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304651.

Li:2021:PRS

[LMW21] Lei Li, Lajos Molnár, and Liguang Wang. On preservers related to the spectral geometric mean. Linear Algebra and its Applications, 610(??):647–672, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304912.

Liu:2023:MIN

[LMX23] Ruifang Liu, Lu Miao, and Jie Xue. Maxima of the Q-index of non-bipartite C_{3}-free graphs. Linear Algebra and its Applications, 673(??):1–13, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001854.

Le:2020:AMT

[LNN^{+}20] Vu A. Le, Tuan A. Nguyen, Tu T. C. Nguyen, Tuyen T. M. Nguyen, and Thieu N. Vo. Applying matrix theory to classify real solvable Lie algebras having 2-dimensional derived ideals. Linear Algebra and its Applications, 588(??):282–303, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305075.

Lev:2021:SFS

[LNP21] Vsevolod F. Lev, János Nagy, and Péter Pál Pach. Sum-full sets are not zerosum-free. Linear Algebra and its Applications, 625 (??):241–247, September 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100210X.

Loewy:2021:SSV

[Loe21]  Raphael Loewy.  On the

smallest singular value in the class of invertible lower triangular (0, 1) matrices. Linear Algebra and its Applications, 608(??):203–213, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304092.

Loewy:2023:PCP

[Loe23] Raphael Loewy. Proof of a conjecture on polynomials preserving nonnegative matrices. Linear Algebra and its Applications, 676 (??):267–276, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002926.

Lin:2020:SSP

[LOS20] Jephian C. H. Lin, Polona Oblak, and Helena Smigoc. The strong spectral property for graphs. Linear Algebra and its Applications, 598(??):68–91, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301592.

Lin:2021:IEP

[LOŠ21a] Jephian C.-H. Lin, Polona Oblak, and Helena Šmigoc. On the inverse eigenvalue problem for block graphs. Linear Algebra and its Applications, 631(??):379–397, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003414.

Losonczi:2021:ESI

[Los21b] László Losonczi. Eigenpairs of some imperfect pentadiagonal Toeplitz matrices. Linear Algebra and its Applications, 608(??):282–298, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304365.

Levene:2022:OSM

[LOŠ22] Rupert H. Levene, Polona Oblak, and Helena Šmigoc. Orthogonal symmetric matrices and joins of graphs. Linear Algebra and its Applications, 652(??):213–238, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002658.

Lin:2023:LSI

[LOŠ23] Jephian C. H. Lin, Polona Oblak, and Helena Šmigoc. The liberation set in the inverse eigenvalue problem of a graph. Linear Algebra and its Applications, 675(??):1–28, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002276.

Liu:2021:SCA

[LP21] Cheng-Kai Liu and Yen-Fu Pu. The structure of ncommuting additive maps on Lie ideals of prime rings. Linear Algebra and its Applications, 631(??):328–361, December 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521002913.

Lopatin:2022:ISP

[LP22] Artem Lopatin and Carlos Arturo Rodriguez Palma. Identities for subspaces of a parametric Weyl algebra. Linear Algebra and its Applications, 654(??):250–266, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003093.

Laface:2023:LSM

[LPS23] Antonio Laface, Elisa Postinghel, and Luis José Santana Sánchez. On linear systems with multiple points on a rational normal curve. Linear Algebra and its Applications, 657(??):197–240, January 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003883.

Liberti:2021:RPC

[LPV21] Leo Liberti, Pierre-Louis Poirion, and Ky Vu. Random projections for conic programs. Linear Algebra and its Applications, 626(??):204–220, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002457.

Li:2021:SRG

[LQ21] Dan Li and Rui Qin. The A_{α}-spectral radius of graphs with a prescribed number of edges for 12 ≤ α ≤ 1. Linear Algebra and its Applications, 628(??):29–41, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002500.

Li:2021:DSL

[LQL21] Yuanjing Li, Rui Qin, and Dan Li. On distance signless Laplacian spectrum of the complements of unicyclic graphs and trees. Linear Algebra and its Applications, 631(??):235–253, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003244.

Levick:2021:PFM

[LR21] Jeremy Levick and Mizanur Rahaman. Positively factorizable maps. Linear Algebra and its Applications, 631(??):282–307, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003360.

Larotonda:2023:WIN

[LR23] Gabriel Larotonda and Ivan Rey. Weakly invariant norms: Geometry of spheres in the space of skew-Hermitian matrices. Linear Algebra and its Applications, 678 (??):136–168, December 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003282.

Lemos:2020:SIK

[LS20a] Rute Lemos and Graça Soares. Spectral inequalities for Kubo–Ando operator means. Linear Algebra and its Applications, 607(??):29–44, December 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303670.

Lima:2020:NSD

[LS20b] Marcos Goulart Lima and Csaba Schneider. Nonsingular derivations of solvable Lie algebras in prime characteristic. Linear Algebra and its Applications, 586(??):170–189, February 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304501.

Lin:2020:RSV

[LS20c] Minghua Lin and Gord Sinnamon. Revisiting a sharpened version of Hadamard’s determinant inequality. Linear Algebra and its Applications, 606(??):192–200, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030361X.

Lockhart:2021:CE

[LS21] Joshua Lockhart and Simone Severini. Combinatorial entanglement. Linear Algebra and its Applications, 619(??):264–327, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000732.

Lavrauw:2022:TRS

[LS22] Michel Lavrauw and John Sheekey. The tensor rank of semifields of order 16 and 81. Linear Algebra and its Applications, 643(??):99–124, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000581.

Lettington:2023:SLR

[LS23] Matthew C. Lettington and Karl Michael Schmidt. On the sum of left and right circulant matrices. Linear Algebra and its Applications, 658(??):62–85, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003895.

Lippold:2024:EM

[LS24a] Steven R. Lippold and Mihai D. Staic. Existence of the map det^{S^{3}}. Linear Algebra and its Applications, 689(??):1–33, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952400079X.

Lledo:2024:GSA

[LS24b] Fernando Lledó and Ignacio Sevillano. Geometric and spectral analysis on weighted digraphs. Linear Algebra and its Applications, 687(??):252–280, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000302.

Lichtenberg:2024:DBA

[LT24] Samuel Lichtenberg and Abiy Tasissa. A dual basis approach to multidimensional scaling. Linear Algebra and its Applications, 682(??):86–95, February 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004184.

Liang:2022:DST

[LTTW22] Xiaoye Liang, Ying-Ying Tan, Hajime Tanaka, and Tao Wang. A duality of scaffolds for translation association schemes. Linear Algebra and its Applications, 638(??):110–124, April 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004535.

Li:2023:LMP

[LTWW23] Chi-Kwong Li, Ming-Cheng Tsai, Ya-Shu Wang, and Ngai-Ching Wong. Linear maps preserving matrices annihilated by a fixed polynomial. Linear Algebra and its Applications, 674(??):46–67, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002057.

Li:2023:CGF

[LTYZ23] Shuchao Li, Bit-Shun Tam, Yuantian Yu, and Qin Zhao. Connected graphs of fixed order and size with maximal A_{α}-index: the onedominating-vertex case. Linear Algebra and its Applications, 662(??):110–135, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000010.

Lubinsky:2022:DET

[Lub22] D. S. Lubinsky. Distribution of eigenvalues of Toeplitz matrices with smooth entries. Linear Algebra and its Applications, 633(??):332–365, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003797.

Lardizabal:2022:MHT

[LV22a] C. F. Lardizabal and L. Velázquez. Mean hitting time formula for positive maps. Linear Algebra and its Applications, 650(??):169–189, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002336.

Laurent:2022:ESS

[LV22b] Monique Laurent and Luis Felipe Vargas. On the exactness of sum-of-squares approximations for the cone of 5 × 5 copositive matrices. Linear Algebra and its Applications, 651(??):26–50, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002373.

Li:2020:ERS

[LW20a] Shuchao Li and Shujing Wang. The energy of random signed graph. Linear Algebra and its Applications, 585(??):227–240, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304343.

Li:2020:ASR

[LW20b] Xueliang Li and Junming Wang. On the ABC spectra radius of unicyclic graphs. Linear Algebra and its Applications, 596(??):71–81, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S002437952030121X.

Lau:2021:PAS

[LW21a] Terry Shue Chien Lau and Kok Bin Wong. On the partitions associated with the smallest eigenvalues of certain Cayley graphs on symmetric group generated by cycles. Linear Algebra and its Applications, 630 (??):179–203, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002925.

Li:2021:TES

[LW21b] Xueliang Li and Zhiqian Wang. Trees with extremal spectral radius of weighted adjacency matrices among trees weighted by degree-based indices. Linear Algebra and its Applications, 620(??):61–75, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000823.

Lu:2021:BRC

[LW21c] Yong Lu and Jingwen Wu. Bounds for the rank of a complex unit gain graph in terms of its maximum degree. Linear Algebra and its Applications, 610(??):73–85, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304638.

Lu:2021:NSG

[LW21d] Yong Lu and Jingwen Wu. No signed graph with the nullity η(G, σ) = |V (G)| − 2m(G) + 2c(G) − 1. Linear Algebra and its Applications, 615(??):175–193, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000045.

Liu:2024:SRG

[LW24] Yuxiang Liu and Ligong Wang. Spectral radius of graphs of given size with forbidden subgraphs. Linear Algebra and its Applications, 689(??):108–125, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000776.

Lei:2020:GWS

[LWB20] Xingyu Lei, Jianfeng Wang, and Maurizio Brunetti. Graphs whose second largest signless Laplacian eigenvalue does not exceed 2 + 2. Linear Algebra and its Applications, 603(??):242–264, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302780.

Lu:2021:CUG

[LWH21] Lu Lu, Jianfeng Wang, and Qiongxiang Huang. Complex unit gain graphs with exactly one positive eigenvalue. Linear Algebra and its Applications, 608(??):270–281, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304389.

Li:2022:HNO

[LWW22] Lei Li, Ya-Shu Wang, and Yau-Lun Wong. How the norm one positive definite functions determine a finite group. Linear Algebra and its Applications, 633(??):71–103, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003694.

Liu:2020:GCG

[LWYL20] Fenjin Liu, Wei Wang, Tao Yu, and Hong-Jian Lai. Generalized cospectral graphs with and without Hamiltonian cycles. Linear Algebra and its Applications, 585(??):199–208, January 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379519304288.

Liang:2023:GTM

[LWZL23] Xin Liang, Li Wang, Lei-Hong Zhang, and Ren-Cang Li. On generalizing trace minimization principles. Linear Algebra and its Applications, 656(??):483–509, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200372X.

Li:2020:SPR

[LY20] Hanyu Li and Shuheng Yin. Single-pass randomized algorithms for LU decomposition. Linear Algebra and its Applications, 595(??):101–122, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301051.

Li:2023:SEG

[LY23] Shuchao Li and Yuantian Yu. On A_{α} spectral extrema of graphs forbidding even cycles. Linear Algebra and its Applications, 668(??):11–27, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001015.

Lyu:2020:MWS

[Lyu20]  Zhen-Hua Lyu.  0–1 ma-

trices whose squares have bounded entries. Linear Algebra and its Applications, 607(??):1–8, December 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303645.

Lin:2023:IUH

[LYZ23] Hongying Lin, Gexin Yu, and Bo Zhou. On the irregularity of uniform hypergraphs. Linear Algebra and its Applications, 678 (??):107–124, December 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003233.

Lin:2020:DSR

[LZ20] Hongying Lin and Bo Zhou. Distance spectral radius of trees with given number of segments. Linear Algebra and its Applications, 600(??):40–59, September 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301920.

Leiter:2021:EMG

[LZ21a] Noam Leiter and Daniel Zelazo. Edge-matching graph contractions and their interlacing properties. Linear Algebra and its Applications, 612(??):289–317, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305280.

Lou:2021:PCE

[LZ21b] Zhenzhen Lou and Mingqing Zhai. Proof of a conjecture on extremal spectral radii of blow-up graphs. Linear Algebra and its Applications, 617(??):168–178, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000392.

Lin:2023:LBS

[LZZ23] Minghua Lin, Jiao Zhang, and Teng Zhang. A lower bound for the spectral radius of complex matrices. Linear Algebra and its Applications, 656(??):1–8, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003299.

Mandal:2022:COP

[MA22] Amrita Mandal and Bibhas Adhikari. A characterization of orthogonal permutative matrices of order 4. Linear Algebra and its Applications, 654(??):102–124, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003172.

Manivannan:2024:PAO

[MA24] Varadha Raj Manivannan and Victor Anandam. A perturbed averaging operator on finite graphs. Linear Algebra and its Applications, 680(??):208–219, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003749.

Marrinan:2021:MEB

[MAG21] Tim Marrinan, P.-A. Absil, and Nicolas Gillis. On a minimum enclosing ball of a collection of linear subspaces. Linear Algebra and its Applications, 625 (??):248–278, September 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002081.

Magajna:2022:OPO

[Mag22] Bojan Magajna. Orthonormal pairs of operators. Linear Algebra and its Applications, 648(??):233–253, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001811.

Mandziuk:2022:ILI

[Mań22] T. Mańdziuk. Identifying limits of ideals of points in the case of projective space. Linear Algebra and its Applications, 634(??):149–178, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003979.

Mao:2020:EHI

[Mao20] Yanling Mao. Extensions of Hartfiel’s inequality to multiple matrices. Linear Algebra and its Applications, 589(??):96–102, March 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305336.

Martin:2021:SGT

[Mar21] William J. Martin. Scaffolds: a graph-theoretic tool for tensor computations related to Bose–Mesner algebras. Linear Algebra and its Applications, 619(??):50–106, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Maroncelli:2024:FSD

[Mar24] Daniel Maroncelli. The Fučík spectrum for discrete systems and some nonlinear existence theorems. Linear Algebra and its Applications, 684(??):63–86, March 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004615.

Mayer:2021:PFG

[May21]  Thorsten Mayer.  Polyhe-

dral faces in Gram spectrahedra of binary forms. Linear Algebra and its Applications, 608(??):133–157, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303980.

Malik:2020:GKF

[MB20] Osman Asif Malik and Stephen Becker. Guarantees for the Kronecker fast Johnson– Lindenstrauss transform using a coherence and sampling argument. Linear Algebra and its Applications, 602(??):120–137, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379520302354.

Meidanis:2021:FMC

[MC21a] Joao Meidanis and Leonid Chindelevitch. Fast median computation for symmetric, orthogonal matrices under the rank distance. Linear Algebra and its Applications, 614(??):394–414, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305139.

Miao:2021:MCP

[MC21b] Xin-He Miao and Jein-Shan Chen. On matrix characterizations for P -property of the linear transformation in second-order cone linear complementarity problems. Linear Algebra and its Applications, 613(??):271–294, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305358.

Mishra:2023:NAE

[MDH23] Satyendra Kumar Mishra, Apurba Das, and Samir Kumar Hazra. Non-abelian extensions of Rota–Baxter Lie algebras and inducibility of automorphisms. Linear Algebra and its Applications, 669(??):147–174, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001325.

Miri:2021:NCM

[ME21] S. M. Miri and S. Effati. New characterizations of the matrix classes P, W and R_{0}. Linear Algebra and its Applications, 621(??):181–192, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001142.

Melman:2023:MWE

[Mel23] A. Melman. Matrices whose eigenvalues are those of a quadratic matrix polynomial. Linear Algebra and its Applications, 676 (??):131–149, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002501.

Meng:2022:TCP

[Men22] Jie Meng. Theoretical and computational properties of semi-infinite quasi-Toeplitz M-matrices. Linear Algebra and its Applications, 653(??):66–85, November 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522002853.

Mezzadri:2021:GID

[MG21] Francesco Mezzadri and Emanuele Galligani. A generalization of irreducibility and diagonal dominance with applications to horizontal and vertical linear complementarity problems. Linear Algebra and its Applications, 621(??):214–234, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001130.

Mojallal:2020:DEG

[MH20] Seyed Ahmad Mojallal and Pierre Hansen. On the difference of energies of a graph and its complement graph. Linear Algebra and its Applications, 595(??):1–12, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300987. See cor-

rigendum [MH21].

Mojallal:2021:CDE

[MH21] Seyed Ahmad Mojallal and Pierre Hansen. Corrigendum to “On the difference of energies of a graph and its complement graph” [Linear Algebra Appl. 595 (2020) 1–12]. Linear Algebra and its Applications, 628(??):228–230, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002755.  See

[MH20].

Mifsud:2021:CDP

[Mif21]  Xandru Mifsud.  λ-Core

distance partitions. Linear Algebra and its Applications, 613(??):170–182, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305668.

Miguel:2021:LCA

[Mig21]  C. Miguel.  Lengths of

cyclic algebras and commutative subalgebras of quaternion matrices. Linear Algebra and its Applications, 628(??):130–139, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002688.

Miller:2024:DPS

[Mil24]  Tomasz Miller.  On the

diagonal product of special unitary matrices. Linear Algebra and its Applications, 691(??):216–232, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001150.

Minh:2022:APM

[Min22]  Hà Quang Minh.  Alpha

Procrustes metrics between positive definite operators: a unifying formulation for the Bures–Wasserstein and Log-Euclidean/Log-Hilbert– Schmidt metrics. Linear Algebra and its Applications, 636(??):25–68, March 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004110.

Mishra:2020:FOS

[Mis20] Hemant Kumar Mishra. First order sensitivity analysis of symplectic eigenvalues. Linear Algebra and its Applications, 604(??):324–345, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303268.

Mishura:2022:CND

[Mis22] Teddy Mishura. Cut norm discontinuity of triangular truncation of graphons. Linear Algebra and its Applications, 650(??):26–41, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

Miyajima:2024:SPC

[Miy24] Shinya Miyajima. Some properties concerning Perron vectors of weakly irreducible nonnegative tensors, and their application to rigorous enclosure. Linear Algebra and its Applications, 687(??):207–228, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000417.

Mao:2024:RTI

[MJ24] Yanling Mao and Guoxing Ji. On Rotfel’d type inequalities for sectorial matrices. Linear Algebra and its Applications, 691(??):123–132, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001010.

Mahato:2023:NDD

[MK23] Iswar Mahato and M. Rajesh Kannan. A note on the distance and distance signless Laplacian spectral radius of complements of trees. Linear Algebra and its Applications, 675(??):344–350, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379523002604.

Morelli:2020:DSI

[MKE^{+}20] Simon Morelli, Claude Klöckl, Christopher Eltschka, Jens Siewert, and Marcus Huber. Dimensionally sharp inequalities for the linear entropy. Linear Algebra and its Applications, 584(??):294–325, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437951930391X.

Min:2022:SUB

[MLH22] Gao Min, Zhenzhen Lou, and Qiongxiang Huang. A sharp upper bound on the spectral radius of C_{5}-free/C_{6}-free graphs with given size. Linear Algebra and its Applications, 640(??):162–178, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000246.

Mlotkowski:2022:QEC

[Mło22] Wojciech Młotkowski. Quadratic embedding constants of path graphs. Linear Algebra and its Applications, 644(??):95–107, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000878.

Miao:2024:USE

[MLZ24] Lu Miao, Ruifang Liu, and Jingru Zhang. The unique spectral extremal graph for intersecting cliques or intersecting odd cycles. Linear Algebra and its Applications, 689(??):176–195, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000788.

Marijuan:2021:CTZ

[MM21] Carlos Marijuán and Julio Moro. A characterization of trace-zero sets realizable by compensation in the SNIEP. Linear Algebra and its Applications, 615(??):42–76, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305887.

uoan:2022:JCM

[– DMM^{+}22] Jacqueline – Doàn, Ján Mináč, Lyle Muller, Tung T. Nguyen, and Federico W. Pasini. Joins of circulant matrices. Linear Algebra and its Applications, 650(??):190–209, October 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200221X.

Milovanovic:2022:ZQP

[MMA22] Gradimir V. Milovanović, Abdullah Mir, and Abrar Ahmad. On the zeros of a quaternionic polynomial with restricted coefficients. Linear Algebra and its Applications, 653(??):231–245, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002877.

Matar:2022:MSR

[MMN22] Nancy Matar, Lon H. Mitchell, and Sivaram K. Narayan. On the minimum semidefinite rank of signed graphs. Linear Algebra and its Applications, 642(??):73–85, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000477.

Masakova:2021:GSS

[MMP21] Zuzana Masáková, Jan Mazáč, and Edita Pelantová. On generalized self-similarities of cut-and-project sets. Linear Algebra and its Applications, 625(??):279–321, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002123.

Mingueza:2020:CEA

[MMR20] David Mingueza, M. Eulàlia Montoro, and Alicia Roca. The centralizer of an endomorphism over an arbitrary field. Linear Algebra and its Applications, 591(??):322–351, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300173.

Marco:2021:EAP

[MMV21] Ana Marco, José-Javier Martínez, and Raquel Viaña. Error analysis, perturbation theory and applications of the bidiagonal decomposition of rectangular totally positive h-Bernstein– Vandermonde matrices. Linear Algebra and its Applications, 613(??):377–392, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305401.

Marco:2022:ACC

[MMV22] Ana Marco, José-Javier Martínez, and Raquel Viaña. Accurate computations with collocation matrices of the Lupaştype (p, q)-analogue of the Bernstein basis. Linear Algebra and its Applications, 651(??):312–331, October 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002464.

Mehl:2021:DPD

[MMW21] C. Mehl, V. Mehrmann, and M. Wojtylak. Distance problems for dissipative Hamiltonian systems and related matrix polynomials. Linear Algebra and its Applications, 623(??):335–366, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302640.

Mao:2022:YAC

[MMW22] Jianxi Mao, Lili Mu, and Yi Wang. Yet another criterion for the total positivity of Riordan arrays. Linear Algebra and its Applications, 634(??):106–111, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003992.

Marino:2020:MCA

[MMZ20] Giuseppe Marino, Maria Montanucci, and Ferdinando Zullo. MRD-codes arising from the trinomial x^{q} + x^{q^{3}} + cx^{q^{5}} ∈ F_{q^{6}}[x]. Linear Algebra and its Applications, 591(??):99–114, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Meier:2024:FRN

[MN24] Maike Meier and Yuji Nakatsukasa. Fast randomized numerical rank estimation for numerically low-rank matrices. Linear Algebra and its Applications, 686(??):1–32, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000016.

Mourrain:2020:MDL

[MO20] Bernard Mourrain and Alessandro Oneto. On minimal decompositions of low rank symmetric tensors. Linear Algebra and its Applications, 607(??):347–377, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303207.

Mohar:2020:NKH

[Moh20]  Bojan Mohar.  A new

kind of Hermitian matrices for digraphs. Linear Algebra and its Applications, 584(??):343–352, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304136.

Mohammadian:2022:RSM

[Moh22]  Ali Mohammadian.  Real

symmetric matrices and their negative eigenvalues. Linear Algebra and its Applications, 640(??):6–11, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000192.

Monterde:2023:FRB

[Mon23]  Hermie Monterde.  Frac-

tional revival between twin vertices. Linear Algebra and its Applications, 676 (??):25–43, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002665.

Morita:2020:RZF

[Mor20] Hideaki Morita. Ruelle zeta functions for finite digraphs. Linear Algebra and its Applications, 603(??):329–358, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303025.

Morales:2022:LSO

[Mor22] John Vincent S. Morales. Linking the special orthogonal algebra ∫≀_{4} and the tetrahedron algebra ⊗. Linear Algebra and its Applications, 637(??):212–239, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004389.

Mori:2023:DMN

[Mor23]  Michiya Mori.  On the

distance from a matrix to nilpotents. Linear Algebra and its Applications, 679 (??):99–103, December 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003476.

Mastnak:2021:LGR

[MORŠ21] Mitja Mastnak, Matjaž Omladič, Heydar Radjavi, and Klemen Šivic. Local and global reducibility of spaces of nilpotent matrices. Linear Algebra and its Applications, 611(??):260–278, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305140.

Mal:2020:CSO

[MP20a] Arpita Mal and Kallol Paul. Characterization of k-smooth operators between Banach spaces. Linear Algebra and its Applications, 586 (??):296–307, February 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304471.

Myskova:2020:SIM

[MP20b] Helena Mysková and Ján Plavka. On the solvability of interval max-min matrix equations. Linear Algebra and its Applications, 590(??):85–96, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305245.

Moon:2021:LUB

[MP21] Sunyo Moon and Seungkook Park. Lower and upper bounds for the number of distinct eigenvalues of a perturbed regular matrix pencil. Linear Algebra and its Applications, 626(??):171–187, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002251.

Martinez-Penas:2022:TAL

[MP22] Umberto Martínez-Peñas. Theory and applications of linearized multivariate skew polynomials. Linear Algebra and its Applications, 637 (??):1–23, March 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521004316.

Marijuan:2023:NSS

[MP23a] C. Marijuán and M. Pisonero. A note for the SNIEP in size 5. Linear Algebra and its Applications, 675(??):338–343, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002616.

Moon:2023:SNF

[MP23b] Sunyo Moon and Seungkook Park. The Smith normal form of the walk matrix of the extended Dynkin graph D̃_{n}. Linear Algebra and its Applications, 678 (??):169–190, December 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003014.

Myskova:2024:RIM

[MP24] Helena Myšková and Ján Plavka. Regularity of interval max-plus matrices. Linear Algebra and its Applications, 680(??):28–44, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300335X.

Mainar:2022:ACW

[MPR22] E. Mainar, J. M. Peña, and B. Rubio. Accurate computations with Wronskian matrices of Bessel and Laguerre polynomials. Linear Algebra and its Applications, 647(??):31–46, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001495.

Mainar:2024:ABF

[MPR24] E. Mainar, J. M. Peña, and B. Rubio. Accurate bidiagonal factorization of quantum Hilbert matrices. Linear Algebra and its Applications, 681(??):131–149, January 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004068.

Morfonios:2021:CPG

[MPRS21] C. V. Morfonios, M. Pyzh, M. Röntgen, and P. Schmelcher. Cospectrality preserving graph modifications and eigenvector properties via walk equivalence of vertices. Linear Algebra and its Applications, 624(??):53–86, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379521001610.

Manjegani:2022:AFI

[MPS22] S. M. Manjegani, A. Peperko, and H. Shokooh Saljooghi. Asymptotic formulae and inequalities for point spectrum in max algebra. Linear Algebra and its Applications, 634(??):112–136, February 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003967.

Mufid:2024:SLE

[MPS24] Muhammad Syifa’ul Mufid, Ebrahim Patel, and Sergeĭ Sergeev. Solving linear equations over maxmin-ω systems. Linear Algebra and its Applications, 681(??):21–46, January 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003865.

Miao:2020:GTF

[MQW20] Yun Miao, Liqun Qi, and Yimin Wei. Generalized tensor function via the tensor singular value decomposition based on the T -product. Linear Algebra and its Applications, 590(??):258–303, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379519305567.

Martinez-Rivera:2021:MWI

[MR21a] Xavier Martínez-Rivera. On 0–1 matrices whose inverses have entries of the same modulus. Linear Algebra and its Applications, 622(??):333–354, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001488.

Monsalve:2021:ESR

[MR21b] Juan Monsalve and Juan Rada. Extremal spectral radius of graphs with rank 4. Linear Algebra and its Applications, 609(??):1–11, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303906.

Miro-Roig:2020:WLP

[MRT20] Rosa M. Miró-Roig and Quang Hoa Tran. The weak Lefschetz property of Gorenstein algebras of codimension three associated to the Apéry sets. Linear Algebra and its Applications, 604(??):346–369, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Mendoza:2020:SMH

[MRV20] Alberto Mendoza, Lázaro Recht, and Alejandro Varela. Supports for minimal Hermitian matrices. Linear Algebra and its Applications, 584(??):458–482, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437951930401X.

Marcoux:2020:DCS

[MRZ20] Laurent W. Marcoux, Heydar Radjavi, and Yuanhang Zhang. Off-diagonal corners of subalgebras of L(C^{n}). Linear Algebra and its Applications, 607(??):58–88, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303657.

Marcoux:2022:DRS

[MRZ22] Laurent W. Marcoux, Heydar Radjavi, and Yuanhang Zhang. Dispersing representations of semi-simple subalgebras of complex matrices. Linear Algebra and its Applications, 642(??):160–220, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S002437952200060X.

McKee:2020:SMQ

[MS20a] James McKee and Chris Smyth. Symmetrizable matrices, quotients, and the trace problem. Linear Algebra and its Applications, 600(??):60–81, September 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301919.

Mitchell:2020:RPC

[MS20b] Lon Mitchell and Irene Sciriha. A reduction procedure for the Colin de Verdière number of a graph. Linear Algebra and its Applications, 596(??):36–48, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301063.

Majdak:2021:CEL

[MS21] Witold Majdak and Laurian Suciu. Convex and expansive liftings close to two-isometries and power bounded operators. Linear Algebra and its Applications, 617(??):1–26, May 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000203.

Matsuda:2022:IGO

[MS22] Takeru Matsuda and Tasuku Soma. Information geometry of operator scaling. Linear Algebra and its Applications, 649(??):240–267, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200180X.

Molnar:2024:SAP

[MS24] Lajos Molnár and Richárd Simon. On some algebraic properties related to Heron type operator means on positive definite cones of C^{⋆}-algebras. Linear Algebra and its Applications, 685(??):214–246, March 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004809.

Mondal:2024:CTP

[MST24] Samir Mondal, K. C. Sivakumar, and Michael Tsatsomeros. The Cayley transform of prevalent matrix classes. Linear Algebra and its Applications, 681(??):1– 20, January 15, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437952300397X.

Momihara:2022:SMM

[MSX22] Koji Momihara, Sho Suda, and Qing Xiang. Square (1, −1)-matrices with large determinants and near-Hadamard matrices. Linear Algebra and its Applications, 654(??):28–55, December 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003123.

Massey:2021:SSH

[MSZ21] Pedro Massey, Demetrio Stojanoff, and Sebastián Zárate. The spectral spread of Hermitian matrices. Linear Algebra and its Applications, 616(??):19–44, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520306030.

Morozov:2023:ORA

[MSZ23] Stanislav Morozov, Matvey Smirnov, and Nikolai Zamarashkin. On the optimal rank-1 approximation of matrices in the Chebyshev norm. Linear Algebra and its Applications, 679(??):4–29, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003373.

Manuilov:2021:KTT

[MT21a] Vladimir Manuilov and Evgenij Troitsky. On Kuiper type theorems for uniform Roe algebras. Linear Algebra and its Applications, 608(??):387–398, January 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304523.

Medina:2021:UBL

[MT21b] Luis Medina and Macarena Trigo. Upper bounds and lower bounds for the spectral radius of Reciprocal Distance, Reciprocal Distance Laplacian and Reciprocal Distance signless Laplacian matrices. Linear Algebra and its Applications, 609(??):386–412, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304468.

McCullough:2022:RFS

[MT22] Scott McCullough and Nicole Tuovila. Reinhardt free spectrahedra. Linear Algebra and its Applications, 640 (??):91–117, May 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379522000234.

Mitamura:2023:MUB

[MT23] Takuma Mitamura and Akira Tanaka. Minimum upper bound of the Cayley transform of an orthogonal matrix multiplied by signed permutation matrices. Linear Algebra and its Applications, 659(??):24–32, February 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004153.

Muranov:2022:PCH

[Mur22] Alexey Muranov. Proof of Cayley–Hamilton theorem using polynomials over the algebra of module endomorphisms. Linear Algebra and its Applications, 645(??):165–169, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001136.

Meulewaeter:2024:UHV

[MV24] Jeroen Meulewaeter and Hendrik Van Maldeghem. 3uniform hypergraphs from vector spaces. Linear Algebra and its Applications, 690 (??):91–111, June 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379524000922.

McDonald:2023:DSS

[MvWC23] André M. McDonald, Michaël A. van Wyk, and Guanrong Chen. A direct solution to the stochastic inverse eigenvalue problem for complexvalued eigenspectra. Linear Algebra and its Applications, 658(??):262–282, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004050.

Moslehian:2020:SNR

[MXZ20] M. S. Moslehian, Q. Xu, and A. Zamani. Seminorm and numerical radius inequalities of operators in semi-Hilbertian spaces. Linear Algebra and its Applications, 591(??):299–321, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300197.

Mroz:2022:WRE

[MZ22] Andrzej Mróz and Katarzyna Zaja̧c. Weyl roots and equivalences of integral quadratic forms. Linear Algebra and its Applications, 650 (??):210–235, October 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002233.

Mulas:2024:TNG

[MZZ24] Raffaella Mulas, Dong Zhang, and Giulio Zucal. There is no going back: Properties of the non-backtracking Laplacian. Linear Algebra and its Applications, 680(??):341–370, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003889.

Nagano:2023:RTR

[Nag23] Atsuhira Nagano. On Riemann type relations for theta functions on bounded symmetric domains of type I. Linear Algebra and its Applications, 673(??):177–199, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001805.

Najafi:2020:SNR

[Naj20] H. Najafi. Some numerical radius inequality for several operators. Linear Algebra and its Applications, 588(??):489–496, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304938.

Nardozza:2023:DPI

[Nar23]  Vincenzo Nardozza.  Dif-

ferential polynomial identities of upper triangular matrices of size three. Linear Algebra and its Applications, 663(??):142–177, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000071.

Nayak:2023:SFY

[Nay23]  Soumyashant Nayak.  A

stronger form of Yamamoto’s theorem on singular values. Linear Algebra and its Applications, 679(??):231– 245, December 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379523003543.

Novanta:2021:QIG

[NdLO21] Anderson Fernandes Novanta, Leonardo de Lima, and Carla Silva Oliveira. Qintegral graphs with at most two vertices of degree greater than or equal to three. Linear Algebra and its Applications, 614(??):144–163, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301555.

Neri:2020:SEG

[Ner20]  Alessandro Neri.  System-

atic encoders for generalized Gabidulin codes and the qanalogue of Cauchy matrices. Linear Algebra and its Applications, 593(??):116–149, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300616.

Nadisic:2023:SSN

[NGK23] Nicolas Nadisic, Nicolas Gillis, and Christophe Kervazo. Smoothed separable nonnegative matrix factorization. Linear Algebra and its Applications, 676(??):174–204, November 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379523002732.

Nica:2020:CRL

[Nic20] Bogdan Nica. Cut ratios and Laplacian eigenvalues. Linear Algebra and its Applications, 593(??):18–28, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300422.

Nica:2024:INR

[Nic24] Bogdan Nica. On the independence number of regular graphs of matrix rings. Linear Algebra and its Applications, 681(??):89–96, January 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004019.

Niezgoda:2024:FRC

[Nie24]  Marek Niezgoda.  Further

results on the concavity of operator means. Linear Algebra and its Applications, 690(??):186–198, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000934.

Niv:2020:OAS

[NMS20] Adi Niv, Marie MacCaig, and Sergei Sergeev. Optimal assignments with supervisions. Linear Algebra and its Applications, 595(??):72–100, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030104X.

Nguyen:2024:PSI

[NN24] Van-Bong Nguyen and Thi Ngan Nguyen. Positive semidefinite interval of matrix pencil and its applications to the generalized trust region subproblems. Linear Algebra and its Applications, 680(??):371–390, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379523003890.

Neumayer:2020:RIN

[NNSS20] Sebastian Neumayer, Max Nimmer, Simon Setzer, and Gabriele Steidl. On the rotational invariant L_{1}-norm PCA. Linear Algebra and its Applications, 587(??):243– 270, February 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379519304756.

Neri:2020:ECL

[NPHT20] Alessandro Neri, Sven Puchinger, and Anna-Lena Horlemann-Trautmann. Equivalence and characterizations of linear rank-metric codes based on invariants. Linear Algebra and its Applications, 603(??):418–469, October 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303050.

Napp:2021:MBL

[NPkT21] Diego Napp, Raquel Pinto, Elif Saçıkara, and Marisa Toste. A matrix based list decoding algorithm for linear codes over integer residue rings. Linear Algebra and its Applications, 614(??):376–393, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304535.

Napp:2022:SRC

[NPR22] Diego Napp, Raquel Pinto, and Conceição Rocha. State representations of convolutional codes over a finite ring. Linear Algebra and its Applications, 640(??):48–66, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004353.

Navarro-Perez:2024:EFC

[NPSE24] Miguel Ángel Navarro-Pérez and Xaro Soler-Escrivà. Equivalence for flag codes. Linear Algebra and its Applications, 690(??):1–26, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000867.

Nechita:2021:GCI

[NS21] Ion Nechita and Satvik Singh. A graphical calculus for integration over random diagonal unitary matrices. Linear Algebra and its Applications, 613(??):46–86, March 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305681.

Needham:2022:AFH

[NS22] Tom Needham and Clayton Shonkwiler. Admissibility and frame homotopy for quaternionic frames. Linear Algebra and its Applications, 645(??):237–255, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001318.

Noferini:2023:CRP

[NV23a] Vanni Noferini and Paul Van Dooren. On computing root polynomials and minimal bases of matrix pencils. Linear Algebra and its Applications, 658(??):86–115, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003901.

Noferini:2023:RVP

[NV23b] Vanni Noferini and Paul Van Dooren. Root vectors of polynomial and rational matrices: Theory and computation. Linear Algebra and its Applications, 656(??):510–540, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003718.

Nishida:2020:CBT

[NWW20] Yuki Nishida, Sennosuke Watanabe, and Yoshihide Watanabe. A characterization of bases of tropical kernels in terms of Cramer’s rule. Linear Algebra and its Applications, 601(??):301–310, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302561.

Neuman:2023:RRG

[NXS23] A. Martina Neuman, Yuying Xie, and Qiang Sun. Restricted Riemannian geometry for positive semidefinite matrices. Linear Algebra and its Applications, 665(??):153–195, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000320.

Nishida:2022:CTP

[NYW22] Yuki Nishida, Akira Yamada, and Yoshihide Watanabe. Characterization of tropical projective quadratic plane curves in terms of the eigenvalue problem. Linear Algebra and its Applications, 644(??):172–191, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522000994.

Nielsen:2024:MFR

[NZ24] Pace P. Nielsen and Yiqiang Zhou. Maximally-fine rings. Linear Algebra and its Applications, 680(??):45–64, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003683.

O:2021:SRM

[O.21] Suil O. Spectral radius and matchings in graphs. Linear Algebra and its Applications, 614(??):316–324, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302950.

O:2022:CDS

[O.22] Suil O. Colorings in digraphs from the spectral radius. Linear Algebra and its Applications, 654(??):1–9, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003056.

Oboudi:2022:SSL

[Obo22] Mohammad Reza Oboudi. On the smallest signless Laplacian eigenvalue of graphs. Linear Algebra and its Applications, 637(??):138–156, March 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004304.

Oboudi:2024:EGN

[Obo24] Mohammad Reza Oboudi. Energy of graphs with no eigenvalue in the interval (−1, 1). Linear Algebra and its Applications, 680(??):126–136, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003622.

Oliveira:2021:LGT

[OdFVM21] Allana S. S. Oliveira, Maria Aguieiras A. de Freitas, Cybele T. M. Vinagre, and Lilian Markenzon. On (k + 1)-line graphs of k-trees and their nullities. Linear Algebra and its Applications, 614(??):244–255, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301907.

OBrien:2020:ASB

[OJQ20] Cian O’Brien, Kevin Jennings, and Rachel Quinlan. Alternating signed bipartite graphs and difference-1 colourings. Linear Algebra and its Applications, 604(??):370–398, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303219.

OLoughlin:2023:SMR

[O’L23] Ryan O’Loughlin. Symmetric matrix representations of truncated Toeplitz operators on finite dimensional spaces. Linear Algebra and its Applications, 666(??):11–28, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000605.

Oliveira:2021:EGA

[Oli21] José Alves Oliveira. Equivalence, group of automorphism and invariants of a family of rank metric codes arising from linearized polynomials. Linear Algebra and its Applications, 630(??):274–292, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003086.

Omidvar:2020:BBN

[OM20] Mohsen Erfanian Omidvar and Hamid Reza Moradi. Better bounds on the numerical radii of Hilbert space operators. Linear Algebra and its Applications, 604 (??):265–277, November 1, 2020. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303128.

Onn:2021:LSO

[Onn21] Shmuel Onn. On line sum optimization. Linear Algebra and its Applications, 610(??):474–479, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304754.

Orera:2021:ADS

[OP21] H. Orera and J. M. Peña. Accurate determinants of some classes of matrices. Linear Algebra and its Applications, 630(??):1–14, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002858.

Ozaslan:2023:MIA

[OPA23] Ibrahim K. Ozaslan, Mert Pilanci, and Orhan Arikan. M-IHS: an accelerated randomized preconditioning method avoiding costly matrix decompositions. Linear Algebra and its Applications, 678(??):57–91, December 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379523003191.

Oh:2022:IGS

[OPPS22] Semin Oh, Jeong Rye Park, Jongyook Park, and Yoshio Sano. On Q-integral graphs with Q-spectral radius 6. Linear Algebra and its Applications, 654(??):267–288, December 1, 2022. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379522003159.

OBrien:2022:ASM

[OQ22] Cian O’Brien and Rachel Quinlan. Alternating sign matrices of finite multiplicative order. Linear Algebra and its Applications, 651 (??):332–358, October 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002178.

Omladic:2023:ACS

[ORŠ23] Matjaž Omladič, Heydar Radjavi, and Klemen Šivic. On approximate commutativity of spaces of matrices. Linear Algebra and its Applications, 676(??):251–266, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002513.

Ouyang:2020:FQC

[OSC20] Yingkai Ouyang, Yi Shen, and Lin Chen. Faster quantum computation with permutations and resonant couplings. Linear Algebra and its Applications, 592(??):270–286, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300240.

Ostermann:2023:PSP

[Ost23]  Maëva Ostermann.  Pseu-

dospectra and simultaneous power control. Linear Algebra and its Applications, 658(??):49–61, February 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003871.

ORourke:2023:QRB

[OW23] Sean O’Rourke and Philip Matchett Wood. Quantitative results for banded Toeplitz matrices subject to random and deterministic perturbations. Linear Algebra and its Applications, 657(??):50–126, January 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003810.

Pankov:2020:OPT

[Pan20]  Mark Pankov.  Orthogo-

nality preserving transformations of Hilbert Grassmannians. Linear Algebra and its Applications, 605 (??):180–189, November 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303426.

Paniello:2021:EOG

[Pan21] Irene Paniello. In-evolution operators in genetic coalgebras. Linear Algebra and its Applications, 614(??):197–207, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030166X.

Pankov:2022:CPT

[Pan22] Mark Pankov. Commutativity preserving transformations on conjugacy classes of compact self-adjoint operators. Linear Algebra and its Applications, 649(??):390–407, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002026.

Panja:2023:ACG

[Pan23] Saikat Panja. Acceptability of classical groups in non-zero characteristic. Linear Algebra and its Applications, 669(??):118–124, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001167.

Pate:2020:IMF

[Pat20] Thomas H. Pate. Independent multistructures, flag structures, and sets of linearly independent polar elements. Linear Algebra and its Applications, 589(??):103–129, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305373.

Petrullo:2021:PRA

[Pet21] P. Petrullo. Palindromic Riordan arrays, classical orthogonal polynomials and Catalan triangles. Linear Algebra and its Applications, 618(??):158–182, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000598.

Pirzada:2020:GDE

[PGRU20] S. Pirzada, Hilal A. Ganie, Bilal A. Rather, and Rezwan Ul Shaban. On generalized distance energy of graphs. Linear Algebra and its Applications, 603(??):1–19, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302603.

Paez-Guillan:2023:SLR

[PGST23] Pilar Páez-Guillán, Salvatore Siciliano, and David A. Towers. On the subalgebra lattice of a restricted Lie algebra. Linear Algebra and its Applications, 660(??):47–65, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004311.

Phung:2024:LNU

[Phu24] Xuan Kien Phung. On linear non-uniform cellular automata: Duality and dynamics. Linear Algebra and its Applications, 688(??):78–103, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000557.

Pikul:2023:JBE

[Pik23]  Piotr Pikul.  Joint back-

ward extension property for weighted shifts on directed trees. Linear Algebra and its Applications, 665(??):253–290, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

Pirzada:2021:DLS

[PK21] S. Pirzada and Saleem Khan. On distance Laplacian spectral radius and chromatic number of graphs. Linear Algebra and its Applications, 625(??):44–54, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001853.

Petkov:2022:NJF

[PK22] Petko H. Petkov and Mihail M. Konstantinov. The numerical Jordan form. Linear Algebra and its Applications, 638(??):1–45, April 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004493.

Poirion:2023:RPL

[PLT23] Pierre-Louis Poirion, Bruno F. Lourenço, and Akiko Takeda. Random projections of linear and semidefinite problems with linear inequalities. Linear Algebra and its Applications, 664(??):24–60, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379523000198.

Poornima:2023:DPC

[PM23] P. Poornima and K. Murugesan. Decomposable pairs and construction of multiwavelets. Linear Algebra and its Applications, 665(??):382–403, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000599.

Polyrakis:2020:NPL

[Pol20] Ioannis A. Polyrakis. The NMF problem and latticesubspaces. Linear Algebra and its Applications, 602(??):1–32, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302147.

Poon:2020:PRU

[Poo20]  Edward Poon.  Preservers

of restricted unitary orbits of projections. Linear Algebra and its Applications, 587(??):111–142, February 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304719.

Pak:2020:BKC

[PP20] Igor Pak and Greta Panova. Bounds on Kronecker coefficients via contingency tables. Linear Algebra and its Applications, 602(??):157–178, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302366.

Pagacz:2020:BNI

[PPW20] Patryk Pagacz, Paweł Pietrzycki, and Michał Wojtylak. Between the von Neumann inequality and the Crouzeix conjecture. Linear Algebra and its Applications, 605 (??):130–157, November 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303256.

Pankov:2021:GGG

[PPŻ21] Mark Pankov, Krzysztof Petelczyc, and Mariusz Żynel. Generalized Grassmann graphs associated to conjugacy classes of finite-rank self-adjoint operators. Linear Algebra and its Applications, 627(??):1–23, October 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002391.

Pankov:2023:AGC

[PPŻ23] Mark Pankov, Krzysztof Petelczyc, and Mariusz Żynel. Automorphisms of graphs corresponding to conjugacy classes of finite-rank selfadjoint operators. Linear Algebra and its Applications, 674(??):192–207, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002136.

Perez:2022:LRM

[PQ22] Javier Pérez and María C. Quintana. Linearizations of rational matrices from general representations. Linear Algebra and its Applications, 647(??):89–126, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001525.

Parzygnat:2022:NCB

[PR22a] Arthur J. Parzygnat and Benjamin P. Russo. A noncommutative Bayes’ theorem. Linear Algebra and its Applications, 644(??):28–94, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000805.

Petek:2022:PUG

[PR22b] Tatjana Petek and Gordana Radić. Preservers of unitary group or equivalence by unitaries. Linear Algebra and its Applications, 633(??):1–22, January 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003578.

Pranesh:2020:BEC

[Pra20] Srikara Pranesh. Backward error and condition number of a generalized Sylvester equation, with application to the stochastic Galerkin method. Linear Algebra and its Applications, 594(??):95–116, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300847.

Preusser:2020:RDU

[Pre20] Raimund Preusser. Reverse decomposition of unipotents over noncommutative rings I: General linear groups. Linear Algebra and its Applications, 601(??):285–300, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302512.

Price:2022:SST

[Pri22] Geoffrey L. Price. On skewsymmetric Toeplitz matrices over finite fields with periodicity conditions. Linear Algebra and its Applications, 642(??):101–117, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000519.

Paras:2020:PIS

[PS20] Agnes T. Paras and Jenny R. Salinasan. The product of an involution and a skewinvolution. Linear Algebra and its Applications, 584(??):431–437, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304240.

Prajapati:2022:ORQ

[PS22] Anshul Prajapati and Punit Sharma. Optimizing the Rayleigh quotient with symmetric constraints and its application to perturbations of structured polynomial eigenvalue problems. Linear Algebra and its Applications, 645(??):256–277, July 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379522001173.

Pavlikova:2023:MPP

[PŠ23a] Soňa Pavlíková and Daniel Ševčovič. On the Moore– Penrose pseudo-inversion of block symmetric matrices and its application in the graph theory. Linear Algebra and its Applications, 673(??):280–303, September 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001994.

Porcelli:2023:NSC

[PS23b] Margherita Porcelli and Valeria Simoncini. Numerical solution of a class of quasilinear matrix equations. Linear Algebra and its Applications, 664(??):349–368, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000307.

Pickmann-Soto:2020:IME

[PSAPNV20] H. Pickmann-Soto, S. Arela-Pérez, Hans Nina, and Elvis Valero. Inverse maximal eigenvalues problems for Leslie and doubly Leslie matrices. Linear Algebra and its Applications, 592(??):93–112, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300239.

Pellegrino:2020:UMF

[PSRS20] Daniel Pellegrino, Diana Serrano-Rodríguez, and Janiely Silva. On unimodular multilinear forms with small norms on sequence spaces. Linear Algebra and its Applications, 595(??):24–32, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300999.

Patnaik:2023:ASI

[PSW23] Sasmita Patnaik, Sanehlata, and Gary Weiss. Automatic selfadjoint-ideal semigroups for finite matrices. Linear Algebra and its Applications, 676(??):211–240, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002677.

Pinasco:2021:OPA

[PSZ21] Damián Pinasco, Ezequiel Smucler, and Ignacio Zalduendo. Orthant probabilities and the attainment of maxima on a vertex of a simplex. Linear Algebra and its Applications, 610(??):785–803, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304961.

Pascoe:2022:MPP

[PTD22] J. E. Pascoe and Ryan Tully-Doyle. Monotonicity of the principal pivot transform. Linear Algebra and its Applications, 643(??):161–165, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000593.

Pitrik:2021:DCI

[PV21a] József Pitrik and Dániel Virosztek. A divergence center interpretation of general symmetric Kubo–Ando means, and related weighted multivariate operator means. Linear Algebra and its Applications, 609(??):203–217, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304286.

Podesta:2021:IEN

[PV21b] Ricardo A. Podestá and Denis E. Videla. Integral equienergetic non-isospectral unitary Cayley graphs. Linear Algebra and its Applications, 612(??):42–74, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S002437952030553X.

Pelantova:2022:PRI

[PV22] Edita Pelantová and Tomáš Vávra. On positional representation of integer vectors. Linear Algebra and its Applications, 633(??):316–331, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003839.

Price:2019:STM

[PW19] Geoffrey Price and Myles Wortham. On sequences of Toeplitz matrices over finite fields. Linear Algebra and its Applications, 561(??):63–80, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379518304464. See cor-

rigendum [PW20].

Price:2020:CST

[PW20] Geoffrey Price and Myles Wortham. Corrigendum to: “On sequences of Toeplitz matrices over finite fields” [Linear Algebra Appl. 561 (2019) 63–80]. Linear Algebra and its Applications, 590(??):330–332, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/

S0024379520300124.  See

[PW19].

Poon:2021:IMC

[PW21] Edward Poon and Hugo J. Woerdeman. Isospectrality and matrices with concentric circular higher rank numerical ranges. Linear Algebra and its Applications, 631(??):174–180, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003220.

Pei:2023:PPT

[PWW23] Yanni Pei, Yaling Wang, and Yi Wang. Positivity problem of three-term recurrence sequences. Linear Algebra and its Applications, 668(??):93–109, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001052.

Polverino:2020:NRS

[PZ20] Olga Polverino and Ferdinando Zullo. On the number of roots of some linearized polynomials. Linear Algebra and its Applications, 601(??):189–218, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379520302408.

Qian:2021:MIE

[QCCS21] Lilong Qian, Lin Chen, Delin Chu, and Yi Shen. A matrix inequality for entanglement distillation problem. Linear Algebra and its Applications, 616(??):139–177, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000173.

Qiu:2023:GSC

[QJMW23] Lihong Qiu, Yizhe Ji, Lihuan Mao, and Wei Wang. Generalized spectral characterizations of regular graphs based on graph-vectors. Linear Algebra and its Applications, 663(??):116–141, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952300006X.

Qiu:2020:TGM

[QJW20] Lihong Qiu, Yizhe Ji, and Wei Wang. On a theorem of Godsil and McKay concerning the construction of cospectral graphs. Linear Algebra and its Applications, 603(??):265–274, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379520302639.

Qin:2020:DEC

[QLCM20] Rui Qin, Dan Li, Yuanyuan Chen, and Jixiang Meng. The distance eigenvalues of the complements of unicyclic graphs. Linear Algebra and its Applications, 598(??):49–67, August 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301683.

Qiu:2024:GSC

[QQM24] Lihong Qiu, Guangkun Qiu, and Lihuan Mao. Generalized spectral characterizations of a new family of noncontrollable graphs. Linear Algebra and its Applications, 683(??):211–234, February 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004561.

Qi:2022:PPN

[QSH22] Xiaofei Qi, Shaoxing Sun, and Jinchuan Hou. Properties and preservers of numerical radius on skew Lie products of operators. Linear Algebra and its Applications, 635(??):24–68, February 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521004079.

Quinlan:2022:LDE

[QSS22] R. Quinlan, M. Shau, and F. Szechtman. Linear Diophantine equations in several variables. Linear Algebra and its Applications, 640(??):67–90, May 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379521004614.

Qiu:2021:OGD

[QWW21] Lihong Qiu, Wei Wang, and Wei Wang. Oriented graphs determined by their generalized skew spectrum. Linear Algebra and its Applications, 622(??):316–332, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001373.

Rai:2023:BML

[Rai23]  Pradeep K. Rai.  Bogo-

molov multiplier of Lie algebras. Linear Algebra and its Applications, 672 (??):122–133, September 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001659.

Rahjoo:2021:BQN

[RAK21] Meysam Rahjoo, Gholamreza Aghamollaei, and Hossein Momenaee Kermani. On the boundedness of quaternionic numerical ranges with respect to nonstandard involutions. Linear Algebra and its Applications, 610(??):59–72, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030464X.

Ramanantoanina:2023:CLO

[Ram23] Andriamanankasina Ramanantoanina. Characterisation of the lattice operations on positive operators with respect to the spectral order. Linear Algebra and its Applications, 656(??):446–462, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003688.

Ran:2022:NSE

[Ran22] A. C. M. Ran. A note on the structure of expansive matrices in indefinite inner product spaces. Linear Algebra and its Applications, 655(??):291–301, December 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379522003445.

Rani:2023:UBS

[Ran23] Sonu Rani. Upper bounds on the smallest positive eigenvalue of trees with at most one zero eigenvalue. Linear Algebra and its Applications, 659(??):42–54, February 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004190.

Rao:2021:BAS

[Rao21] T. S. S. R. K. Rao. Best approximation in spaces of compact operators. Linear Algebra and its Applications, 627(??):72–79, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100241X.

Roy:2022:LEQ

[RB22] Nandita Roy and Shreemayee Bora. Locating eigenvalues of quadratic matrix polynomials. Linear Algebra and its Applications, 649(??):452–490, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002014.

Rahim:2023:RCF

[RC23] Fahim Rahim and Nicholas J. Cavenagh. Row–column factorial designs with strength at least 2. Linear Algebra and its Applications, 667(??):44–70, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000733.

Reyes:2023:STG

[RDFM23] M. A. Reyes, C. Dalfó, M. A. Fiol, and A. Messegué. On the spectra of token graphs of cycles and other graphs. Linear Algebra and its Applications, 679(??):38–66, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003348.

Reis:2023:CDF

[Rei23] Lucas Reis. Counting distinct functional graphs from linear finite dynamical systems. Linear Algebra and its Applications, 656(??):409–420, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003706.

Reese:2022:EEP

[RFP22] Tyler M. Reese, Joseph D. Fehribach, and Randy C. Paffenroth. Equitable edge partitions and Kirchhoff graphs. Linear Algebra and its Applications, 639(??):225–242, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200009X.

Rizzo:2023:DCE

[Riz23]  Carla Rizzo.  Differential

codimensions and exponential growth. Linear Algebra and its Applications, 675(??):294–311, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002525.

Roy:2021:APB

[RKBA21] Nandita Roy, Michael Karow, Shreemayee Bora, and Gorka Armentia. Approximation of pseudospectra of block triangular matrices. Linear Algebra and its Applications, 623(??):398–419, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303815.

Rubiano-Murcia:2024:HOD

[RMG24] Jorge I. Rubiano-Murcia and Juan Galvis. Higher order derivatives of the adjugate matrix and the Jordan form. Linear Algebra and its Applications, 680(??):137–155, January 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003804.

Ramos:2021:MID

[RMP21] C. Correia Ramos, N. Martins, and P. R. Pinto. Markov invariant dynamics. Linear Algebra and its Applications, 620(??):268–296, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000951.

Robol:2020:RKA

[Rob20] Leonardo Robol. Rational Krylov and ADI iteration for infinite size quasi-Toeplitz matrix equations. Linear Algebra and its Applications, 604(??):210–235, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303049.

Rocha:2020:BCH

[Roc20]  Israel Rocha.  Brouwer’s

conjecture holds asymptotically almost surely. Linear Algebra and its Applications, 597(??):198–205, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301476.

Rodriguez:2021:RAS

[Rod21] Jorge Tomás Rodríguez. On the rank and the approximation of symmetric tensors. Linear Algebra and its Applications, 628(??):72–102, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002603.

Rosenfeld:2022:IUW

[Ros22]  Matthieu Rosenfeld.  It

is undecidable whether the growth rate of a given bilinear system is 1. Linear Algebra and its Applications, 651(??):131–143, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002452.

Rowlinson:2020:PCG

[Row20] Peter Rowlinson. A problem concerning graphs with just three distinct eigenvalues. Linear Algebra and its Applications, 592(??):260–269, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300343.

Roy:2024:WDN

[Roy24] Saikat Roy. The weak differentiability of norm and a generalized Bhatia–Šemrl theorem. Linear Algebra and its Applications, 685(??):46–65, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004755.

Reza-Rahmati:2023:ETF

[RRF23] Mohammad Reza-Rahmati and Gerardo Flores. An extended trace formula for vertex operators. Linear Algebra and its Applications, 661(??):1–19, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004281.

Reitzner:2024:CML

[RRvW24] Matthias Reitzner, Tim Römer, and Mandala von Westenholz. Covariance matrices of length power functionals of random geometric graphs — an asymptotic analysis. Linear Algebra and its Applications, 691(??):151–181, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001289.

Roitberg:2021:IBD

[RS21a] Inna Roitberg and Alexander Sakhnovich. On the inversion of the block doublestructured and of the triplestructured Toeplitz matrices and on the corresponding reflection coefficients. Linear Algebra and its Applications, 610(??):506–528, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030478X.

Rowlinson:2021:SGT

[RS21b] Peter Rowlinson and Zoran Stanić. Signed graphs with three eigenvalues: Biregularity and beyond. Linear Algebra and its Applications, 621(??):272–295, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001154.

Roy:2021:OBF

[RSS21] Saikat Roy, Tanusri Senapati, and Debmalya Sain. Orthogonality of bilinear forms and application to matrices. Linear Algebra and its Applications, 615(??):104–111, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379521000033.

Rumynin:2021:RRG

[RT21] Dmitriy Rumynin and James Taylor. Real representations of C_{2}-graded groups: the antilinear theory. Linear Algebra and its Applications, 610(??):135–168, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304687.

Raposo:2023:VVK

[RT23] Anselmo Raposo and Katiuscia B. Teixeira. Vector-valued Kahane–Salem–Zygmund inequalities with asymptotically bounded constants. Linear Algebra and its Applications, 672(??):134–146, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001349.

Rubei:2021:IMR

[Rub21] Elena Rubei. Interval matrices: Realization of ranks by rational matrices. Linear Algebra and its Applications, 608(??):299–321, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304390.

Rubei:2022:ASM

[Rub22] Elena Rubei. Affine subspaces of matrices with constant rank. Linear Algebra and its Applications, 644(??):259–269, July 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000908.

Rump:2022:ELU

[Rum22] S. M. Rump. Entrywise lower and upper bounds for the Perron vector. Linear Algebra and its Applications, 653(??):314–319, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002889.

Ruzickova:2021:DRM

[Ruž21] Viera Štoudková Ružičková. Discrete Riccati matrix equation and the order preserving property. Linear Algebra and its Applications, 618(??):58–75, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000409.

Rosa:2021:CSR

[RW21a] Kennett L. Dela Rosa and Hugo J. Woerdeman. Continuity of submatrices and Ritz sets associated to a point in the numerical range. Linear Algebra and its Applications, 624(??):1–13, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100135X.

Rozanski:2021:HOI

[RW21b] Michał Różański and Roman Wituła. Hadamard’s optimized inequality. Linear Algebra and its Applications, 620(??):109–123, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000963.

Rue:2022:NLD

[RW22] Juanjo Rué and Maximilian Wötzel. Normal limiting distributions for systems of linear equations in random sets. Linear Algebra and its Applications, 649(??):122–151, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001902.

Roy:2021:SPI

[RX21] Sandip Roy and Mengran Xue. Sign patterns of inverse doubly-nonnegative matrices. Linear Algebra and its Applications, 610 (??):480–487, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304766.

Reich:2024:CMA

[RZ24] Simeon Reich and Rafał Zalas. Comparing the methods of alternating and simultaneous projections for two subspaces. Linear Algebra and its Applications, 683(??):235–263, February 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004573.

Sanchez:2021:LIB

[Sán21] Javier Cabello Sánchez. Linearity of isometries between convex Jordan curves. Linear Algebra and its Applications, 621(??):1–17, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000744.

Sano:2023:MPM

[San23] Takashi Sano. Matrix perspectives and matrix convexity. Linear Algebra and its Applications, 666(??):29–36, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000654.

Sarty:2021:NRC

[Sar21] Gordon E. Sarty. Natural reconstruction coordinates for imperfect TRASE MRI. Linear Algebra and its Applications, 611(??):94–117, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305553.

Sasane:2021:SEB

[Sas21] Amol Sasane. The Sylvester equation in Banach algebras. Linear Algebra and its Applications, 631(??):1–9, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003128.

Satake:2021:RIP

[Sat21] Shohei Satake. On the restricted isometry property of the Paley matrix. Linear Algebra and its Applications, 631(??):35–47, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003153.

Sarkar:2020:JHT

[SB20] Amitesh Sarkar and Anirban Banerjee. Joins of hypergraphs and their spectra. Linear Algebra and its Applications, 603(??):101–129, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302676.

Sheriff:2020:CCO

[SC20a] Mohammed Rayyan Sheriff and Debasish Chatterjee. A complete characterization of optimal dictionaries for least squares representation. Linear Algebra and its Applications, 601(??):219–264, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302421.

Somodi:2020:IZF

[SC20b] Marius Somodi and Danny Cheuk. Ihara zeta functions of coronae. Linear Algebra and its Applications, 601(??):1–16, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301828.

Sun:2023:ENL

[SC23] Shaowei Sun and Mengsi Chen. Extremal normalized Laplacian spectral radii of graphs. Linear Algebra and its Applications, 679 (??):261–274, December 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003658.

Schonlein:2022:FEU

[Sch22]  Michael Schönlein.  Feed-

back equivalence and uniform ensemble reachability. Linear Algebra and its Applications, 646(??):175–194, August 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001343.

Schweitzer:2023:IRH

[Sch23]  Marcel Schweitzer.  Inte-

gral representations for higherorder Fréchet derivatives of matrix functions: Quadrature algorithms and new results on the level-2 condition number. Linear Algebra and its Applications, 656(??):247–276, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003573.

Schulz:2024:BAM

[Sch24] François Schulz. Banach algebra mappings preserving the invertibility of linear pencils. Linear Algebra and its Applications, 691(??):109–122, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524001009.

Sun:2020:CSR

[SD20] Shaowei Sun and Kinkar Chandra Das. A conjecture on the spectral radius of graphs. Linear Algebra and its Applications, 588(??):74–80, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930504X.

Sun:2021:CPC

[SD21a] Shaowei Sun and Kinkar Chandra Das. Corrigendum to “Proof of a conjecture on distance energy change of complete multipartite graph due to edge deletion” [Linear Algebra Appl. 611 (2021) 253– 259]. Linear Algebra and its Applications, 618(??):203–204, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100104X.  See

[SD21d].

Sun:2021:MNL

[SD21b] Shaowei Sun and Kinkar Chandra Das. On the multiplicities of normalized Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 609(??):365–385, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304444.

Sun:2021:PDC

[SD21c] Shaowei Sun and Kinkar Chandra Das. Proof and disproof of conjectures on spectral radii of coclique extension of cycles and paths. Linear Algebra and its Applications, 618(??):1–11, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000380.

Sun:2021:PCD

[SD21d] Shaowei Sun and Kinkar Chandra Das. Proof of a conjecture on distance energy change of complete multipartite graph due to edge deletion. Linear Algebra and its Applications, 611(??):253–259, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305127. See cor-

rigendum [SD21a].

Sorensen:2021:BFS

[SDS21] Mikael Sørensen, Lieven De Lathauwer, and Nicholaos D.

Sidiropoulos. Bilinear factorizations subject to monomial equality constraints via tensor decompositions. Linear Algebra and its Applications, 621(??):296–333, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001191.

Sahu:2023:CAA

[SDTR23] Atal Narayan Sahu, Aritra Dutta, Aashutosh Tiwari, and Peter Richtárik. On the convergence analysis of asynchronous SGD for solving consistent linear systems. Linear Algebra and its Applications, 663(??):1–31, April 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004517.

Schork:2020:RRG

[SG20] Lukas Schork and Jacek Gondzio. Rank revealing Gaussian elimination by the maximum volume concept. Linear Algebra and its Applications, 592(??):1–19, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305580.

Shao:2021:DBE

[SGGZ21] Yanling Shao, Yubin Gao, Wei Gao, and Xuan Zhao. Degreebased energies of trees. Linear Algebra and its Applications, 621(??):18–28, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001063.

Salehi:2020:HOS

[SH20] Maryam Salehi and Karim Hedayatian. On higher order selfadjoint operators. Linear Algebra and its Applications, 587(??):358–386, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304859.

Sepitka:2021:CIL

[ŠH21a] Peter Šepitka and Roman Šimon Hilscher. Comparative index and Lidskii angles for symplectic matrices. Linear Algebra and its Applications, 624(??):174–197, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001695.

Sepitka:2021:DNF

[ŠH21b] Peter Šepitka and Roman Šimon Hilscher. Distribution and number of focal points for linear Hamiltonian systems. Linear Algebra and its Applications, 611(??):26–45, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305437.

Shah:2022:DBM

[Sha22] Rikhav Shah. Determinants of binary matrices achieve every integral value up to Ω(2^{n}/n). Linear Algebra and its Applications, 645(??):229–236, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952200132X.

Shi:2020:CSO

[Shi20a] Guanghua Shi. Concavity of some operator functions and Fréchet differential mappings. Linear Algebra and its Applications, 590(??):224–234, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305452.

Shitov:2020:RRS

[Shi20b] Yaroslav Shitov. A remark on ranks of sign patterns. Linear Algebra and its Applications, 593(??):150–151, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300550.

Shi:2021:VRR

[Shi21a] Guanghua Shi. Variational representations related to quantum Rényi relative entropies. Linear Algebra and its Applications, 610 (??):257–273, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304729.

Shitov:2021:CSS

[Shi21b] Yaroslav Shitov. Column subset selection is NP-complete. Linear Algebra and its Applications, 610(??):52–58, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304377.

Shitov:2021:DCT

[Shi21c] Yaroslav Shitov. The determinants of certain (0, 1) Toeplitz matrices. Linear Algebra and its Applications, 618(??):150–157, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000549.

Shitov:2022:EPF

[Shi22]  Yaroslav Shitov.  On the

extreme points of a family of matrices related to a theorem of Birkhoff. Linear Algebra and its Applications, 632(??):233–240, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003499.

Siconolfi:2021:RCG

[Sic21] Viola Siconolfi. Ricci curvature, graphs and eigenvalues. Linear Algebra and its Applications, 620(??):242–267, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000938.

Simson:2020:CTC

[Sim20] Daniel Simson. A computational technique in Coxeter spectral study of symmetrizable integer Cartan matrices. Linear Algebra and its Applications, 586 (??):190–238, February 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304495.

Simson:2021:CSC

[Sim21] Daniel Simson. A Coxeter spectral classification of positive edge-bipartite graphs. II. Dynkin type D_{n}. Linear Algebra and its Applications, 612(??):223–272, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305255.

Singla:2021:GDN

[Sin21] Sushil Singla. Gateaux derivative of C^{⋆} norm. Linear Algebra and its Applications, 629(??):208–218, November 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002846.

Saha:2021:CNN

[SK21a] Tanay Saha and Swanand R. Khare. Computation of the nearest non-prime polynomial matrix: Structured low-rank approximation approach. Linear Algebra and its Applications, 616(??):84–124, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000082.

Samanta:2021:BEC

[SK21b] Aniruddha Samanta and M. Rajesh Kannan. Bounds for the energy of a complex unit gain graph. Linear Algebra and its Applications, 612(??):1–29, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305449.

Skripka:2020:TBM

[Skr20] Anna Skripka. Tracial bounds for multilinear Schur multipliers. Linear Algebra and its Applications, 590(??):62–84, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305476.

Skripka:2021:MBE

[Skr21] Anna Skripka. MSE bounds for estimators of matrix functions. Linear Algebra and its Applications, 609 (??):231–252, January 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304158.

Sammour:2022:GAN

[SKS22] Samah Abu Sammour, Fuad Kittaneh, and Mohammad Sababheh. A geometric approach to numerical radius inequalities. Linear Algebra and its Applications, 652(??):1–17, November 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002592.

Sanchez:2020:EUD

[SL20a] Andrés David Báez Sánchez and Carlile Lavor. On the estimation of unknown distances for a class of Euclidean distance matrix completion problems with interval data. Linear Algebra and its Applications, 592(??):287–305, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030046X.

Sun:2020:SPZ

[SL20b] Wanting Sun and Shuchao Li. A short proof of Zhou, Wong and Sun’s conjecture. Linear Algebra and its Applications, 589(??):80–84, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305348.

Shi:2021:TCP

[SL21] Yuming Shi and Yan Liu. Trace class perturbation of closed linear relations. Linear Algebra and its Applications, 610(??):837–858, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520304997.

Sun:2022:NES

[SLL22] Gaoxing Sun, Feng Liu, and Kaiyang Lan. A note on eigenvalues of signed graphs. Linear Algebra and its Applications, 652(??):125–131, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002683.

Slowik:2020:SCE

[Sło20] R. Słowik. Some (counter)examples on totally positive Riordan arrays. Linear Algebra and its Applications, 594(??):117–123, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300938. See cor-

rigendum [Sło21a].

Slowik:2021:CSC

[Sło21a] R. Słowik. Corrigendum to “Some (counter)examples on totally positive Riordan arrays” [Linear Algebra Appl. 594 (2020) 117–123]. Linear Algebra and its Applications, 619(??):338–339, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001129.  See

[Sło20].

Slowik:2021:MAI

[Sło21b] R. Słowik. More about involutions in the group of almost-Riordan arrays. Linear Algebra and its Applications, 624(??):247–258, September 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001804.

Schopfer:2022:ERK

[SLTW22] Frank Schöpfer, Dirk A. Lorenz, Lionel Tondji, and Maximilian Winkler. Extended randomized Kaczmarz method for sparse least squares and impulsive noise problems. Linear Algebra and its Applications, 652 (??):132–154, November 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002610.

Sergeev:2022:ECS

[SlW22] Sergeĭ Sergeev and Hui li Wang. Extremality criteria for the supereigenvector space in maxplus algebra. Linear Algebra and its Applications, 653(??):116–134, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002828.

Shaked-Monderer:2023:SPI

[SM23] Naomi Shaked-Monderer. Sign patterns of inverse doubly nonnegative matrices and inverse completely positive matrices. Linear Algebra and its Applications, 659(??):10–23, February 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004165.

Sarkar:2020:PLQ

[SMA20] Rohit Sarma Sarkar, Amrita Mandal, and Bibhas Adhikari. Periodicity of lively quantum walks on cycles with generalized Grover coin. Linear Algebra and its Applications, 604(??):399–424, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303293.

Saravanan:2021:GFL

[SMA21] M. Saravanan, S. P. Murugan, and G. Arunkumar. A generalization of Fiedler’s lemma and the spectra of H-join of graphs. Linear Algebra and its Applications, 625(??):20–43, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001798.

Sosa:2024:FOS

[SMM24] Fredy Sosa, Julio Moro, and Christian Mehl. First order structure-preserving perturbation theory for eigenvalues of symplectic matrices: Part II. Linear Algebra and its Applications, 689(??):196–229, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000636.

Sain:2024:LPO

[SMP24] Debmalya Sain, Jayanta Manna, and Kallol Paul. On local preservation of orthogonality and its application to isometries. Linear Algebra and its Applications, 690(??):112–131, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000879.

Singh:2022:NSM

[SNK22] Gambheer Singh, S. K. Neogy, and Promila Kumar. A new subclass of Q_{0}-matrix in linear complementarity theory. Linear Algebra and its Applications, 647(??):64–77, August 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001628.

Soltan:2020:PSB

[Soł20] Piotr M. Sołtan. Podle’s spheres for the braided quantum SU(2). Linear Algebra and its Applications, 591(??):169–204, April 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030015X.

Sossa:2023:FTD

[Sos23]  David Sossa.  A Fiedler-

type determinantal inequality in Euclidean Jordan algebras. Linear Algebra and its Applications, 667(??):151–164, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000952.

Sra:2021:MIJ

[Sra21] Suvrit Sra. Metrics induced by Jensen–Shannon and related divergences on positive definite matrices. Linear Algebra and its Applications, 616(??):125–138, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305905.

Stampach:2020:HMC

[SS20] Frantisek Stampach and Pavel Stovícek. On Hankel matrices commuting with Jacobi matrices from the Askey scheme. Linear Algebra and its Applications, 591(??):235–267, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300203.

Saltenberger:2021:SHM

[SS21a] Philip Saltenberger and Michel-Niklas Senn. On skew-Hamiltonian matrices and their Krylov-Lagrangian subspaces. Linear Algebra and its Applications, 611 (??):66–81, February 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305528.

Seeger:2021:CCS

[SS21b] Alberto Seeger and David Sossa. On cardinality of complementarity spectra of connected graphs. Linear Algebra and its Applications, 614(??):5–23, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304884.

Sharma:2021:MIE

[SS21c] Debashish Sharma and Mausumi Sen. The minimax inverse eigenvalue problem for matrices whose graph is a generalized star of depth 2. Linear Algebra and its Applications, 621(??):334–344, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100118X.

Schmitt:2022:SRS

[SS22] Philipp Schmitt and Matthias Schötz. Symmetry reduction of states II: a noncommutative Positivstellensatz for CP^{n}. Linear Algebra and its Applications, 649(??):326–353, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001987.

Samanta:2023:BCO

[SS23a] Amit Samanta and Santanu Sarkar. Boundedness of composition operator on several variable Paley–Wiener space. Linear Algebra and its Applications, 660(??):66–79, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200430X.

Solymosi:2023:CMS

[SS23b] József Solymosi and Endre Szabó. Classification of maps sending lines into translates of a curve. Linear Algebra and its Applications, 668(??):161–172, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000940.

Sun:2023:HDD

[SS23c] Cheng Sun and Yidong Sun. On the halves of double and 3-dimensional Riordan arrays. Linear Algebra and its Applications, 679(??):194–219, December 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003397.

Shapiro:2024:RGP

[SS24a] Louis W. Shapiro and Minho Song. A Riordan group poset. Linear Algebra and its Applications, 689(??):60–85, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000582.

Stochel:2024:HCT

[SS24b] Jan Stochel and Jerzy Bartłomiej Stochel. The hyperbolic cosine transform and its applications to composition operators. Linear Algebra and its Applications, 691(??):1–36, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000880.

Saha:2022:LSU

[SSP22] S. S. Saha, K. Sharma, and S. K. Panda. On the Laplacian spectrum of kuniform hypergraphs. Linear Algebra and its Applications, 655(??):1–27, December 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003226.

Song:2021:CCB

[SSS21] Guanbang Song, Guifu Su, and Huichao Shi. A complete characterization of bidegreed split graphs with four distinct signless Laplacian eigenvalues. Linear Algebra and its Applications, 629 (??):232–245, November 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002974.

Szigeti:2020:PCH

[SSvW20] Jenő Szigeti, Szilvia Szilágyi, and Leon van Wyk. A power Cayley–Hamilton identity for n×n matrices over a Lie nilpotent ring of index k. Linear Algebra and its Applications, 584(??):153–163, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519303994.

Seeger:2020:SMI

[ST20] Alberto Seeger and Mounir Torki. On spectral maps induced by convex cones. Linear Algebra and its Applications, 592(??):65–92, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300264.

Sano:2022:KMM

[ST22] Takashi Sano and Kazuki Takeuchi. Kraus matrices and matrix convex functions. Linear Algebra and its Applications, 649(??):268– 272, September 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437952200194X.

Stanic:2020:LBL

[Sta20a] Zoran Stanić. Lower bounds for the least Laplacian eigenvalue of unbalanced blocks. Linear Algebra and its Applications, 584(??):145–152, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519303921.

Stanic:2020:OGW

[Sta20b] Zoran Stanić. Oriented graphs whose skew spectral radius does not exceed 2. Linear Algebra and its Applications, 603(??):359–367, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303098.

Starcic:2021:HCF

[Sta21] Tadej Starčič. Hong’s canonical form of a Hermitian matrix with respect to orthogonal *congruence. Linear Algebra and its Applications, 630(??):241–251, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002986.

Stanic:2023:DPD

[Sta23a]  Zoran Stanić.  Determi-

nation of particular double starlike trees by the Laplacian spectrum. Linear Algebra and its Applications, 672(??):182–194, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001696.

Stanic:2023:RBS

[Sta23b] Zoran Stanić. Relations between the skew spectrum of an oriented graph and the spectrum of an associated signed graph. Linear Algebra and its Applications, 676(??):241–250, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002756.

Steinerberger:2021:SAS

[Ste21a] Stefan Steinerberger. A spectral approach to the shortest path problem. Linear Algebra and its Applications, 620(??):182–200, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000720.

Stewart:2021:LOR

[Ste21b] Michael Stewart. A locally optimal rank revealing product decomposition. Linear Algebra and its Applications, 623(??):420–453, August 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000185.

Stechlinski:2022:GDE

[Ste22a] Peter Stechlinski. Generalized derivatives of eigenvalues of a symmetric matrix. Linear Algebra and its Applications, 649(??):63–95, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001707.

Ster:2022:NCI

[Šte22b] Janez Šter. Nil-clean index of M_{n}(F_{2}). Linear Algebra and its Applications, 632(??):294–307, January 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003669.

Steinerberger:2022:EPR

[StW22] Stefan Steinerberger and Hau tieng Wu. Eigenvector Phase Retrieval: Recovering eigenvectors from the absolute value of their entries. Linear Algebra and its Applications, 652(??):239–252, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002798.

Suda:2022:PCC

[Sud22] Sho Suda. Q-polynomial coherent configurations. Linear Algebra and its Applications, 643(??):166–195, June 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379522000520.

Shaul:2021:PPO

[SV21] Liran Shaul and Victor Vinnikov. Pole placement for overdetermined 2D systems. Linear Algebra and its Applications, 610(??):529–559, February 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379520304791.

Stefan:2021:SPS

[SW21] Anthony Stefan and Aaron Welters. A short proof of the symmetric determinantal representation of polynomials. Linear Algebra and its Applications, 627(??):80–93, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002421.

Szymanski:2023:SMP

[SW23] Oskar Jakub Szymański and Michał Wojtylak. Stability of matrix polynomials in one and several variables. Linear Algebra and its Applications, 670(??):42–67, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001313.

Shan:2021:SSR

[SWW21] Haiying Shan, Zhiyi Wang, and Feifei Wang. The smallest spectral radius of bicyclic uniform hypergraphs with a given size. Linear Algebra and its Applications, 622(??):166–188, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001348.

Shan:2022:LMU

[SWW22] Haiying Shan, Zhiyi Wang, and Feifei Wang. (α, β)- labelling method for k-uniform hypergraph and its applications. Linear Algebra and its Applications, 642(??):30–49, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000489.

Sykes:2020:CFP

[SZ20] David Sykes and Igor Zelenko. A canonical form for pairs consisting of a Hermitian form and a self-adjoint antilinear operator. Linear Algebra and its Applications, 590(??):32–61, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305324.

Simson:2021:MGR

[SZ21] Daniel Simson and Katarzyna Zaja̧c. On mesh geometries of root Coxeter orbits and mesh algorithms for corank two edge-bipartite signed graphs. Linear Algebra and its Applications, 610(??):698–765, February 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304936.

Simson:2022:AMA

[SZ22a] Daniel Simson and Katarzyna Zaja̧c. Applications of mesh algorithms and self-dual mesh geometries of root Coxeter orbits to a Horn–Sergeichuk type problem. Linear Algebra and its Applications, 632(??):79–152, January 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003384.

Sriwongsa:2022:AGI

[SZ22b] Songpon Sriwongsa and Yi Ming Zou. On automorphism groups of idempotent evolution algebras. Linear Algebra and its Applications, 641(??):143–155, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379522000532.

Shao:2023:NPB

[SZ23a] Meiyue Shao and Sizhe Zhang. A note on “Perturbation bounds for Williamson’s symplectic normal form”. Linear Algebra and its Applications, 677(??):194–199, November 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002987.  See

[IGW17].

Shi:2023:FNI

[SZ23b] Shuo Shi and Yun Zhang. Further norm inequalities for sums of operators. Linear Algebra and its Applications, 658(??):250–261, February 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004013.

Szczygielski:2021:FAC

[Szc21] Krzysztof Szczygielski. On the Floquet analysis of commutative periodic Lindbladians in finite dimension. Linear Algebra and its Applications, 609(??):176–202, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304201.

Song:2022:SRD

[SZH22] Xiaoyu Song, Baodong Zheng, and Riguang Huang. The symmetric rank and decomposition of m-order n-dimensional (n = 2, 3, 4) symmetric tensors over the binary field. Linear Algebra and its Applications, 653(??):1–32, November 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002774.

Singh:2020:NVW

[SZSMB20] Ranveer Singh, Cheng Zheng, Naomi Shaked-Monderer, and Abraham Berman. Nonsingular (vertex-weighted) block graphs. Linear Algebra and its Applications, 602(??):138–156, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030238X.

Tanaka:2023:CGF

[Tan23] Hajime Tanaka. Characterizing graphs with fully positive semidefinite Q-matrices. Linear Algebra and its Applications, 671(??):59–66, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001532.

Taranenko:2020:APP

[Tar20] Anna A. Taranenko. Algebraic properties of perfect structures. Linear Algebra and its Applications, 607(??):286–306, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303852.

Tsai:2020:TSN

[TBH20] Ming-Cheng Tsai, Meaza Bogale, and Huajun Huang. On triangular similarity of nilpotent triangular matrices. Linear Algebra and its Applications, 596(??):1–35, July 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301154.

Tian:2021:EIG

[TCC21] Gui-Xian Tian, Ya-Xue Chen, and Shu-Yu Cui. The extremal α-index of graphs with no 4-cycle and 5-cycle. Linear Algebra and its Applications, 619(??):160–175, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000811.

Tian:2022:PWP

[TD22] Xiaoyi Tian and Chunyuan Deng. On the properties of weakly parallel summable operators. Linear Algebra and its Applications, 635 (??):139–158, February 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521004250.

Teixeira:2022:RHM

[Tei22] Katiuscia B. Teixeira. Remarks on Hadamard matrices and a theorem of Macphail. Linear Algebra and its Applications, 640(??):1–5, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000118.

Taghvaee:2020:NSE

[TFT20] Fatemeh Taghvaee and Gholam Hossein Fath-Tabar. The number of the skeweigenvalues of digraphs and their relationship with optimum skew energy. Linear Algebra and its Applications, 605(??):190–205, November 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303281.

Thorne:2022:TC

[Tho22] Michael A. S. Thorne. Tipping cycles. Linear Algebra and its Applications, 646(??):43–53, August 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001203.

terHorst:2021:LMM

[tHvdM21] S. ter Horst and A. van der Merwe. Linear matrix maps for which positivity and complete positivity coincide. Linear Algebra and its Applications, 628(??):140–181, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002652.

Toyonaga:2021:EPD

[TJ21] Kenji Toyonaga and Charles R. Johnson. The effect of perturbation of an off-diagonal entry pair on the geometric multiplicity of an eigenvalue. Linear Algebra and its Applications, 615(??):112–142, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305875.

Tasissa:2021:LRM

[TL21] Abiy Tasissa and Rongjie Lai. Low-rank matrix completion in a general nonorthogonal basis. Linear Algebra and its Applications, 625(??):81–112, September 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001956.

Tian:2020:CDE

[TLC20] Gui-Xian Tian, Yuan Li, and Shu-Yu Cui. The change of distance energy of some special complete multipartite graphs due to edge deletion. Linear Algebra and its Applications, 584(??):438–457, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304173.

Tomasevic:2023:NPP

[Tom23] Mateo Tomašević. A nonlinear preserver problem on positive matrices. Linear Algebra and its Applications, 666(??):96–113, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000770.

Thanwerdas:2023:IRM

[TP23] Yann Thanwerdas and Xavier Pennec. O(n)-invariant Riemannian metrics on SPD matrices. Linear Algebra and its Applications, 661(??):163–201, March 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004360.

Tsakiris:2023:DCH

[Tsa23] Manolis C. Tsakiris. Determinantal conditions for homomorphic sensing. Linear Algebra and its Applications, 656(??):210–223, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522003512.

Tarcsay:2021:OAD

[TT21] Zsigmond Tarcsay and Tamás Titkos. Operators on antidual pairs: Generalized Schur complement. Linear Algebra and its Applications, 614(??):125–143, April 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301038.

Torres:2024:BCC

[TT24] Guilherme Simon Torres and Vilmar Trevisan. Brouwer’s Conjecture for the Cartesian product of graphs. Linear Algebra and its Applications, 685(??):66–76, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004767.

Tatakis:2023:MCB

[TTV23] Christos Tatakis, Apostolos Thoma, and Marius Vladoiu. The Markov complexity of book graphs. Linear Algebra and its Applications, 661(??):51–78, March 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004396.

Tura:2021:CET

[Tur21] Fernando C. Tura. A conjecture on the eigenvalues of threshold graphs. Linear Algebra and its Applications, 612(??):345–356, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305322.

Tisseur:2021:MME

[TV21] Françoise Tisseur and Marc Van Barel. Min-max elementwise backward error for roots of polynomials and a corresponding backward stable root finder. Linear Algebra and its Applications, 623(??):454–477, August 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520305176.

Tian:2021:FCG

[TW21] Fenglei Tian and Yiju Wang. Full characterization of graphs having certain normalized Laplacian eigenvalue of multiplicity n − 3. Linear Algebra and its Applications, 630(??):69–83, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002895.

Tian:2022:MPE

[TW22] Fenglei Tian and Yiju Wang. On the multiplicity of positive eigenvalues of a graph. Linear Algebra and its Applications, 652(??):105–124, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200266X.

Tian:2020:PSO

[TWD20] Xiaoyi Tian, Shuaijie Wang, and Chunyuan Deng. On parallel sum of operators. Linear Algebra and its Applications, 603(??):57–83, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302627.

Tian:2020:CGS

[TWW20] Fenglei Tian, Dein Wong, and Sai Wang. Characterization of graphs with some normalized Laplacian eigenvalue of multiplicity n − 3. Linear Algebra and its Applications, 606(??):127–143, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303505.

Tian:2024:FDP

[TXF24] Xiaoyi Tian, Qingxiang Xu, and Chunhong Fu. The Frobenius distances from projections to an idempotent matrix. Linear Algebra and its Applications, 688(??):21–43, May 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000545.

Terwilliger:2023:CBP

[TŽ23] Paul Terwilliger and Arjana Žitnik. Circular bidiagonal pairs. Linear Algebra and its Applications, 669(??):24–56, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001039.

Uchiyama:2021:OMM

[Uch21]  Mitsuru Uchiyama.  Op-

erator means and matrix quadratic equations. Linear Algebra and its Applications, 609(??):163–175, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304195.

Uchyama:2023:SMM

[Uch23]  Mitsuru Uchyama.  Sym-

metric matrix means. Linear Algebra and its Applications, 656(??):112–130, January 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003482.

Urure:2022:SSS

[Uru22] Ronald Ismael Quispe Urure. Symmetric and skew-symmetric polynomial identities with involution for the upper triangular matrix algebras of even order. Linear Algebra and its Applications, 641(??):98–114, May 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000453.

Urschel:2021:DTT

[UZ21] John C. Urschel and Ludmil T. Zikatanov. Discrete trace theorems and energy minimizing spring embeddings of planar graphs.

Linear Algebra and its Applications, 609(??):73–107, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304146.

VanMieghem:2024:ECS

[Van24a] P. Van Mieghem. Eigenvector components of symmetric, graph-related matrices. Linear Algebra and its Applications, 692(??):91–134, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001319.

Vannieuwenhoven:2024:CNS

[Van24b] Nick Vannieuwenhoven. The condition number of singular subspaces, revisited. Linear Algebra and its Applications, 687(??):157–180, April 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000399.

Varga:2021:RCC

[Var21]  Andreas Varga.  On re-

cursive computation of coprime factorizations of rational matrices. Linear Algebra and its Applications, 623(??):478–502, August 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300409.

Vos:2020:DAU

[VB20] Alexis De Vos and Stijn De Baerdemacker. The decomposition of an arbitrary 2^{w} × 2^{w} unitary matrix into signed permutation matrices. Linear Algebra and its Applications, 606(??):23–40, December 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303402.

Vu:2021:PEE

[VCR21] Trung Vu, Evgenia Chunikhina, and Raviv Raich. Perturbation expansions and error bounds for the truncated singular value decomposition. Linear Algebra and its Applications, 627(??):94–139, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002366.

Vermeersch:2022:TCB

[VD22] Christof Vermeersch and Bart De Moor. Two complementary block Macaulay matrix algorithms to solve multiparameter eigenvalue problems. Linear Algebra and its Applications, 654 (??):177–209, December 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002919.

vanderMerwe:2024:PIM

[vdMvSW24] Alma van der Merwe, Madelein van Straaten, and Hugo J. Woerdeman. Partial isospectrality of a matrix pencil and circularity of the c-numerical range. Linear Algebra and its Applications, 689(??):247–259, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000661.

vomEnde:2022:MP

[vED22] Frederik vom Ende and Gunther Dirr. The d-Majorization Polytope. Linear Algebra and its Applications, 649 (??):152–185, September 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001926.

Vitas:2021:IMP

[Vit21] Daniel Vitas. Images of multilinear polynomials in the algebra of finitary matrices contain trace zero matrices. Linear Algebra and its Applications, 626(??):221–233, October 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002263.

Mieghem:2024:CEG

[VJ24] Piet Van Mieghem and Ivan Jokić. Co-eigenvector graphs. Linear Algebra and its Applications, 689(??):34–59, May 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379524000521.

Vill:2022:ISD

[VMB22] Julian Vill, Mateusz Michałek, and Alexander Taveira Blomenhofer. Ideals of spaces of degenerate matrices. Linear Algebra and its Applications, 648(??):56–69, September 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001677.

Volkov:2023:AAT

[Vol23] Yury Volkov. Anticommutative algebras of the third level. Linear Algebra and its Applications, 662(??):18–38, April 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004554.

vanRensburg:2021:SRN

[vRvSTT21] D. B. Janse van Rensburg, M. van Straaten, F. Theron, and C. Trunk. Square roots of H-nonnegative matrices. Linear Algebra and its Applications, 621(??):29–49, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001038.

Vaia:2020:PJM

[VS20] Ruggero Vaia and Lidia Spadini. Persymmetric Jacobi matrices with square-integer eigenvalues and dispersionless mass–spring chains. Linear Algebra and its Applications, 585(??):164–177, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437951930429X.

Verde-Star:2021:UCA

[VS21] Luis Verde-Star. A unified construction of all the hypergeometric and basic hypergeometric families of orthogonal polynomial sequences. Linear Algebra and its Applications, 627(??):242–274, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002470.

Verde-Star:2023:LCC

[VS23] Luis Verde-Star. Linearization and connection coefficients of polynomial sequences: a matrix approach. Linear Algebra and its Applications, 672(??):195–209, September 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001726.

Vinagre:2020:CTG

[VTBC20] Cybele T. M. Vinagre, Vilmar Trevisan, Johann Bolckau, and Rodrigo Chimelli. Characterizing threshold graphs with k main signless Laplacian eigenvalues. Linear Algebra and its Applications, 602(??):33–45, October 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302160.

Valencia:2022:AAA

[VV22] Carlos E. Valencia and Ralihe R. Villagrán. Algorithmic aspects of arithmetical structures. Linear Algebra and its Applications, 640(??):191–208, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000350.

Vavilov:2020:MCE

[VZ20] Nikolai Vavilov and Zuhong Zhang. Multiple commutators of elementary subgroups: end of the line. Linear Algebra and its Applications, 599(??):1–17, August 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301798.

Wang:2021:SNF

[Wan21] Wei Wang. On the Smith normal form of walk matrices. Linear Algebra and its Applications, 612(??):30–41, March 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305565.

Wang:2023:SNF

[Wan23] Wei Wang. The Smith normal form of the walk matrix of the Dynkin graph D_{n} for n ≡ 0 (mod 4). Linear Algebra and its Applications, 671(??):121–134, August 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001568.

Wang:2024:STP

[Wan24]  Yongang Wang.  Spec-

tral Turán problem for K_{5}^{−}- free signed graphs. Linear Algebra and its Applications, 691(??):96–108, June 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001022.

Wang:2021:ESR

[WCF21] Bing Wang, Wenwen Chen, and Longfei Fang. Extremal spectral radius of K_{3,3}/K_{2,4}- minor free graphs. Linear Algebra and its Applications, 628(??):103–114, November 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002615.

Wang:2020:PCM

[WCKQ20] Dan Wang, Wei Chen, Sei Zhen Khong, and Li Qiu. On the phases of a complex matrix. Linear Algebra and its Applications, 593(??):152–179, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300458.

Wen:2020:TSI

[WCS20] Chao-Tao Wen, Xiao Shan Chen, and Hai-Wei Sun. A two-step inexact Newton– Chebyshev-like method for inverse eigenvalue problems. Linear Algebra and its Applications, 585(??):241–262, January 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304318.

Weinstein:2022:RSM

[Wei22] Madeleine Weinstein. Real symmetric matrices with partitioned eigenvalues. Linear Algebra and its Applications, 633(??):281–289, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003815.

Werner:2022:NIS

[Wer22] Nicholas J. Werner. Null ideals of subsets of matrix rings over fields. Linear Algebra and its Applications, 642(??):50–72, June 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000507.

Waki:2021:RSI

[WES21] Hayato Waki, Yoshio Ebihara, and Noboru Sebe. Reduction of SISO H-infinity output feedback control problem. Linear Algebra and its Applications, 610(??):321–378, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520304626.

Wang:2020:MAEb

[WFGT20] Long Wang, Xianwen Fang, Xianya Geng, and Fenglei Tian. On the multiplicity of an arbitrary A_{α}-eigenvalue of a connected graph. Linear Algebra and its Applications, 589(??):28–38, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930535X.

Wei:2020:RBI

[WFL20] Wei Wei, Zhimin Feng, and Shuchao Li. Relations between the inertia indices of a mixed graph and those of its underlying graph. Linear Algebra and its Applications, 588(??):19–53, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437951930494X.

Watanabe:2020:WMP

[WFSS20] Sennosuke Watanabe, Akiko Fukuda, Etsuo Segawa, and Iwao Sato. A walk on max-plus algebra. Linear Algebra and its Applications, 598(??):29–48, August 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Wu:2024:IDS

[WFY24] Yongjiang Wu, Lihua Feng, and Guihai Yu. The integrality of distance spectra of quasiabelian 2-Cayley graphs. Linear Algebra and its Applications, 681(??):97–130, January 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004032.

Wang:2020:SUBa

[WG20a] Zhi-Wen Wang and Ji-Ming Guo. A sharp upper bound of the nullity of a connected graph in terms of order and maximum degree. Linear Algebra and its Applications, 584(??):287–293, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519303908.

Wang:2020:SUBb

[WG20b] Zhi-Wen Wang and Ji-Ming Guo. Some upper bounds on the spectral radius of a graph. Linear Algebra and its Applications, 601 (??):101–112, September 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379520302184.

Wang:2020:ISS

[WH20] Dijian Wang and Yaoping Hou. Integral signed subcubic graphs. Linear Algebra and its Applications, 593(??):29–44, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300549.

Wang:2024:ERF

[WHL24] Dijian Wang, Yaoping Hou, and Deqiong Li. Extremal results for C_{3}^{−}-free signed graphs. Linear Algebra and its Applications, 681(??):47–65, January 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004044.

Wang:2023:SEG

[WHM23] Hongyu Wang, Xinmin Hou, and Yue Ma. Spectral extrema of graphs with bounded clique number and matching number. Linear Algebra and its Applications, 669(??):125–135, July 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001179.

Wiertel:2022:IRH

[Wie22]  Magdalena Wiertel.  Ir-

reducible representations of Hecke–Kiselman monoids. Linear Algebra and its Applications, 640(??):12–33, May 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522000076.

Winter:2021:GTS

[Win21] Martin Winter. Geometry and topology of symmetric point arrangements. Linear Algebra and its Applications, 611(??):1–25, February 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305425.

Wang:2024:DCN

[WJLW24] Hongxing Wang, Tianhe Jiang, Qiuli Ling, and Yimin Wei. Dual core-nilpotent decomposition and dual binary relation. Linear Algebra and its Applications, 684(??):127–157, March 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004640.

Wang:2023:SGA

[WLL23] Yongang Wang, Dan Li, and Huiqiu Lin. The signed graphs with all but at most three eigenvalues equal to −1. Linear Algebra and its Applications, 664(??):314–323, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000277.

Wu:2022:CSG

[WLT22] Qi Wu, Yong Lu, and Bit-Shun Tam. On connected signed graphs with rank equal to girth. Linear Algebra and its Applications, 651(??):90–115, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002427.

Wang:2023:PSS

[WMCQ23] Dan Wang, Xin Mao, Wei Chen, and Li Qiu. On the phases of a semi-sectorial matrix and the essential phase of a Laplacian. Linear Algebra and its Applications, 676(??):441–458, November 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002744.

Wojcik:2020:SD

[Wój20] Paweł Wójcik. On seminorm and its decomposition. Linear Algebra and its Applications, 588(??):419–427, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437951930518X.

Wojcik:2022:AON

[Wój22a]  Paweł Wójcik.  Approxi-

mate orthogonality in normed spaces and its applications II. Linear Algebra and its Applications, 632(??):258–267, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100344X.

Wojcik:2022:VUB

[Wój22b] Paweł Wójcik. Vertices of the unit ball of subspaces in L(H) and strong unicity of best approximation in L(l_{2}^{2}). Linear Algebra and its Applications, 632(??):281–293, January 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003438.

Wu:2023:GSL

[WQP23] Xiaoxia Wu, Jianguo Qian, and Haigen Peng. Graphs with second largest eigenvalue less than 1/2. Linear Algebra and its Applications, 665(??):339–353, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000460.

Wang:2022:WHM

[WS22] Wen-Huan Wang and Rui Sun. The weighted hypergraph with the maximum spectral radius. Linear Algebra and its Applications, 649(??):47–62, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001653.

Wang:2020:SPS

[WSW20] Feifei Wang, Haiying Shan, and Zhiyi Wang. On some properties of the αspectral radius of the kuniform hypergraph. Linear Algebra and its Applications, 589(??):62–79, March 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305178.

Willemsen:2024:UCA

[WvdVvdH24] Rick S. H. Willemsen, Michel van de Velden, and Wilco van den Heuvel. On the uniqueness of correspondence analysis solutions. Linear Algebra and its Applications, 690(??):162–185, June 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524000971.

Wang:2020:KPJ

[WW20] Yunjie Wang and Gang Wu. On the Kahan–Parlett–Jiang theorem — a globally optimal backward perturbation error for two-sided invariant subspaces. Linear Algebra and its Applications, 602(??):73–92, October 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302342.

Wei:2022:NIM

[WW22] Chun Wei and Fan Wen. On norms of iterations of {0, 1}-matrices. Linear Algebra and its Applications, 649(??):375–389, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001938.

Wang:2022:WMD

[WWG22] Wei Wang, Chuanming Wang, and Songlin Guo. On the walk matrix of the Dynkin graph D_{n}. Linear Algebra and its Applications, 653 (??):193–206, November 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002920.

Wang:2020:MAEa

[WWJ20] Long Wang, Liangli Wei, and Yongrong Jin. The multiplicity of an arbitrary eigenvalue of a graph in terms of cyclomatic number and number of pendant vertices. Linear Algebra and its Applications, 584(??):257–266, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519303969.

Wang:2020:BLS

[WWT20] Sai Wang, Dein Wong, and Fenglei Tian. Bounds for the largest and the smallest A_{α} eigenvalues of a graph in terms of vertex degrees. Linear Algebra and its Applications, 590(??):210–223, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305609.

Wong:2023:RBM

[WWY23] Dein Wong, Jiao Wang, and Jiahui Yin. A relation between multiplicity of nonzero eigenvalues of trees and their matching numbers. Linear Algebra and its Applications, 660(??):80–88, March 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004323.

Wu:2020:RQD

[WX20] Nianci Wu and Hua Xiang. Randomized QLP decomposition. Linear Algebra and its Applications, 599(??):18–35, August 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301762.

Wang:2024:MOR

[WXL24] Xiaolong Wang, Kejia Xu, and Li Li. Model order reduction for discrete time-delay systems based on Laguerre function expansion. Linear Algebra and its Applications, 692(??):160–184, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001393.

Wang:2024:IHM

[WXY24] Yunpeng Wang, Ao Xu, and Jizhen Yang. A t-intersecting Hilton–Milner theorem for vector spaces. Linear Algebra and its Applications, 680(??):220–238, January 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523003816.

Wang:2020:VCC

[WYF^{+}20] Long Wang, Chunyu Yan, Xianwen Fang, Xianya Geng, and Fenglei Tian. Vertexconnectivity, chromatic number, domination number, maximum degree and Laplacian eigenvalue distribution. Linear Algebra and its Applications, 607(??):307–318, December 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520303840.

Wang:2021:ECN

[WYQ21] Wei Wang, Zhidan Yan, and Jianguo Qian. Eigenvalues and chromatic number of a signed graph. Linear Algebra and its Applications, 619(??):137–145, June 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952100077X.

Wang:2023:PCD

[WZ23a] Yanna Wang and Bo Zhou. A proof of a conjecture on the distance spectral radius. Linear Algebra and its Applications, 674(??):124–154, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002082.

Wei:2023:ICI

[WZ23b] Yangjiang Wei and Yi Ming Zou. Isomorphism classes of idempotent evolution algebras. Linear Algebra and its Applications, 675(??):142–178, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002185.

Wang:2023:SRC

[WZC23] Junjie Wang, Jiaxin Zheng, and Yonglei Chen. Spectral radius conditions for fractional [a, b]-covered graphs. Linear Algebra and its Applications, 666(??):1–10, June 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000629.

Wu:2023:DAS

[WZFW23] Yongjiang Wu, Xiaoqian Zhang, Lihua Feng, and Tingzeng Wu. Distance and adjacency spectra and eigenspaces for three (di)graph lifts: a unified approach. Linear Algebra and its Applications, 672(??):147–181, September 1, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379523001672.

Wang:2021:IPU

[WZL21] Yu Wang, Jia Zhou, and Yingyu Luo. The image of polynomials on 2×2 upper triangular matrix algebras. Linear Algebra and its Applications, 610(??):560–570, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304808.

Wong:2020:MET

[WZT20] Dein Wong, Qi Zhou, and Fenglei Tian. On the multiplicity of eigenvalues of trees. Linear Algebra and its Applications, 593(??):180–187, May 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030063X.

Wang:2021:MLE

[WZW^{+}21] Xinlei Wang, Qi Zhou, Dein Wong, Chi Zhang, and Fenglei Tian. The multiplicity of Laplacian eigenvalue two in a connected graph with a perfect matching. Linear Algebra and its Applications, 609(??):152–162, January 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0024379520304183.

Xu:2023:DCM

[XBSF23] Wei-Ru Xu, Natália Bebiano, Qian-Yu Shu, and Ting-Ting Feng. A divide-and-conquer method for constructing a pseudo-Jacobi matrix from mixed given data. Linear Algebra and its Applications, 674(??):256–281, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002094.

Xu:2023:ESR

[XHW23] Shao-Han Xu, Fu-Tao Hu, and Yi Wang. On extremal spectral radius of blow-up uniform hypergraphs. Linear Algebra and its Applications, 667(??):71–87, June 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000939.

Xie:2021:TCM

[Xie21] Yangxinyu Xie. On 2 × 2 tropical commuting matrices. Linear Algebra and its Applications, 620(??):92–108, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000835.

Xu:2023:MRD

[XJLL23] Kang-Li Xu, Yao-Lin Jiang, Zhen Li, and Li Li. Model reduction of discrete timedelay systems based on Charlier polynomials and highorder Krylov subspaces. Linear Algebra and its Applications, 661(??):222–246, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004542.

Xue:2020:SLD

[XLS20] Jie Xue, Huiqiu Lin, and Jinlong Shu. On the second largest distance eigenvalue of a block graph. Linear Algebra and its Applications, 591(??):284–298, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300185.

Xue:2020:DNL

[XLYS20] Jie Xue, Ruifang Liu, Guanglong Yu, and Jinlong Shu. Domination number and Laplacian eigenvalue of trees. Linear Algebra and its Applications, 592(??):210–227, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300355.

Xia:2023:ASG

[XTLK23] Tian-Tian Xia, Ying-Ying Tan, Xiaoye Liang, and Jack H. Koolen. On association schemes generated by a relation or an idempotent. Linear Algebra and its Applications, 670(??):1–18, August 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523001271.

Xiao:2021:EAS

[XW21] Peng Xiao and Ligong Wang. The effect on the adjacency and signless Laplacian spectral radii of uniform hypergraphs by grafting edges. Linear Algebra and its Applications, 610(??):591–607, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304882.

Xu:2020:MEM

[XWT20] Feng Xu, Dein Wong, and Fenglei Tian. On the multiplicity of α as an eigenvalue of the A_{α} matrix of a graph in terms of the number of pendant vertices. Linear Algebra and its Applications, 594(??):193–204, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0024379520300975.

Xie:2020:TLA

[XX20] Shui-Lian Xie and Hong-Ru Xu. A two-level additive Schwarz method for a kind of tensor complementarity problem. Linear Algebra and its Applications, 584(??):394–408, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304148.

Xi:2020:CCM

[XY20a] Changchang Xi and Shujun Yin. Cellularity of centrosymmetric matrix algebras and Frobenius extensions. Linear Algebra and its Applications, 590(??):317–329, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300069.

Xu:2020:HPH

[XY20b] Qingxiang Xu and Guanjie Yan. Harmonious projections and Halmos’ two projections theorem for Hilbert C^{⋆}-module operators. Linear Algebra and its Applications, 601(??):265–284, September 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

Xu:2020:LSD

[XY20c] Yuyang Xu and Jianfeng Yao. On Laplacian spectrum of dendrite trees. Linear Algebra and its Applications, 591(??):215–234, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300045.

Xi:2021:SCM

[XZ21] Changchang Xi and Jinbi Zhang. Structure of centralizer matrix algebras. Linear Algebra and its Applications, 622(??):215–249, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521001385.

Xu:2023:PCD

[XZ23] Leyou Xu and Bo Zhou. Proof of a conjecture on distribution of Laplacian eigenvalues and diameter, and beyond. Linear Algebra and its Applications, 678(??):92–106, December 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523003178.

Xu:2020:CUG

[XZWT20] Feng Xu, Qi Zhou, Dein Wong, and Fenglei Tian. Complex unit gain graphs of rank 2. Linear Algebra and its Applications, 597(??):155–169, July 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301518.

Yamazaki:2021:IAT

[Yam21]  Takeaki Yamazaki.  The

induced Aluthge transformations. Linear Algebra and its Applications, 628(??):1–28, November 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002494.

Yang:2021:ACE

[Yan21] Zhao Yang. Accurate computations of eigenvalues of quasi-Cauchy–Vandermonde matrices. Linear Algebra and its Applications, 622(??):268–293, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001464.

Yang:2022:TMA

[Yan22] Guiyu Yang. Triangular matrix algebras over affine quasihereditary algebras. Linear Algebra and its Applications, 633(??):213–226, January 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003748.

Yasumura:2023:UEG

[Yas23] Felipe Yukihide Yasumura. Universal enveloping of a graded Lie algebra. Linear Algebra and its Applications, 674(??):208–229, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002112.

Yazawa:2021:EHM

[Yaz21] Akiko Yazawa. The eigenvalues of the Hessian matrices of the generating functions for trees with k components. Linear Algebra and its Applications, 631(??):48–66, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003141.

Yang:2024:SMD

[YBW24] Xiuwen Yang, Hajo Broersma, and Ligong Wang. The A_{α} spectral moments of digraphs with a given dichromatic number. Linear Algebra and its Applications, 685(??):77–103, March 15, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004810.

Yuan:2021:FTM

[YD21] Yan Yuan and Zhibin Du. The first two maximum ABC spectral radii of bicyclic graphs. Linear Algebra and its Applications, 615(??):28–41, April 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520305991.

Yang:2020:IEP

[YGW20] Lu Yang, Yongxia Guo, and Guangsheng Wei. Inverse eigenvalue problems for a damped Stieltjes string with mixed data. Linear Algebra and its Applications, 601(??):55–71, September 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302172.

Yang:2021:ORG

[YK21] Jae Young Yang and Jack H. Koolen. On the order of regular graphs with fixed second largest eigenvalue. Linear Algebra and its Applications, 610(??):29–39, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304420.

Yao:2021:NBT

[YLCB21] Hongmei Yao, Zhen Li, Lixiang Chen, and Changjiang Bu. New Brualdi-type eigenvalue inclusion sets for tensors. Linear Algebra and its Applications, 614(??):343–355, April 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303372.

Yu:2021:NGI

[YLZH21] Qian Yu, Fenjin Liu, Hao Zhang, and Ziling Heng. Note on graphs with irreducible characteristic polynomials. Linear Algebra and its Applications, 629 (??):72–86, November 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002780.

Yoshida:2021:MDS

[Yos21] Yuuya Yoshida. Maximum dimension of subspaces with no product basis. Linear Algebra and its Applications, 620(??):228–241, July 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521000926.

Yamamoto:2021:CEE

[YOV21] Yusaku Yamamoto, Gabriel Okša, and Marián Vajteršic. On convergence to eigenvalues and eigenvectors in the block-Jacobi EVD algorithm with dynamic ordering. Linear Algebra and its Applications, 622(??):19–45, August 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001312.

Yu:2020:ERN

[Yu20] Luyan Yu. Extreme rays of the l^{∞}-nearest ultrametric tropical polytope. Linear Algebra and its Applications, 587(??):23–44, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304665.

Yan:2022:DCM

[YW22] Zhidan Yan and Wei Wang. Disproof of a conjecture on the main spectrum of generalized Bethe trees. Linear Algebra and its Applications, 639(??):18–23, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.

Yang:2023:LGT

[YW23] Jing Yang and Long Wang. Line graphs of trees with the largest eigenvalue multiplicity. Linear Algebra and its Applications, 676 (??):56–65, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002641.

Yuan:2021:LEM

[YWF21] Bo-Jun Yuan, Yi Wang, and Yi-Zheng Fan. On the largest eigenvalue of a mixed graph with partial orientation. Linear Algebra and its Applications, 627(??):150–161, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100238X.

Yang:2021:RAL

[YY21] Lin Yang and Sheng-Liang Yang. Riordan arrays, Łukasiewicz paths and Narayana polynomials. Linear Algebra and its Applications, 622 (??):1–18, August 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379521001099.

Yang:2023:SHR

[YY23] Lin Yang and Sheng-Liang Yang. The skew halves of a Riordan array. Linear Algebra and its Applications, 664(??):216–235, May 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523000113.

Yu:2021:SRM

[YYS^{+}21] Guanglong Yu, Chao Yan, Lin Sun, Yarong Wu, and Hailiang Zhang. Spectral radius and matching number of the unicyclic hypergraph. Linear Algebra and its Applications, 610(??):571–590, February 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952030481X.

Yang:2020:HTI

[YZ20] Chaojun Yang and Fuzhen Zhang. Harnack type inequalities for matrices in majorization. Linear Algebra and its Applications, 588(??):196–209, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519305014.

Yang:2024:HDM

[YZY24] Lin Yang, Yu-Yuan Zhang, and Sheng-Liang Yang. The halves of Delannoy matrix and Chung–Feller properties of the m-Schröder paths. Linear Algebra and its Applications, 685(??):138–161, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004780.

Zakharov:2021:ZFE

[Zak21]  Dmitry Zakharov.  Zeta

functions of edge-free quotients of graphs. Linear Algebra and its Applications, 629(??):40–71, November 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002676.

Zalar:2022:TMP

[Zal22] Aljaž Zalar. The truncated moment problem on the union of parallel lines. Linear Algebra and its Applications, 649(??):186–239, September 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001951.

Zamani:2023:WHS

[Zam23] Ali Zamani. The weighted Hilbert–Schmidt numerical radius. Linear Algebra and its Applications, 675 (??):225–243, October 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002495.

Zuo:2021:FCC

[ZBCI21] Kezheng Zuo, Oskar Maria Baksalary, and Dragana Cvetković-Ilić. Further characterizations of the co-EP matrices. Linear Algebra and its Applications, 616(??):66–83, May 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520306029.

Zniyed:2020:HOT

[ZBdAF20] Yassine Zniyed, Rémy Boyer, André L. F. de Almeida, and Gérard Favier. High-order tensor estimation via trains of coupled third-order CP and Tucker decompositions. Linear Algebra and its Applications, 588(??):304–337, March 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519304744.

Zhou:2020:SSR

[ZBWL20] Qiannan Zhou, Hajo Broersma, Ligong Wang, and Yong Lu. On sufficient spectral radius conditions for Hamiltonicity of k-connected graphs. Linear Algebra and its Applications, 604(??):129–145, November 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520303037.

Zhu:2023:NRM

[ZC23] Shouchong Zhu and Wei Chen. On numerical ranges of matrices with Perron–Frobenius properties and some negative entries. Linear Algebra and its Applications, 661(??):20–34, March 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004347.

Zemanek:2022:RSD

[Zem22]  Petr Zemánek.  Resolvent

and spectrum for discrete symplectic systems in the limit point case. Linear Algebra and its Applications, 634(??):179–209, February 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003955.

Zhou:2020:HCP

[ZFZ20] Anwa Zhou, Jinyan Fan, and Guangming Zhou. Hermitian completely positive matrices. Linear Algebra and its Applications, 604 (??):187–209, November 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302986.

Zheng:2021:EDT

[ZG21] Junling Zheng and Hanpeng Gao. The extension dimension of triangular matrix algebras. Linear Algebra and its Applications, 624(??):44–52, September 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521001580.

Zhang:2022:MLS

[ZG22] Rong Zhang and Shu-Guang Guo. Maxima of the Laplacian spectral radius of (minimally) 2-connected graphs with fixed size. Linear Algebra and its Applications, 651(??):390–406, October 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522002506.

Zhang:2021:GFD

[Zha21] Xiaoling Zhang. Graphs with few distinct D-eigenvalues determined by their Dspectra. Linear Algebra and its Applications, 628 (??):42–55, November 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002585.

Zhan:2022:FDT

[Zha22] Hanmeng Zhan. Factoring discrete-time quantum walks on distance regular graphs into continuous-time quantum walks. Linear Algebra and its Applications, 648 (??):88–103, September 1, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522001689.

Zhang:2023:PMF

[Zha23] Xue Zhang. Periodic multivariate formal power series. Linear Algebra and its Applications, 657(??):32–49, January 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003822.

Zhang:2024:FTE

[Zha24a] Shengtong Zhang. On the first two eigenvalues of regular graphs. Linear Algebra and its Applications, 686(??):102–110, April 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379524000028.

Zhang:2024:SPR

[Zha24b] Yueheng Zhang. On the stability of the principal ratio. Linear Algebra and its Applications, 685(??):104–124, March 15, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523004779.

Zheng:2020:SCS

[Zhe20]  Ya-Nan Zheng.  Simplify-

ing the computation of the spectrum of the complete kuniform hypergraph. Linear Algebra and its Applications, 595(??):145–156, June 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520301142.

Zheng:2021:CPC

[Zhe21] Ya-Nan Zheng. The characteristic polynomial of the complete 3-uniform hypergraph. Linear Algebra and its Applications, 627(??):275–286, October 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002512.

Zhao:2021:SLS

[ZHG21] Yanhua Zhao, Xueyi Huang, and Hangtian Guo. The signless Laplacian spectral radius of graphs with no intersecting triangles. Linear Algebra and its Applications, 618(??):12–21, June 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521000379.

Zheng:2020:CCA

[ZHSX20] Baodong Zheng, Riguang Huang, Xiaoyu Song, and Jinli Xu. On Comon’s conjecture over arbitrary fields. Linear Algebra and its Applications, 587(??):228–242, February 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304860.

Zhu:2021:VWB

[Zhu21] Lin Zhu. A vertex weighted Bartholdi zeta function for a graph. Linear Algebra and its Applications, 621(??):254–271, July 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952100121X.

Zhao:2021:SRP

[ZHW21] Yanhua Zhao, Xueyi Huang, and Zhiwen Wang. The A_{α}- spectral radius and perfect matchings of graphs. Linear Algebra and its Applications, 631(??):143–155, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521003256.

Zheng:2020:SGB

[ZJCA20] Yanpeng Zheng, Xiaoyu Jiang, Xiaoting Chen, and Fawaz Alsaadi. On some generalizations of the Brunn– Minkowski inequality. Linear Algebra and its Applications, 586(??):103–110, February 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519304628.

Zaitsev:2021:MES

[ZK21] Vasilii Zaitsev and Inna Kim. Matrix eigenvalue spectrum assignment for linear control systems by static output feedback. Linear Algebra and its Applications, 613(??):115–150, March 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520305838.

Zhou:2020:EPS

[ZKLS20] Yacong Zhou, Liying Kang, Lele Liu, and Erfang Shan. Extremal problems for the p-spectral radius of Berge hypergraphs. Linear Algebra and its Applications, 600(??):22–39, September 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520301865.

Zhang:2022:MIG

[ZL22] Yanting Zhang and Zhenzhen Lou. Maxima of the Q-index: Graphs with no K_{1,t}-minor. Linear Algebra and its Applications, 653 (??):135–150, November 15, 2022. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002816.

Zhang:2023:RBS

[ZL23a] Yanjun Zhang and Hanyu Li. Randomized block subsampling Kaczmarz–Motzkin method. Linear Algebra and its Applications, 667 (??):133–150, June 15, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379523000800.

Zhao:2023:CAS

[ZL23b] Jing Zhao and Huiqing Liu. On a conjecture about the spectral radius of block graphs. Linear Algebra and its Applications, 659(??):1– 9, February 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379522004177.

Zhou:2023:TSC

[ZL23c] Sizhong Zhou and Hongxia Liu. Two sufficient conditions for odd [1, b ]- factors in graphs. Linear Algebra and its Applications, 661(??):149–162, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522004529.

Zhang:2020:UHF

[ZLG20] Jianbin Zhang, Jianping Li, and Haiyan Guo. Uniform hypergraphs with the first two smallest spectral radii. Linear Algebra and its Applications, 594(??):71–80, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300835.

Zhang:2022:DSF

[ZLLZ22] Yuke Zhang, Huiqiu Lin, Qinghai Liu, and Jinfeng Zheng. Distance spectrum, 1-factor and vertex-disjoint cycles. Linear Algebra and its Applications, 654(??):10–27, December 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522003184.

Zarai:2021:ESC

[ZM21] Yoram Zarai and Michael Margaliot. On the exponent of several classes of oscillatory matrices. Linear Algebra and its Applications, 608(??):363–386, January 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520304432.

Zerbo:2023:LPQ

[ZMP23] Santiago Gonzalez Zerbo, Alejandra Maestripieri, and Francisco Martínez Pería. Linear pencils and quadratic programming problems with a quadratic constraint. Linear Algebra and its Applications, 665(??):12–35, May 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523000356.

Zhao:2022:LPR

[ZRQK22] Di Zhao, Axel Ringh, Li Qiu, and Sei Zhen Khong. Low phase-rank approximation. Linear Algebra and its Applications, 639(??):177–204, April 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522000040.

Zadeh:2022:IDV

[ZS22a] Pourya Habib Zadeh and Suvrit Sra. Introducing discrepancy values of matrices with application to bounding norms of commutators. Linear Algebra and its Applications, 651(??):359–384, October 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200249X.

Zhu:2022:ENR

[ZS22b] Li Zhu and Huaqing Sun. On essential numerical ranges and essential spectra of Hamiltonian systems with one singular endpoint. Linear Algebra and its Applications, 645(??):9–29, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S002437952200101X.

Zhang:2023:FCL

[ZT23] He Zhang and Chunming Tang. Further constructions of large cyclic subspace codes via Sidon spaces. Linear Algebra and its Applications, 661(??):106–115, March 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522004384.

Zhang:2024:MOD

[ZTX24] Xiaofeng Zhang, Xiaoyi Tian, and Qingxiang Xu. The maximum operator distance from an idempotent to the set of projections. Linear Algebra and its Applications, 692(??):62–70, July 1, 2024. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379524001320.

Zhou:2021:SRC

[ZvD21] Jiang Zhou and Edwin R. van Dam. Spectral radius and clique partitions of graphs. Linear Algebra and its Applications, 630(??):84–94, December 1, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379521002901.

Zhang:2023:MED

[ZvD23] Yuke Zhang and Edwin R. van Dam. Matching extension and distance spectral radius. Linear Algebra and its Applications, 674(??):244–255, October 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0024379523002148.

Zhang:2020:RJM

[ZW20] Jing Zhang and Guangsheng Wei. Reconstruction of Jacobi matrices with mixed spectral data. Linear Algebra and its Applications, 591(??):44–60, April 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520300033.

Zamani:2021:AGN

[ZW21] Ali Zamani and Paweł Wójcik. Another generalization of the numerical radius for Hilbert space operators. Linear Algebra and its Applications, 609(??):114–128, January 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520304110.

Zhou:2022:GEM

[ZW22] Qi Zhou and Dein Wong. Graphs with eigenvalue −1 of multiplicity 2θ(G) + ρ(G) − 1. Linear Algebra and its Applications, 645(??):137–152, July 15, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379522001112.

Zhai:2020:STP

[ZWF20] Mingqing Zhai, Bing Wang, and Longfei Fang. The spectral Turán problem about graphs with no 6-cycle. Linear Algebra and its Applications, 590(??):22–31, April 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379519305464.

Zhou:2020:SLS

[ZWL20a] Qiannan Zhou, Ligong Wang, and Yong Lu. Signless Laplacian spectral conditions for Hamilton-connected graphs with large minimum degree. Linear Algebra and its Applications, 592(??):48–64, May 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520300252.

Zhou:2020:SCH

[ZWL20b] Qiannan Zhou, Ligong Wang, and Yong Lu. Sufficient conditions for Hamilton-connected graphs in terms of (signless Laplacian) spectral radius. Linear Algebra and its Applications, 594(??):205–225, June 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002437952030094X.

Zhou:2021:CGO

[ZWT21] Qi Zhou, Dein Wong, and Bit-Shun Tam. On connected graphs of order n with girth g and nullity n − g. Linear Algebra and its Applications, 630(??):56–68, December 1, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521002639.

Zhang:2023:MSR

[ZWW23] Wenqian Zhang, Jianfeng Wang, and Weifan Wang. The matchings and spectral radius of graphs involving minimum degree. Linear Algebra and its Applications, 668(??):149–160, July 1, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523001076.

Zhang:2020:VHP

[ZX20] Lin Zhang and Hua Xiang. A variant of Horn’s problem and the derivative principle. Linear Algebra and its Applications, 584(??):79–106, January 1, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379519303945.

Zhai:2020:SLS

[ZXL20] Mingqing Zhai, Jie Xue, and Zhenzhen Lou. The signless Laplacian spectral radius of graphs with a prescribed number of edges. Linear Algebra and its Applications, 603(??):154–165, October 15, 2020. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379520302822.

Zhou:2023:RBT

[ZXZW23] Qi Zhou, Feng Xu, Yuanshuai Zhang, and Dein Wong. Relation between the trace norm of an oriented graph and its rank. Linear Algebra and its Applications, 675(??):244–255, October 15, 2023. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523002483.

Zhai:2024:SEF

[ZY24] Yanni Zhai and Xiying Yuan. Spectral extrema of {K_{k+1}, L_{s}}- free graphs. Linear Algebra and its Applications, 682(??):309–322, February 1, 2024. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379523004299.

Zhang:2020:DRM

[ZZ20] Liming Zhang and Xiqiang Zhao. q-double Riordan matrices. Linear Algebra and its Applications, 603 (??):212–225, October 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379520302652.

Zhang:2021:LBS

[ZZ21] Peng-Li Zhang and Xiao-Dong Zhang. Lower bounds for the A_{α}-spectral radius of uniform hypergraphs. Linear Algebra and its Applications, 631(??):308–327, December 15, 2021. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379521003189.

Zhang:2023:HDR

[ZZJ23] Liming Zhang, Xiqiang Zhao, and Hua Jiang. The halves of a 3-dimensional Riordan array. Linear Algebra and its Applications, 676 (??):374–394, November 1, 2023. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0024379523002823.

Zhang:2022:NME

[ZZW22] Yuanshuai Zhang, Qi Zhou, and Dein Wong. A note on the multiplicities of the eigenvalues of a tree. Linear Algebra and its Applications, 652(??):97–104, November 1, 2022. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379522002646.